### v1

# Quantum Polymorphisms and Commutativity Gadgets

Abstract.
We introduce the concept of quantum polymorphisms to the complexity theory of non-local games. We use this notion to give a full characterisation of the existence of commutativity gadgets for relational structures, introduced by Ji as a method for achieving quantum soundness of classical CSP reductions. Prior to our work, a classification was only known in the Boolean case [Culf–Mastel, STOC’25]. As an application of our framework, we prove that the entangled CSP parameterised by odd cycles is undecidable. Furthermore, we establish a quantum version of Galois connection for entangled CSPs in the case of non-oracular quantum homomorphisms.
1. Introduction
In 1971, Cook’s introduction of $\mathsf{NP}$-completeness [cook1971complexity] provided a way to mathematically assess the inherent hardness of computational problems. Boolean satisfiability, graph colouring, linear equations over finite fields, and many other computatational problems of a seemingly distinct nature systematically began to fall into either the polynomial-time solvable or the $\mathsf{NP}$-complete buckets [Karp72], while intermediate complexity turned out to be a rarer phenomenon [ladner1975structure]. This fact soon sparked the hypothesis that the broad, natural class of computational problems involving satisfaction of constraints over finite alphabets—collectively known as (finite-domain) CSPs and encompassing the aforementioned problems and many others—could never exhibit the unnatural intermediate behaviour. $\operatorname{P}$ vs. $\mathsf{NP}$-complete’ dichotomies for fragments of CSPs started to emerge, primary examples being Schaefer’s [Schaefer78:stoc] and Hell–Nešetřil’s [HellN90] dichotomies for Boolean and graph CSPs, respectively. However, a unified framework for approaching the CSP Dichotomy Conjecture [Feder98:monotone] only emerged at the turn of the century, with the foundational line of work initiated by Jeavons, Cohen, and Gyssens that became known as the (universal-)algebraic approach to CSPs [Jeavons97:closure, Jeavons98:algebraic]. By lifting CSP theory from the realm of Turing machines and reductions to that of algebra, the study of polymorphism clones—universal-algebraic objects capturing high-order invariance properties of CSP solutions—provided a clearer viewpoint on the CSP complexity landscape, and it eventually culminated with the positive resolution of the conjecture by Bulatov [Bulatov17:focs] and, independently, Zhuk [Zhuk17_FOCS, Zhuk20:jacm] in 2017.
Seven years before the publication of Cook’s result, the work of Bell [bell1964einstein] originating from the Einstein–Podolsky–Rosen paradox [einstein1935can] demonstrated that quantum systems can display correlations unattainable within classical physics, thus ruling out local hidden-variable theories as realistic explanations for quantum mechanics. In its mathematical essence, Bell’s result can be formulated as a property of CSPs: There exist specific CSP languages for which the classical value and the entangled value can differ, where the latter is defined as the optimal success probability in a multiplayer interactive proof of the satisfiability of the CSP instance where the players have access to quantum resources. A primary question arising from Bell’s theorem concerns the computational complexity of determining the entangled value of a CSP. In particular, it is known that determining whether an instance of SAT is perfectly quantum satisfiable is undecidable (in fact, even the gap version of the problem is known to be undecidable, as a consequence of the celebrated $\operatorname{MIP^{*}=RE}$ theorem [ji2021mip]). On the other hand, examples of tractable CSP languages for which quantum and classical satisfiability coincide, which implies polynomial-time solvability of the corresponding entangled CSPs, are known [atserias_generalized_2019, ciardo_quantum_minion, bulatov2025satisfiability].
Nevertheless, unlike in the classical case, the complexity landscape of entangled CSPs remains largely unexplored. A natural strategy to shed some light on it is to transfer the algebraic machinery that allowed proving the (classical) CSP Dichotomy to the quantum setting. In particular, for classical CSPs, the concept of polymorphisms turned out to be the right tool for exploring CSP complexity at the algebraic level. The root of this phenomenon is a Galois connection between relational clones (expressing the CSP language) and operation clones (expressing the family of polymorphisms). Via this connection, interreducibility between CSPs associated with different languages can be explained via inclusion (or weaker forms of the latter, in particular, so called height-1 clone homomorphisms) between their polymorphism clones.
To what extent can this machinery by transferred to the setting of entangled CSPs, and thus provide some understanding of their complexity and interreducibility? This question was first addressed at this level of generality by Ji in [ji2013binary]. In particular, that work identified a substantial obstacle against the prospect of quantising all types of algebraic CSP reductions known as gadgets. Ultimately, the obstacle is caused by the non-classical phenomenon of contextuality: Replacing constraints by gadgets increases the size of the contexts—which, in general, destroys simultaneous measurability. In the same paper, Ji provided a method to solve this issue for specific CSP languages, via the new notion of commutativity gadgets. On a high level, these are specific types of gadgets having no classical counterpart, that can be plugged into classical gadgets to enforce simultaneous measurability of the final instance resulting from the reduction. For example, by showing that a specific triangular prism construction yields a commutativity gadgets for the 3-clique, Ji showed that the classical polynomial-time reduction from 3-SAT to 3-colouring can be lifted to a reduction between the corresponding entangled CSPs, thus proving undecidability of quantum 3-colouring.
Later works generalised this idea to more general types of CSPs. In particular, commutativity gadgets for certain Boolean CSPs were constructed in [atserias_generalized_2019]. In a subsequent recent breakthrough [culf2024re], Culf and Mastel gave a complete characterisation of Boolean CSPs admitting commutativity gadgets. In the same work, the authors identified a sufficient property of CSP languages—which they called two-variable (non-)falsifiability—guaranteeing the existence of commutativity gadgets for the corresponding CSPs, and thus allows quantising classical reductions. Very recently, Culf, de Bruyn, Vernooij, and Zeman [Zeman] gave a different, necessary condition for the existence of commutativity gadgets, by showing that CSPs with non-classical quantum endomorphisms do not admit commutativity gadgets.
Despite this recent progress, a full characterisation of CSPs exhibiting commutativity gadgets is not available. In particular, [Zeman] identified odd cycles as a challenging open case for which the existence or non-existence of commutativity gadgets cannot be established via available methods.
Contributions
We give a complete characterisation of CSPs exhibiting commutativity gadgets. To that end, we introduce the notion of quantum polymorphisms, which lift classical polymorphisms to the setting of entangled CSPs. We show that quantum polymorphisms completely capture the existence of commutativity gadgets, by proving that a commutativity gadget exists if, and only if, the minion of quantum polymorphisms coincides with the quantum closure of the clone of standard polymorphisms. Through this characterisation, we prove that all odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable. This answers the question raised in [Zeman]. In addition, we use quantum polymorphisms to build commutativity gadgets for Boolean structures in a unified manner, thus giving a different proof of the result of [culf_re-completeness_2025]. Our characterisation works for both oracular and non-oracular CSPs. In the non-oracular setting, we additionally establish a quantum version of Galois connection that parallels the corresponding classical connection.
Structure of the paper
In Section 2, we give formal definitions of classical and entangled CSPs, and we list some basic results about quantum homomorphisms. Section 3 presents the main definition of this work—quantum polymorphims—as well as their use to characterise the existence of oracular commutativity gadgets for entangled CSPs. In Section 4 we show that the quantum polymorphisms of cliques are non-contextual, which gives a different proof of the undecidability of the corresponding entangled CSPs due to [ji2013binary, Zeman]. In section 5, we prove a general result on the consequence of contextuality in quantum homomorphisms, by showing that it causes the emergence of a specific non-orthogonality pattern in the underlying structures. We make use of this pattern in section 6, to prove that odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable. In Section 7, we address the non-oracular case, by showing that a similar characterisation to the one in Section 3 captures commutativity gadgets in the non-oracular setting. In addition, we establish therein a quantum version of Galois connection for the non-oracular quantum analogue of pp-definitions. Finally, in Section 8, we investigate the existence of commutativity gadgets for structures with a 2-element domain. We obtain that structures that do not admit the ternary majority as a polymorphism must have a commutativity gadget, and that apart from trivial cases, structures that do have the majority operation as a polymorphism do not have a commutativity gadget.
2. Preliminaries
2.1. Relational structures
For $\ell\in\mathbb{N}$, $\ell$] is the set $\{1,2,\dots,\ell\}$. A signature $\sigma$ is a finite set of relation symbols $R$, each with its arity $\operatorname{ar}(R)\in\mathbb{N}$. A relational structure $\mathbb{X}$ with signature $\sigma$ (in short, a $\sigma$-structure or simply a structure when $\sigma$ is implicit) consists of a finite set $X$ and a relation $R^{\mathbb{X}}\subseteq X^{\operatorname{ar}(R)}$ for each symbol $R\in\sigma$. All relational structures in this work shall be implicitly assumed to be finite. A (classical) homomorphism $f:\mathbb{X}\to\mathbb{Y}$ between two relational structures $\mathbb{X}$ and $\mathbb{Y}$ on a common signature $\sigma$ is a map from the domain $X$ of $\mathbb{X}$ to the domain $Y$ of $\mathbb{Y}$ that preserves all relations; i.e., $f(\mathbf{x})\in R^{\mathbb{Y}}$ for each symbol $R\in\sigma$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$ (where $f(\mathbf{x})$ is the entrywise application of $f$ the the entries of $\mathbf{x}$). We denote the existence of a classical homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ by the notation $\mathbb{X}\to\mathbb{Y}$. Given a $\sigma$-structure $\mathbb{A}$, a positive integer $r\in\mathbb{N}$, and a set $S\subseteq A^{r}$, we define the structure $\mathbb{A}_{S}$ as follows. The signature of $\mathbb{A}_{S}$ contains a unique $r$-ary symbol $R$ (which does not belong to $\sigma$), and the interpretation of $R$ in $\mathbb{A}_{S}$ is $R^{\mathbb{A}_{S}}=S$. For typographical reasons, we shall sometimes denote $\mathbb{A}_{S}$ by $(A;S)$. For a digraph $\mathbb{X}$, we let $\mathbb{X}^{u}$ be the undirected graph on the same vertex set of $\mathbb{X}$, having an undirected edge $\{x,y\}$ whenever $(x,y)$ or $(y,x)$ form an edge in $\mathbb{X}$. We let the Gaifman graph of a relational structure $\mathbb{X}$ be the loopless directed graph $\operatorname{Gaif}(\mathbb{X})$ on vertex set $V(\operatorname{Gaif}(\mathbb{X}))=X$ and edge set $E(\operatorname{Gaif}(\mathbb{X}))$ containing a pair $(x,y)$ if, and only if, $x$ and $y$ appear together in some tuple in some relation of $\mathbb{X}$. Given a $\sigma$-structure $\mathbb{Y}$, the constraint satisfaction problem parameterised by $\mathbb{Y}$ (in short, $\operatorname{CSP}(\mathbb{Y})$) is the following computational problem: [^1] This is the decision version of $\operatorname{CSP}(\mathbb{Y})$. In the search version, the problem is to find an explicit homomorphism $f:\mathbb{X}\to\mathbb{Y}$ for any input $\mathbb{X}$ such that $\mathbb{X}\to\mathbb{Y}$. The two versions of $\operatorname{CSP}$ are equivalent up to polynomial-time reductions [bulatov2005classifying]. Given as input a $\sigma$-structure $\mathbb{X}$, output $\mathrm{YES}$ if $\mathbb{X}\to\mathbb{Y}$, and $\mathrm{NO}$ if $\mathbb{X}\not\to\mathbb{Y}$. For $\ell\in\mathbb{N}$, we denote by $\mathbb{Y}^{\ell}$ the $\ell$-th direct power of $\mathbb{Y}$; i.e., $\mathbb{Y}^{\ell}$ is the $\sigma$-structure whose domain is $Y^{\ell}$ and whose relations are defined as follows: Given $R\in\sigma$ and any $\ell\times\operatorname{ar}(R)$ matrix $M$ whose rows are tuples in $R^{\mathbb{Y}}$, the columns of $M$ form a tuple in $R^{\mathbb{Y}^{\ell}}$. A polymorphism of $\mathbb{Y}$ (of arity $\ell$) is a homomorphism from $\mathbb{Y}^{\ell}$ to $\mathbb{Y}$. By $\operatorname{Pol}(\mathbb{Y})$ (resp. $\operatorname{Pol}^{(\ell)}(\mathbb{Y})$) we denote the set of all polymorphisms (resp. all polymorphisms of arity $\ell$) of $\mathbb{Y}$. The core of a $\sigma$-structure $\mathbb{X}$—in symbols, $\operatorname{core}(\mathbb{X})$—is a minimal (under inclusion) induced substructure of $\mathbb{X}$ that is homomorphically equivalent to $\mathbb{X}$. It is easy to check that $\operatorname{core}(\mathbb{X})$ is unique up to isomorphism, and any endomorphism is an automorphism (see e.g. [HellNesetrilCore]).
The primitive positive (pp) fragment of first-order logic consists of the formulas that are built out of atomic formulas using conjunctions and existential quantifiers. If $\mathbb{A}$ is a relational structure and $\varphi(x_{1},\dots,x_{n})$ is a primitive positive formula in the signature of $\mathbb{A}$ with free variables $x_{1},\dots,x_{n}$, then the relation defined by $\varphi$ is the set of $n$-tuples that satisfy the formula in $\mathbb{A}$. If $R\subseteq A^{n}$ is a relation, we say that it is pp-definable over $\mathbb{A}$ if there exists a pp-formula defining it. A classical fact about pp-formulas is the following duality with gadgets: every pp-formula can be understood as a relational structure $\mathbb{X}_{\varphi}$ with distinguished vertices $x_{1},\dots,x_{n}$ such that for every structure $\mathbb{A}$, $(a_{1},\dots,a_{n})$ is in the relation defined by $\varphi$ in $\mathbb{A}$ if, and only if, there exists a homomorphism $\mathbb{X}_{\varphi}\to\mathbb{A}$ that maps $x_{i}$ to $a_{i}$.
2.2. Quantum CSPs
Let $H$ be a non-trivial finite-dimensional Hilbert space. We say that a linear endomorphism $p$ of $H$ is an orthogonal projector (or simply a projector) if it is self-adjoint and idempotent (i.e., if $p^{2}=p=p^{*}$). We denote by $\operatorname{id}_{H}$ (or simply by $\operatorname{id}$ when the Hilbert space is clear) the identity projector over $H$. A projective measurement (PVM) over a Hilbert space $H$ and a set $B$ and is a family $(Q_{b})_{b\in B}$ of projectors on $H$ with the property that $\sum_{b\in B}Q_{b}=\operatorname{id}$. A quantum function from $A$ to $B$ is a family of PVMs $(Q_{a,b})_{b\in B}$, one PVM for each $a\in A$. We shall denote this by the notation $Q\colon A\to_{q}B$ or $Q\colon A\to_{H}B$ if we want to emphasize the Hilbert space. We say that two PVMs $(Q_{b})_{b\in B}$ and $(Q^{\prime}_{b^{\prime}})_{b^{\prime}\in B^{\prime}}$ over the same Hilbert space commute if $Q_{b},Q^{\prime}_{b^{\prime}}]=0$ for each $b\in B,b^{\prime}\in B^{\prime}$. Given a quantum function $Q:A\to_{q}B$ for two sets $A,B$, we say that $Q$ is non-contextual if $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for each $a,a^{\prime}\in A$ and each $b,b^{\prime}\in B$. We say that $Q$ is contextual if it is not non-contextual. Note that each classical function $f:A\to B$ can be seen as a non-contextual quantum function.
Let $\mathbb{A},\mathbb{B}$ be relational structures with the same signature $\sigma$. A quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$, first defined at this level of generality in [abramsky2017quantum], is given by a non-trivial Hilbert space $H$ and a quantum function $Q\colon A\to_{q}B$ such that the following conditions hold:
  • ($\mathbf{QH_{1}}$)
    $\prod_{i\in[\operatorname{ar}(R)]}Q_{a_{i},b_{i}}=0\quad\forall R\in\sigma,\mathbf{a}\in R^{\mathbb{A}},\mathbf{b}\in B^{\operatorname{ar}(R)}\setminus R^{\mathbb{B}}$;
  • ($\mathbf{QH_{2}}$)
    $[Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0\quad\forall(a,a^{\prime})\in E(\operatorname{Gaif}(\mathbb{A})),(b,b^{\prime})\in B^{2}$.
Note: Condition $($$\mathbf{QH_{2}}$)) means that $Q$ corresponds to perfect oracular correlations. In Section 7 we shall extend our results to the non-oracular case. We also point out that the notion of a quantum homomorphism can be phrased equivalently as a representation of a certain finitely generated *-algebra [MastelSlofstra2024].
Definition 1 (Entangled CSP).
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP^{*}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{A}$.
2.3. Operations on quantum homomorphisms
We describe here certain constructions taking as input two quantum homomorphisms and giving another quantum homomorphism as the output.
Recall that if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\oplus Q^{\prime}$ is the operator on $H\oplus H^{\prime}$ defined by $(Q\oplus Q^{\prime})(v+v^{\prime})=Q(v)+Q(v^{\prime})$. We extend this definition to quantum functions $Q\colon X\to_{H}A$, $Q^{\prime}:X\to_{H^{\prime}}A$ by letting $Q\oplus Q^{\prime}:X\to_{H\oplus H^{\prime}}A$ be defined by $(Q\oplus Q^{\prime})_{x,a}:=Q_{x,a}\oplus Q^{\prime}_{x,a}$. Note that this is well defined as the set $\{Q_{x,a}\oplus Q^{\prime}_{x,a}\}_{a\in A}$ is a PVM over the Hilbert space $H\oplus H^{\prime}$. Indeed, the fact that its members are projectors onto $H\oplus H^{\prime}$ immediately follows from the definition. Moreover, it is easy to verify that the projector $\sum_{a\in A}(Q_{x,a}\oplus Q^{\prime}_{x,a})$ acts like the identity on both $H$ and $H^{\prime}$—and, hence, on the whole $H\oplus H^{\prime}$ by linearity.
Similarly, if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\otimes Q^{\prime}$ is the linear operator on $H\otimes H^{\prime}$ defined by letting $(Q\otimes Q^{\prime})(v\otimes v^{\prime})=Q(v)\otimes Q^{\prime}(v^{\prime})$ for all $v\in H$ and $v^{\prime}\in H^{\prime}$ and extending by linearity. As for the direct sum, we extend this definition to quantum functions $Q\colon X\to_{H}A$ and $Q^{\prime}\colon X\to_{H^{\prime}}B$ by $(Q\otimes Q^{\prime})_{x,(a,b)}:=Q_{x,a}\otimes Q^{\prime}_{x,b}$ for all $a\in A$ and $b\in B$. This is a quantum function $Q\otimes Q^{\prime}\colon X\to_{H\otimes H^{\prime}}A\times B$.
Lemma 2.
Let $Q\colon\mathbb{X}\to_{H}\mathbb{A}$ and $Q^{\prime}\colon\mathbb{X}\to_{H^{\prime}}\mathbb{A}$ be quantum homomorphisms. Then $Q\oplus Q^{\prime}$ is a quantum homomorphism $\mathbb{X}\to_{H\oplus H^{\prime}}\mathbb{A}$.
Proof.
Let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$ and $(a_{1},\dots,a_{r})\not\in R^{\mathbb{A}}$. Then $R_{x_{1},a_{1}}\ldots R_{x_{r},a_{r}}=Q_{x_{1},a_{1}}\ldots Q_{x_{r},a_{r}}\oplus Q^{\prime}_{x_{1},a_{1}}\ldots Q^{\prime}_{x_{r},a_{r}}=0$, so $($$\mathbf{QH_{1}}$)) holds.
For $($$\mathbf{QH_{2}}$)), take a symbol $R\in\sigma$ of arity $r$ and let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$, $i,j\in\{1,\dots,r\}$, and $a,a^{\prime}\in A$. Then $R_{x_{i},a},R_{x_{j},a^{\prime}}]=[Q_{x_{i},a},Q_{x_{j},a^{\prime}}]\oplus[Q^{\prime}_{x_{i},a},Q^{\prime}_{x_{j},a^{\prime}}]=0$, so $($$\mathbf{QH_{2}}$)) holds. ∎
Observe that classical homomorphisms $h\colon A\to B$ can be identified with quantum homomorphisms $h\colon A\to_{H}B$ over a Hilbert space $H$ of dimension $1$. We will implicitly use this identification in the following. Now, if $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is such that $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for all $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$, then $Q=h_{1}\oplus\dots\oplus h_{d}$ for some $d\geq 1$ and $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$ classical homomorphisms (where $d$ can be taken to be the dimension of $Q$). This is a simple consequence of the spectral theorem.
Definition 3.
Let $\mathbb{A},\mathbb{B}$ be two $\sigma$-structures. The quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ is the set $\oplus{\operatorname{Hom}(\mathbb{A},\mathbb{B})}$ of all quantum homomorphisms of the form $h_{1}\oplus\dots\oplus h_{d}$, where $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$.
[…]
(1) $Q_{a,b}(x)=\sum_{i=1}^{d}\mathbf{1}[h_{i}(a)=b]\langle x,e_{i}\rangle e_{i}$
for all $x\in H$, where $\mathbf{1}{[h_{i}(a)=b]}$ is $1$ if $h_{i}(a)=b$ and 0 otherwise.
Lemma 4.
Let $Q\colon\mathbb{X}\to_{H}\mathbb{A}$ and $Q^{\prime}\colon\mathbb{X}\to_{H^{\prime}}\mathbb{B}$ be quantum homomorphisms. Then $Q\otimes Q^{\prime}$ is a quantum homomorphism $\mathbb{X}\to_{H\otimes H^{\prime}}\mathbb{A}\times\mathbb{B}$.
Proof.
Let $R=Q\otimes Q^{\prime}$. Let $x\in X,a\in A,b\in B$. Then
$R_{x,(a,b)}(R_{x,(a,b)}(v\otimes v^{\prime}))$ $=R_{x,(a,b)}(Q_{x,a}(v)\otimes Q^{\prime}_{x,b}(v^{\prime}))$
$=Q_{x,a}(Q_{x,a}(v))\otimes Q^{\prime}_{x,b}(Q^{\prime}_{x,b}(v^{\prime}))$
$=Q_{x,a}(v)\otimes Q^{\prime}_{x,b}(v^{\prime})$
$=R_{x,(a,b)}(v\otimes v^{\prime})$
so that each $R_{x,(a,b)}$ is a projector. Moreover, if $x\in X$, then
$\sum_{a,b}R_{x,(a,b)}$ $=\sum_{a,b}Q_{x,a}\otimes Q_{x^{\prime},b}$
$=\sum_{a}Q_{x,a}\otimes(\sum_{b}Q_{x^{\prime},b})$
$=\sum_{a}Q_{x,a}\otimes\operatorname{id}_{H^{\prime}}$
$=(\sum_{a}Q_{x,a})\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}}$
so $R$ is a quantum function. The verification that $($$\mathbf{QH_{1}}$)) holds is similar.
For $($$\mathbf{QH_{2}}$)), let $x,y$ appear together in a tuple in a relation of $\mathbb{X}$, and let $a,a^{\prime}\in A,b,b^{\prime}\in B$. Then
$R_{x,(a,b)}R_{y,(a^{\prime},b^{\prime})}$ $=(Q_{x,a}\otimes Q^{\prime}_{x,b})(Q_{y,a^{\prime}}\otimes Q^{\prime}_{y,b^{\prime}})$
$=(Q_{x,a}Q_{y,a^{\prime}})\otimes(Q^{\prime}_{x,b}Q^{\prime}_{y,b^{\prime}})$
$=(Q_{y,a^{\prime}}Q_{x,a})\otimes(Q^{\prime}_{y,b^{\prime}}Q^{\prime}_{x,b})$
$=R_{y,(a^{\prime},b^{\prime})}R_{x,(a,b)}$
so $($$\mathbf{QH_{2}}$)) holds. ∎
We shall also make use of the following composition operation for quantum homomorphisms.
Definition 5.
Let $Q\colon A\to_{H}B$ and $R\colon B\to_{H^{\prime}}C$ be quantum functions. We define
$R\bullet Q\colon A\to_{H\otimes H^{\prime}}C$
by $(R\bullet Q)_{a,c}=\sum_{b\in B}R_{b,c}\otimes Q_{a,b}$.
Note that the sum in the definition is finite (even if $B$ is infinite), since $H$ has finite dimension and therefore only finitely many of the $Q_{a,b}$ are non-zero.
The fact that the composition of quantum homomorphisms is a quantum homomorphism was proved in [abramsky2017quantum] using that this operation can be expressed as the composition of Kleisli arrows in the Kleisli category for the quantum monad. For the convenience of the reader, we give here a self-contained proof following [abramsky2017quantum].
Lemma 6 ([abramsky2017quantum]).
The composition of quantum homomorphisms $Q\colon\mathbb{A}\to_{H}\mathbb{B}$ and $R\colon\mathbb{B}\to_{H^{\prime}}\mathbb{C}$ is a quantum homomorphism $\mathbb{A}\to_{H\otimes H^{\prime}}\mathbb{C}$.
Proof.
We first verify that $S=R\bullet Q$ is a quantum function. For $a\in A,c\in C$, we get
$S_{a,c}S_{a,c}$ $=\sum_{b,b^{\prime}}(R_{b,c}\otimes Q_{a,b})(R_{b^{\prime},c}\otimes Q_{a,b^{\prime}})$
$=\sum_{b,b^{\prime}}(R_{b,c}R_{b^{\prime},c})\otimes(Q_{a,b}Q_{a,b^{\prime}})$
$=\sum_{b}R_{b,c}\otimes Q_{a,b}$ (since $Q_{a,b}Q_{a,b^{\prime}}=0$ for $b\neq b^{\prime}$)
$=S_{a,c}.$
Similarly,
$S_{a,c}^{\top}=(\sum_{b}R_{b,c}\otimes Q_{a,b})^{\top}=\sum_{b}R_{b,c}^{\top}\otimes Q_{a,b}^{\top}=\sum_{b}R_{b,c}\otimes Q_{a,b}=S_{a,c},$
so $S_{a,c}$ is a projector. We also have
$\sum_{c}S_{a,c}=\sum_{b,c}R_{b,c}\otimes Q_{a,b}=\sum_{b}\operatorname{id}\otimes Q_{a,b}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}}.$
This means that $S$ is indeed a quantum function. Let us prove it is in fact a quantum homomorphism.
If $(a_{1},\dots,a_{r})\in R^{\mathbb{A}}$ and $(c_{1},\dots,c_{r})\not\in R^{\mathbb{C}}$ for some symbol $R\in\sigma$ of arity, say, $r$, then
$S_{a_{1},c_{1}}\dots S_{a_{r},c_{r}}$ $=\sum_{b_{1},\dots,b_{r}\in B}(R_{b_{1},c_{1}}\otimes Q_{a_{1},c_{1}})\dots(R_{b_{r},c_{r}}\otimes Q_{a_{r},b_{r}})$
$=\sum_{b_{1},\dots,b_{r}\in B}(R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}})\otimes(Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}).$
Note that for every $(b_{1},\dots,b_{r})$, either $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$—in which case $R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}}=0$ since $R$ satisfies $($$\mathbf{QH_{1}}$))—or $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$—and then $Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}=0$ since $Q$ satisfies $($$\mathbf{QH_{1}}$)). Thus, the sum above evaluates to 0 and $($$\mathbf{QH_{1}}$)) holds.
If $a,a^{\prime}$ appear together in a tuple in a relation of $\mathbb{A}$ and $c,c^{\prime}\in C$, then
$S_{a,c}S_{a^{\prime},c^{\prime}}$ $=\sum_{b,b^{\prime}}(R_{b,c}\otimes Q_{a,b})(R_{b^{\prime},c^{\prime}}\otimes Q_{a^{\prime},b^{\prime}})$
$=\sum_{b,b^{\prime}}(R_{b,c}R_{b^{\prime},c^{\prime}})\otimes(Q_{a,b}Q_{a^{\prime},b^{\prime}})$
$=\sum_{b,b^{\prime}}(R_{b^{\prime},c^{\prime}}R_{b,c})\otimes(Q_{a^{\prime},b^{\prime}}Q_{a,b})$ by $($$\mathbf{QH_{2}}$))
$=S_{a^{\prime},c^{\prime}}S_{a,c}.$
so $($$\mathbf{QH_{2}}$)) holds. ∎
We will also make use of the following simple result.
Lemma 7.
Let $\mathbb{X}$ be a core and let $Q\colon\mathbb{X}\to_{q}\mathbb{X}$ be a quantum homomorphism. If $Q$ is non-contextual, then
$\sum_{x\in X}Q_{x,y}=\operatorname{id}$
for each vertex $y$ of $\mathbb{X}$.
Proof.
Since $Q$ is non-contextual, we can write it as $Q=\bigoplus_{i\in[d]}h_{i}$ for some $d\in\mathbb{N}$ and some classical homomorphisms $h_{i}:\mathbb{X}\to\mathbb{X}$. By the definition of a core, each $h_{i}$ is an automorphism of $\mathbb{X}$. In particular, for each $i\in[d]$ there exists a unique $x_{i}$ such that $h_{i}(x_{i})=y$. It follows that
$\sum_{x\in X}Q_{x,y}=\bigoplus_{i\in[d]}\sum_{x\in X}(h_{i})_{x,y}=\operatorname{id}.$
3. Quantum polymorphisms and reductions
In this section, we show that reductions between entangled CSPs can be obtained via a quantised notion of pp-definitions that we call q-definitions. In order to prove that q-definitions exist, we make use of the notion of commutativity gadgets as defined in [Ji]. Then, we prove that the existence of commutativity gadgets is entirely determined by a closure property of the quantum polymorphisms of the given language. We begin by defining q-definitions.
Definition 8.
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
  • $(\mathbf{q_{1}})$
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
  • $(\mathbf{q_{2}})$
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$.
We can then define a notion of q-definability, as follows.
Definition 9.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{B}$ is q-definable from $\mathbb{A}$ if there exists a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
Observe, in ($(\mathbf{q_{1}})$), that a quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ is necessarily non-contextual, since all the vertices of $\mathbb{R}$ lie in a common constraint. (This will not be true anymore in the non-oracular setting, see Section 7.) Classical gadgets for CSPs that can be defined one from the other via pp-definitions provide logspace reductions between CSPs of different languages. The next result shows that the same is true for $\operatorname{CSP^{*}}$, provided that pp-definitions are replaced by the stronger q-definitions.
Proposition 10.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures such that $\mathbb{B}$ is q-definable from $\mathbb{A}$. Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{*}}(\mathbb{A})$ in logspace.
We prove Proposition 10 later in this section. For now, we focus on the following problem: How to construct q-definitions guaranteeing that a given $\operatorname{CSP^{*}}$ can be reduced to a different one? It is not hard to show that every q-definition yields a classical pp-definition. Indeed, considering ($(\mathbf{q_{1}})$) for quantum homomorphisms with a Hilbert space of dimension 1 (which corresponds to classical homomorphisms), we find that the tuples in $R$ are in one-to-one correspondence with the restrictions of homomorphisms $\mathbb{G}\to\mathbb{A}$ to $\{g_{1},\dots,g_{r}\}$. However, not every pp-definition is a q-definition. It turns out that, in order for a pp-definition to ensure the existence of a q-definition, we need to pair it with an extra construction making sure that the process of replacing constraints with gadgets preserves the existence of quantum homomorphisms. This is the notion of commutativity gadgets we define below, generalising [Ji].
Definition 11.
Fix a $\sigma$-structure $\mathbb{A}$. A commutativity gadget for $\mathbb{A}$ is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
  • $(\mathbf{c_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ extending $Q$.
  • $(\mathbf{c_{2}})$
    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
Note that a commutativity gadget as in Definition 11 is not the same thing as a q-definition for $\mathbb{A}$ and $R=A^{2}$ in the sense of Definition 8.
Remark 12.
In [Zeman], an algebraic commutativity gadget is defined by replacing ($(\mathbf{c_{1}})$) by the following condition:
  • $(\mathbf{c_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
Note that this is a special case of $(\mathbf{c_{1}})$ and, thus, any commutativity gadget is an algebraic commutativity gadget. While this appears to be a strictly weaker requirement than the one in Definition 11, we shall prove below that the existence of an algebraic commutativity gadget in the sense of [Zeman] is equivalent to the existence of a commutativity gadget in our sense.
If $\mathbb{A}$ has a commutativity gadget, then we have a natural way of obtaining a sound reduction from $\operatorname{CSP^{*}}(\mathbb{B})$ to $\operatorname{CSP^{*}}(\mathbb{A})$ out of a classical gadget: For every pair of variables $x,y$ in a relation in an instance of $\operatorname{CSP^{*}}(\mathbb{B})$, glue to $x,y$ a copy of the commutativity gadget, identifying $x$ with $u$ and $y$ with $v$. The second condition in Definition 11 forces the associated projectors to commute in every quantum homomorphism to $\mathbb{A}$, and the first condition tells us that no additional constraint is forced upon $x,y$ by gluing the commutativity gadget. Formally, we prove the following result.
Proposition 13.
Let $\mathbb{A},\mathbb{B}$ be structures such that $(i)$ $\mathbb{A}$ has a commutativity gadget, and $(ii)$ $\mathbb{B}$ is pp-definable from $\mathbb{A}$. Then $\mathbb{B}$ is q-definable from $\mathbb{A}$.
As a consequence of what we have discussed so far, the crucial step in order to quantise classical reductions from CSP theory is to build commutativity gadgets. The main result of this section classifies the languages for which commutativity gadgets exist, in terms of a property of their quantum polymorphisms, which we now define.
Definition 14.
A quantum polymorphism of arity $n\in\mathbb{N}$ for a structure $\mathbb{A}$ is a quantum homomorphism $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$.
We denote by $\operatorname{qPol}(\mathbb{A})$ the set of quantum polymorphisms of $\mathbb{A}$. Moreover, we define the quantum closure of $\operatorname{Pol}(\mathbb{A})$, denoted by $\operatorname{\oplus Pol}(\mathbb{A})$, as the union of the quantum closures of the set of $n$-ary polymorphisms $\operatorname{Hom}(\mathbb{A}^{n},\mathbb{A})$, for all $n\in\mathbb{N}$.
Remark 15.
Observe that $\operatorname{qPol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are both abstract minions (see the definition in [BBKO21, ciardo_quantum_minion]). Furthermore, it is not hard to check that $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are homomorphically equivalent, and $\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{A})$. In particular, it follows that, if $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$, then we have a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{A})$.
The main result of this section is the following.
Theorem 16.
Let $\mathbb{A}$ be a finite structure. The following are equivalent:
  1. (1)
    $\mathbb{A}$ has a commutativity gadget.
  2. (2)
    $\mathbb{A}$ has an algebraic commutativity gadget.
  3. (3)
    $\mathbb{A}^{n}$ is an (algebraic) commutativity gadget for all large enough $n\in\mathbb{N}$.
  4. (4)
    $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
In the rest of this section, we provide the proofs of the results we stated.
Proof of Proposition 10.
For every symbol $R\in\rho$ of some arity $r$, let $\mathbb{G}_{R}$ be a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ (with its distinguished vertices $g_{1},\dots,g_{r}$). Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{*}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$. For $R\in\rho$ of some arity $r$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$, and we glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to_{q}\mathbb{A}^{\prime}$. Then, ($(\mathbf{q_{1}})$) guarantees the existence of a quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$ such that
$Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
for each $i\in[r]$ and each $a\in A$. Overall, this is a quantum homomorphism $\mathbb{Y}\to_{q}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\to_{q}\mathbb{A}$ is a quantum homomorphism. Consider the quantum function $Q^{\prime}\colon X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Observe that the restriction of $Q$ to $\mathbb{G}_{R}^{(\mathbf{x})}$ is a quantum homomorphism $\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$. By ($(\mathbf{q_{2}})$), this means that the assignment $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$). In particular, recalling that $x_{i}$ and $g_{i}^{(\mathbf{x})}$ are identified in $\mathbb{Y}$, it follows that the PVMs $Q^{\prime}_{x_{i}}=Q_{g_{i}^{(\mathbf{x})}}$ and $Q^{\prime}_{x_{j}}=Q_{g_{j}^{(\mathbf{x})}}$ commute for each choice of $i,j\in[r]$. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
(2) $\prod_{i\in[r]}Q^{\prime}_{x_{i},b_{i}}=0,$
which means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. ∎
Proof of Proposition 13.
Let $S=R^{\mathbb{B}}$ be a relation of $\mathbb{B}$ of some arity $r$, and let $\mathbb{G}$ be a classical gadget for $R^{\mathbb{B}}$ in $\mathbb{A}$, with distinguished vertices $g_{1},\dots,g_{r}$. Let $\mathbb{H}$ be a commutativity gadget for $\mathbb{A}$, with distinguished vertices $u,v$. We let $\mathbb{G}^{\prime}$ be the structure that initially consists of the disjoint union of $\mathbb{G}$ and a copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$ for each pair of distinct vertices $x,y$ in $G$, and in which we identify each such pair $x,y$ with the vertices $u,v$ from the copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$.
We show that $\mathbb{G}^{\prime}$ (with distinguished vertices $g_{1},\dots,g_{r}$) is a q-definition for $S$ in $\mathbb{A}$. Let $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ be a quantum homomorphism. In particular, $Q$ is non-contextual and can therefore be written as $h_{1}\oplus\dots\oplus h_{d}$ for some homomorphisms $h_{k}\colon\mathbb{R}\to\mathbb{A}_{S}$. Since $\mathbb{G}$ is a gadget for $S$ in $\mathbb{A}$, it follows that there exist homomorphisms $f_{k}\colon\mathbb{G}\to\mathbb{A}$ such that $f_{k}(g_{i})=h_{k}(i)$ for all $i\in[r]$. Then $Q^{\prime}=\bigoplus_{k\in[d]}f_{k}$ is a non-contextual quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$. In particular, $Q^{\prime}_{x}$ and $Q^{\prime}_{y}$ commute for all $x,y\in G$, and it follows from ($(\mathbf{c_{1}})$) that for all $x,y$ there exists a quantum homomorphism $Q^{(x,y)}\colon\mathbb{H}\to_{q}\mathbb{A}$ such that $Q^{(x,y)}_{u}=Q^{\prime}_{x}$ and $Q^{(x,y)}_{v}=Q^{\prime}_{y}$. We thus obtained a quantum homomorphism $\mathbb{G}^{\prime}\to_{q}\mathbb{A}$, which proves ($(\mathbf{q_{1}})$).
Let now $Q\colon\mathbb{G}^{\prime}\to_{q}\mathbb{A}$ be a quantum homomorphism. For each $x,y\in G$, we have by ($(\mathbf{c_{2}})$) that $Q_{x}$ and $Q_{y}$ commute. Thus, $Q$ restricts to a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ that is non-contextual and can thus be written as $h_{1}\oplus\dots\oplus h_{d}$, each $h_{k}$ being a classical homomorphism $\mathbb{G}\to\mathbb{A}$. Since $\mathbb{G}$ is a classical gadget for $R$, we obtain that $(h_{k}(g_{1}),\dots,h_{k}(g_{r}))\in R^{\mathbb{B}}$ for all $k\in[d]$. Thus, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}=\bigoplus_{k}(h_{k})_{g_{i},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$. ∎
Next, we show that q-definitions are compatible with quantum homomorphisms.
Proposition 17.
Let $\mathbb{A},\mathbb{B}$ be structures, $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively. Then every quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is a quantum homomorphism $(A;R_{A})\to_{q}(B;R_{B})$.
Proof.
Let $g_{1},\dots,g_{r}$ be the distinguished vertices of $\mathbb{G}$.
We prove $($$\mathbf{QH_{1}}$)). Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$. By ($(\mathbf{q_{1}})$), there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in\{1,\dots,r\}$. Let $S=Q\bullet Q^{\prime}$, which is a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{B}$. By ($(\mathbf{q_{2}})$), we have that the quantum function defined by $Q^{\prime\prime}_{i,b_{i}}:=S_{g_{i},b_{i}}$ for all $i\in\{1,\dots,r\}$ is a quantum homomorphism $Q^{\prime\prime}\colon\mathbb{R}\to(B,R_{B})$. Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$. Note that $S_{g_{i},b_{i}}$ is by definition $\sum_{a\in A}Q_{a,b_{i}}\otimes Q^{\prime}_{g_{i},a}=Q_{a_{i},b_{i}}\otimes\operatorname{id}$. Thus, $(\prod_{i=1}^{r}Q_{a_{i},b_{i}})\otimes\operatorname{id}=0$, from which we obtain that $\prod_{i=1}^{r}Q_{a_{i},b_{i}}=0$.
We now turn to $($$\mathbf{QH_{2}}$)). Let $a_{1},a_{2}$ appear together in a tuple in $R_{A}$ at positions $i$ and $j$ and let $b_{1},b_{2}\in B$ be arbitrary. Let $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ be such that $Q^{\prime}_{g_{i},a_{1}}=Q^{\prime}_{g_{j},a_{2}}=\operatorname{id}$, whose existence is asserted by ($(\mathbf{q_{1}})$). Again, let $S\colon\mathbb{G}\to_{q}\mathbb{B}$ be the composition $Q\bullet Q^{\prime}$ and $Q^{\prime\prime}\colon\mathbb{R}\to_{q}(B,R_{B})$ be the quantum homomorphism given by ($(\mathbf{q_{2}})$). We get $0=[R_{i,b_{1}},R_{j,b_{2}}]=[S_{g_{i},b_{1}},S_{g_{j},b_{2}}]=[Q_{a_{1},b_{1}}\otimes\operatorname{id},Q_{a_{2},b_{2}}\otimes\operatorname{id}]=[Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]\otimes\operatorname{id}$ and therefore $Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]=0$. ∎
In the rest of this section, we give a proof of Theorem 16. We start by observing the following fact.
Proposition 18.
A quantum polymorphism of a structure $\mathbb{A}$ is non-contextual if and only if it belongs to $\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
This is a simple consequence of the spectral theorem. ∎
Definition 19.
Let $\mathbb{A}$ be a structure. The pairs $(a_{1},b_{1}),\dots,(a_{n},b_{n})\in A^{2}$ are called generators of $A^{2}$ in $\operatorname{Pol}(\mathbb{A})$ if for every pair $(c,d)\in A^{2}$, there exists $g\in\operatorname{Pol}(\mathbb{A})$ of arity $n$ such that $g(a_{1},\dots,a_{n})=c$ and $g(b_{1},\dots,b_{n})=d$.
Note that, if $A$ is finite (as we always assume), $A^{2}$ admits a family of $n=|A|^{2}$ generators. Indeed, we can simply let $g$ be the $i$-th projection $\mathbb{A}^{n}\to\mathbb{A}$, with $i\in[n]$ being the index of the pair $(c,d)$. However, $n$ could be smaller for certain $\mathbb{A}$. We point out that there are many cases of infinite structures $\mathbb{A}$ where $A^{2}$ is finitely generated under $\operatorname{Pol}(\mathbb{A})$; this is for example the case for the class of $\omega$-categorical structures, which has been widely studied in the context of classical constraint satisfaction problems.
Proposition 20.
Let $\mathbb{A}$ be a structure such that $A^{2}$ admits a family of $n$ generators under $\operatorname{Pol}(\mathbb{A})$. Suppose also that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Then $\mathbb{A}^{n}$ is a commutativity gadget for $\mathbb{A}$.
Proof.
Let $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ be generators for $A^{2}$ under $\operatorname{Pol}(\mathbb{A})$. Let $\mathbb{G}$ be $\mathbb{A}^{n}$ and the two distinguished elements be $\mathbf{a}=(a_{1},\dots,a_{n})$ and $\mathbf{b}=(b_{1},\dots,b_{n})$.
Every quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ is a quantum polymorphism of $\mathbb{A}$ and therefore ($(\mathbf{c_{2}})$) holds by the assumption that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$. Moreover, by the spectral theorem, every quantum function $Q\colon\{\mathbf{a},\mathbf{b}\}\to A$ such that $Q(\mathbf{a})$ commutes with $Q(\mathbf{b})$ is of the form $h_{1}\oplus\cdots\oplus h_{d}$, where $h_{i}\colon\{\mathbf{a},\mathbf{b}\}\to A$ is a classical function. Using that $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ are generators for $A^{2}$, we know that, for each $i$, there exists $g_{i}\in\operatorname{Pol}(\mathbb{A})$ such that $g_{i}(\mathbf{a})=h_{i}(\mathbf{a})$ and $g_{i}(\mathbf{b})=h_{i}(\mathbf{b})$. Then $Q$ extends to a quantum homomorphism $Q^{\prime}:=g_{1}\oplus\cdots\oplus g_{d}$, thus proving ($(\mathbf{c_{1}})$). ∎
We now prove that the existence of a commutativity gadget guarantees that all quantum polymorphisms are classical. In fact, we prove that the same fact holds even for the algebraic commutativity gadgets from Remark 12.
Proposition 21.
Let $\mathbb{A}$ be a structure. If $\mathbb{A}$ has an algebraic commutativity gadget, then $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Let $\mathbb{G}$ be an algebraic commutativity gadget with distinguished elements $u,v\in G$. Let $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$ be a quantum polymorphism, and choose two tuples $\mathbf{a},\mathbf{b}\in A^{n}$ and two elements $c,d\in A$. We want to show that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, which is enough to conclude by Proposition 18.
Using ($(\mathbf{c_{1}^{\prime}})$) we know that, for each $i\in[n]$, there exists a quantum homomorphism $W^{(i)}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that
$W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}.$
Consider now the quantum homomorphism $W\colon\mathbb{G}\to_{q}\mathbb{A}^{n}$ given by the tensor product
$W=\bigotimes_{i\in[n]}W^{(i)}$
and observe that $W_{u,\mathbf{a}}=W_{v,\mathbf{b}}=\operatorname{id}$. Take now the composition
$U=Q\bullet W:\mathbb{G}\to_{q}\mathbb{A}.$
Using ($(\mathbf{c_{2}})$), we observe that $U$ satisfies $U_{u,c},U_{v,d}]=0$. However, expanding the composition as per Definition 5, we have
$U_{u,c}=\sum_{\mathbf{t}\in A^{n}}Q_{\mathbf{t},c}\otimes W_{u,\mathbf{t}}=Q_{\mathbf{a},c}\otimes\operatorname{id}$
and, similarly, $U_{v,d}=Q_{\mathbf{b},d}\otimes\operatorname{id}$. Thus,
$Q_{\mathbf{a},c},Q_{\mathbf{b},d}]\otimes\operatorname{id}=[Q_{\mathbf{a},c}\otimes\operatorname{id},Q_{\mathbf{b},d}\otimes\operatorname{id}]=[U_{u,c},U_{v,d}]=0,$
which implies that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we are done. ∎
Proof of Theorem 16.
The result is a direct consequence of Propositions 20 and 21 and the fact that any commutativity gadget is an algebraic commutativity gadget (as observed in Remark 12). ∎
4. Warm-up: Quantum polymorphisms of cliques
In this section, we give a first simple example of using quantum polymorphisms to obtain commutativity gadgets, in the setting of cliques. In particular, we prove the following.
Theorem 22.
For all $m\geq 3$, $\operatorname{qPol}(\mathbb{K}_{m})=\operatorname{\oplus Pol}(\mathbb{K}_{m})$.
Proof.
The proof goes by induction on $m\geq 3$.
For the case of $m=3$, it follows from [Ji, Lemma 4] that $\mathbb{K}_{3}$ has a commutativity gadget. Then, Theorem 16 implies the result.
Take now $m\geq 4$. Let $Q\colon(\mathbb{K}_{m})^{n}\to_{q}\mathbb{K}_{m}$ be a quantum homomorphism over a Hilbert space $H$. Pick $\mathbf{u},\mathbf{v}\in[m]^{n}$ and $i,j\in[m]$. Our goal is to show that $Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=0$.
Choose $\mathbf{w}\in[m]^{n}$ in such a way that $(\mathbf{u},\mathbf{w})$ and $(\mathbf{v},\mathbf{w})$ are both edges of $(\mathbb{K}_{m})^{n}$. (Note that this is possible as long as $m\geq 3$.) Choose also $k\in[m]\setminus\{i,j\}$. We first claim that
(3) $[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},k}]=0.$
This is clearly the case if $Q_{\mathbf{w},k}=0$, so assume this is not the case. Define $\mathbb{X}$ to be the subgraph of $(\mathbb{K}_{m})^{n}$ induced by the vertices $\mathbf{z}$ such that $z_{p}\neq w_{p}$ for all $p\in\{1,\dots,n\}$, namely the subgraph induced by the neighborhood of $\mathbf{w}$. Note that $\mathbb{X}$ is isomorphic to $(\mathbb{K}_{m-1})^{n}$.
Define $R\colon\mathbb{X}\to_{q}\mathbb{K}_{m-1}$ by setting $R_{\mathbf{z},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}$ over the (non-trivial) space $Q_{\mathbf{w},k}(H)$. We check that this is a quantum homomorphism. Every $R_{\mathbf{z},\ell}$ is a projector since $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell}]=0$. For $\mathbf{z}\in X$, we have $\sum_{\ell}R_{\mathbf{z},\ell}=\sum_{\ell}Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}=Q_{\mathbf{w},k}(\sum_{\ell}Q_{\mathbf{z},\ell})=Q_{\mathbf{w},k}$, which is indeed the identity function on $Q_{\mathbf{w},k}(H)$ since $Q_{\mathbf{w},k}$ is a projector. Thus, $R$ is a quantum function. If $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$, then $\mathbf{z},\mathbf{z}^{\prime},\mathbf{w}$ form a triangle in $(\mathbb{K}_{m})^{n}$ and therefore the projectors $Q_{\mathbf{w},k},Q_{\mathbf{z},\ell},Q_{\mathbf{z}^{\prime},\ell^{\prime}}$ pairwise commute for $\ell,\ell^{\prime}$ distinct from $k$. Thus, $($$\mathbf{QH_{2}}$)) holds. Finally, if $\mathbf{z},\mathbf{z}^{\prime}$ are adjacent in $\mathbb{X}$ and $\ell\neq k$, then $R_{\mathbf{z},\ell}R_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{w},k}Q_{\mathbf{z}^{\prime},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}Q_{\mathbf{z}^{\prime},\ell}=0$. Thus, $($$\mathbf{QH_{1}}$)) holds.
By the inductive hypothesis, we have $\operatorname{qPol}(\mathbb{K}_{m-1})=\operatorname{\oplus Pol}(\mathbb{K}_{m-1})$. Therefore, $R_{\mathbf{z},\ell},R_{\mathbf{z}^{\prime},\ell^{\prime}}]=0$ for all $\mathbf{z},\mathbf{z}^{\prime}\in X$ and all $\ell,\ell^{\prime}\in[m]\setminus\{k\}$. For $\mathbf{z}=\mathbf{u}$ and $\mathbf{z}^{\prime}=\mathbf{v}$, this gives the claimed eq. 3.
We can now already conclude by noting that
$Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=[Q_{\mathbf{u},i}(\sum_{k}Q_{\mathbf{w},k}),Q_{\mathbf{v},j}(\sum_{\ell}Q_{\mathbf{w},\ell})]=\sum_{k,\ell}[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}],$
that all the terms with $k=\ell$ in the sum vanish by eq. 3, and that all the terms with $k\neq\ell$ in the sum vanish, since
$[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}]$ $=Q_{\mathbf{u},i}Q_{\mathbf{w},k}Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}-Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}Q_{\mathbf{u},i}Q_{\mathbf{w},k}$
$=Q_{\mathbf{u},i}Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}Q_{\mathbf{v},j}-Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}Q_{\mathbf{u},i}$
$=0,$
where the second equality is due to the fact that $Q_{\mathbf{w},k}$ and $Q_{\mathbf{w},\ell}$ commute with $Q_{\mathbf{u},i}$ and $Q_{\mathbf{v},j}$, and the third holds since $Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}=0=Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}$. ∎
Note that $\mathbb{K}_{2}$ has quantum polymorphisms that are not in the quantum closure of $\operatorname{Pol}(\mathbb{K}_{2})$. Indeed, let $n=2$ and $H=\mathbb{C}^{2}$, and define $Q\colon(\mathbb{K}_{2})^{2}\to_{q}\mathbb{K}_{2}$ by the following matrices:
  • $Q_{(0,0),0}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},Q_{(0,0),1}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$
  • $Q_{(1,1),0}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},Q_{(1,1),0}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$
  • $Q_{(1,0),0}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix},Q_{(1,0),1}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$
  • $Q_{(0,1),0}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix},Q_{(0,1),1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$.
It is easily checked that this is a quantum polymorphism and that $Q_{(0,0),0},Q_{(1,0),0}]\neq 0$. By Theorem 16, we obtain that $\mathbb{K}_{2}$ does not have a commutativity gadget.
Proposition 23.
Let $m\geq 3$. Then $\mathbb{K}_{m}$ has a commutativity gadget. In fact, for every $n\geq 2$ and $\mathbf{u},\mathbf{v}\in(\mathbb{K}_{m})^{n}$ such that $u_{i}=v_{i}$ for at least one $i\in[n]$ and $u_{j}\neq v_{j}$ for at least one $j\in[n]$, $(\mathbb{K}_{m})^{n}$ with the distinguished vertices $\mathbf{u},\mathbf{v}$ is a commutativity gadget for $\mathbb{K}_{m}$.
Proof.
We apply Proposition 20. The assumptions on $\mathbf{u},\mathbf{v}$ imply that $(u_{1},v_{1}),\dots,(u_{n},v_{n})$ generate $m]^{2}$. Indeed, any pair $(c,c)$ is the image of $(u_{i},v_{i})$ under a permutation (which is in particular a unary polymorphism of $\mathbb{K}_{m}$), while any pair $(c,d)$ with $c\neq d$ is the image of $(u_{j},v_{j})$ under a permutation. ∎
A consequence of Proposition 23 is the undecidability of the entangled $m$-coloring problem for $m\geq 3$. The case of $m=3$ was proved in [culf_re-completeness_2025], and the case for $m>3$ was recently and independently proved in [Zeman, Proposition 5.2, Lemma 4.6]. We include a simple proof here using commutativity gadgets for the convenience of the reader.
Corollary 24.
$\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable for all $m\geq 3$.
Proof.
We shall reduce from $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, whose undecidability was proved in [Ji, culf2024re]. Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, which we can assume does not contain isolated vertices. Let $\mathbb{Y}$ be the graph obtained by adjoining m3m-3 new vertices $y_{4},\dots,y_{m}$ to $\mathbb{X}$ and such that, whenever $x$ is a vertex in $\mathbb{X}$, then $\{x,y_{4},\dots,y_{m}\}$ forms a clique in $\mathbb{Y}$. Finally, adjoin a commutativity gadget between every $y_{i}$ and every other vertex in $\mathbb{Y}$. Such a commutativity gadget exists by Proposition 23. We let $\mathbb{Z}$ be the resulting structure.
If $\mathbb{X}$ has a quantum homomorphism $Q$ to $\mathbb{K}_{3}$, then $Q^{\prime}\colon\mathbb{Y}\to_{q}\mathbb{K}_{m}$ defined by $Q^{\prime}_{y_{i},i}=\operatorname{id}$ for all $i\in\{4,\dots,m\}$ and $Q_{x,j}=Q^{\prime}_{x,j}$ for all $x\in X$ is a quantum homomorphism. This $Q^{\prime}$ can be extended to a quantum homomorphism $\mathbb{Z}\to\mathbb{K}_{m}$ by ($(\mathbf{c_{1}})$).
Conversely, suppose that there exists a quantum homomorphism $Q\colon\mathbb{Z}\to_{q}\mathbb{K}_{m}$. By ($(\mathbf{c_{2}})$), it follows that for every $x\in X$, the restriction of $Q$ to $\{x,y_{2},\dots,y_{4}\}$ is non-contextual. Thus, this restriction can be written as $h_{1}\oplus\dots\oplus h_{d}$ for some classical homomorphisms $h_{1},\dots,h_{d}$. Up to composing the $h_{i}$ with an automorphism of $\mathbb{K}_{m}$, we can assume without loss of generality that $h_{i}(y_{j})=j$ for all $i\in\{1,\dots,d\}$ and $j\in\{4,\dots,m\}$. Thus, for each $i\in\{1,\dots,d\}$ and $x\in X$, we have that $h_{i}(x)\in\{1,2,3\}$. It follows that $\sum_{i=1}^{3}Q_{x,i}=\operatorname{id}$ and that $Q$ restricts to a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{K}_{3}$. ∎
5. Intermezzo: Contextuality and bifurcations
In this section, we present a general result on the emergence of a certain non-orthogonality structure in any contextual quantum homomorphism. While in the current work we shall only use this result to analyse quantum polymorphisms of odd cycles (see Section 6), we believe this result will be useful in a broader setting.
$Q_{x_{0}}$$Q_{x_{1}}$$Q_{x_{2}}$$a_{0}$$a_{0}^{\prime}$$a_{1}$$a_{2}^{\prime}$$a_{2}$
Figure 1. A bifurcation of length $2$. Solid lines mean non-orthogonality.
The high-level idea is that, whenever some non-commutativity occurs in a quantum homomorphism $Q:\mathbb{X}\to_{q}\mathbb{Y}$, a bifurcation must appear in the product $X\times A$, in a sense that we formalise below. For simplicity, we restrict the analysis to binary structures (i.e., digraphs), although the same argument applies to quantum homomorphisms between arbitrary structures.
Definition 25.
Let $\mathbb{X},\mathbb{A}$ be digraphs, let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a quantum homomorphism, and let $\omega$ be a positive integer. A bifurcation for $Q$ is a set
$V=\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}\subseteq X\times A$
such that
  • $x_{0},x_{1},\dots,x_{\omega}$ is a path in $\mathbb{X}$;
  • $a_{0}\neq a_{0}^{\prime}$ and $a_{\omega}\neq a_{\omega}^{\prime}$;
  • $Q_{\alpha,\beta}Q_{\delta,\gamma}\neq 0$ for each $((\alpha,\beta),(\delta,\gamma))\in V^{2}\setminus\{((x_{0},a_{0}^{\prime}),(x_{\omega},a_{\omega}^{\prime})),((x_{\omega},a_{\omega}^{\prime}),(x_{0},a_{0}^{\prime}))\}$ such that $\alpha\neq\delta$. We call $\omega$ the length of the bifurcation.
Recall that $\mathbb{X}^{u}$ denotes the undirected graph underlying a digraph $\mathbb{X}$.
Theorem 26.
Let $\mathbb{X}$ and $\mathbb{A}$ be digraphs such that $\mathbb{X}^{u}$ is connected, and let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism. Then there exists a bifurcation for $Q$ whose length is at most the diameter of $\mathbb{X}^{u}$.
We will need the following result from [manvcinska2016quantum], of which we give an alternative proof making use of the machinary of Section 3.
Proposition 27 ([manvcinska2016quantum]).
Let $Q\colon\mathbb{X}\to_{q}\mathbb{Y}$ be a quantum homomorphism between two digraphs $\mathbb{X}$ and $\mathbb{Y}$. If $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ are such that there exists a directed $\ell$-walk from $x$ to $x^{\prime}$ in $\mathbb{X}$ but there is no directed $\ell$-walk from $y$ to $y^{\prime}$ in $\mathbb{Y}$, then $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$.
Proof.
Consider the structure $\mathbb{G}$ consisting of a path of length $\ell$, and call $u,v$ the two extremities of this path. Since this structure is a tree, Proposition 47 implies that $(\mathbb{G},u,v)$ is a q-definition for certain binary relations $R_{X}\subseteq X^{2},R_{Y}\subseteq Y^{2}$ in $\mathbb{X}$ and $\mathbb{Y}$, respectively. It follows from Proposition 17 that $Q$ is a quantum homomorphism $(X,R_{X})\to_{q}(Y,R_{Y})$. By assumption, $R_{X}$ contains $(x,x^{\prime})$ and $R_{Y}$ does not contain $(y,y^{\prime})$, and therefore $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$ by $($$\mathbf{QH_{1}}$)). ∎
We shall repeatedly use a simple fact about projectors, stated below without proof.
Fact 28.
Let $P,Q$ be projectors onto some Hilbert space $H$. Then $[P,Q]=0$ if and only if $PQP=PQ=QP$.
Proof of Theorem 26.
Let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism over some Hilbert space $H$ of dimension $h$. Let $S$ be the set of pairs $(x,x^{\prime})\in X^{2}$ for which there exist $(a,a^{\prime})\in A^{2}$ such that $Q_{x,a},Q_{x^{\prime},a^{\prime}}]\neq 0$. Since $Q$ is contextual, $S$ is non-empty. Choose $(x,x^{\prime})\in S$ in a way that $\operatorname{dist}(x,x^{\prime})$ is minimum over all pairs in $S$ (where the distance is computed in $\mathbb{X}^{u}$), and call such distance $\omega$. Note that $\omega\geq 2$, as $Q$ satisfies $($$\mathbf{QH_{2}}$)). Let $x=x_{0},x_{1},\dots,x_{\omega}=x^{\prime}$ be a shortest oriented path in $\mathbb{X}$ connecting $x$ and $x^{\prime}$. By the minimality of $\operatorname{dist}(x,x^{\prime})$, observe that the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\beta}}$ commute for each $\alpha\leq\beta\in\{0,\dots,\omega\}$ such that $\alpha\neq 0$ or $\beta\neq\omega$. In particular, we can choose an o.n. basis $\mathcal{E}=(e_{1},\dots,e_{h})$ for $H$ that diagonalises all PVMs $Q_{x_{\alpha}}$ for each $\alpha\in\{0,\dots,\omega-1\}$. On the other hand, since $Q_{x_{0}}$ and $Q_{x_{\omega}}$ do not commute, there exist $a_{0},a_{\omega}\in A$ such that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
Take the set $Z=\{i\in[h]:Q_{x_{\omega},b}e_{i}\in\{0,e_{i}\}\}$, and consider the projectors $P_{1}$ onto the space $\operatorname{span}(\{e_{i}:i\in Z\})$ and $P_{2}$ onto the space $\operatorname{span}(\{e_{i}:i\in[h]\setminus Z\})$. Observe that $P_{1}P_{2}=0$ and $P_{1}+P_{2}=\operatorname{id}$. Moreover, we have
(4) $[P_{1},M]=[P_{2},M]=0\quad\forall M\in\{Q_{x_{\alpha},d}:\alpha\in\{0,\dots,\omega-1\},d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}\}.$
Indeed, $P_{1}$ is diagonal in the same basis $\mathcal{E}$ that diagonalises each $Q_{x_{\alpha}}$, so it commutes with all projectors in such PVMs. Furthermore,
$Q_{x_{\omega},a_{\omega}}P_{1}=\sum_{i\in Z}Q_{x_{\omega},a_{\omega}}e_{i}e_{i}^{*}=\sum_{i\in Z}e_{i}e_{i}^{*}Q_{x_{\omega},a_{\omega}}=P_{1}Q_{x_{\omega},a_{\omega}}$
since, for $i\in Z$, $e_{i}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}$. Hence, $P_{1}$ also commutes with $Q_{x_{\omega},a_{\omega}}$. Finally, $P_{1},M]+[P_{2},M]=[\operatorname{id},M]=0$ for each $M$ because of the bilinearity of the commutator, so eq. 4 holds. In particular, the operators $Q_{x_{0},a_{0}}P_{1}$, $Q_{x_{0},a_{0}}P_{2}$, $Q_{x_{\omega},a_{\omega}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{2}$ are all projectors. Observe now that
$[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]$ $=[Q_{x_{0},a_{0}}P_{1}+Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}+Q_{x_{\omega},a_{\omega}}P_{2}]$
$=[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{1}]+[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{2}]+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]$
(5) $+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}].$
Combining eq. 4 and 28, we find that the second summand in the right-hand side of section 5 is
$[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{2}]=[P_{1}Q_{x_{0},a_{0}}P_{1},P_{2}Q_{x_{\omega},a_{\omega}}P_{2}]=0$
since $P_{1}P_{2}=0$. Similarly, we obtain $Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]=0$. Consider now the first summand in the right-hand side of section 5. Both the projectors $Q_{x_{0},a_{0}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{1}$ are diagonal in the basis $\mathcal{E}$. Hence, they commute, too. We deduce that
(6) $[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0.$
We now claim that the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ pairwise commute. The fact that $Q_{x_{\alpha},d}P_{2},Q_{x_{\alpha}^{\prime},d^{\prime}}P_{2}]=0$ for each $\alpha,\alpha^{\prime}\in[\omega-1],d,d^{\prime}\in A$ is trivial, since $\mathcal{E}$ diagonalises both. Moreover,
$Q_{x_{\alpha},d}P_{2}Q_{x_{\omega},a_{\omega}}P_{2}=Q_{x_{\alpha},d}Q_{x_{\omega},a_{\omega}}P_{2}=Q_{x_{\omega},a_{\omega}}Q_{x_{\alpha},d}P_{2}=Q_{x_{\omega},a_{\omega}}P_{2}Q_{x_{\alpha},d}P_{2},$
where the first and third equalities come from eq. 4 and 28, while the second holds since the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\omega}}$ are assumed to commute by the minimality of $\operatorname{dist}(x,x^{\prime})$. So, the claim is true, and we can simultaneously diagonalise the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ in some o.n. basis $\mathcal{F}=(f_{1},\dots,f_{h})$ of $H$.
For the next step, it will be useful to define the sets $L,M\subseteq[h]$ such that $Q_{x_{0},a_{0}}P_{2}=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}$ and $Q_{x_{\omega},a_{\omega}}P_{2}=\sum_{m\in M}f_{m}f_{m}^{*}$. For $m\in M$, consider also the set $A_{m}=\{i\in[h]:{\langle e_{i},f_{m}\rangle}\neq 0\}$. Observe that, for any $m\in M$,
$Q_{x_{0},a_{0}}P_{2}f_{m}f_{m}^{*}$ $=Q_{x_{0},a_{0}}P_{2}\left(\sum_{i\in[h]}e_{i}e_{i}^{*}\right)f_{m}f_{m}^{*}\left(\sum_{j\in[h]}e_{j}e_{j}^{*}\right)=\sum_{i,j\in[h]}Q_{x_{0},a_{0}}P_{2}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}$
$=\sum_{\begin{subarray}{c}i\in L\cap A_{m}\\ j\in A_{m}\end{subarray}}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}$
and
$f_{m}f_{m}^{*}Q_{x_{0},a_{0}}P_{2}$ $=\left(\sum_{i\in[h]}e_{i}e_{i}^{*}\right)f_{m}f_{m}^{*}\left(\sum_{j\in[h]}e_{j}e_{j}^{*}\right)Q_{x_{0},a_{0}}P_{2}=\sum_{i,j\in[h]}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}Q_{x_{0},a_{0}}P_{2}$
$=\sum_{\begin{subarray}{c}i\in A_{m}\\ j\in L\cap A_{m}\end{subarray}}e_{i}e_{i}^{*}f_{m}f_{m}^{*}e_{j}e_{j}^{*}.$
Hence, if $A_{m}\subseteq L$ or if $L\cap A_{m}=\emptyset$, it holds that $Q_{x_{0},a_{0}}P_{2},f_{m}f_{m}^{*}]=0$. On the other hand, we know from eq. 6 that
$0\neq[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=\sum_{m\in M}[Q_{x_{0},a_{0}}P_{2},f_{m}f_{m}^{*}],$
whence we deduce that there must exist some $\bar{m}\in M$ for which $A_{\bar{m}}\not\subseteq L$ and $A_{\bar{m}}\not\subseteq[h]\setminus L$.
We now choose $\xi_{1}\in A_{\bar{m}}\cap L$ and $\xi_{2}\in A_{\bar{m}}\cap([h]\setminus L)$. We claim that
(7) $\xi_{i}\in[h]\setminus Z\quad\mbox{ for both }i=1\mbox{ and }i=2.$
Indeed, suppose for the sake of contradiction that $\xi_{i}\in Z$ for some $i\in[2]$. This would imply that $P_{2}e_{\xi_{i}}=0$, whence in turn we could deduce that
$\sum_{m\in M}f_{m}f_{m}^{*}e_{\xi_{i}}=Q_{x_{\omega},a_{\omega}}P_{2}e_{\xi_{i}}=0.$
It would follow that ${\langle e_{\xi_{i}},f_{m}\rangle}=0$ for each $m\in M$. In particular, $\xi_{i}\not\in A_{\bar{m}}$, which is a contradiction. Hence, (7) is true.
Let now $a_{0}^{\prime}\in A$ be such that $Q_{x_{0},a_{0}^{\prime}}e_{\xi_{2}}=e_{\xi_{2}}$. Notice that there exists exactly one such $a_{0}^{\prime}$, since $Q_{x_{0}}$ is a PVM diagonalised by $\mathcal{E}$. We claim that $a_{0}^{\prime}\neq a_{0}$. Suppose, for the sake of contradiction, that $a_{0}^{\prime}=a_{0}$. Then, it would hold that $Q_{x_{0},a_{0}}e_{\xi_{2}}=e_{\xi_{2}}$. On the other hand, we have
(8) $0=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}e_{\xi_{2}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{2}}=P_{2}Q_{x_{0},a_{0}}e_{\xi_{2}}=P_{2}e_{\xi_{2}}=e_{\xi_{2}},$
where the first equality is true since $\xi_{2}\not\in L$, the second is true by definition of $L$, the third by eq. 4, the fourth by the observation above, and the fifth by eq. 7. This is a contradiction, so the claimed disequality $a_{0}^{\prime}\neq a_{0}$ must hold.
Fix now some $\alpha\in[\omega-1]$, and recall that the PVM $Q_{x_{\alpha}}$ is diagonalised by $\mathcal{E}$. Hence, there exists exactly one element $c_{1}\in A$ such that $Q_{x_{\alpha},c_{1}}e_{\xi_{1}}=e_{\xi_{1}}$, and exactly one element $c_{2}\in A$ such that $Q_{x_{\alpha},c_{2}}e_{\xi_{2}}=e_{\xi_{2}}$. We claim that $c_{1}=c_{2}$. Suppose this is not the case. It is time to remember that we chose $\mathcal{F}$ in such a way that it diagonalises both $Q_{x_{\omega},a_{\omega}}P_{2}$ and all the matrices $Q_{x_{\alpha},d}P_{2}$ for $d\in A$. Hence, we can write $Q_{x_{\alpha},c_{1}}P_{2}=\sum_{p\in V_{1}}f_{p}f_{p}^{*}$ and $Q_{x_{\alpha},c_{2}}P_{2}=\sum_{p\in V_{2}}f_{p}f_{p}^{*}$ for some $V_{1},V_{2}\subseteq[h]$. Since $Q_{x_{\alpha}}$ is a PVM, we have $Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}=0$. Using eq. 4 and 28, we deduce that
$0=Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}P_{2}=Q_{x_{\alpha},c_{1}}P_{2}Q_{x_{\alpha},c_{2}}P_{2},$
so $V_{1}\cap V_{2}=\emptyset$. On the other hand, using eq. 7, we have
$e_{\xi_{1}}$ $=Q_{x_{\alpha},c_{1}}e_{\xi_{1}}=Q_{x_{\alpha},c_{1}}P_{2}e_{\xi_{1}}=\sum_{p\in V_{1}}f_{p}f_{p}^{*}e_{\xi_{1}}\quad\mbox{and}$
$e_{\xi_{2}}$ $=Q_{x_{\alpha},c_{2}}e_{\xi_{2}}=Q_{x_{\alpha},c_{2}}P_{2}e_{\xi_{2}}=\sum_{p\in V_{2}}f_{p}f_{p}^{*}e_{\xi_{2}}.$
Since we are assuming that $\{\xi_{1},\xi_{2}\}\subseteq A_{\bar{m}}$, we must have $f_{\bar{m}}^{*}e_{\xi_{1}}\neq 0$ and $f_{\bar{m}}^{*}e_{\xi_{2}}\neq 0$. But this means that $\bar{m}\in V_{1}\cap V_{2}$, a contradiction. This proves the claimed equality $c_{1}=c_{2}$. Call such element $a_{\alpha}$.
Observe now that
$Q_{x_{0},a_{0}}e_{\xi_{1}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{1}}=e_{\xi_{1}},$
where the first equality is due to eq. 7 and the second to the fact that $\xi_{1}\in L$ and, for each $\ell\in L$, $e_{\ell}$ is an eigenvector of $Q_{x_{0},a_{0}}P_{2}$ with eigenvalue $1$. It follows that, for each $i\in[\omega-1]$,
$Q_{x_{0},a_{0}}Q_{x_{i},a_{i}}e_{\xi_{1}}=Q_{x_{0},a_{0}}e_{\xi_{1}}=e_{\xi_{1}},$
which means that $Q_{x_{0},a_{0}}Q_{x_{i},a_{i}}\neq 0$. Similarly,
$Q_{x_{0},a_{0}^{\prime}}Q_{x_{i},a_{i}}e_{\xi_{2}}=Q_{x_{0},a_{0}^{\prime}}e_{\xi_{2}}=e_{\xi_{2}},$
so $Q_{x_{0},a_{0}^{\prime}}Q_{x_{i},a_{i}}\neq 0$. Moreover, clearly, $Q_{x_{i},a_{i}}Q_{x_{j},a_{j}}\neq 0$ for each $i\neq j\in[\omega-1]$ since $e_{\xi_{1}}$ is an eigenvector of both with eigenvalue $1$.
Observe now that
$Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=f_{\bar{m}}$
since $\bar{m}\in M$ and, for each $m\in M$, $f_{m}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}P_{2}$ with eigenvalue $1$. Hence, for each $i\in[\omega-1]$,
$e_{\xi_{1}}^{*}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=e_{\xi_{1}}^{*}f_{\bar{m}}\neq 0$
since $\xi_{1}\in A_{\bar{m}}$. In particular, this means that $Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}\neq 0$. Analogously,
$e_{\xi_{2}}^{*}Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=e_{\xi_{2}}^{*}f_{\bar{m}}\neq 0$
since $\xi_{2}\in A_{\bar{m}}$. Hence, $Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}\neq 0$. Note also that the $Q_{x_{0},a_{0}}Q_{x_{\omega},a_{\omega}}=0$ is trivial from the fact that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
To conclude, we claim that there exists some $a_{\omega}^{\prime}\in A$ such that $a_{\omega}^{\prime}\neq a_{\omega}$ and
(9) $\left(\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\right)Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0.$
Note that this would be enough to finish the proof. Indeed, it follows from the minimality of $\operatorname{dist}(x,x^{\prime})$ that (9) guarantees that $Q_{x_{\omega},a_{\omega}^{\prime}}$ commutes with $Q_{x_{i},a_{i}}$ for each $0\leq i\leq\omega-1$.
Suppose, for the sake of contradiction, that the claim is false. Then, the fact that $Q_{x_{\omega}}$ is a PVM would imply that
$\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}$ $=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\operatorname{id}=\sum_{d\in A}\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},d}$
$=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}.$
But then, we would have
$\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}e_{\xi_{1}}=e_{\xi_{1}}$
so, in particular,
$\|\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|=1.$
However, we know from eq. 7 that $\xi_{1}\not\in Z$, which by the definition of $Z$ implies that $Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\neq e_{\xi_{1}}$. Since $Q_{x_{\omega},a_{\omega}}$ is a projector, this means that $\|Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|<1$. Since
$\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}$
is also a projector, it follows that
$\|\prod_{i\in[\omega-1]}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|<1,$
a contradiction. Hence, the claim is true.
Putting all together, we conclude that the vertices
$\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}$
form a bifurcation for $Q$ of length $\omega$, and the proof is complete. ∎
[…]
Figure 2. A bifurcation of length $\omega$. Solid lines mean non-orthogonality.
6. Quantum polymorphisms of odd cycles
The goal of this section is to prove the following.
Theorem 29.
For each odd $m\geq 3$, it holds that $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$.
We begin with a simple consequence of Theorem 26, which handles the case of quantum polymorphisms of arity 11—i.e., quantum endomorphisms.
Corollary 30.
$\operatorname{qHom}(\mathbb{C}_{m},\mathbb{C}_{m})=\operatorname{\oplus Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ for each odd $m\geq 3$.
Proof.
Suppose, for the sake of contradiction, that $\operatorname{qHom}(\mathbb{C}_{m},\mathbb{C}_{m})\neq\operatorname{\oplus Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$, and let $Q$ be a contextual quantum homomorphism $\mathbb{C}_{m}\to_{q}\mathbb{C}_{m}$. Applying Theorem 26, we deduce that there exists a bifurcation for $Q$ of length $\omega\leq\frac{m-1}{2}$. Consider, in particular, the initial elements $(x_{0},a_{0}),(x_{0},a_{0}^{\prime}),(x_{1},a_{1}),(x_{2},a_{2})$ of such bifurcation. By Definition 25 we know that $Q_{x_{0},a_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{0},a^{\prime}_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$, $Q_{x_{0},a_{0}}Q_{x_{2},a_{2}}\neq 0$, and $Q_{x_{0},a_{0}^{\prime}}Q_{x_{2},a_{2}}\neq 0$. Using $($$\mathbf{QH_{1}}$)), we deduce that $(a_{0},a_{1}),(a_{0}^{\prime},a_{1})$, and $(a_{1},a_{2})$ are all edges of $\mathbb{C}_{m}$. Note also that $x_{0}$ and $x_{2}$ lie on an (m2)(m-2)-walk in $\mathbb{X}$. Since no walk of such length connects a vertex to itself in $\mathbb{C}_{m}$Proposition 27 ensures that $a_{0}\neq a_{2}$ and $a_{0}^{\prime}\neq a_{2}$. Finally $a_{0}\neq a_{0}^{\prime}$ again by Definition 25. We deduce that $a_{1}$ has degree at least three in $\mathbb{C}_{m}$, a contradiction. ∎
We point out that a different proof of Corollary 30 was obtained in [Zeman] via a result from [banica2007quantum] on the quantum symmetry of vertex-transitive graphs. In the following, we fix an odd integer $m\in\mathbb{N}$. We start with the following observation.
Lemma 31.
Let $Q$ be a quantum endomorphism of $\mathbb{C}_{m}$ and let $x,y,z$ be a $2$-walk in $\mathbb{C}_{m}$. Then $Q_{x,a}Q_{y,b}Q_{z,c}=Q_{y,b}Q_{z,c}$ for each $a,b,c\in\mathbb{Z}_{m}$.
Proof.
If $x=z$ the result is trivial so we may assume that $x\neq z$. We know from Corollary 30 that there exists $d\in\mathbb{N}$ and $h_{1},\dots h_{d}\in\operatorname{Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ such that $Q=\bigoplus_{i\in[d]}h_{i}$. Using eq. 1, we can write
$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\left(\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right)\left(\sum_{\begin{subarray}{c}j\in[d]\\ h_{j}(y)=b\end{subarray}}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\right)\left(\sum_{\begin{subarray}{c}k\in[d]\\ h_{k}(z)=c\end{subarray}}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)$
$=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}.$
Let $i\in[d]$ give a non-trivial contribution to the sum above. Since $x,y,z$ form a path in $\mathbb{C}_{m}$, any choice of $(b,c)\in\mathbb{Z}_{m}^{2}$ for which $h_{i}(y)=b$ and $h_{i}(z)=c$ gives precisely one choice of $a\in\mathbb{Z}_{m}$ such that $h_{i}(x)=a$. It follows that
$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}=Q_{y,b}Q_{z,c},$
as needed. ∎
Next, we focus on quantum polymorphisms of arity at least two. Henceforth in this section, we fix some arity $n\geq 2$. Furthermore, whenever a plus or minus sign is applied to an element of $\mathbb{Z}_{m}$ or to a tuple of such elements, it shall be meant modulo $m$.
Lemma 32.
Fix two vertices $\mathbf{x},\mathbf{y}$ of $\mathbb{C}_{m}^{n}$ such that $\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})=2$. For any $a\in\mathbb{Z}_{m}$, it holds that
$[Q_{\mathbf{x},a},Q_{\mathbf{y},a}]=0.$
Proof.
In this proof, we shall abbreviate $Q_{\mathbf{a},a}$ to $Q_{\mathbf{a}}$ for each $\mathbf{a}\in[m]^{n}$. Let $\mathbf{x}\to\mathbf{z}\to\mathbf{y}$ be a $2$-path connecting $\mathbf{x}$ and $\mathbf{y}$ in $\mathbb{C}_{m}^{n}$. Let also $\mathbf{u}=\mathbf{z}-\mathbf{x}$ and $\mathbf{v}=\mathbf{y}-\mathbf{z}$, and notice that $\mathbf{u},\mathbf{v}\in\{\pm 1\}^{n}$. Observe that the vertices $\mathbf{y},\mathbf{y}+\mathbf{v},\mathbf{y}+2\mathbf{v},\dots,\mathbf{y}+(m-1)\mathbf{v}$ induce a $\mathbb{C}_{m}$ within $\mathbb{C}_{m}^{n}$. The restriction of $Q$ to such vertices is a quantum endomorphism of $\mathbb{C}_{m}$. Therefore, combining Corollary 30 and Lemma 7, we obtain that
$\sum_{i\in\mathbb{Z}_{m}}Q_{\mathbf{y}+i\mathbf{v}}=\operatorname{id}.$
Consider the two paths
$\mathbf{x}\to\mathbf{x}+\mathbf{u}\to\mathbf{y}\to\mathbf{y}+\mathbf{v}\to\mathbf{y}+2\mathbf{v}\to\dots\to\mathbf{y}+(m-4)\mathbf{v},$
$\mathbf{x}\to\mathbf{x}+\mathbf{u}\to\mathbf{y}-2\mathbf{v}\to\mathbf{y}-3\mathbf{v}\to\dots\to\mathbf{y}-(m-2)\mathbf{v},$
and notice that both have length m2m-2. Since there is no closed walk in $\mathbb{C}_{m}$ of odd length less than $m$, it follows from Proposition 27 that $Q_{\mathbf{x}}Q_{\mathbf{y}+i\mathbf{v}}=0$ for each $i\in\mathbb{Z}_{m}\setminus\{0,-2\}$. Therefore,
$0$ $=Q_{\mathbf{y}+\mathbf{v}}Q_{\mathbf{x}}=\left(\operatorname{id}-\sum_{i\in\mathbb{Z}_{m}\setminus\{1\}}Q_{\mathbf{y}+i\mathbf{v}}\right)Q_{\mathbf{x}}=Q_{\mathbf{x}}-(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}.$
Hence, $Q_{\mathbf{x}}=(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}$. Since projectors are Hermitian, it follows that
$[Q_{\mathbf{x}},Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}}]=0.$
Consider now the automorphism $\xi$ of $\mathbb{C}_{m}^{n}$ that acts as the reflection around $x_{\alpha}$ for each coordinate $\alpha\in[n]$ for which $u_{\alpha}=v_{\alpha}$, and acts as the identity for all other coordinates. Note that both $\mathbf{x}$ and $\mathbf{x}+\mathbf{u}-\mathbf{v}=\mathbf{y}-2\mathbf{v}$ are fixed points of $\xi$, while $\xi(\mathbf{y})=2\mathbf{x}-\mathbf{y}$. Since the previous part of the argument is invariant under (classical) automorphisms of $\mathbb{C}_{m}^{n}$, we deduce that
$[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}}]=0.$
By the bilinearity of the commutator, it follows that
$[Q_{\mathbf{x}},Q_{\mathbf{y}}]=[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}].$
We now apply the same argument after applying the automorphism of $\mathbb{C}_{m}^{n}$ given by $\mathbf{a}\mapsto\mathbf{a}+\mathbf{y}-\mathbf{x}$. We obtain
$[Q_{\mathbf{x}},Q_{\mathbf{y}}]=[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}]$ $=[Q_{3\mathbf{x}-2\mathbf{y}},Q_{2\mathbf{x}-\mathbf{y}}]=[Q_{3\mathbf{x}-2\mathbf{y}},Q_{4\mathbf{x}-3\mathbf{y}}]=\dots$
$=[Q_{m\mathbf{x}-(m-1)\mathbf{y}},Q_{(m+1)\mathbf{x}-m\mathbf{y}}]=[Q_{\mathbf{y}},Q_{\mathbf{x}}],$
whence we conclude that $Q_{\mathbf{x}},Q_{\mathbf{y}}]=0$, as required. ∎
Lemma 33.
Fix two vertices $\mathbf{x},\mathbf{y}$ of $\mathbb{C}_{m}^{n}$ such that $\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})=2$. For any $a,b\in\mathbb{Z}_{m}$, it holds that
$[Q_{\mathbf{x},a},Q_{\mathbf{y},b}]=0.$
Proof.
We can assume without loss of generality that $b=a\pm 2$ $\mod m$, as otherwise $Q_{\mathbf{x},a}Q_{\mathbf{y},b}=0$ by Proposition 27. Up to symmetry, we can further suppose that $b=a+2$ $\mod m$. Let $\mathbf{u}$ and $\mathbf{v}$ be as in the proof of Lemma 32, and observe that
$Q_{\mathbf{x},a}Q_{\mathbf{y},a+2}$ $=Q_{\mathbf{x},a}\left(\sum_{b\in\mathbb{Z}_{m}}Q_{\mathbf{y}-\mathbf{v},b}\right)\left(\sum_{c\in\mathbb{Z}_{m}}Q_{\mathbf{y}-2\mathbf{v},c}\right)Q_{\mathbf{y},a+2}$
$=\sum_{b,c\in\mathbb{Z}_{m}}Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},b}Q_{\mathbf{y}-2\mathbf{v},c}Q_{\mathbf{y},a+2}.$
Since $Q_{\mathbf{y}-\mathbf{v},b},Q_{\mathbf{y}-2\mathbf{v},c}]=0$, for a product in the summation above to be non-zero, it must be the case that $b$ is a neighbour of both $a$ and $a+2$; i.e., $b=a+1$. Consider now $Q_{\mathbf{y}-2\mathbf{v},c}$. For the product to be non-zero, since $\mathbf{y}-2\mathbf{v}$ is a $\mathbb{C}_{m}^{n}$-neighbour of $\mathbf{y}-\mathbf{v}$, it must hold that $c$ is a $\mathbb{C}_{m}$-neighbour of $b$; i.e., $c=a$ or $c=a+2$. Suppose $Q_{\mathbf{y}-2\mathbf{v},a+2}\neq 0$. We know from Lemma 32 that $Q_{\mathbf{y}-2\mathbf{v},a+2},Q_{\mathbf{y},a+2}]=0$ and, hence, the two projectors can be diagonalised simultaneously. Hence, if it was the case that $Q_{\mathbf{y}-2\mathbf{v},a+2}Q_{\mathbf{y},a+2}\neq 0$, the ranges of $Q_{\mathbf{y}-2\mathbf{v},a+2}$ and of $Q_{\mathbf{y},a+2}$ would have non-trivial intersection. But this would imply that $\sum_{i\in\mathbb{Z}_{m}}Q_{\mathbf{y}+i\mathbf{v},a+2}\neq\operatorname{id}$, contradicting Corollary 30 and Lemma 7. It follows that the only choice for $c$ giving a non-trivial contribution is $c=a$, so
$Q_{\mathbf{x},a}Q_{\mathbf{y},a+2}=Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y},a+2}.$
By a similar argument, we have
$Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}$ $=Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y},a+2}.$
Hence, we obtain
$Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}Q_{\mathbf{y},a+2}$ $=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y},a+2}$
$=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}$
$=Q_{\mathbf{y},a+2}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{x},a}$
(where we are using that $Q_{\mathbf{x},a},Q_{\mathbf{y}-\mathbf{v},a+1}]=0$ as $\mathbf{x}\sim\mathbf{y}-\mathbf{v}$ and $Q_{\mathbf{x},a},Q_{\mathbf{y}-2\mathbf{v},a}]=0$ by Lemma 32)
$=Q_{\mathbf{y},a+2}Q_{\mathbf{y}-2\mathbf{v},a}Q_{\mathbf{y}-\mathbf{v},a+1}Q_{\mathbf{x},a}$
$=(Q_{\mathbf{x},a}Q_{\mathbf{y},a+2})^{*}=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}.$
In particular, we deduce that $Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}$ is a Hermitian matrix, so
$[Q_{\mathbf{x},a},Q_{\mathbf{y},a+2}]=0,$
as needed. ∎
We can now finalise the proof of the main result of this section.
Proof of Theorem 29.
Let $\mathbf{x}$ and $\mathbf{y}$ be two vertices of $\mathbb{C}_{m}^{n}$ and let $a,b$ be two vertices of $\mathbb{C}_{m}$. We need to show that $Q_{\mathbf{x},a},Q_{\mathbf{y},b}]=0$. We shall use induction on $\delta=\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})$. If $\delta\in\{0,1\}$, the result is clear from the definition of quantum homomorphisms. If $\delta=2$, the result is Lemma 33. Let then $\delta\geq 3$, and suppose that the result holds for all pairs of vertices at mutual distance at most $\delta-1$. Let $\mathbf{x},\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{\delta-1},\mathbf{y}$ be a shortest path from $\mathbf{x}$ to $\mathbf{y}$ in $\mathbb{C}_{m}^{n}$. Since $\delta\geq 3$, we can assume without loss of generality that the vectors $\mathbf{z}_{1}-\mathbf{x}$ and $\mathbf{z}_{2}-\mathbf{z}_{1}$ are equal. [^2] Note that this is false in general when $\operatorname{dist}_{\mathbb{C}_{m}^{n}}(\mathbf{x},\mathbf{y})=2$. For instance, consider the case that $n=2$, $m\geq 5$, $\mathbf{x}=(0,0)$, and $\mathbf{y}=(0,2)$. We now expand the product $Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ as follows:
$Q_{\mathbf{x},a},Q_{\mathbf{y},b}$ $=Q_{\mathbf{x},a}\left(\sum_{c_{1}\in\mathbb{Z}_{m}}Q_{\mathbf{z}_{1},c_{1}}\right)\left(\sum_{c_{2}\in\mathbb{Z}_{m}}Q_{\mathbf{z}_{2},c_{2}}\right)\dots\left(\sum_{c_{\delta-1}\in\mathbb{Z}_{m}}Q_{\mathbf{z}_{\delta-1},c_{\delta-1}}\right)Q_{\mathbf{y},b}$
$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}Q_{\mathbf{y},b}.$
Note now that $\operatorname{dist}(\mathbf{z}_{i},\mathbf{y})\leq\delta-1$ for each $i\in[\delta-1]$. Hence, the inductive hypothesis allows us to freely permute the corresponding projectors in the expression above, thus yielding
(10) $Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$
Multiplying both terms by $Q_{\mathbf{x},a}$ on the right and further rearranging gives
$Q_{\mathbf{x},a}Q_{\mathbf{y},b}Q_{\mathbf{x},a}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}Q_{\mathbf{x},a}$
$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}$
Observe now that $\mathbf{x}$, $\mathbf{z}_{1}$, and $\mathbf{z}_{2}$ lie on a common $m$-cycle within $\mathbb{C}_{m}^{n}$ since $\mathbf{z}_{1}-\mathbf{x}=\mathbf{z}_{2}-\mathbf{z}_{1}$. By Lemma 31, it holds that $Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}=Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}$ for each $c_{1},c_{2}\in\mathbb{Z}_{m}$. We deduce that
$Q_{\mathbf{x},a}Q_{\mathbf{y},b}Q_{\mathbf{x},a}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}=Q_{\mathbf{x},a}Q_{\mathbf{y},b},$
where the second equality is eq. 10. From this, we conclude that $Q_{\mathbf{x},a}$ and $Q_{\mathbf{y},b}$ commute, as required.
Corollary 34.
Let $m\geq 3$ be an odd natural number. Then $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$. In particular, $\mathbb{C}_{m}$ has a commutativity gadget and $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable.
Proof.
$\mathbb{C}_{m}$ has a commutativity gadget by Theorem 29 and Theorem 16.
Note that $\mathbb{K}_{m}$ is pp-definable in $\mathbb{C}_{m}$ with the gadget consisting of a simple path of length m2m-2. Since $\mathbb{C}_{m}$ has a commutativity gadget, it follows that $\mathbb{K}_{m}$ is q-definable in $\mathbb{C}_{m}$ by Proposition 13. By Proposition 10, $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ reduces to $\operatorname{CSP^{*}}(\mathbb{C}_{m})$, and $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable by Corollary 24. Thus, $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable. ∎
We conclude with the proof that quantum homomorphisms between different odd cycles can be contextual.
Proposition 35.
There is a quantum homomorphism $\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space that is not in $\oplus{\operatorname{Hom}(\mathbb{C}_{7},\mathbb{C}_{5})}$.
Proof.
Consider for this proof $\mathbb{C}_{5}$ to have domain $\{0,1,2,5,6\}$ with edges between adjacent numbers, and $\mathbb{C}_{7}$ to have domain $[7]$. Consider the two classical endomorphisms $h,g\colon\mathbb{C}_{5}\to\mathbb{C}_{5}$ such that $h(2)=6,h(5)=0$ and $g(5)=0,g(2)=1$. Then $h\oplus g$ is a quantum homomorphism $Q\colon\mathbb{C}_{5}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space. Extend $Q$ by setting $Q_{3,a}=Q_{5,a}$ for $a\in\{0,1,2,5,6\}$. Choose a basis in which all the projectors are diagonal, and in this basis define $Q_{4,1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$ and $Q_{4,6}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$. Then $Q\colon\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ and $Q_{4,0},Q_{2,1}]\neq 0$. ∎
7. The non-oracular setting
In this section, we consider the non-oracular version of the results in Section 3. A non-oracular quantum homomorphism between two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, denoted by $Q\colon\mathbb{A}\operatorname{\to_{qno}}\mathbb{B}$, is a quantum function $Q\colon A\to_{q}B$ satisfying the condition $($$\mathbf{QH_{1}}$)) but possibly not $($$\mathbf{QH_{2}}$)).
The following is the non-oracular version of the entangled CSP.
Definition 36 (Non-oracular entangled CSP).
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP^{**}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a non-oracular quantum homomorphism $\mathbb{X}\operatorname{\to_{qno}}\mathbb{A}$.
We now define quantum polymorphisms in the non-oracular setting.
Definition 37.
Let $\mathbb{A}$ be a relational structure. A non-oracular quantum polymorphism is a non-oracular quantum homomorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$. We denote by $\operatorname{qPol^{no}}(\mathbb{A})$ the set of non-oracular quantum polymorphisms of $\mathbb{A}$.
We will obtain a quantisation of the Galois connection between classical polymorphisms and relations. The following is the main result of this section.
Theorem 38.
Let $\mathbb{A},\mathbb{B}$ be structures on potentially different signatures such that
$\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B}).$
Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
We begin by defining the non-oracular version of q-definitions.
Definition 39.
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A non-oracular q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
  1. $(\mathbf{noq_{1}})$
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
  2. $(\mathbf{noq_{2}})$
    For every non-oracular quantum homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}_{S}$.
Note that, in the first item of the definition above, $Q$ is required to be oracular (i.e., to satisfy $($$\mathbf{QH_{2}}$))), while $Q^{\prime}$ need not.
Definition 40.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{B}$ is q-no-definable from $\mathbb{A}$ if there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
In the following, it shall be useful to observe that Lemma 2, Lemma 7Lemma 4, and Lemma 6 also hold in the non-oracular setting. Furthermore, the definition of quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ (Definition 3) does not change in the non-oracular setting.
The following result is the key to the quantisation of the relation/operation Galois connection in the non-oracular quantum setting.
Theorem 41.
Let $\mathbb{A}$ be a finite structure and let $S\subseteq A^{r}$ for some $r\in\mathbb{N}$. Then there exists a non-oracular q-definition for $\mathbb{A}$ and $S$ if, and only if, $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{A}_{S})$.
Proof.
Henceforth in this proof, we shall denote $\mathbb{A}_{S}$ by $\mathbb{D}$ for the sake of readability.
  • $(\Rightarrow)$
    Let $\mathbb{G}$ be a non-oracular q-definition for $\mathbb{A}$ and $S$. This means that every quantum homomorphism $\mathbb{R}\to_{q}\mathbb{D}$ extends to a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, and every non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ restricts to a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{D}$.
    Fix now a non-oracular quantum polymorphism $Q$ of $\mathbb{A}$; i.e., $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$. We want to prove that $Q\in\operatorname{qPol^{no}}(\mathbb{D})$. Note that $Q$ is a quantum function $A^{n}\to_{q}A$, so we only need to prove $($$\mathbf{QH_{1}}$)) applied to the relation of $\mathbb{D}^{n}$; i.e., to the relation $R^{\mathbb{D}^{n}}$ containing all $r$-tuples $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})$ where each $\mathbf{b}^{(i)}$ is an element of $A^{n}$, and for each $j\in[n]$ it holds that $(b^{(1)}_{j},\dots,b^{(r)}_{j})\in R^{\mathbb{D}}=S$.
    Fix then such a tuple $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})\in R^{\mathbb{D}^{n}}$, and choose a tuple $\mathbf{c}\in A^{r}\setminus S$. For every $i\in[n]$, consider the deterministic homomorphism $f_{i}:\mathbb{R}\to\mathbb{D}$ defined by $f_{i}(j)=b^{(j)}_{i}$ for each $j\in[r]$. In particular, $f_{i}$ is a quantum homomorphism. Hence, ($(\mathbf{noq_{1}})$) implies that there exists a non-oracular quantum homomorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $f_{i}$. Consider now the tensor product
    $W=\bigotimes_{i\in[n]}W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}^{n},$
    and compose it with $Q$ to obtain $Q\bullet W:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$. Fix $\ell\in[r]$ and $a\in A$. Expanding the composition and the tensor product yields
    $(Q\bullet W)_{g_{\ell},a}=\sum_{\mathbf{a}\in A^{n}}Q_{\mathbf{a},a}\otimes W_{g_{\ell},\mathbf{a}}=\sum_{\mathbf{a}\in A^{n}}Q_{\mathbf{a},a}\otimes\bigotimes_{i\in[n]}W^{(i)}_{g_{\ell},a_{i}}$.
    We know from ($(\mathbf{noq_{1}})$) that $W^{(i)}_{g_{\ell},a_{i}}=(f_{i})_{\ell,a_{i}}$. Since $f_{i}$ is deterministic, this means that in order for the $\mathbf{a}$-th summand in the sum above to be non-zero it must be that $a_{i}=f_{i}(\ell)=b_{i}^{(\ell)}$ for each $i\in[n]$; i.e., $\mathbf{a}=\mathbf{b}^{(\ell)}$. Hence,
    $(Q\bullet W)_{g_{\ell},a}=Q_{\mathbf{b}^{(\ell)},a}\otimes\operatorname{id}.$
    On the other hand, ($(\mathbf{noq_{2}})$) implies that the assignment $(\ell,a)\mapsto(Q\bullet W)_{g_{\ell},a}$ yields a non-oracular quantum homomorphism from $\mathbb{R}$ to $\mathbb{D}$. Since we are assuming that $\mathbf{c}\not\in S=R^{\mathbb{D}}$, we deduce that
    $0=\prod_{\ell\in[r]}(Q\bullet W)_{g_{\ell},c_{\ell}}=\prod_{\ell\in[r]}(Q_{\mathbf{b}^{(\ell)},a}\otimes\operatorname{id})=\left(\prod_{\ell\in[r]}Q_{\mathbf{b}^{(\ell)},a}\right)\otimes\operatorname{id}.$
    Therefore, we find $\prod_{\ell\in[r]}Q_{\mathbf{b}^{(\ell)},a}=0$. It follows that $Q$ is a non-oracular quantum homomorphism from $\mathbb{D}^{n}$ to $\mathbb{D}$; i.e., an element of $\operatorname{qPol^{no}}(\mathbb{D})$, as desired.
  • $(\Leftarrow)$
    Suppose now that $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{D})$. Let us enumerate the tuples in $S$ as $\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}$, where $n=|S|$. We will show that $\mathbb{A}^{n}$ is a non-oracular quantum gadget for $\mathbb{A}$ and $S$, with distinguished vertices $g_{1},\dots,g_{r}$ defined by $g_{\ell}=(a^{(1)}_{\ell},\dots,a^{(n)}_{\ell})$ for $\ell\in[r]$. We show that the two conditions in Items $(\mathbf{noq_{1}})$ and $(\mathbf{noq_{2}})$ hold.
    Let $Q\colon\mathbb{R}\to_{q}\mathbb{D}$ be a quantum homomorphism over some Hilbert space $H$, and let $d$ be the dimension of $H$. Since $Q$ satisfies $($$\mathbf{QH_{2}}$)), $Q$ must be classical, so by the spectral theorem we can decompose it as $Q=\bigoplus_{i\in[d]}h_{i}$, where $h_{i}\colon\mathbb{R}\to\mathbb{D}$ is a deterministic homomorphism for each $i\in[d]$. By the definition of $\mathbb{R}$, it must be that $(h_{i}(1),\dots,h_{i}(r))\in S$ for each $i\in[d]$. In other words, there exists a function $f:[d]\to[n]$ satisfying
    (11) $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{(f(i))}$
    for each $i\in[d]$. Consider now the deterministic dictator polymorphisms $\pi^{n}_{j}:\mathbb{A}^{n}\to\mathbb{A}$ defined by $\pi^{n}_{j}(\mathbf{a})=a_{j}$ for each $j\in[n]$ and each $\mathbf{a}\in A^{n}$. Define
    $P=\bigoplus_{i\in[d]}\pi^{n}_{f(i)}.$
    Note that $P$ is a classical (a fortiori, non-oracular quantum) homomorphism from $\mathbb{A}^{n}$ to $\mathbb{A}$. Take now $\ell\in[r]$ and $a\in A$, and observe that
    $P_{g_{\ell},a}=\bigoplus_{i\in[d]}(\pi^{n}_{f(i)})_{g_{\ell},a}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ \pi^{n}_{f(i)}(g_{\ell})=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ a^{f(i)}_{\ell}=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ h_{i}(\ell)=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\end{subarray}}(h_{i})_{\ell,a}=Q_{\ell,a}$
    (where the fourth equality follows from Equation 11). Hence, $P$ witnesses that ($(\mathbf{noq_{1}})$) holds.
    Finally, let $Q\colon\mathbb{\mathbb{A}}^{n}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism, and consider the quantum function $Q^{\prime}$ given by $Q^{\prime}_{\ell,a}=Q_{g_{\ell},a}$ for $\ell\in[r]$ and $a\in A$. Recall now that each $\mathbf{a}^{(i)}$ is an element of $S=R^{\mathbb{D}}$ and, thus, the tuples $(g_{1},\dots,g_{r})$ belongs to $R^{\mathbb{D}^{n}}$ by the definition of categorical power of structures. Since $Q$ is a non-oracular quantum polymorphism of $\mathbb{A}$, by assumption it is also an $n$-ary non-oracular quantum polymorphism of $\mathbb{D}$. Hence, for each $\mathbf{c}\in A^{r}\setminus S$, we must have
    $0=\prod_{\ell\in[r]}Q_{g_{\ell},c_{\ell}}=\prod_{\ell\in[r]}Q^{\prime}_{\ell,c_{\ell}},$
    which means that $Q^{\prime}$ is a non-oracular quantum homomorphism from $\mathbb{R}$ to $\mathbb{D}$. This shows that ($(\mathbf{noq_{2}})$) is also satisfied, and concludes the proof of the theorem.∎
The following result is an immediate consequence of Theorem 41.
Corollary 42.
Two structures $\mathbb{A},\mathbb{B}$ satisfy $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ if and only if $\mathbb{B}$ is q-no-definable from $\mathbb{A}$.
Unlike in the oracular setting, non-oracular q-gadgets are not strong enough to guarantee reductions. To that end, they need to be strengthened via a non-oracular commutativity gadget, which we now define.
Definition 43.
Fix a $\sigma$-structure $\mathbb{A}$. A non-oracular commutativity gadget for $\mathbb{A}$ is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
  1. $(\mathbf{noc_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum non-oracular homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $Q$.
  2. $(\mathbf{noc_{2}})$
    For every quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
We now give the non-oracular analogue of Theorem 16. Just like in the oracular setting, the characterisation of commutativity gadgets via quantum polymorphisms (stated in the result below) also holds for the weaker algebraic non-oracular commutativity gadgets of [Zeman], which are defined by replacing ($(\mathbf{noc_{1}})$) in Definition 43 with the weaker condition stated below:
  1. $(\mathbf{noc_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
Theorem 44.
Let $\mathbb{A}$ be a finite structure. The following are equivalent:
  1. (1)
    $\mathbb{A}$ has a non-oracular commutativity gadget.
  2. (2)
    $\mathbb{A}$ has an algebraic non-oracular commutativity gadget.
  3. (3)
    $\mathbb{A}^{n}$ is an (algebraic) non-oracular commutativity gadget for all large enough $n\in\mathbb{N}$.
  4. (4)
    $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
We prove the following implications.
  • $(4.\Rightarrow 3.)$
    The proof follows from exactly the same argument as for Proposition 20.
  • $(3.\Rightarrow 1.)$
    Obvious.
  • $(1.\Rightarrow 2.)$
    This is clear, since each non-oracular commutativity gadget is in particular an algebraic non-oracular commutativity gadget (via the same argument as in Remark 12).
  • $(2.\Rightarrow 4.)$
    Take an algebraic non-oracular commutativity gadget $\mathbb{G}$ with distinguished elements $u,v$, a non-oracular quantum polymorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$, two tuples $\mathbf{a},\mathbf{b}\in A^{n}$, and two elements $c,d\in A$. We proceed as in Proposition 21. We first build a non-oracular quantum polymorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ for each $i\in[n]$ with the property that $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}$, using ($(\mathbf{noc_{1}^{\prime}})$). Then, we take the tensor product $W=\bigotimes_{i\in[n]}W^{(i)}$ of such homomorphisms, and we define $U:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ as the composition of $Q$ and $W$. On the one hand, ($(\mathbf{noq_{2}})$) guarantees that $U_{u,c},U_{v,d}]=0$. On the other hand, expanding the composition yields $U_{u,c}=Q_{\mathbf{a},c}\otimes\operatorname{id}$ and $U_{v,d}=Q_{\mathbf{b},d}\otimes\operatorname{id}$. Combining the two facts, we conclude that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we deduce that $Q$ is classical.∎
The next result shows that non-oracular q-definitions, combined with non-oracular commutativity gadgets, give logspace reductions from $\operatorname{CSP^{*}}$s to $\operatorname{CSP^{**}}$s.
Proposition 45.
Let $\mathbb{A}$ be a $\sigma$-structure and let $\mathbb{B}$ be a $\rho$-structure having the same domain (but potentially distinct signatures $\sigma$ and $\rho$), and suppose that
  • $\mathbb{A}$ has a non-oracular commutativity gadget, and
  • for every $R\in\rho$, there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$.
Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
Proof.
For every symbol $R\in\rho$ of some arity $r$, let $\mathbb{G}_{R}$ be a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ (with its distinguished vertices $g_{1},\dots,g_{r}$). Let also $\mathbb{H}$ be a non-oracular commutativity gadget for $\mathbb{A}$ with its distinguished vertices $u,v$.
Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$. Recall that $\operatorname{Gaif}(\mathbb{X})$ denotes the Gaifman digraph of $\mathbb{X}$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{**}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$, as well as a copy $\mathbb{H}^{(x,y)}$ of $\mathbb{H}$ for each edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$. For $R\in\rho$ of some arity $r\in\mathbb{N}$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$. Similarly, we denote by $u^{(x,y)}$ (resp. $v^{(x,y)}$) the copy of $u$ (resp. $v$) in $\mathbb{H}^{(x,y)}$. For every edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$, glue $u^{(x,y)}$ with $x$ and glue $v^{(x,y)}$ with $y$. For each $R\in\rho$ of some arity $r$, each $\mathbf{x}\in R^{\mathbb{X}}$, and each $i\in[r]$, glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to\mathbb{A}^{\prime}$. (Observe that this would not be guaranteed if $Q$ was a non-oracular quantum homomorphism.) Then, ($(\mathbf{noq_{1}})$) guarantees the existence of a non-oracular quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\operatorname{\to_{qno}}\mathbb{A}$ such that
$Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
for each $i\in[r]$ and each $a\in A$. Fix now an edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$. Using again the fact that $Q$ is oracular, we know that $Q_{x,a},Q_{y,b}]=0$ for every $a,b\in A$. Hence, we can apply ($(\mathbf{noc_{1}})$) to find a non-oracular quantum homomorphism
$Q^{(x,y)}\colon\mathbb{H}^{(x,y)}\operatorname{\to_{qno}}\mathbb{A}$
extending $Q$ in $x$ and $y$. Overall, this is a non-oracular quantum homomorphism $\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$ is a non-oracular quantum homomorphism. Consider the quantum function $Q^{\prime}:X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is an (oracular) quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Take $i,j\in[r]$. We claim that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute. This is easily true if $x_{i}=x_{j}$ (as in this case the PVMs coincide), so suppose that $x_{i}\neq x_{j}$. In this case, $(x_{i},x_{j})$ is an edge in $\operatorname{Gaif}(\mathbb{X})$. Note now that, by construction of $\mathbb{Y}$, the vertices $x_{i}$, $u^{(x_{i},x_{j})}$, and $g_{i}^{(\mathbf{x})}$ are identified in $Y$. Similarly, the vertices $x_{j}$, $v^{(x_{i},x_{j})}$, and $g_{j}^{(\mathbf{x})}$ are identified in $Y$. Since $Q$ induces a non-oracular quantum homomorphism over $\mathbb{H}^{(x_{i},x_{j})}$, we deduce from ($(\mathbf{noc_{2}})$) that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute also in this case. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). We now apply ($(\mathbf{noq_{2}})$) to deduce that the quantum function $U$ defined by $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}=Q^{\prime}_{x_{i},a}$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$). Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
(12) $\prod_{i\in[r]}Q_{x_{i},b_{i}}=0.$
(Note that for this to be true we do not need $U$ to be oracular.) This means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum (oracular) homomorphism from $\mathbb{X}$ to $\mathbb{B}$. ∎
We can finally complete the proof of the main result of this section.
Proof of Theorem 38.
Since $\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})$Theorem 44 guarantees that $\mathbb{A}$ has a non-oracular commutativity gadget. Moreover, we know from $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ and Corollary 42 that all the relations of $\mathbb{B}$ have a q-no-definition in $\mathbb{A}$. Note also that $\mathbb{A}$ and $\mathbb{B}$ have the same domain since $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$. Hence, the result directly follows from Proposition 45. ∎
We now show that tree gadgets can be used to produce q-no-definitions. A relational structure $\mathbb{X}$ is a tree if for every sequence of tuples $\mathbf{t}^{1},\dots,\mathbf{t}^{m}$ that all belong to (potentially different) relations $R_{i}$ of $\mathbb{X}$ with arity $r_{i}$, the total number of elements appearing in $\mathbf{t}^{1},\dots,\mathbf{t}^{m}$ is at least $1+\sum_{i=1}^{m}(r_{i}-1)$.
Lemma 46.
Let $\mathbb{T}$ be a tree structure, and let $x,y$ be elements of $\mathbb{T}$. Let $Q\colon\mathbb{T}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism. If $Q_{x,a}Q_{y,b}\neq 0$, there exists a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ such that $h(x)=a$ and $h(y)=b$.
Proof.
We prove this by induction on the size of $\mathbb{T}$. Assume first that $\mathbb{T}$ consists of a single constraint $(z_{1},\dots,z_{r})\in R^{\mathbb{T}}$ with $x=z_{i}$ and $y=z_{j}$ for some $i,j\in\{1,\dots,r\}$. For the argument below, we can assume without loss of generality that $i<j$ up to swapping $x$ and $y$. Then we have
$Q_{x,a}Q_{y,b}$ $=\sum_{c_{1},\dots,c_{r}\in A:c_{i}=a,c_{j}=b}\prod_{k=1}^{r}Q_{z_{k},c_{k}}$
$=\sum_{(c_{1},\dots,c_{r})\in R^{\mathbb{A}}:c_{i}=a,c_{j}=b}\prod_{k=1}^{r}Q_{z_{k},c_{k}}$ by $($$\mathbf{QH_{1}}$))
and since $Q_{x,a}Q_{y,b}\neq 0$, it must be that there exists $(c_{1},\dots,c_{r})\in R^{\mathbb{A}}$ such that $c_{i}=a$ and $c_{j}=b$. We obtain a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ by setting $h(z_{k})=c_{k}$ for all $k\in\{1,\dots,r\}$.
Otherwise, $\mathbb{T}$ consists of the union of some trees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$ that share a single vertex $z$. Since $\sum_{c}Q_{x,a}Q_{z,c}Q_{z,c}Q_{y,b}=Q_{x,a}(\sum_{c}Q_{z,c})Q_{y,b}=Q_{x,a}Q_{y,b}\neq 0$, there exists a $c$ such that $Q_{x,a}Q_{z,c}\neq 0$ and $Q_{z,c}Q_{y,b}\neq 0$. By induction hypothesis on the subtrees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$, there exist homomorphisms $h_{i}\colon\mathbb{T}_{i}\to\mathbb{A}$ such that $h_{i}(z)=c$ for all $i$, and such that if $x\in T_{i}$ (resp. $y\in\mathbb{T}_{i}$) then $h_{i}(x)=a$ (resp. $h_{i}(y)=b$). The union of these homomorphisms gives a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ with $h(x)=a$ and $h(y)=b$. ∎
Proposition 47.
Let $\mathbb{A}$ be a structure, and let $R\subseteq A^{2}$ be a relation with a classical gadget $\mathbb{G}$ that is a tree. Then $\mathbb{G}$ is a non-oracular q-definition for $R$.
Proof.
Suppose that $Q\colon\mathbb{R}\to_{q}(A;R)$ is a quantum homomorphism. Then we have $Q=h_{1}\oplus\dots\oplus h_{d}$, where each $h_{i}$ is a map $\{1,\dots,r\}\to A$ such that $(h_{i}(1),h_{i}(2))\in R$. Thus, each $h_{i}$ can be extended to a homomorphism $g_{i}\colon\mathbb{G}\to\mathbb{A}$ since $\mathbb{G}$ is a classical gadget for $R$. It follows that $Q^{\prime}:=g_{1}\oplus\dots\oplus g_{d}$ is a (non-oracular) quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ satisfying the conclusion of ($(\mathbf{noq_{1}})$).
To prove ($(\mathbf{noq_{2}})$), let $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism. Let $x_{1},x_{2}$ be the distinguished vertices of $\mathbb{G}$. Suppose that $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$. Since $\mathbb{G}$ is a tree, there exists by Lemma 46 a homomorphism $h\colon\mathbb{G}\to\mathbb{A}$ such that $h(x_{i})=a_{i}$ for $i\in\{1,2\}$. Thus, $(a_{1},a_{2})\in R$, and therefore the quantum function $Q^{\prime}\colon\{1,2\}\to A$ defined by $Q^{\prime}_{i,a}=Q_{x_{i},a}$ satisfies $($$\mathbf{QH_{1}}$)) and is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}(A;R)$. ∎
We conclude this section by showing that also the non-oracular version of the CSP parameterised by odd cycles is undecidable. The result is proved in same way as its oracular counterpart.
Corollary 48.
For every odd $m\geq 3$, $\operatorname{CSP^{**}}(\mathbb{C}_{m})$ is undecidable.
Proof.
Simply observe that the proof of Theorem 29 does not make use of the fact that quantum polymorphisms of $\mathbb{C}_{m}$ are oracular. Hence, the same argument establishes that $\operatorname{qPol^{no}}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$ for each odd $m\geq 3$. Then, the result follows from Theorem 38 applied to $\mathbb{A}=\mathbb{B}=\mathbb{C}_{m}$ and Corollary 34. ∎
8. Quantum polymorphisms of Boolean structures
In this section, we provide some results regarding the quantum polymorphisms of Boolean structures. If a quantum homomorphism $Q$ has range $\{0,1\}$, then $Q_{x,0}$ is determined by $Q_{x,1}$ since $Q_{x,0}=\operatorname{id}-Q_{x,1}$. Moreover, the set $\{0,1\}^{n}$ is naturally in bijection with the powerset of $\{1,\dots,n\}$. Given $S\subseteq[n]$, we let $s\in\{0,1\}^{n}$ be the tuple such that $s_{i}=1$ iff $i\in S$. In the following, if $\mathbb{A}$ is a Boolean structure, we write the projectors of a quantum homomorphism $\mathbb{A}^{n}\to_{q}\mathbb{A}$ by $Q_{S}$ where $S\subseteq[n]$.
8.1. The case of structures without a majority polymorphism
In this subsection we show that any Boolean relational structure which is not invariant under the majority polymorphism has a commutativity gadget. This specific statement has previously been proven in [culf_re-completeness_2025, Section 7], our contribution is to simplify the proof by using the characterization in Theorem 44 to assert the non-existence of commutativity gadgets through the study of quantum polymorphisms. This allows us to make use of standard CSP methods in the proof.
We focus our study on the following class of Boolean relational structures:
Definition 49.
Let $t\in\{0,1\}^{k}$ be an arbitrary Boolean tuple. Then the $t$-translate of 11-in-$k$ is the relational structure $\mathbb{O}_{t}=(\{0,1\};\{x\oplus_{2}t\mid x\in R_{1/k}\})$, where
  1. (1)
    $R_{1/k}=\{x\in\{0,1\}^{k}\mid\sum_{i=1}^{k}x_{i}=1\}$ is the relation of 11-in-$k$ SAT and,
  2. (2)
    $\oplus_{2}$ is componentwise addition modulo 2.
The relation of $\mathbb{O}_{t}$ is named $R_{t}$.
One can easily see that the ternary majority function on $\{0,1\}$ does not preserve $R_{t}$; the majority of any 3 distinct tuples in the relation yields $t\not\in R_{t}$, witnessing the non-preservation.
We show that if we translate 11-in-$k$ by a tuple $t\in R_{1/k}$, then $\operatorname{\oplus Pol}(\mathbb{O}_{t})=\operatorname{qPol^{no}}(\mathbb{O}_{t})$. First consider the case where $k=3$ and $t=(1,0,0)$:
Proposition 50.
For every non-oracular polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$, the following properties hold:
  1. (1)
    If $S\cap T=\emptyset$, then $Q_{S}Q_{T}=0$.
  2. (2)
    If $S\cap T=\emptyset$, then $Q_{S\cup T}=Q_{S\cup T}(Q_{S}+Q_{T})$.
  3. (3)
    $Q_{S}=Q_{S\cup T}Q_{S}$ for all $S$
  4. (4)
    $Q_{S}=\sum_{i\in S}Q_{\{i\}}$.
Proof.
Consider $S,T\subseteq[n]$ with $S\cap T=\emptyset$ and let $R:=S\cup T$. Then $(r,s,t)$ is in $R_{(1,0,0)}$, which gives the following identities by property $($$\mathbf{QH_{1}}$)) of $Q$:
  • $Q_{S\cup T}Q_{S}Q_{T}=0$, since $(1,1,1)$ is not in $R_{(1,0,0)}$,
  • $(\operatorname{id}-Q_{S\cup T})Q_{S}Q_{T}=0$, since $(0,1,1)$ is not in $R_{(1,0,0)}$,
  • $Q_{S\cup T}(\operatorname{id}-Q_{S})(\operatorname{id}-Q_{T})=0$, since $(1,0,0)$ is not in $R_{(1,0,0)}$.
  • $(\operatorname{id}-Q_{S\cup T})Q_{S}(\operatorname{id}-Q_{T})=0$, since $(0,1,0)$ is not in $R_{(1,0,0)}$.
We can add the first two identities to get $Q_{S}Q_{t}=0$ so, (1) holds. Moreover, adding the first and third identities yields (2).
To derive $Q_{S}=Q_{S\cup T}Q_{S}$ we add the first and fourth identities to get $Q_{S}=Q_{S\cup T}Q_{S}+Q_{S}Q_{T}$. Application of (1) then yields (3).
Finally, from the equalities
$Q_{S\cup T}=Q_{S\cup T}(Q_{S}+Q_{T})=Q_{S\cup T}Q_{S}+Q_{S\cup T}Q_{T}=Q_{S}+Q_{T}$
we can derive (4) by inductively building any $Q_{S}$ from sums of the individual elements of $S$. ∎
In particular, it follows that in every non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$ and every $S,T\subseteq[n]$, we have $Q_{S},Q_{T}]=0$, which implies that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})$ is the quantum closure of $\operatorname{Pol}(\mathbb{O}_{(1,0,0)})$.
Lemma 51.
For any natural number $k\geq 3$, the translation $\mathbb{O}_{(1,0,\dots,0)}$ of 11-in-$k$ by $(1,0,\dots,0)$ has the property that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,\dots,0)})=\operatorname{\oplus Pol}(\mathbb{O}_{(1,0,\dots,0)})$.
Proof.
Since $(0,0,\dots,0)=(1,0,\dots,0)\oplus_{2}(1,0,\dots,0)\in R_{(1,0,\dots,0)}$ and $(1,1,\dots,1)\notin R_{(1,0,\dots,0)}$, we can pp-define 0 in $\mathbb{O}_{(1,0,\dots,0)}$ by
$a=0\Leftrightarrow R_{(1,0,\dots,0)}(a,\dots,a)$
This relation is preserved by every non-oracular quantum polymorphism of $R_{(1,0,\dots,0)}$ since,
$Q_{\mathbf{a},1}=Q_{\mathbf{a},1}\dots Q_{\mathbf{a},1}=0$
for every $\mathbf{a}\in\{0,1\}^{n}$ and $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to\mathbb{O}_{(1,0,\dots,0)}$. We then proceed to define $R_{(1,0,0)}$ in $(\{0,1\},R_{(1,0,\dots,0)},0)$ by
$R_{(1,0,0)}(x,y,z)\Leftrightarrow R_{(1,0,\dots,0)}(x,y,z,0,\dots,0)$
We prove that $R_{(1,0,0)}$ is preserved by every non-oracular $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to\mathbb{O}_{(1,0,\dots,0)}$. Assume we have tuples $\mathbf{a},\mathbf{b},\mathbf{c}\in\{0,1\}^{n}$ with $(\mathbf{a},\mathbf{b},\mathbf{c})$ in the $n$th direct power of $R_{(1,0,0)}$ and variables $x,y,z\in\{0,1\}$ such that $(x,y,z)\notin R_{(1,0,0)}$. Then $(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{0},\dots,\mathbf{0})$ is in the $n$th power $R_{(1,0,\dots,0)}$ and $(x,y,z,0,\dots,0)$ is not an element of $R_{(1,0,\dots,0)}$
Now assume we have some non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to_{q}\mathbb{O}_{(1,0,\dots,0)}$ and consider:
$Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}$ $=Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}\sum_{w\in\{0,1\}}Q_{\mathbf{0},w}$
$=\sum_{w\in\{0,1\}}Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},w}$
$=\sum_{w\in\{0,1\}}(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},w}\dots Q_{\mathbf{0},w})(Q_{\mathbf{0},w}\dots Q_{\mathbf{0},w})$
$=(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})(Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})$
$=0(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})0=0$
where the penultimate equality follows from a combination of $($$\mathbf{QH_{1}}$)) and the fact that $(1,\dots,1)\notin R_{(1,0,\dots,0)}$.
This proves that every non-oracular quantum polymorphism of $\mathbb{O}_{(1,0,\dots,0)}$ preserves $R_{(1,0,0)}$ and thus
$\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,\dots,0)})\subseteq\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})=\operatorname{\oplus Pol}(\mathbb{O}_{(1,0,0)}).$
It follows that every non-oracular quantum polymorphism of $\mathbb{O}_{(1,0,\dots,0)}$ is non-contextual, which completes the proof. ∎
For a generalisation of Lemma 51 to an arbitrary $t\in R_{1/k}$, we can pp-define $R_{t}$ in $\mathbb{O}_{(1,0,0\dots,0)}$ by
$R_{t}(x_{1},\dots,x_{k})\Leftrightarrow R_{(1,0,0\dots,0)}(x_{l},x_{1}\dots,x_{l-1},x_{l+1},\dots,x_{k})$
where $1\leq l\leq k$ is the singular non-zero component of $t$.
Corollary 52.
If $k\geq 3$ and $t\in R_{1/k}$, the non-oracular polymorphisms of $\mathbb{O}_{t}$ are non-contextual.
Now for the when $t\notin R_{1/k}$. We reduce it to the case from Corollary 52.
Lemma 53.
Assume $k\geq 3$, and neither $t$ nor $\overline{t}$ ($(1,\dots,1)\oplus_{2}t$) are in $R_{1/k}$. Then $\operatorname{qPol^{no}}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$.
Proof.
Without loss of generality, assume that $t=t^{l}$ is the tuple with $l$ 1-entries at the start, one can do this from an arbitrary $t$ by just permuting the entries.
If $l\notin\{1,k-1\}$ (i.e., neither $t$ nor $\overline{t}$ are in $R_{1/k}$) then $(0,\dots,0)=t\oplus_{2}t$ and $(1,\dots,1)=t\oplus_{2}\overline{t}$ are not in $R_{t}$. We first pp-define $\neq$ in $\mathbb{O}_{t}$ via
$x\neq y\Leftrightarrow R_{t}(x,\dots,x,y,\dots,y)$
where we fill the first $l+1$ entries with $x$’s. This relation is preserved by every non-oracular $Q\colon\mathbb{O}_{t}^{n}\to\mathbb{O}_{t}^{n}$ since for every $\mathbf{a}\neq\mathbf{b}\in\{0,1\}^{n}$ and $x\in\{0,1\}$ we have
$Q_{\mathbf{a},x}Q_{\mathbf{b},x}=Q_{\mathbf{a},x}\dots Q_{\mathbf{a},x}Q_{\mathbf{b},x}\dots Q_{\mathbf{b},x}=0.$
Let $t^{\prime}=(0,1\dots,1,0,\dots,0)$ ($t^{l}$ with the first entry switched). We define $R_{t^{\prime}}$ by the formula
$R_{t^{\prime}}(s_{1},\dots,s_{k})\Leftrightarrow\exists q\left(R_{t}(q,s_{2},\dots,s_{k})\wedge q\neq s_{1}\right).$
Assume we have tuples $\mathbf{s}_{1},\dots,\mathbf{s}_{k}\in\{0,1\}^{n}$ with $(\mathbf{s}_{1},\dots,\mathbf{s}_{k})$ in the $n$th direct power of $R_{t^{\prime}}$ and variables $x_{1},\dots,x_{k}\in\{0,1\}$ such that $(x_{1},\dots,x_{k})\notin R_{t^{\prime}}$. Then we have $\mathbf{q}\in\{0,1\}^{n}$ such that $(\mathbf{q},\mathbf{s}_{2},\dots,\mathbf{s}_{k})$ are elements of the $n$th power of $R_{t}$ and $\mathbf{q}\neq\mathbf{s}_{1}$ . Also, there is no $y\in\{0,1\}$ such that both $(y,x_{2},\dots,x_{k})\in R_{t}$ and $y\neq x_{1}$, i.e., we know that $(\neg x_{1},x_{2},\dots,x_{k})\notin R_{t}$.
Now let $Q\colon\mathbb{O}_{t}^{n}\to_{q}\mathbb{O}_{t}$ be a non-oracular polymorphism of $\mathbb{O}_{t}$. Then we have
$Q_{\mathbf{s}_{1},x_{1}}\dots Q_{\mathbf{s}_{k},x_{k}}$ $=Q_{\mathbf{s}_{1},x_{1}}\left(\sum_{y\in\{0,1\}}Q_{\mathbf{q},y}\right)Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}}$
$=\sum_{y\in\{0,1\}}Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},y}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}}$
$=\sum_{y\in\{0,1\}}(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},y})(Q_{\mathbf{q},y}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})$
$=(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},x_{1}})(Q_{\mathbf{q},x_{1}}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})+(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},\neg x_{1}})(Q_{\mathbf{q},\neg x_{1}}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})$
$=(0)(Q_{\mathbf{q},x_{1}}Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}})+(Q_{\mathbf{s}_{1},x_{1}}Q_{\mathbf{q},\neg x_{1}})(0)=0$
where the penultimate equality follows from $($$\mathbf{QH_{1}}$)). This proves that every non-oracular quantum polymorphism of $\mathbb{O}_{t^{l}}$ preserves $R_{t^{\prime}}$ and thus
$\operatorname{qPol^{no}}(\mathbb{O}_{t^{l}})\subseteq\operatorname{qPol^{no}}(\mathbb{O}_{t^{\prime}})=\operatorname{qPol^{no}}(\mathbb{O}_{t^{l-1}})$
by induction, we then get that
$\operatorname{qPol^{no}}(\mathbb{O}_{t})\subseteq\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,\dots,0)})$
completing the proof after an application of Lemma 51.
Note that this theorem also covers the case when $t=(0,\dots,0)$ i.e. $R_{t}=R_{1/k}$ (or $l=0$). In this case $R_{t^{\prime}}$ as pp-defined above is itself $R_{(1,0,\dots,0)}$ and we are done immediately. ∎
Next, we argue that the only real cases we need to consider in the full theorem is the $\mathbb{O}_{t}$’s.
Lemma 54.
Every relational structure $\mathbb{A}$ with a relation $R$ that has any full binary projection has the property that $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Consider a relational structure with a relation $R$ that has a full binary projection on the components $\{i,j\}$. If this is the case, then in any power, any two elements appear together in some tuple of the direct power of RR.∎
Theorem 55.
For any relation $R\subseteq\{0,1\}^{k}$ the following statements are equivalent:
  1. (1)
    $R$ is not invariant under majority, majority preserves every proper projection of $R$, and every binary projection of $R$ is not full.
  2. (2)
    $R=R_{t}$ for some tuple $t$.
Proof.
2. implies 1. Note that translating a relation by a tuple does not change the size of the binary projections and does not change preservation by majority, since majority is a self-dual operation. It follows that it suffices to prove the implication in the case of $R_{1/k}$, for which all the properties in 1. are clear.
For 1. implies 2. suppose $R\subseteq\{0,1\}^{k}$ is not invariant under majority but majority does preserve every proper projection and no binary projection is full. Suppose $t\in\{0,1\}^{k}$ is a witness to the fact that $R$ is not invariant under majority, we first show that $\{x\oplus_{2}t\mid x\in R_{1/k}\}\subseteq R$.
For $i\in[k]$, let $x\in R_{1/k}$ be the tuple with a 1 in the $i$th position. We know that $\pi_{[k]\setminus\{i\}}(R)$ is invariant under majority so there is an $r_{i}\in R$ with $\pi_{[k]\setminus\{i\}}(r_{i})=\pi_{[k]\setminus\{i\}}(t)$. Now, since $t\notin R$; we can conclude that $\pi_{\{i\}}(r_{i})\neq\pi_{\{i\}}(t)$ and therefore $x\oplus_{2}t=r_{i}$, so $x\oplus_{2}t\in R$.
Finally, we use the non-fullness of binary projections to argue that any $r\in R$ must match $\overline{t}=(1,\dots,1)\oplus_{2}t$ in exactly one argument, placing it in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$.
Assume that $r$ matches $\overline{t}$ in two arguments $1\leq i<j\leq k$. We know that $\pi_{i,j}(R)$ contains $\pi_{i,j}(t)$, $(\pi_{i}(t),\pi_{j}(\overline{t}))$ and $(\pi_{i}(\overline{t}),\pi_{j}(t))$; then $\pi_{i,j}(R)$ is full, a contradiction. Additionally, $r$ cannot match $\overline{t}$ on zero arguments since then it would equal $t$, which is not in $R$. Thus every $r$ is in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$, completing the proof. ∎
We can now easily prove the general result.
Theorem 56.
If $\mathbb{A}$ is a Boolean relational structure such that $\operatorname{Pol}(\mathbb{A})$ does not contain majority, then $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Consider a relational structure $\mathbb{A}$; without loss of generality, every projection of a relation of $\mathbb{A}$ is also a relation of $\mathbb{A}$, since this does not change the hypothesis or the conclusion of the statement. Let $R$ be a relation of $\mathbb{A}$ that is not preserved by majority and of minimal arity. By Lemma 54, if a binary projection of $R$ is full then we are done, so assume that this is not the case. Then Theorem 55 implies that $R=R_{t}$ for some $k$-tuple $t$.
Finally, we know that $k\geq 3$ since every binary relation of arity 2 is preserved under majority and so, most of the result follows from Corollary 52 and Lemma 53. All that remains is for us to argue that Lemma 52 also covers the case where $\overline{t}\in R_{1/k}$.
This is since, $\mathbb{O}_{t}$ has ‘essentially’ the same polymorphisms as $\mathbb{O}_{\overline{t}}$ for any $t$. The formal fact we use is that $\{Q_{S}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{t}$ if and only if $\{\operatorname{id}-Q_{\overline{S}}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{\overline{t}}$. In particular $\mathbb{O}_{t}$ has contextual polymorphisms if and only if $\mathbb{O}_{\overline{t}}$ does. Thus we can assume that $\overline{t}\notin R_{1/k}$, completing the proof. ∎
8.2. The minimal contextual clone
In this section, we provide a clone with contextual polymorphisms that is contained within the quantum polymorphism clone of any Boolean relational structure with contextual polymorphisms.
In the previous section, we have seen that every boolean relational structure with contextual polymorphisms has non-full binary projections and contains majority in its (classical) polymorphism clone.
Let $\mathbb{B}$ be the structure $\mathbb{B}:=(\{0,1\};S_{00},S_{11},S_{10})$ where $S_{ab}=\{0,1\}^{2}\setminus\{(a,b)\}$. In the classical setting, $\operatorname{Pol}(\mathbb{B})$ is the smallest clone that contains the majority operation. We first show that this property extends to oracular quantum polymorphisms.
Lemma 57.
Let $\mathbb{A}$ be a Boolean structure that admits majority as a polymorphism, and such that no binary projection of a relation of $\mathbb{A}$ is full. Then $\operatorname{qPol}(\mathbb{B})\subseteq\operatorname{qPol}(\mathbb{A})$.
Proof.
Consider a relational structure $\mathbb{A}:=(\{0,1\};R)$ that is preserved by majority, then it is well known that $R$ (of arity $r$) can be pp-defined in $\mathbb{A}_{2}:=(\{0,1\};\{\pi_{i,j}(R)\mid 1\leq i,j\leq r\})$ via the conjunction of its binary projections $\bigwedge_{1\leq i,j\leq r}\pi_{\{i,j\}}(R)(x_{i},x_{j})$. We show that $R$ is preserved by every quantum polymorphism $Q\colon\mathbb{A}_{2}^{n}\to\mathbb{A}_{2}$. Assume that $(\mathbf{a}_{1},\dots,\mathbf{a}_{r})$ is in the $n$th direct power of $R$ while in $(x_{1},\dots,x_{r})$, we have indices $i,j$ such that $(x_{i},x_{j})\notin\pi_{\{i,j\}}(R)$. Then
$Q_{\mathbf{a}_{1},x_{1}}\dots Q_{\mathbf{a}_{r},x_{r}}=(Q_{\mathbf{a}_{i},x_{i}}Q_{\mathbf{a}_{j},x_{j}})Q_{\mathbf{a}_{1},x_{1}}\dots Q_{\mathbf{a}_{r},x_{r}}=(0)Q_{\mathbf{a}_{1},x_{1}}\dots Q_{\mathbf{a}_{r},x_{r}}=0$
where the second equality is by duplicating the $Q_{\mathbf{a}_{i},x_{i}}$ and $Q_{\mathbf{a}_{j},x_{j}}$ projectors and then using property $($$\mathbf{QH_{2}}$)) to move them to the front. Moreover, if $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}^{n}))$ then $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}_{2}^{n}))$ so, for any $x,x^{\prime}\in\{0,1\}$ $Q_{\mathbf{a},x},Q_{\mathbf{a}^{\prime},x^{\prime}}]=0$ by property $($$\mathbf{QH_{2}}$)). $Q$ is then an oracular quantum polymorphism of $\mathbb{A}$ and so $\operatorname{qPol}(\mathbb{A}_{2})\subseteq\operatorname{qPol}(\mathbb{A})$.
By assumption, no binary projection of a relation of $\mathbb{A}$ is full, thus, each relation $R$ in $\mathbb{A}_{2}$ can be defined as the (non-empty!) conjunction $\bigwedge_{(c,d)\not\in R}S_{cd}(x,y)$, where we use the shorthand $S_{01}(x,y)$ for the formula $S_{10}(y,x)$. We show that each relation $R$ of $\mathbb{A}_{2}$ is preserved by the quantum polymorphisms of $\mathbb{B}$.
Let $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ be a quantum polymorphism, $(\mathbf{a},\mathbf{b})$ be in the $n$th direct power of $R$ and $(c,d)\notin R$. We then have $Q_{\mathbf{a},c}Q_{\mathbf{b},d}=0$ by $($$\mathbf{QH_{1}}$)) since $(c,d)\not\in S_{cd}$ and $(\mathbf{a},\mathbf{b})$ is in the relation $S_{cd}$ in the $n$th power of $\mathbb{B}$. Thus, $($$\mathbf{QH_{1}}$)) holds.
Finally, if $(\mathbf{a},\mathbf{b})$ is a pair in a relation $R$ in the $n$th power of $\mathbb{A}_{2}$, then it is necessarily a pair in one of the relations $S_{cd}$ in the $n$th power of $\mathbb{B}$. Thus, $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute, so that $($$\mathbf{QH_{2}}$)) holds. ∎
Now we show that this clone actually contains some contextual polymorphisms.
Theorem 58.
For $n\leq 3$, every quantum polymorphism $Q\colon\mathbb{B}^{n}\to\mathbb{B}$ is non-contextual. For all $n\geq 4$, $\mathbb{B}$ has contextual quantum polymorphisms of arity $n$.
Proof.
We start by proving that every quantum polymorphism $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ is non-contextual for $n\leq 3$. For this, it suffices by $($$\mathbf{QH_{2}}$)) to prove that for every pair $S,T$ of subsets of $\{1,\dots,n\}$, either $(S,T)$ or $(T,S)$ belongs to one of the relations of $\mathbb{B}^{n}$. If $(S,T)$ does not belong to the relation $S_{00}$ in $\mathbb{B}^{n}$, then there exists $i\in\{1,\dots,n\}$ such that $i\in\overline{S}\cap\overline{T}$. Similarly, for it to not belong to the relation $S_{11}$ in $\mathbb{B}^{n}$, there must exist $j\in\{1,\dots,n\}$ such that $j\in S\cap T$. If $n\leq 3$, there remains at one most other element $k$ in $\{1,\dots,n\}\setminus\{i,j\}$. If $k$ is not in $S\cap T$ nor in $\overline{S}\cap\overline{T}$, then it is the case that $(S,T)$ or $(T,S)$ belongs to the relation $S_{10}$ of $\mathbb{B}^{n}$, so we are done.
For the $n\geq 4$ case we define a concrete quantum polymorphism for $n=4$. We provide Figure 3 as a visual aid.
0000100001000010000111001010100101100101001111101101101101111111
Figure 3. Simplified illustration of $\mathbb{B}^{4}$. Black lines indicate the poset $S_{10}$. The blue (resp. red) dotted lines show parts of $S_{00}$ (resp. $S_{11}$); this relation additionally relates $(0,0,0,0)$ (resp. $(1,1,1,1)$) to everything and relates two nodes when they are a bit-flips of each other (see coloring).
We define $Q\colon\mathbb{B}^{4}\to_{q}\mathbb{B}$ by listing the projectors. Let $T\subseteq[4]$. If $|T|\geq 3$, define $Q_{T}=\operatorname{id}$. Pick now arbitrary matrices $A,B,C$. Define $Q_{\{1,2\}}:=A$, $Q_{\{1,3\}}=B$, and $Q_{\{1,4\}}=C$. In all the other cases, define $Q_{T}=\operatorname{id}-Q_{[4]\setminus T}$. Note that if at least two of the matrices $A,B,C$ do not commute, then $Q$ is contextual.
It remains to prove that $Q$ preserves the relations of $\mathbb{B}^{4}$. Assume we have $S,T\subseteq[4]$ with $(s,t)$ in the fourth power of $S_{11}$ and $x,y\in\{0,1\}$ with $(x,y)\notin S_{11}$. If $S\neq\overline{T}$, we are in one of 2 cases; either $S$ or $T$ is the empty set or $|S|=|T|=1$. In both cases at least $Q_{S}$ or $Q_{T}$ equals 0 and we are done immediately. If $S=\overline{T}$, then $Q_{S}Q_{T}=Q_{S}Q_{\overline{S}}=Q_{S}(\operatorname{id}-Q_{S})=0$ and we are done. The case of $S_{00}$ is symmetric.
Now assume we have $S,T\subseteq[4]$ such that the corresponding pair of tuples $(s,t)$ in the fourth power of $S_{10}$. If both $|S|$ and $|T|$ do not equal $2$ we know $Q_{S},Q_{T},(\operatorname{id}-Q_{S}),(\operatorname{id}-Q_{T})\in\{0,\operatorname{id}\}$ and we immediately get that $Q_{S},{\operatorname{id}-Q_{T}}]=0$. The final case is when one of $|S|$ or $|T|$ equal $2$, here we get that either
$Q_{S}(\operatorname{id}-Q_{T})=0(\operatorname{id}-Q_{T})=0$
or,
$Q_{S}(\operatorname{id}-Q_{T})=Q_{S}(\operatorname{id}-\operatorname{id})=0$
showing preservation of the relations. ∎
9. Acknowledgments
The first and last author are thankful for the attentive listening and comments by Peter Zeman, to whom they explained the definition of quantum polymorphisms and the main results in this paper concerning the characterisation of the existence of commutativity gadgets on the occasion of the CSP World Congress in September 2025. The second and last author acknowledge funding from the project “Hamburg Quantum Computing”, co-financed by ERDF of the European Union and by the Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB).

### v2

@@
-Quantum Polymorphisms and Commutativity Gadgets
-Abstract.
-We introduce the concept of quantum polymorphisms to the complexity theory of non-local games.
-We use this notion to give a full characterisation of the existence of commutativity gadgets for relational structures, introduced by Ji as a method for achieving quantum soundness of classical CSP reductions.
-Prior to our work, a classification was only known in the Boolean case [Culf–Mastel, STOC’25].
-As an application of our framework, we prove that the entangled CSP parameterised by odd cycles is undecidable.
-Furthermore, we establish a quantum version of Galois connection for entangled CSPs in the case of non-oracular quantum homomorphisms.
-1. Introduction
-In 1971, Cook’s introduction of $\mathsf{NP}$-completeness [cook1971complexity] provided a way to mathematically assess the inherent hardness of computational problems.
-Boolean satisfiability, graph colouring, linear equations over finite fields, and many other computatational problems of a seemingly distinct nature systematically began to fall into either the polynomial-time solvable or the $\mathsf{NP}$-complete buckets [Karp72], while intermediate complexity turned out to be a rarer phenomenon [ladner1975structure].
-This fact soon sparked the hypothesis that the broad, natural class of computational problems involving satisfaction of constraints over finite alphabets—collectively known as (finite-domain) CSPs and encompassing the aforementioned problems and many others—could never exhibit the unnatural intermediate behaviour.
-$\operatorname{P}$ vs.
-$\mathsf{NP}$-complete’ dichotomies for fragments of CSPs started to emerge, primary examples being Schaefer’s [Schaefer78:stoc] and Hell–Nešetřil’s [HellN90] dichotomies for Boolean and graph CSPs, respectively.
-However, a unified framework for approaching the CSP Dichotomy Conjecture [Feder98:monotone] only emerged at the turn of the century, with the foundational line of work initiated by Jeavons, Cohen, and Gyssens that became known as the (universal-)algebraic approach to CSPs [Jeavons97:closure, Jeavons98:algebraic].
-By lifting CSP theory from the realm of Turing machines and reductions to that of algebra, the study of polymorphism clones—universal-algebraic objects capturing high-order invariance properties of CSP solutions—provided a clearer viewpoint on the CSP complexity landscape, and it eventually culminated with the positive resolution of the conjecture by Bulatov [Bulatov17:focs] and, independently, Zhuk [Zhuk17_FOCS, Zhuk20:jacm] in 2017.
-Seven years before the publication of Cook’s result, the work of Bell [bell1964einstein] originating from the Einstein–Podolsky–Rosen paradox [einstein1935can] demonstrated that quantum systems can display correlations unattainable within classical physics, thus ruling out local hidden-variable theories as realistic explanations for quantum mechanics.
-In its mathematical essence, Bell’s result can be formulated as a property of CSPs: There exist specific CSP languages for which the classical value and the entangled value can differ, where the latter is defined as the optimal success probability in a multiplayer interactive proof of the satisfiability of the CSP instance where the players have access to quantum resources.
-A primary question arising from Bell’s theorem concerns the computational complexity of determining the entangled value of a CSP.
-In particular, it is known that determining whether an instance of SAT is perfectly quantum satisfiable is undecidable (in fact, even the gap version of the problem is known to be undecidable, as a consequence of the celebrated $\operatorname{MIP^{*}=RE}$ theorem [ji2021mip]).
-On the other hand, examples of tractable CSP languages for which quantum and classical satisfiability coincide, which implies polynomial-time solvability of the corresponding entangled CSPs, are known [atserias_generalized_2019, ciardo_quantum_minion, bulatov2025satisfiability].
-Nevertheless, unlike in the classical case, the complexity landscape of entangled CSPs remains largely unexplored.
-A natural strategy to shed some light on it is to transfer the algebraic machinery that allowed proving the (classical) CSP Dichotomy to the quantum setting.
-In particular, for classical CSPs, the concept of polymorphisms turned out to be the right tool for exploring CSP complexity at the algebraic level.
-The root of this phenomenon is a Galois connection between relational clones (expressing the CSP language) and operation clones (expressing the family of polymorphisms).
-Via this connection, interreducibility between CSPs associated with different languages can be explained via inclusion (or weaker forms of the latter, in particular, so called height-1 clone homomorphisms) between their polymorphism clones.
-To what extent can this machinery by transferred to the setting of entangled CSPs, and thus provide some understanding of their complexity and interreducibility?
-This question was first addressed at this level of generality by Ji in [ji2013binary].
-In particular, that work identified a substantial obstacle against the prospect of quantising all types of algebraic CSP reductions known as gadgets.
-Ultimately, the obstacle is caused by the non-classical phenomenon of contextuality: Replacing constraints by gadgets increases the size of the contexts—which, in general, destroys simultaneous measurability.
-In the same paper, Ji provided a method to solve this issue for specific CSP languages, via the new notion of commutativity gadgets.
-On a high level, these are specific types of gadgets having no classical counterpart, that can be plugged into classical gadgets to enforce simultaneous measurability of the final instance resulting from the reduction.
-For example, by showing that a specific triangular prism construction yields a commutativity gadgets for the 3-clique, Ji showed that the classical polynomial-time reduction from 3-SAT to 3-colouring can be lifted to a reduction between the corresponding entangled CSPs, thus proving undecidability of quantum 3-colouring.
-Later works generalised this idea to more general types of CSPs.
-In particular, commutativity gadgets for certain Boolean CSPs were constructed in [atserias_generalized_2019].
-In a subsequent recent breakthrough [culf2024re], Culf and Mastel gave a complete characterisation of Boolean CSPs admitting commutativity gadgets.
-In the same work, the authors identified a sufficient property of CSP languages—which they called two-variable (non-)falsifiability—guaranteeing the existence of commutativity gadgets for the corresponding CSPs, and thus allows quantising classical reductions.
-Very recently, Culf, de Bruyn, Vernooij, and Zeman [Zeman] gave a different, necessary condition for the existence of commutativity gadgets, by showing that CSPs with non-classical quantum endomorphisms do not admit commutativity gadgets.
-Despite this recent progress, a full characterisation of CSPs exhibiting commutativity gadgets is not available.
-In particular, [Zeman] identified odd cycles as a challenging open case for which the existence or non-existence of commutativity gadgets cannot be established via available methods.
-Contributions
-We give a complete characterisation of CSPs exhibiting commutativity gadgets.
-To that end, we introduce the notion of quantum polymorphisms, which lift classical polymorphisms to the setting of entangled CSPs.
-We show that quantum polymorphisms completely capture the existence of commutativity gadgets, by proving that a commutativity gadget exists if, and only if, the minion of quantum polymorphisms coincides with the quantum closure of the clone of standard polymorphisms.
-Through this characterisation, we prove that all odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable.
-This answers the question raised in [Zeman].
-In addition, we use quantum polymorphisms to build commutativity gadgets for Boolean structures in a unified manner, thus giving a different proof of the result of [culf_re-completeness_2025].
-Our characterisation works for both oracular and non-oracular CSPs.
-In the non-oracular setting, we additionally establish a quantum version of Galois connection that parallels the corresponding classical connection.
+Quantum Polymorphisms and the Complexity of Quantum Constraint Satisfaction
+Abstract
+We introduce the concept of quantum polymorphisms to the complexity theory of quantum constraint satisfaction.
+Via this notion, we build an algebraic framework of reductions between quantum CSPs, and we establish a Galois connection between quantum polymorphism minions and quantum relational constructions.
+By leveraging a contextuality property of quantum polymorphisms, we fully characterise the existence of commutativity gadgets for relational structures, introduced by Ji as a method for achieving quantum soundness of classical CSP reductions.
+Prior to our work, only a partial classification was known for a subclass of Boolean languages and for non-Boolean languages meeting specific structural conditions [Culf–Mastel, FOCS’25].
+As an application of our framework, we prove that the quantum CSPs parameterised by odd cycles and the quantum CSP expressing quantum satisfiability of Siggers clauses are undecidable.
+1 Introduction
+The quantum constraint satisfaction problem is the decision problem of testing for the existence of a winning quantum strategy for a 2-player 1-round game where the two (cooperating, computationally unbounded) players aim to prove to a classical verifier that a system of constraints can be simultaneously satisfied.
+Here, a quantum strategy consists of a bipartite finite-dimensional quantum state shared by the players and a collection of measurements that the players perform on the state.
+Upon receiving from the verifier questions involving local constraints of the system, the players measure their part of the state and answer with partial assignments based on the measurement outcomes.
+The origins of such game—often known as a non-local game—come from physics: Bell’s[12] and Kochen–Specker’s [53] proofs of the incompatibility of quantum mechanics with local hidden-variable theories, stemming from the Einstein–Podolsky–Rosen paradox [35], depend on the fact that, for certain types of constraints, quantum strategies generate correlations in the player’s answers that are unattainable in classical physics.
+[…]
+However, a unified framework for approaching the CSP Dichotomy Conjecture [36] only emerged at the turn of the century, with the foundational line of work initiated by Jeavons, Cohen, and Gyssens that became known as the (universal-)algebraic approach to CSPs [44, 45].
+By lifting CSP theory from the realm of Turing machines and reductions to that of algebra, the study of polymorphism clones—universal-algebraic objects capturing high-order invariance properties of CSP solutions—provided a clearer viewpoint on the CSP complexity landscape, and it eventually culminated with the positive resolution of the conjecture by Bulatov [21] and, independently, Zhuk [68, 69] in 2017.
+In the case of quantum CSPs, the complexity landscape is much less understood.
+While there are various examples of constraint classes for which the quantum CSP is undecidable [67, 46, 41, 2, 33, 63, 34] or polynomial-time solvable [50, 47, 2, 30, 63, 20, 56], the complexity is unknown for most classes of languages, due to the lack of a general approach for studying them with unified tools.
+The main conceptual contribution of this work is the introduction of the notion of quantum polymorphisms as a tool to lift the study of algebraic invariants of classical CSP languages to the quantum setting.
+Via this notion, we initiate an algebraic complexity theory of quantum CSPs modelled on the algebraic approach initiated in [44, 45] and culminated 20 years later with the proof of the (classical) Dichotomy Theorem.
+Classical polymorphisms turned out to be the right tool for exploring CSP complexity at the algebraic level.
+The root of this phenomenon is a connection between relational constructions (expressing the CSP language) and operation systems (expressing the family of polymorphisms).
+Via this connection, reducibility between CSPs associated with different languages can be explained via inclusion (or algebraic forms of the latter, in particular, so called minion homomorphisms) between their polymorphism sets.
+In this work, we prove that quantum polymorphisms completely determine the complexity of quantum CSPs, thus enabling the use of algebra in the complexity analysis of such problems.
+Moreover, we link quantum polymorphisms to a quantised version of classical relational constructions that we name q-definitions and q-constructions, obtained by lifting the classical model-theoretic counterparts to the setting of quantum strategies.
+Broadly, computable reductions in the literature on non-local games and quantum CSP complexity typically follow a common conceptual pattern: taking a classical CSP reduction (in particular, a so-called subdivision of constraints or a more general pp-definition), and trying to achieve quantum completeness and soundness.
+It is well known that this approach does not work in general.
+A substantial obstacle to the prospect of using classical gadgets as reductions between quantum CSPs is ultimately caused by the following, non-classical phenomenon: replacing constraints by gadgets increases the size of the contexts—which, in general, destroys simultaneous measurability (see, e.g., [46, 63, 2]).
+On the other hand, for some families of constraint languages, the issue can be solved by pairing classical gadgets with an extra type of gadget known as a commutativity gadget.
+On a high level, these are specific types of gadgets that can be plugged into classical gadgets to enforce simultaneous measurability of the final instance resulting from the reduction.
+For example, by showing that a specific triangular prism construction yields a commutativity gadget for the 3-clique, Ji showed that the classical polynomial-time reduction from 3-SAT to 3-colouring can be lifted to a reduction between the corresponding quantum CSPs [47].
+These commutativity gadgets have no counterpart in the classical setting, as they enforce (by design) no additional constraints on the variables in their scope apart from commutativity.
+Therefore, they are useless from the classical point of view.
+Later works generalised this idea to other types of quantum CSPs.
+In particular, commutativity gadgets for certain Boolean CSPs were constructed in [2].
+In a subsequent recent breakthrough [33], Culf and Mastel designed commutativity gadgets for all Boolean constraint languages whose classical CSP is NP-complete (or, more generally, is not preserved under majority [66]).
+In the same work, the authors identified a sufficient property of CSP languages—which they called two-variable (non-)falsifiability—that guarantees the existence of commutativity gadgets for the corresponding CSPs and thus allows quantising classical reductions.
+Very recently, Culf, de Bruyn, Vernooij, and Zeman [34] gave a different, necessary condition for the existence of commutativity gadgets, by showing that CSPs with non-classical quantum endomorphisms do not admit commutativity gadgets.
+Despite this recent progress, a full characterisation of CSPs admitting commutativity gadgets was hitherto not available.
+We prove that Ji’s reductions obtained by pairing classical gadgets with commutativity gadgets are captured by (and are a strict subset of) the reductions coming from quantum polymorphisms.
+Hence, as a second main contribution of this work, we give a complete characterisation of the class of CSP languages admitting commutativity gadgets, in terms of a contextuality property of their quantum polymorphisms.
+In particular, we show that a commutativity gadget exists if, and only if, the minion of quantum polymorphisms coincides with the quantum closure of the clone of standard polymorphisms.
+Through this characterisation, we prove that all odd cycles, as well as the Siggers digraph [49], admit commutativity gadgets and, thus, the corresponding quantum CSP is undecidable.
+[^2] This answers an open question raised in [34].
+In addition, we show that whenever a language admits a commutativity gadget, this can be chosen to be a high-enough categorical power of the language itself.
+We use this to build commutativity gadgets for Boolean structures in a unified manner, and we extend the result in [33] to a full classification of Boolean structures admitting commutativity gadgets.
+As a consequence, we obtain a complexity dichotomy for Boolean quantum CSPs, by showing that they are either in P or undecidable.
+2 Overview of results and techniques
+The gist of the algebraic approach to CSPs is a correspondence between $(i)$ algebraic properties of invariants of solution sets of CSPs—namely, polymorphisms—and $(ii)$ relational constructions that enable reductions between different CSPs and, thus, determine tractability or hardness.
+In the next three subsections of this overview, we lift both objects to the quantum setting, and we use them to characterise the existence of commutativity gadgets.
+We then apply the polymorphic framework to prove undecidability for classes of quantum CSPs by using a result on the emergence of certain non-orthogonality patterns in contextual quantum polymorphisms.
+This section contains an informal overview of the main ideas we use to establish these results.
+Full details are in the body of the paper (from Section 3 on).
+2.1 Quantum algebraic reductions
+Consider a 2-prover 1-round game as illustrated in the Introduction.
+Suppose that the admitted constraints in the protocol are encoded in a relational structure $\mathbb{A}$; i.e., the goal of the provers is to convince the verifier that a given system $\mathbb{X}$ (called the instance) of $\mathbb{A}$-clauses applied to variables of $\mathbb{X}$ is simultaneously satisfiable.
+The problem of deciding whether a winning quantum strategy exists for such game is the quantum CSP parameterised by the relational structure $\mathbb{A}$, and it is denoted by $\operatorname{CSP_{q}}(\mathbb{A})$.
+We define a quantum polymorphism of $\mathbb{A}$ as a winning quantum strategy for such protocol, where the instance $\mathbb{X}$ is constrained to be some $n$-ary categorical power $\mathbb{A}^{n}$ of $\mathbb{A}$.
+[^3] If $\mathbb{A}$ is a digraph (i.e., it has a single, binary relation), then $\mathbb{A}^{n}$ is the standard direct power of digraphs.
+In the language of quantum homomorphisms between relational structures [1], this can be rephrased as follows.
+Definition 1.
+A quantum polymorphism of arity $n\in\mathbb{N}$ for a structure $\mathbb{A}$ is a quantum homomorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$.
+We let $\operatorname{qPol}(\mathbb{A})$ be the set of all quantum polymorphisms of $\mathbb{A}$.
+By forcing the admitted strategies in the game to be classical and deterministic, we retrieve the classical polymorphisms of $\mathbb{A}$; i.e., $\operatorname{Pol}(\mathbb{A})\subseteq\operatorname{qPol}(\mathbb{A})$.
+The role of classical polymorphisms in CSP complexity is well understood: $\operatorname{Pol}(\mathbb{A})$ completely determines the complexity of $\operatorname{CSP}(\mathbb{A})$, in the sense that, given two langauges $\mathbb{A}$ and $\mathbb{B}$ such that $\operatorname{Pol}(\mathbb{A})=\operatorname{Pol}(\mathbb{B})$, the two problems $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{CSP}(\mathbb{B})$ are polynomial-time equivalent.
+Moreover, not all information contained in $\operatorname{Pol}(\mathbb{A})$ is relevant complexity-wise.
+In fact, the complexity of $\operatorname{CSP}(\mathbb{A})$ is fully determined by the identities of a restricted form satisfied by $\operatorname{Pol}(\mathbb{A})$, namely the so-called minor identities.
+This leads to the notion of a function minion (or minor-closed function class), which provides an algebraic tool for comparing the complexity of CSPs across different languages.
+Function minions capture precisely the minor identities satisfied by polymorphisms, thereby isolating the complexity-relevant structure of $\operatorname{Pol}(\mathbb{A})$.
+Theorem 2 ([10]).
+Let $\mathbb{A}$ and $\mathbb{B}$ be relational structures.
+If there exists a minion homomorphism from $\operatorname{Pol}(\mathbb{A})$ to $\operatorname{Pol}(\mathbb{B})$, then $\operatorname{CSP}(\mathbb{B})$ is logspace-reducible to $\operatorname{CSP}(\mathbb{A})$.
+Quantum polymorphisms are not operations over the domain $A$ and, thus, cannot be directly equipped with the structure of a function minion.
+Nevertheless, we show that they behave algebraically like classical polymorphisms, by endowing them with the structure of abstract minions.
+These objects, originally introduced in the contexts of relaxations of promise CSPs [5, 27, 28, 29, 26, 17, 55], are designed to retain the algebraic structure of function minions while dispensing from the requirement that their elements be concrete functions.
+To establish that $\operatorname{qPol}(\mathbb{A})$ is an abstract minion, we exploit the fact that quantum homomorphisms admit a specific composition operation, described in [1] using the framework of Kleisli categories (see Section 4.1).
+This structure enables an algebraic comparison of quantum polymorphisms across different languages.
+In particular, we lift Theorem 2 to the following result.
+Theorem 3.
+Let $\mathbb{A}$ and $\mathbb{B}$ be relational structures.
+If there exists a minion homomorphism from $\operatorname{qPol}(\mathbb{A})$ to $\operatorname{qPol}(\mathbb{B})$, then $\operatorname{CSP_{q}}(\mathbb{B})$ is logspace-reducible to $\operatorname{CSP_{q}}(\mathbb{A})$.
+The proof of Theorem 3 (whose full details are given in Section 4.2) is based on a reduction coming from the Long-Code testing for classical constraint systems [13] (see also [5, §3.3]).
+We establish quantum completeness and soundness of this reduction using the minor-preservation property induced by the minion homomorphism.
+As a consequence of the above result, quantum polymorphisms—and, more precisely, the minor identities they satisfy—completely characterise the complexity of quantum CSPs.
+Moreover, Theorem 3 enables the transfer of known undecidability results for specific quantum CSPs (for instance, quantum 3SAT [46, 60] or quantum XOR [67]) to a much broader class of problems via the analysis of the corresponding quantum polymorphisms.
+This draws a nice parallel with the classical setting, where the algebraic study of polymorphisms enabled the use of tools from universal algebra, ultimately playing a central role in the proof of the CSP dichotomy theorem.
+The next step is to look at the relational side of the picture.
+2.2 Quantum relational constructions
+The fact that a homomorphism $\xi$ between (classical) polymorphism minions $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{Pol}(\mathbb{B})$ gives rise to reductions can be viewed in two distinct ways.
+At the level of the instance, $\xi$ allows a complete and sound reduction from $\operatorname{CSP}(\mathbb{B})$ to $\operatorname{CSP}(\mathbb{A})$ by ensuring that satisfying assignment to the first problem correspond to satisfying assignments of the second.
+At the level of the language, $\xi$ captures the fact that $\mathbb{A}$ can simulate $\mathbb{B}$ via a specific type of uniform construction.
+These two sides are tied together via strong categorical links (see, for example, [40]).
+We now look at the latter side, in the case of quantum CSPs.
+Classically, the right notion of simulation for CSPs is based on primitive positive (pp) formulae—first-order formulae using relations in the language, equality, conjunction, and existential quantification—a concept that originates in model theory and universal algebra.
+Lifting this concept directly to the case of quantum CSPs presents a seriour obstacle: It is well known that, in general, reductions between classical CSPs based on pp-formulae are not complete and sound in the quantum setting.
+This is due to the fact that they alter the commutativity contexts of the instance, see e.g.
+[47, 2, 33].
+For this reason, we lift the notion of pp-formula to a quantum counterpart that provably works as a reduction between quantum CSPs.
+This results in the notion of q-definition, which we formally describe in Section 5.1.
+At a high level, this notion should be viewed as the “minimal set of conditions required to quantise classical pp-definitions of constraint languages”.
+One limitation of q- (and pp-) definitions is that any structure that is definable from $\mathbb{A}$ must have domain $A$.
+In order to be able to reduce between quantum CSPs having distinct domains, we introduce the more general notion of q-constructions, modelled around the classical pp-constructions.
+The next step is to show that q-definitions and the more expressive q-constructions perform the required tasks of simulating a quantum CSP within another.
+To that end, we prove that q-constructions induce minion homomorphisms of the corresponding quantum polymorphism minions (and thus, via Theorem 3, reductions between the corresponding quantum CSPs).
+Theorem 4.
+If $\mathbb{A}$ q-constructs $\mathbb{B}$, then there exists a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$.
+Note that, in the statement above, neither the domains nor the signatures of $\mathbb{A}$ and $\mathbb{B}$ are required to coincide.
+The key to proving the above result is to show that it is possible to perform a “vectorisation step” identifying tuples of tuples with vectors (see condition $(ii)$ in Definition 44) at the level of minion homomorphisms of quantum polymorphism minions (and, thus, of reductions between quantum CSPs).
+While for classical pp-constructions this fact is straightforward, it is not a priori clear that this operation does not affect completeness and soundness in the quantum case.
+One might expect that an extra commutativity gadget (see later) is needed specifically to quantise this part of the classical construction.
+We prove that, in fact, no extra gadget is necessary for this particular step to work.
+To this end, we make use of a minion-theoretic observation on the extension of partial minion homomorphisms defined on elements having only essential coordinates to global minion homomorphisms.
+In the case of q-definitions, we additionally prove a quantum counterpart of the classical operation/relation Galois connection for pp-definitions [15, 38].
+Theorem 5.
+Let $\mathbb{A}$ and $\mathbb{B}$ be relational structures over the same domain.
+Then $\mathbb{A}$ q-defines $\mathbb{B}$ if, and only if, $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(\mathbb{B})$.
+2.3 Non-contextual quantum polymorphisms and commutativity gadgets
+We have, at this point, two perspectives on reductions between quantum CSPs—one algebraic, and one relational.
+The next step is to obtain a direct link between such reductions and their classical counterparts, with the goal of achieving a systematic method for lifting classical hardness reductions to quantum undecidability reductions.
+To that end, we make use of the notion of commutativity gadgets.
+At a high level, the idea is that, while every q-definition is in particular a pp-definition, the converse is not true in general, as the commutativity contexts of the measurements of the quantum strategy for the reduced instance are in general unrelated to those for the initial instance.
+Commutativity gadgets were introduced in [47] to overcome this obstacle for certain constraint languages, by artificially introducing the required commutativity relations between the PVMs.
+Recall that quantum strategies for a non-local game consist of projection-valued measurements (PVMs) with outcome set given by the language $\mathbb{A}$.
+In general, the correlation associated with the measurements of a quantum strategy may locally satisfy all constraints with perfect probability even if such correlation cannot be associated with any system of measurements simultaneously performed on all constraints.
+This is ultimately a phenomenon of quantum contextuality, occurring whenever the PVMs of the strategy do not give rise to a global context on the whole instance—i.e., whenever some of the PVMs do not commute with each other.
+This phenomenon can also occur for quantum polymorphisms: we say that a quantum polymorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ is contextual if at least one pair $(Q_{1},Q_{2})$ of PVMs in $Q$ does not commute.
+If, on the other hand, all PVMs commute, we say that $Q$ is non-contextual.
+As a straightforward application of the spectral theorem, non-contextual quantum polymorphisms can be equivalently viewed as probability distributions over classical deterministic polymorphisms.
+A main conceptual contribution of our work is to show that the existence of a commutativity gadget for a language can be fully characterised by the non-contextuality of its quantum polymorphisms, as stated next.
+Theorem 6.
+Let $\mathbb{A}$ be a relational structure. The following are equivalent:
+  1. 1.
+    $\mathbb{A}$ has a commutativity gadget.
+  2. 2.
+    $\mathbb{A}^{n}$ is a commutativity gadget for all large enough $n\in\mathbb{N}$.
+  3. 3.
+    All quantum polymorphisms of $\mathbb{A}$ are non-contextual.
+In order to establish the $3.\Rightarrow 2.\Rightarrow 1.$ implications in the theorem above, we use the fact that any pair of elements in $A$ can be generated as the image of some polymorphism of high-enough arity $n$.
+Then, we prove that, for such $n$, the $n$-th categorical power of $\mathbb{A}$ yields a commutativity gadget for $\mathbb{A}$ provided that all quantum homomorphisms $\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ (i.e., all $n$-ary quantum polymorphisms of $\mathbb{A}$) are non-contextual.
+In order to prove the $1.\Rightarrow 3.$ implication, the high-level idea is to start with a quantum polymorphism of $\mathbb{A}$, and compose it with a tensor product of quantum homomorphisms from the commutativity gadget to $\mathbb{A}$.
+To make this step work, we make use of the composition operation introduced in [1] to prove that quantum homomorphisms can be expressed as Kleisli morphisms for a graded monad.
+Via the characterisation of commutativity gadgets in Theorem 6, we are able to connect classical pp-constructions to both the algebraic reductions in Section 2.1, and the relational constructions from Section 2.2.
+As a consequence, we prove that if $\mathbb{A}$ $p$-constructs $\mathbb{B}$ and all quantum polymorphisms of $\mathbb{A}$ are non-contextual, then $\mathbb{A}$ q-constructs $\mathbb{B}$ and, thus, $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$.
+In particular, this gives the following.
+Theorem 7.
+Let $\mathbb{A}$ be a structure such that $\operatorname{CSP}(\mathbb{A})$ is not in $\mathsf{P}$ and all quantum polymorphisms of $\mathbb{A}$ are non-contextual.
+Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+This result yields a broadly applicable undecidability criterion for quantum CSPs.
+[^4] We point out that the condition $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$ in Theorem 7 can be replaced by the algebraic condition of $\operatorname{Pol}(\mathbb{A})$ being trivial (or, equivalently, by $\mathbb{A}$ pp-constructing, say, 3SAT).
+In this way, the result does not trivialise if $\mathsf{P}=\mathsf{NP}$; see Corollary 56.
+In particular, it captures and extends the known undecidability results established in [2] and in [33]: We show in Section 5.4 that any relational structure that does not meet the so-called two-variable falsifiability (TVF) property (and, thus, is captured by the undecidability criteria of [2, 33]) cannot have contextual quantum polymorphisms and, thus, is captured by Theorem 7.
+As we shall see, the latter result is strictly more expressive, as it applies to various classes of CSP languages that are not captured by [2, 33].
+2.4 Contextuality and bifurcations
+A consequence of the machinery developed so far is that we are now able to systematically investigate the complexity landscape of quantum CSPs by studying the properties of their quantum polymorphisms—in particular, whether they are non-contextual.
+Recall that we defined a quantum homomorphism to be contextual when at least two of the measurements of the corresponding quantum strategy do not commute and, thus, cannot be simultaneously performed in at least one reference system.
+Before proceeding further, it shall be useful to address the following question: What information can we derive from the fact that a quantum homomorphism is contextual?
+Informally, the idea here is to “trade non-commutativity for non-orthogonality”.
+At first glance, this trade appears unfavourable as any two non-commuting projectors are, in particular, non-orthogonal.
+However, it turns out that a single instance of non-commutativity among projectors arising in a perfect quantum strategy generates a structured pattern of multiple non-orthogonality relations.
+We call such non-orthogonality patterns (contextuality) bifurcations.
+(For the formal definition, see Section 7; two illustrations of such patterns are given in Figure 3 and Figure 4.) Specifically, we prove the following.
+Theorem 8.
+Let $\mathbb{X}$ and $\mathbb{A}$ be relational structures such that $\operatorname{Gaif}(\mathbb{X})$ is a connected graph, and let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a contextual quantum homomorphism.
+Then there exists a bifurcation for $Q$ whose length is at most the diameter of $\operatorname{Gaif}(\mathbb{X})$.
+In practice, the strategy for applying Theorem 7 to show that the quantum polymorphisms of a structure $\mathbb{A}$ are non-contextual (and hence that $\mathbb{A}$ admits commutativity gadgets) proceeds as follows.
+First, one observes that the support of any bifurcation defines a substructure $\mathbb{G}\subseteq\mathbb{A}^{n}\times\mathbb{A}$, since $Q$ is a perfect strategy.
+Next, one shows that the existence of a contextual quantum polymorphism would yield sufficiently many bifurcations (or bifurcations of a suitable form) to force $\mathbb{G}$ to contain edges not present in $\mathbb{A}^{n}\times\mathbb{A}$, leading to a contradiction.
+Figure 1: A reduction proving undecidability of $\operatorname{CSP_{q}}(\mathbb{C}_{5})$ by a reduction from the ternary Boolean 1-in-3 predicate.
+Each variable of an instance of 1-in-3-SAT is replaced by a cloud of $5^{2}$ vertices as shown, isomorphic to the graph $(\mathbb{C}_{5})^{2}$.
+Each constraint $C(x,y,z)$ is then replaced by a cloud of $5^{3}$ vertices inducing the graph $(\mathbb{C}_{5})^{3}$, and the variable-clouds are identified with subsets of the constraint-cloud as depicted in the image.
+2.5 Quantum polymorphisms for Siggers and odd cycles
+The machinery developed so far allows transferring undecidability across classes of quantum CSPs, in a similar way as NP-hardness propagates across classical CSPs via gadget reductions associated with pp-definitions and pp-constructions.
+To showcase this strategy, we prove undecidability of the quantum CSPs parameterised by two simple types of languages.
+In Section 9, we prove the following.
+Theorem 9.
+Let $\mathbb{C}_{m}$ be the undirected cycle of size $m$. Then $\operatorname{CSP_{q}}(\mathbb{C}_{m})$ is undecidable for each odd $m\geq 3$.
+This resolves an open question raised in [34], where it was mentioned that the known techniques for proving undecidability of quantum CSPs do not capture this class of graphs.
+The Siggers digraph is the $3$-vertex digraph obtained by making one edge of a directed triangle undirected.
+The prominence of this language in algebraic CSP theory lies in the fact that the corresponding loop condition gives rise to Siggers-identities [49, Theorem 2.2], which are known to exactly capture the borderline between tractability and hardness for classical CSPs [9] (like other types of similar identities, see e.g.
+[54]).
+In Section 8, we prove the following.
+Theorem 10.
+Let $\mathbb{A}$ be the Siggers digraph. Then $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+The proofs of these results, given in full detail in the body of the paper, are based on establishing that the quantum polymorphisms of such languages are non-contextual, via a combination of the bifurcation result (Theorem 8) and ad-hoc arguments exploiting the particular structures of the languages.
+Then, undecidability follows from Theorem 7.
+2.6 Dichotomy for Boolean languages
+As a third application of our framework to concrete classes of languages, we consider the case of Boolean structures, and we obtain a complete classification of the Boolean structures admitting a commutativity gadget.
+Theorem 11.
+Let $\mathbb{A}$ be a Boolean relational structure. Then the following are equivalent:
+  • $(i)$
+    $\mathbb{A}$ admits a commutativity gadget;
+  • $(ii)$
+    $\mathbb{A}$ has no majority polymorphism or is not two-variable falsifiable.
+We point out that the $(ii)\Rightarrow(i)$ implication was shown in [33]; here, we provide a different proof based on the analysis of quantum polymorphisms of Boolean structures.
+We show that the polymorphisms of 1-in-3-SAT, are non-contextual and then use quantum compatible pp-definitions to show this for other Boolean structures.
+We note that some of these arguments make use of non-oracular quantum polymorphisms, see Section 2.7.
+To obtain the converse implication $(i)\Rightarrow(ii)$, we identify a structure $\operatorname{\mathbb{M}}$ whose quantum polymorphism minion $\operatorname{qPol}(\operatorname{\mathbb{M}})$ has the property of being contained in the quantum polymorphism minion of any Boolean structure that is invariant under majority and is two-variable falsifiable, see Figure 2 and Theorem 115.
+Then, we prove that the minion $\operatorname{qPol}(\operatorname{\mathbb{M}})$ is contextual, as it contains contextual quantum polymorphisms of any arity $\geq 4$.
+On the other hand, we show that $\operatorname{\mathbb{M}}$ does not admit contextual quantum polymorphisms of any arity $n\leq 3$ (see Theorem 116).
+In particular, this means that all quantum endomorphisms of $\operatorname{\mathbb{M}}$ are non-contextual.
+Hence, the necessary condition for the existence of commutativity gadgets given in [34] is provably insufficient to establish that Boolean TVF structures that are preserved by majority cannot admit a commutativity gadget, and the full power of our quantum polymorphism characterisation is needed.
+0000100001000010000111001010100101100101001111101101101101111111
+Figure 2: Illustration of the $4$-th power of the structure $\operatorname{\mathbb{M}}$ yielding the minimal contextual minion.
+The proof of the $(i)\Rightarrow(ii)$ implication of Theorem 11 relies on showing that this structure admits a contextual quantum homomorphism to $\operatorname{\mathbb{M}}$; see Section 11.2 for full details.
+Combining Theorem 11 with results from [67, 2], we obtain the following complexity dichotomy for quantum CSPs parameterised by Boolean languages.
+We denote by $\mathbb{XOR}$ the structure encoding Boolean linear systems of equations.
+Corollary 12.
+Let $\mathbb{A}$ be a Boolean relational structure.
+If $\mathbb{A}$ $p$-defines $\mathbb{XOR}$, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable; otherwise, $\operatorname{CSP_{q}}(\mathbb{A})$ is solvable in polynomial time.
+We refer the reader to Remark 107 for a comparison of this result with the Boolean dichotomies given in [63, Theorem 5.11 (a)] and in [2, §6.1].
+2.7 Non-oracular quantum polymorphisms
+So far, we have only considered the oracular setting for quantum homomorphisms.
+At the level of two-player games, this corresponds to the so-called constraint-constraint or constraint-variable formulations of the CSP game (see, for example, as opposed to the variable-variable variant where the cooperating players both receive as questions two variables of the instance).
+At the level of quantum homomorphisms, this corresponds to requiring that the projectors in a winning strategy associated with variables in the same constraint (or context) should commute.
+We believe this is the natural setting for establishing a theory of reductions between quantum CSPs, for two reasons.
+First, categorical interpretations of quantum CSPs—primarily, the quantum-monad formulation of [1]—appear to naturally capture the oracular setting for CSPs of arbitrary arities.
+Second, in the gapped version of quantum CSPs, the definition of the quantum value of a game in terms of the expected value of the trace of products of projectors in a randomly sampled constraint requires commutativity between such projectors in order for the trace to be a real number, if the arity is at least $3$; see, for example, the discussion in [62, §3.1].
+On the other hand, in the case of binary CSPs, both notions are natural, and they are known to sometimes give rise to different behaviours.
+For example, the non-oracular Unique Games are provably in P [50], while it is plausible that the oracular version is undecidable [62]; see also [48].
+See also [1, §4.3], where non-oracular quantum graph homomorphisms in the sense of [59] are shown to be captured via oracular homomorphisms for specific Boolean constraint systems.
+We show that the machinery illustrated so far extends to a large extent to the non-oracular setting.
+In particular, we define the non-oracular version of quantum polymorphisms, and we show that it precisely captures the existence of commutativity gadgets for non-oracular quantum CSPs in exactly the same way as in the oracular case; see Theorem 97.
+Furthermore, we extend the notion of q-definition to the non-oracular case, and we prove that the relation/operation Galois connection of Theorem 5 extends to the non-oracular setting.
+Theorem 13.
+Two structures $\mathbb{A},\mathbb{B}$ satisfy $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ if, and only if, $\mathbb{A}$ q-no-defines $\mathbb{B}$.
+We refer to Section 10 for more details on the properties of quantum non-oracular polymorphisms, and the reductions that are captured by them.
+2.8 Outlook
+Quantum polymorphisms provide a bridge between two types of methods: $(i)$ reductions from the classical algebraic CSP theory, and $(ii)$ the quantum-specific techniques used in the complexity analysis of non-local games—in particular, commutativity gadgets.
+It is, in our view, quite surprising that both $(i)$ and $(ii)$ can be captured within a single formalism, given the different contexts from which they originate.
+Establishing this link is a key conceptual contribution of the paper and plays a crucial role in our undecidability results for concrete classes of languages.
+In more detail, quantum polymorphisms are ultimately encodings of specific quantum strategies; as such, they exhibit, in general, non-classical behaviours—specifically, non-commutativity of measurements across different contexts.
+This is precisely what enables them to capture (and characterise the existence of) commutativity gadgets.
+At the same time, quantum polymorphisms admit the same algebraic structure as their classical counterparts, namely that of minions, and can thus be related to other such structures on a purely algebraic ground.
+It is this dual nature that makes quantum polymorphisms able to transfer undecidability across different problems over different languages—simulating the way NP-completeness is transferred across classical CSPs.
+Among the open directions arising from this paper, we mention the prospect of obtaining a stronger contextuality bifurcation theorem than Theorem 8, resulting in more complex non-orthogonality patterns generated by contextual homomorphisms.
+This would allow applying our undecidability criterion Theorem 7 to new classes of languages.
+Similarly, it is an interesting prospect to adapt and extend the techniques used in Section 11 for proving the existence of contextual quantum polymorphisms for non-Boolean languages, thus ruling out the existence of commutativity gadgets.
+While in the current paper we exclusively deal with the exact version of quantum CSPs—i.e., the decision problem asking to decide whether a perfect quantum strategy for the corresponding 2-player game exists—techniques based on the weighted-algebra formalism introduced by Mastel–Slofstra [60] appear to be suited for a translation of our results to the setting of imperfect soundness (see also [34], where it is shown that Ji’s commutativity gadgets are robust, in the sense that they can be used in the imperfect-soundness version of the quantum CSP).
+Finally, in the current work we only consider quantum strategies for non-local games, described by measurements of a finite-dimensional bipartite quantum state.
+An interesting direction for follow-up work is to translate our results to the case of quantum-approximable strategies (i.e., limits of quantum strategies) and commuting-operator strategies (i.e., possibly infinite-dimensional quantum strategies where the two players’ measurements do not act on separate tensor factors, but are merely required to commute with each other); see for example [31, 57].
 Structure of the paper
-In Section 2, we give formal definitions of classical and entangled CSPs, and we list some basic results about quantum homomorphisms.
-Section 3 presents the main definition of this work—quantum polymorphims—as well as their use to characterise the existence of oracular commutativity gadgets for entangled CSPs.
-In Section 4 we show that the quantum polymorphisms of cliques are non-contextual, which gives a different proof of the undecidability of the corresponding entangled CSPs due to [ji2013binary, Zeman].
-In section 5, we prove a general result on the consequence of contextuality in quantum homomorphisms, by showing that it causes the emergence of a specific non-orthogonality pattern in the underlying structures.
-We make use of this pattern in section 6, to prove that odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable.
-In Section 7, we address the non-oracular case, by showing that a similar characterisation to the one in Section 3 captures commutativity gadgets in the non-oracular setting.
-In addition, we establish therein a quantum version of Galois connection for the non-oracular quantum analogue of pp-definitions.
-Finally, in Section 8, we investigate the existence of commutativity gadgets for structures with a 2-element domain.
-We obtain that structures that do not admit the ternary majority as a polymorphism must have a commutativity gadget, and that apart from trivial cases, structures that do have the majority operation as a polymorphism do not have a commutativity gadget.
-2. Preliminaries
-2.1. Relational structures
+The remaining part of the paper is organised as follows.
+In Section 3, we formally define classical and quantum CSPs, and we give a list of some basic results about quantum homomorphisms and operations thereof that shall we useful throughout the paper.
+Section 4 presents the algebraic side of our result: we show in Section 4.1 that quantum polymorphisms are closed under minor operations and, thus, form an abstract minion, and in Section 4.2 we prove that minion homomorphism between quantum polymorphism minions determine complexity reductions between the corresponding quantum CSPs.
+We then turn to the relational side of classical/quantum bridge in Section 5, by formally describing q-definitions and q-constructions in Section 5.1, as well as establishing a Galois connection with quantum polymorphism.
+In Section 5.2 and Section 5.3, we show how the contextuality of quantum polymorphisms of a language fully captures commutativity gadgets, which translates into a fully polymorphic, concrete undecidability criterion for quantum CSPs (Theorem 7).
+This provably extends the known undecidability criteria in the literature on non-local games, as discussed in Section 5.4.
+In Section 6, we describe the quantum polymorphisms of cliques and obtain a simple proof of the undecidability of the corresponding CSPs, first established in [47, 33, 34].
+Next, in Section 7, we describe non-orthogonality patterns for systems of orthogonal projectors that we call bifurcations, and we show that they occur in any contextual quantum homomorphism (or quantum polymorphism).
+We then specialise our analysis to particular classes of quantum CSPs.
+In particular, by using the polymorphic framework, contextuality bifurcations, and ad-hoc arguments based on the specific types of constraints, we establish in Section 8 and Section 9 the undecidability of the problems encoding quantum satisfiability of Siggers clauses and quantum homomorphisms to odd cycles, respectively.
+In Section 10, we consider the case of non-oracular quantum CSPs, we define the non-oracular version of quantum polymorphisms, and we show that they capture non-oracular commutativity gadgets in a similar way as in the oracular setting.
+Finally, Section 11 address quantum CSPs parameterised by Boolean languages.
+By describing the quantum polymorphisms of such languages, we obtain a full classification of the existence of commutativity gadgets, as well as a complete $\mathsf{P}$ vs.
+undecidable complexity dichotomy for Boolean quantum CSPs.
+3 Preliminaries
+In this section, we give formal definitions of most of the notions used throughout the paper, and we list a few simple facts on quantum homomorphisms and their operations.
+3.1 Relational structures
+We let $\mathbb{N}=\{1,2,\dots\}$ denote the set of positive integers.
 For $\ell\in\mathbb{N}$, $\ell$] is the set $\{1,2,\dots,\ell\}$.
@@
 All relational structures in this work shall be implicitly assumed to be finite.
-A (classical) homomorphism $f:\mathbb{X}\to\mathbb{Y}$ between two relational structures $\mathbb{X}$ and $\mathbb{Y}$ on a common signature $\sigma$ is a map from the domain $X$ of $\mathbb{X}$ to the domain $Y$ of $\mathbb{Y}$ that preserves all relations; i.e., $f(\mathbf{x})\in R^{\mathbb{Y}}$ for each symbol $R\in\sigma$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$ (where $f(\mathbf{x})$ is the entrywise application of $f$ the the entries of $\mathbf{x}$).
+A (classical) homomorphism $f\colon\mathbb{X}\to\mathbb{Y}$ between two relational structures $\mathbb{X}$ and $\mathbb{Y}$ on a common signature $\sigma$ is a map from the domain $X$ of $\mathbb{X}$ to the domain $Y$ of $\mathbb{Y}$ that preserves all relations; i.e., $f(\mathbf{x})\in R^{\mathbb{Y}}$ for each symbol $R\in\sigma$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$ (where $f(\mathbf{x})$ is the entrywise application of $f$ to the entries of $\mathbf{x}$).
 We denote the existence of a classical homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ by the notation $\mathbb{X}\to\mathbb{Y}$.
-Given a $\sigma$-structure $\mathbb{A}$, a positive integer $r\in\mathbb{N}$, and a set $S\subseteq A^{r}$, we define the structure $\mathbb{A}_{S}$ as follows.
-The signature of $\mathbb{A}_{S}$ contains a unique $r$-ary symbol $R$ (which does not belong to $\sigma$), and the interpretation of $R$ in $\mathbb{A}_{S}$ is $R^{\mathbb{A}_{S}}=S$.
-For typographical reasons, we shall sometimes denote $\mathbb{A}_{S}$ by $(A;S)$.
-For a digraph $\mathbb{X}$, we let $\mathbb{X}^{u}$ be the undirected graph on the same vertex set of $\mathbb{X}$, having an undirected edge $\{x,y\}$ whenever $(x,y)$ or $(y,x)$ form an edge in $\mathbb{X}$.
-We let the Gaifman graph of a relational structure $\mathbb{X}$ be the loopless directed graph $\operatorname{Gaif}(\mathbb{X})$ on vertex set $V(\operatorname{Gaif}(\mathbb{X}))=X$ and edge set $E(\operatorname{Gaif}(\mathbb{X}))$ containing a pair $(x,y)$ if, and only if, $x$ and $y$ appear together in some tuple in some relation of $\mathbb{X}$.
+Moreover, we let $\operatorname{Hom}(\mathbb{X},\mathbb{Y})$ be the set of all homomorphisms $f:\mathbb{X}\to\mathbb{Y}$.
+Given a $\sigma$-structure $\mathbb{X}$, a positive integer $r\in\mathbb{N}$, and a set $S\subseteq X^{r}$, we let $(X;S)$ be the structure with domain $X$ and a single relation $S$.
+We let the Gaifman graph of a relational structure $\mathbb{X}$ be the loopless undirected graph $\operatorname{Gaif}(\mathbb{X})$ on vertex set $V(\operatorname{Gaif}(\mathbb{X}))=X$ and edge set $E(\operatorname{Gaif}(\mathbb{X}))$ containing an edge $\{x,y\}$ if, and only if, $x$ and $y$ appear together in a tuple in some relation of $\mathbb{X}$.
 Given a $\sigma$-structure $\mathbb{Y}$, the constraint satisfaction problem parameterised by $\mathbb{Y}$ (in short, $\operatorname{CSP}(\mathbb{Y})$) is the following computational problem: [^5] This is the decision version of $\operatorname{CSP}(\mathbb{Y})$.
 In the search version, the problem is to find an explicit homomorphism $f:\mathbb{X}\to\mathbb{Y}$ for any input $\mathbb{X}$ such that $\mathbb{X}\to\mathbb{Y}$.
-The two versions of $\operatorname{CSP}$ are equivalent up to polynomial-time reductions [bulatov2005classifying].
+The two versions of $\operatorname{CSP}$ are equivalent up to polynomial-time reductions [22].
 Given as input a $\sigma$-structure $\mathbb{X}$, output $\mathrm{YES}$ if $\mathbb{X}\to\mathbb{Y}$, and $\mathrm{NO}$ if $\mathbb{X}\not\to\mathbb{Y}$.
 For $\ell\in\mathbb{N}$, we denote by $\mathbb{Y}^{\ell}$ the $\ell$-th direct power of $\mathbb{Y}$; i.e., $\mathbb{Y}^{\ell}$ is the $\sigma$-structure whose domain is $Y^{\ell}$ and whose relations are defined as follows: given $R\in\sigma$ and any $\ell\times\operatorname{ar}(R)$ matrix $M$ whose rows are tuples in $R^{\mathbb{Y}}$, the columns of $M$ form a tuple in $R^{\mathbb{Y}^{\ell}}$.
-A polymorphism of $\mathbb{Y}$ (of arity $\ell$) is a homomorphism from $\mathbb{Y}^{\ell}$ to $\mathbb{Y}$.
+A (classical) polymorphism of $\mathbb{Y}$ (of arity $\ell$) is a homomorphism from $\mathbb{Y}^{\ell}$ to $\mathbb{Y}$.
 By $\operatorname{Pol}(\mathbb{Y})$ (resp.
@@
 all polymorphisms of arity $\ell$) of $\mathbb{Y}$.
-The core of a $\sigma$-structure $\mathbb{X}$—in symbols, $\operatorname{core}(\mathbb{X})$—is a minimal (under inclusion) induced substructure of $\mathbb{X}$ that is homomorphically equivalent to $\mathbb{X}$.
-It is easy to check that $\operatorname{core}(\mathbb{X})$ is unique up to isomorphism, and any endomorphism is an automorphism (see e.g.
-[HellNesetrilCore]).
-The primitive positive (pp) fragment of first-order logic consists of the formulas that are built out of atomic formulas using conjunctions and existential quantifiers.
-If $\mathbb{A}$ is a relational structure and $\varphi(x_{1},\dots,x_{n})$ is a primitive positive formula in the signature of $\mathbb{A}$ with free variables $x_{1},\dots,x_{n}$, then the relation defined by $\varphi$ is the set of $n$-tuples that satisfy the formula in $\mathbb{A}$.
-If $R\subseteq A^{n}$ is a relation, we say that it is pp-definable over $\mathbb{A}$ if there exists a pp-formula defining it.
-A classical fact about pp-formulas is the following duality with gadgets: every pp-formula can be understood as a relational structure $\mathbb{X}_{\varphi}$ with distinguished vertices $x_{1},\dots,x_{n}$ such that for every structure $\mathbb{A}$, $(a_{1},\dots,a_{n})$ is in the relation defined by $\varphi$ in $\mathbb{A}$ if, and only if, there exists a homomorphism $\mathbb{X}_{\varphi}\to\mathbb{A}$ that maps $x_{i}$ to $a_{i}$.
-2.2. Quantum CSPs
+The core of a $\sigma$-structure $\mathbb{X}$ is a minimal (under inclusion) induced substructure of $\mathbb{X}$ that is homomorphically equivalent to $\mathbb{X}$.
+It is easy to check that the core of a structure is unique up to isomorphism, and any endomorphism is an automorphism (see e.g.
+[43]).
+The primitive positive (pp) fragment of first-order logic consists of the formulae that are built out of atomic formulae using conjunctions and existential quantifiers.
+If $\mathbb{Y}$ is a relational structure and $\varphi(x_{1},\dots,x_{n})$ is a primitive positive formula in the signature of $\mathbb{Y}$ with free variables $x_{1},\dots,x_{n}$, then the relation defined by $\varphi$ is the set of $n$-tuples that satisfy the formula in $\mathbb{Y}$.
+If $R\subseteq Y^{n}$ is a relation, we say that it is pp-definable over $\mathbb{Y}$ if there exists a pp-formula defining it.
+A classical fact about pp-formulae is the following duality with gadgets: every pp-formula can be understood as a relational structure $\mathbb{X}_{\varphi}$ with distinguished vertices $x_{1},\dots,x_{n}$ such that for every structure $\mathbb{Y}$, $(y_{1},\dots,y_{n})$ is in the relation defined by $\varphi$ in $\mathbb{Y}$ if, and only if, there exists a homomorphism $\mathbb{X}_{\varphi}\to\mathbb{Y}$ that maps $x_{i}$ to $y_{i}$ for each $i\in[n]$.
+Suppose $\mathbb{X}$ and $\mathbb{Y}$ are relational structures with possible distinct domains and signatures.
+We say that $\mathbb{Y}$ has a pp-construction in $\mathbb{X}$ (or that $\mathbb{X}$ $p$-constructs $\mathbb{Y}$) if $\mathbb{Y}$ is homomorphically equivalent to a structure having domain $X^{d}$ for some $d\in\mathbb{N}$, and relations $R_{1},\dots,R_{k}$, where each $R_{i}\subseteq(X^{d})^{r_{i}}$ is pp-definable over $\mathbb{X}$, when seen as a relation $\tilde{R_{i}}\subseteq X^{dr_{i}}$ by “vectorising” tuples of tuples.
+3.2 Quantum CSPs
 Let $H$ be a non-trivial finite-dimensional Hilbert space.
@@
 We denote by $\operatorname{id}_{H}$ (or simply by $\operatorname{id}$ when the Hilbert space is clear) the identity projector over $H$.
-A projective measurement (PVM) over a Hilbert space $H$ and a set $B$ and is a family $(Q_{b})_{b\in B}$ of projectors on $H$ with the property that $\sum_{b\in B}Q_{b}=\operatorname{id}$.
+A projective measurement (PVM) over a Hilbert space $H$ with outcome set $B$ is a family $(Q_{b})_{b\in B}$ of projectors on $H$ with the property that $\sum_{b\in B}Q_{b}=\operatorname{id}$.
 A quantum function from $A$ to $B$ is a family of PVMs $(Q_{a,b})_{b\in B}$, one PVM for each $a\in A$.
-We shall denote this by the notation $Q\colon A\to_{q}B$ or $Q\colon A\to_{H}B$ if we want to emphasize the Hilbert space.
-We say that two PVMs $(Q_{b})_{b\in B}$ and $(Q^{\prime}_{b^{\prime}})_{b^{\prime}\in B^{\prime}}$ over the same Hilbert space commute if $Q_{b},Q^{\prime}_{b^{\prime}}]=0$ for each $b\in B,b^{\prime}\in B^{\prime}$.
-Given a quantum function $Q:A\to_{q}B$ for two sets $A,B$, we say that $Q$ is non-contextual if $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for each $a,a^{\prime}\in A$ and each $b,b^{\prime}\in B$.
-We say that $Q$ is contextual if it is not non-contextual.
-Note that each classical function $f:A\to B$ can be seen as a non-contextual quantum function.
+We shall denote this by the notation $Q\colon A\to_{\operatorname{q}}B$ or $Q\colon A\to_{H}B$ if we want to emphasise the Hilbert space.
+We say that two PVMs $(Q_{b})_{b\in B}$ and $(Q^{\prime}_{b^{\prime}})_{b^{\prime}\in B^{\prime}}$ over the same Hilbert space commute if $Q_{b},Q^{\prime}_{b^{\prime}}]=0$ for each $b\in B$ and each $b^{\prime}\in B^{\prime}$.
+Given a quantum function $Q\colon A\to_{\operatorname{q}}B$ for two sets $A,B$, we say that $Q$ is non-contextual if $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for each $a,a^{\prime}\in A$ and each $b,b^{\prime}\in B$.
+Otherwise, we say that $Q$ is contextual.
+Note that each classical function $f\colon A\to B$ can be seen as a non-contextual quantum function by setting $Q_{a,b}=\operatorname{id}$ if $f(a)=b$ and $Q_{a,b}=0$ otherwise.
 Let $\mathbb{A},\mathbb{B}$ be relational structures with the same signature $\sigma$.
-A quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$, first defined at this level of generality in [abramsky2017quantum], is given by a non-trivial Hilbert space $H$ and a quantum function $Q\colon A\to_{q}B$ such that the following conditions hold:
+A quantum homomorphism $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$, first defined at this level of generality in [1], is given by a non-trivial finite-dimensional Hilbert space $H$ and a quantum function $Q\colon A\to_{\operatorname{q}}B$ such that the following conditions hold:
   • ($\mathbf{QH_{1}}$)
@@
     $[Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0\quad\forall(a,a^{\prime})\in E(\operatorname{Gaif}(\mathbb{A})),(b,b^{\prime})\in B^{2}$.
-Note: Condition $($$\mathbf{QH_{2}}$)) means that $Q$ corresponds to perfect oracular correlations.
-In Section 7 we shall extend our results to the non-oracular case.
-We also point out that the notion of a quantum homomorphism can be phrased equivalently as a representation of a certain finitely generated *-algebra [MastelSlofstra2024].
-Definition 1 (Entangled CSP).
+It was proved in [1] that a quantum homomorphism $\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ exists if and only if the constraint-variable game associated with $\mathbb{A}$ and $\mathbb{B}$ admits a winning finite-dimensional tensor-product strategy.
+In the literature on non-local games, this is often expressed in terms of the existence of a perfect correlation in the set $C_{q}$ of quantum correlation matrices, see for example [67, 41, 46, 56].
+Equivalently, the notion of a quantum homomorphism can be phrased as a representation of a certain finitely generated *-algebra [31, 64, 52, 60, 63].
+Condition $($$\mathbf{QH_{2}}$)) is sometimes phrased as $Q$ expressing perfect oracular correlations; in Section 10 we shall extend our results to the non-oracular case.
+Definition 14 (Quantum CSP).
 Fix a $\sigma$-structure $\mathbb{A}$.
-$\operatorname{CSP^{*}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{A}$.
-2.3. Operations on quantum homomorphisms
-We describe here certain constructions taking as input two quantum homomorphisms and giving another quantum homomorphism as the output.
-Recall that if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\oplus Q^{\prime}$ is the operator on $H\oplus H^{\prime}$ defined by $(Q\oplus Q^{\prime})(v+v^{\prime})=Q(v)+Q(v^{\prime})$.
-We extend this definition to quantum functions $Q\colon X\to_{H}A$, $Q^{\prime}:X\to_{H^{\prime}}A$ by letting $Q\oplus Q^{\prime}:X\to_{H\oplus H^{\prime}}A$ be defined by $(Q\oplus Q^{\prime})_{x,a}:=Q_{x,a}\oplus Q^{\prime}_{x,a}$.
+$\operatorname{CSP_{q}}(\mathbb{A})$ is the following computational problem: given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a quantum homomorphism $\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$.
+We will use the next observation on the length of walks in $\operatorname{Gaif}(\mathbb{A})$ and $\operatorname{Gaif}(\mathbb{B})$ for two structures $\mathbb{A}$ and $\mathbb{B}$ that are quantum homomorphic.
+This result is a slight modification of [59, Lemma 4.12] (which concerns the case where $\mathbb{A}$ and $\mathbb{B}$ are undirected graphs), and its proof closely follows that in [59].
+Definition 15.
+Let $\mathbb{A}$ be a relational structure with signature $\sigma$.
+We say that $\mathbb{A}$ is essentially binary if, for every $R\in\sigma$ and every $\mathbf{a}\in R^{\mathbb{A}}$, $\mathbf{a}$ has at most two distinct entries.
+Proposition 16.
+Let $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ be a quantum homomorphism between two structures $\mathbb{A}$ and $\mathbb{B}$ such that $\mathbb{A}$ is essentially binary.
+If $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$ are such that there exists an $\ell$-walk from $a$ to $a^{\prime}$ in $\operatorname{Gaif}(\mathbb{A})$ but there is no $\ell$-walk from $b$ to $b^{\prime}$ in $\operatorname{Gaif}(\mathbb{B})$, then $Q_{a,b}Q_{a^{\prime},b^{\prime}}=0$.
+Proof.
+We use induction over $\ell$.
+For $\ell=0$, the result is clear.
+Suppose $\ell=1$; i.e., $\{a,a^{\prime}\}\in E(\operatorname{Gaif}(\mathbb{A}))$ but $\{b,b^{\prime}\}\not\in E(\operatorname{Gaif}(\mathbb{B}))$.
+Assume, by contradiction, that $Q_{a,b}Q_{a^{\prime},b^{\prime}}\neq 0$.
+Since $\mathbb{A}$ is essentially binary, there exist some symbol $R\in\sigma$ and some tuple $\mathbf{z}=(z_{1},\dots,z_{r})\in R^{\mathbb{A}}$ such that $\{z_{1},\dots,z_{r}\}=\{a,a^{\prime}\}$.
+Consider now the tuple $\mathbf{w}=(w_{1},\dots,w_{r})\in B^{r}$ such that, for each $i\in[r]$, $w_{i}=b$ if $z_{i}=a$, and $w_{i}=b^{\prime}$ if $z_{i}=a^{\prime}$.
+Using $($$\mathbf{QH_{2}}$)) and the idempotency of projectors, we find that
+$\prod_{i\in[r]}Q_{z_{i},w_{i}}=Q_{a,b}Q_{a^{\prime},b^{\prime}}\neq 0.$
+We then deduce from $($$\mathbf{QH_{1}}$)) that $\mathbf{w}\in R^{\mathbb{B}}$, so $\{b,b^{\prime}\}\in E(\operatorname{Gaif}(\mathbb{B}))$; a contradiction.
+Take now $\ell\geq 2$, and suppose the result is true for all smaller values of $\ell$.
+Let $a_{0},a_{1},\dots,a_{\ell}$ be an $\ell$-walk in $\operatorname{Gaif}(\mathbb{A})$ such that $a_{0}=a$ and $a_{\ell}=a^{\prime}$.
+We have
+$Q_{a,b}Q_{a^{\prime},b^{\prime}}=\sum_{z\in B}Q_{a,b}Q_{a_{\ell-1},z}Q_{a^{\prime},b^{\prime}}.$
+Applying the inductive hypothesis to the $(\ell-1)$-walk $a_{0},\dots,a_{\ell-1}$ and the $1$-walk $a_{\ell-1},a_{\ell}$ in $\operatorname{Gaif}(\mathbb{A})$, we deduce that, in order for $z$ to yield a nonzero contribution in the summation above, $z$ must be connected to $b$ via an $(\ell-1)$-walk in $\operatorname{Gaif}(\mathbb{B})$, and to $b^{\prime}$ via a $1$-walk in $\operatorname{Gaif}(\mathbb{B})$.
+Concatenating the two walks yields an $\ell$-walk in $\operatorname{Gaif}(\mathbb{B})$ connecting $b$ to $b^{\prime}$.
+Hence, if no such walk exists, the product $Q_{a,b}Q_{a^{\prime},b^{\prime}}$ must be zero.
+∎
+3.3 Operations on quantum homomorphisms
+In this subsection, we describe certain constructions taking as input two quantum homomorphisms and returning another quantum homomorphism as the output.
+Given a linear operator $Q$ on $H$ and a linear operator $Q^{\prime}$ on $H^{\prime}$, we let $Q\oplus Q^{\prime}$ be the linear operator on $H\oplus H^{\prime}$ defined by $(Q\oplus Q^{\prime})(v+v^{\prime})=Q(v)+Q(v^{\prime})$.
+We extend this definition to quantum functions $Q\colon X\to_{H}A$, $Q^{\prime}\colon X\to_{H^{\prime}}A$ by letting $Q\oplus Q^{\prime}\colon X\to_{H\oplus H^{\prime}}A$ be defined by $(Q\oplus Q^{\prime})_{x,a}:=Q_{x,a}\oplus Q^{\prime}_{x,a}$.
 Note that this is well defined as the set $\{Q_{x,a}\oplus Q^{\prime}_{x,a}\}_{a\in A}$ is a PVM over the Hilbert space $H\oplus H^{\prime}$.
@@
 Moreover, it is easy to verify that the projector $\sum_{a\in A}(Q_{x,a}\oplus Q^{\prime}_{x,a})$ acts like the identity on both $H$ and $H^{\prime}$—and, hence, on the whole $H\oplus H^{\prime}$ by linearity.
-Similarly, if $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, then $Q\otimes Q^{\prime}$ is the linear operator on $H\otimes H^{\prime}$ defined by letting $(Q\otimes Q^{\prime})(v\otimes v^{\prime})=Q(v)\otimes Q^{\prime}(v^{\prime})$ for all $v\in H$ and $v^{\prime}\in H^{\prime}$ and extending by linearity.
-As for the direct sum, we extend this definition to quantum functions $Q\colon X\to_{H}A$ and $Q^{\prime}\colon X\to_{H^{\prime}}B$ by $(Q\otimes Q^{\prime})_{x,(a,b)}:=Q_{x,a}\otimes Q^{\prime}_{x,b}$ for all $a\in A$ and $b\in B$.
-This is a quantum function $Q\otimes Q^{\prime}\colon X\to_{H\otimes H^{\prime}}A\times B$.
+We now show that this operation preserves quantum homomorphisms.
 Lemma 17.
@@
 Then $R_{x_{1},a_{1}}\ldots R_{x_{r},a_{r}}=Q_{x_{1},a_{1}}\ldots Q_{x_{r},a_{r}}\oplus Q^{\prime}_{x_{1},a_{1}}\ldots Q^{\prime}_{x_{r},a_{r}}=0$, so $($$\mathbf{QH_{1}}$)) holds.
-For $($$\mathbf{QH_{2}}$)), take a symbol $R\in\sigma$ of arity $r$ and let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$, $i,j\in\{1,\dots,r\}$, and $a,a^{\prime}\in A$.
+For $($$\mathbf{QH_{2}}$)), take a symbol $R\in\sigma$ of arity $r$ and let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$, $i,j\in[r]$, and $a,a^{\prime}\in A$.
 Then $R_{x_{i},a},R_{x_{j},a^{\prime}}]=[Q_{x_{i},a},Q_{x_{j},a^{\prime}}]\oplus[Q^{\prime}_{x_{i},a},Q^{\prime}_{x_{j},a^{\prime}}]=0$, so $($$\mathbf{QH_{2}}$)) holds.
@@
 We will implicitly use this identification in the following.
-Now, if $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is such that $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for all $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$, then $Q=h_{1}\oplus\dots\oplus h_{d}$ for some $d\geq 1$ and $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$ classical homomorphisms (where $d$ can be taken to be the dimension of $Q$).
+Now, if $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ is such that $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for all $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$, then $Q=h_{1}\oplus\dots\oplus h_{d}$ for some $d\geq 1$ and $h_{1},\dots,h_{d}\colon\mathbb{A}\to\mathbb{B}$ classical homomorphisms (where $d$ can be taken to be the dimension of $Q$).
 This is a simple consequence of the spectral theorem.
@@
 […]
-(1) $Q_{a,b}(x)=\sum_{i=1}^{d}\mathbf{1}[h_{i}(a)=b]\langle x,e_{i}\rangle e_{i}$
-for all $x\in H$, where $\mathbf{1}{[h_{i}(a)=b]}$ is $1$ if $h_{i}(a)=b$ and 0 otherwise.
+$Q_{a,b}(x)=\sum_{i=1}^{d}\mathbf{1}_{[h_{i}(a)=b]}\langle x,e_{i}\rangle e_{i}$ (1)
+for all $x\in H$, where $\mathbf{1}_{[h_{i}(a)=b]}$ is $1$ if $h_{i}(a)=b$ and 0 otherwise.
+Suppose now that $Q$ is a linear operator on $H$ and $Q^{\prime}$ is a linear operator on $H^{\prime}$, and let $Q\otimes Q^{\prime}$ be the linear operator on $H\otimes H^{\prime}$ defined by setting $(Q\otimes Q^{\prime})(v\otimes v^{\prime})=Q(v)\otimes Q^{\prime}(v^{\prime})$ for all $v\in H$ and $v^{\prime}\in H^{\prime}$ and extending by linearity.
+As for the direct sum, we extend this definition to quantum functions $Q\colon X\to_{H}A$ and $Q^{\prime}\colon X\to_{H^{\prime}}B$ by $(Q\otimes Q^{\prime})_{x,(a,b)}:=Q_{x,a}\otimes Q^{\prime}_{x,b}$ for all $a\in A$ and $b\in B$.
+This is a quantum function $Q\otimes Q^{\prime}\colon X\to_{H\otimes H^{\prime}}A\times B$.
+Like for the direct sum, we now show that the tensor product preserves quantum homomorphisms.
 Lemma 19.
@@
 $=Q_{x,a}(v)\otimes Q^{\prime}_{x,b}(v^{\prime})$
-$=R_{x,(a,b)}(v\otimes v^{\prime})$
-so that each $R_{x,(a,b)}$ is a projector. Moreover, if $x\in X$, then
-$\sum_{a,b}R_{x,(a,b)}$ $=\sum_{a,b}Q_{x,a}\otimes Q_{x^{\prime},b}$
-$=\sum_{a}Q_{x,a}\otimes(\sum_{b}Q_{x^{\prime},b})$
-$=\sum_{a}Q_{x,a}\otimes\operatorname{id}_{H^{\prime}}$
-$=(\sum_{a}Q_{x,a})\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}}$
+$=R_{x,(a,b)}(v\otimes v^{\prime}),$
+so each $R_{x,(a,b)}$ is a projector onto $H\otimes H^{\prime}$. Moreover, if $x\in X$, then
+$\sum_{a,b}R_{x,(a,b)}$ $=\sum_{a,b}Q_{x,a}\otimes Q_{x^{\prime},b}=\sum_{a}Q_{x,a}\otimes\sum_{b}Q_{x^{\prime},b}=\operatorname{id}_{H}\otimes\operatorname{id}_{H^{\prime}}=\operatorname{id}_{H\otimes H^{\prime}},$
 so $R$ is a quantum function. The verification that $($$\mathbf{QH_{1}}$)) holds is similar.
@@
 $=(Q_{y,a^{\prime}}Q_{x,a})\otimes(Q^{\prime}_{y,b^{\prime}}Q^{\prime}_{x,b})$
-$=R_{y,(a^{\prime},b^{\prime})}R_{x,(a,b)}$
+$=R_{y,(a^{\prime},b^{\prime})}R_{x,(a,b)},$
 so $($$\mathbf{QH_{2}}$)) holds. ∎
@@
 $R\bullet Q\colon A\to_{H\otimes H^{\prime}}C$
-by $(R\bullet Q)_{a,c}=\sum_{b\in B}R_{b,c}\otimes Q_{a,b}$.
-Note that the sum in the definition is finite (even if $B$ is infinite), since $H$ has finite dimension and therefore only finitely many of the $Q_{a,b}$ are non-zero.
-The fact that the composition of quantum homomorphisms is a quantum homomorphism was proved in [abramsky2017quantum] using that this operation can be expressed as the composition of Kleisli arrows in the Kleisli category for the quantum monad.
-For the convenience of the reader, we give here a self-contained proof following [abramsky2017quantum].
-Lemma 6 ([abramsky2017quantum]).
+by $(R\bullet Q)_{a,c}=\sum_{b\in B}R_{b,c}\otimes Q_{a,b}$ for each $a\in A$, $c\in C$.
+The fact that the composition of quantum homomorphisms is a quantum homomorphism was proved in [1], and it corresponds to the fact that this operation can be expressed as the composition of Kleisli arrows in the Kleisli category for the quantum monad.
+(See also [48] for a recent extension of the Kleisli framework to different quantum relaxations of classical morphisms.) For the convenience of the reader, we give here a self-contained proof following [1].
+Lemma 21 ([1]).
 The composition of quantum homomorphisms $Q\colon\mathbb{A}\to_{H}\mathbb{B}$ and $R\colon\mathbb{B}\to_{H^{\prime}}\mathbb{C}$ is a quantum homomorphism $\mathbb{A}\to_{H\otimes H^{\prime}}\mathbb{C}$.
@@
 Similarly,
-$S_{a,c}^{\top}=(\sum_{b}R_{b,c}\otimes Q_{a,b})^{\top}=\sum_{b}R_{b,c}^{\top}\otimes Q_{a,b}^{\top}=\sum_{b}R_{b,c}\otimes Q_{a,b}=S_{a,c},$
+$S_{a,c}^{*}=(\sum_{b}R_{b,c}\otimes Q_{a,b})^{*}=\sum_{b}R_{b,c}^{*}\otimes Q_{a,b}^{*}=\sum_{b}R_{b,c}\otimes Q_{a,b}=S_{a,c},$
 so $S_{a,c}$ is a projector. We also have
@@
 $=\sum_{b_{1},\dots,b_{r}\in B}(R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}})\otimes(Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}).$
-Note that for every $(b_{1},\dots,b_{r})$, either $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$—in which case $R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}}=0$ since $R$ satisfies $($$\mathbf{QH_{1}}$))—or $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$—and then $Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}=0$ since $Q$ satisfies $($$\mathbf{QH_{1}}$)).
+Note that for every $(b_{1},\dots,b_{r})\in B^{r}$, either $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$—in which case $R_{b_{1},c_{1}}\dots R_{b_{r},c_{r}}=0$ since $R$ satisfies $($$\mathbf{QH_{1}}$))—or $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$—and then $Q_{a_{1},b_{1}}\dots Q_{a_{r},b_{r}}=0$ since $Q$ satisfies $($$\mathbf{QH_{1}}$)).
 Thus, the sum above evaluates to 0 and $($$\mathbf{QH_{1}}$)) holds.
@@
 $=\sum_{b,b^{\prime}}(R_{b^{\prime},c^{\prime}}R_{b,c})\otimes(Q_{a^{\prime},b^{\prime}}Q_{a,b})$ by $($$\mathbf{QH_{2}}$))
-$=S_{a^{\prime},c^{\prime}}S_{a,c}.$
+$=S_{a^{\prime},c^{\prime}}S_{a,c},$
 so $($$\mathbf{QH_{2}}$)) holds. ∎
+We also generalise the composition of quantum functions to functions with arity greater than 1.
+Definition 22.
+Fix positive integers $k_{1},\dots,k_{n}$, and let $W_{1},\dots,W_{n}$ be quantum functions, where $W_{i}\colon A^{k_{i}}\to_{\operatorname{q}}B$ for all $i\in[n]$.
+Take also a quantum function $Q\colon B^{n}\to_{\operatorname{q}}C$.
+We define
+$Q*(W_{1},\dots,W_{n})\colon A^{\sum k_{i}}\to_{\operatorname{q}}C$
+by $(Q*(W_{1},\dots,W_{n}))_{(\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}),c}=\sum_{\mathbf{b}\in B^{n}}Q_{\mathbf{b},c}\otimes\bigotimes_{i=1}^{n}(W_{i})_{\mathbf{a}^{(i)},b_{i}}$.
+Note that this is a well-defined quantum function, since
+$\sum_{c\in C}(Q*(W_{1},\dots,W_{n}))_{(\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}),c}$ $=\sum_{c\in C}\sum_{\mathbf{b}\in B^{n}}Q_{\mathbf{b},c}\otimes\bigotimes_{i=1}^{n}(W_{i})_{\mathbf{a}^{(i)},b_{i}}$
+$=\operatorname{id}\otimes\bigotimes_{i=1}^{n}\sum_{b\in B}(W_{i})_{\mathbf{a}^{(i)},b}=\operatorname{id}.$
+Lemma 23.
+If $W_{1},\dots,W_{n}$ are homomorphisms $\mathbb{A}^{k_{i}}\to_{H_{i}}\mathbb{B}$ and $Q\colon\mathbb{B}^{n}\to_{H}\mathbb{C}$, then $Q*(W_{1},\dots,W_{n})$ is a homomorphism $\mathbb{A}^{\sum k_{i}}\to_{H\otimes\bigotimes H_{i}}\mathbb{C}$.
+Proof.
+Consider the quantum function $W:A^{\sum k_{i}}\to_{\operatorname{q}}B^{n}$ defined as follows.
+For $\mathbf{a}^{(1)}\in A^{k_{1}},\dots,\mathbf{a}^{(n)}\in A^{k_{n}}$ and $\mathbf{b}\in B^{n}$, we set
+$W_{((\mathbf{a}^{(1)},\dots,\mathbf{a}^{(n)}),\mathbf{b})}=\bigotimes_{i\in[n]}(W_{i})_{\mathbf{a}^{(i)},b_{i}}.$
+Reasoning as in the proof of Lemma 19, we readily check that $W$ is a quantum homomorphism $\mathbb{A}^{\sum k_{i}}\to_{\bigotimes H_{i}}\mathbb{B}^{n}$, and that
+$Q*(W_{1},\dots,W_{n})=Q\bullet W.$
+Then, the result follows from Lemma 21. ∎
 We will also make use of the following simple result.
 Lemma 24.
-Let $\mathbb{X}$ be a core and let $Q\colon\mathbb{X}\to_{q}\mathbb{X}$ be a quantum homomorphism. If $Q$ is non-contextual, then
+Let $\mathbb{X}$ be a core and let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{X}$ be a quantum homomorphism. If $Q$ is non-contextual, then
 $\sum_{x\in X}Q_{x,y}=\operatorname{id}$
@@
 It follows that
-$\sum_{x\in X}Q_{x,y}=\bigoplus_{i\in[d]}\sum_{x\in X}(h_{i})_{x,y}=\operatorname{id}.$
-3. Quantum polymorphisms and reductions
-In this section, we show that reductions between entangled CSPs can be obtained via a quantised notion of pp-definitions that we call q-definitions.
-In order to prove that q-definitions exist, we make use of the notion of commutativity gadgets as defined in [Ji].
+$\sum_{x\in X}Q_{x,y}=\bigoplus_{i\in[d]}\sum_{x\in X}(h_{i})_{x,y}=\operatorname{id},$
+as required. ∎
+4 Quantum algebraic reductions
+In this section, we start our analysis of reductions between quantum CSPs by lifting Theorem 2 to the quantum setting via the notion of quantum polymorphisms.
+First of all, in the next subsection, we collect information on the algebraic structure of the quantum polymorphism set of a constraint language.
+4.1 The algebraic structure of quantum polymorphisms
+The set of classical polymorphisms of a relational structure $\mathbb{A}$ can be endowed with the algebraic structure of a function minion over $A$.
+A function minion over some set $S$ is a subset $\mathscr{M}$} of the set of functions $\{f:S^{n}\to S:n\in\mathbb{N}\}$ that is closed under identifying variables, permuting variables, and adding dummy variables.
+This can be expressed compactly as follows: For each $n,m\in\mathbb{N}$, each map $\pi:[n]\to[m]$, and each $f\in{\mathscr{M}}$ of arity $n$ (meaning that $f$ is a function $S^{n}\to S$), the function $f_{/\pi}:S^{m}\to S$ defined by
+$f_{/\pi}(x_{1},\dots,x_{m})=f(x_{\pi(1)},x_{\pi(2)},\dots,x_{\pi(n)})$
+belongs to $\mathscr{M}$}.
+It is clear from the definition of classical polymorphisms that the set $\operatorname{Pol}(\mathbb{A})$ is invariant under this type of operations (known as minors), so $\operatorname{Pol}(\mathbb{A})$ is indeed a function minion over $A$.
+In fact, one can endow $\operatorname{Pol}(\mathbb{A})$ with the richer structure of a (function) clone, due to the fact that polymorphisms are closed under composition (and composition is associative).
+Nevertheless, it was observed in [10] that the complexity of CSPs is entirely determined by the height-one or minor identities holding in $\operatorname{Pol}(\mathbb{A})$; i.e., identities of the form $f=g_{/\pi}$.
+In particular, in order to capture reducibility of CSPs via pp-definitions and pp-constructions, it is sufficient to compare their polymorphism clones in terms of minor-preserving maps; i.e., so-called minion homomorphisms.
+The minion framework proved robust enough to capture reducibility and complexity of variants of the CSP; in particular, the promise CSP [3, 19, 5], whose polymorphisms do not compose and, thus, do not form clones.
+In addition, minions were shown to be well-suited to capture the power of relaxations of (promise) CSPs via multiple algorithmic models [5, 27, 28, 29, 25, 17, 30, 11, 55]; see also [16] for a recent application to CSP sparsification and non-redundancy.
+In these works, the notion of function minions was gradually lifted to the categorical concept of abstract minions, retaining the algebraic structure of function minions without the restriction of being applied to sets of functions; see, in particular, [18, 6].
+Definition 25.
+An abstract minion (or simply a minion) $\mathscr{M}$} is a functor from the skeleton category of non-empty finite sets to the category of non-empty sets.
+[^6] The fact that abstract minions can be described as functors was first observed by Barto; see also [40] for a categorical treatment of the algebraic approach to constraint satisfaction.
+In other words, $\mathscr{M}$} is the disjoint union of non-empty sets ${\mathscr{M}}^{(n)}$ for $n\in\mathbb{N}$, equipped with operations $(\cdot)_{/\pi}\colon{\mathscr{M}}^{(n)}\to{\mathscr{M}}^{(m)}$ (the minor operations) for each $n,m\in\mathbb{N}$ and each map $\pi\colon[n]\to[m]$, such that
+  • $(i)$
+    $M_{/\operatorname{id}}=M$ for each $n\in\mathbb{N}$ and $M\in{\mathscr{M}}^{(n)}$, where $\operatorname{id}$ is the identity map on the set $[n]$;
+  • $(ii)$
+    $(M_{/\pi})_{/\tilde{\pi}}=M_{/\tilde{\pi}\circ\pi}$ for each $n,m,p\in\mathbb{N}$, $M\in{\mathscr{M}}^{(n)}$, $\pi\colon[n]\to[m]$, and $\tilde{\pi}\colon[m]\to[p]$.
+Recall now the notion of quantum polymorphisms (Definition 1).
+Quantum polymorphisms are not functions over a set, but rather quantum functions, i.e., systems of projective measurements.
+Hence, in order to lift the algebraic structure of $\operatorname{Pol}(\mathbb{A})$ to the set $\operatorname{qPol}(\mathbb{A})$ of quantum polymorphisms of $\mathbb{A}$, it is crucial to adopt the formalism of abstract, as opposed to function, minions.
+To prove the following result, we define the minor operations in $\operatorname{qPol}(\mathbb{A})$ in terms of compositions and tensor products of quantum homomorphisms, as described in Section 3.3.
+Proposition 26.
+For each relational structure $\mathbb{A}$, $\operatorname{qPol}(\mathbb{A})$ is an abstract minion.
+Proof.
+First, observe that $\operatorname{qPol}(\mathbb{A})$ is the disjoint union of the non-empty sets $\operatorname{qPol}(\mathbb{A})^{(n)}$ for $n\in\mathbb{N}$, where $\operatorname{qPol}(\mathbb{A})^{(n)}$ is the set of $n$-ary quantum polymorphisms of $\mathbb{A}$; i.e., quantum homomorphisms $\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$.
+For $1\leq i\leq k$, we let $\pi_{i}^{k}$ be the $k$-ary projection $\pi_{i}^{k}\colon\mathbb{A}^{k}\to\mathbb{A}$ defined by $\mathbf{a}\mapsto a_{i}$ for each $\mathbf{a}=(a_{1},\dots,a_{k})\in A^{k}$.
+Note that this is a classical polymorphism and, thus, a quantum polymorphism.
+We define minor operations as follows.
+For $n,m\in\mathbb{N}$, $\tau\colon[n]\to[m]$, and $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$, we set
+$Q_{/\tau}=Q\bullet\bigotimes_{j\in[n]}\pi_{\tau(j)}^{m}.$ (2)
+We need to check that this definition satisfies the two conditions in Definition 25.
+Unfolding the definitions of composition and tensor product of quantum homomorphisms given in Section 3.3, we see that, for each $\mathbf{a}\in A^{m}$ and $b\in A$,
+$(Q_{/\tau})_{\mathbf{a},b}$ $=\left(Q\bullet\bigotimes_{j\in[n]}\pi_{\tau(j)}^{m}\right)_{\mathbf{a},b}=\sum_{\mathbf{c}\in A^{n}}Q_{\mathbf{c},b}\otimes\left(\bigotimes_{j\in[n]}\pi_{\tau(j)}^{m}\right)_{\mathbf{a},\mathbf{c}}=\sum_{\mathbf{c}\in A^{n}}Q_{\mathbf{c},b}\otimes\bigotimes_{j\in[n]}(\pi_{\tau(j)}^{m})_{\mathbf{a},c_{j}}$
+$=Q_{\tau(\mathbf{a}),b}\otimes\operatorname{id}^{\otimes n},$
+where $\tau(\mathbf{a})=(a_{\tau(1)},\dots,a_{\tau(n)})$.
+Therefore, up to the canonical identification of the Hilbert space $H$ associated with $Q$ with $H\otimes\mathbb{K}^{\otimes n}$ (where $\mathbb{K}$ is the base field), we obtain that $(Q_{/\tau})_{\mathbf{a},b}=Q_{\tau(\mathbf{a}),b}$.
+From this, it follows immediately that $Q_{/\operatorname{id}}=Q$ and that $(Q_{/\tau})_{/\tau^{\prime}}=Q_{\tau^{\prime}\circ\tau}$.
+Therefore, the minor operations in eq.
+2 meet the conditions $(i)$ and $(ii)$ of Definition 25, and it follows that $\operatorname{qPol}(\mathbb{A})$ is a well-defined abstract minion.
+∎
+We now introduce a notion that shall be important for the classification of languages admitting commutativity gadgets (see Section 5.3).
+Definition 27.
+Let $\mathbb{A}$ be a relational structure.
+The quantum closure of $\operatorname{Pol}(\mathbb{A})$, denoted by $\operatorname{\oplus Pol}(\mathbb{A})$, is the union of the quantum closures of the set of $n$-ary classical polymorphisms $\operatorname{Hom}(\mathbb{A}^{n},\mathbb{A})$, for all $n\in\mathbb{N}$.
+We observe the following simple fact.
+Proposition 28.
+A quantum polymorphism of a structure $\mathbb{A}$ belongs to $\operatorname{\oplus Pol}(\mathbb{A})$ if, and only if, it is non-contextual.
+Proof.
+This is an immediate consequence of the spectral theorem. ∎
+Since every classical polymorphism is non-contextual, and every non-contextual polymorphism is in particular a quantum polymorphism, we have the following inclusions for each structure $\mathbb{A}$:
+$\operatorname{Pol}(\mathbb{A})\subseteq\operatorname{\oplus Pol}(\mathbb{A})\subseteq\operatorname{qPol}(\mathbb{A}).$ (3)
+Note also that, since the projection maps $\pi_{i}^{k}\colon\mathbb{A}^{k}\to\mathbb{A}$ used in the proof of Proposition 26 are classical and, thus, non-contextual, $\operatorname{\oplus Pol}(\mathbb{A})$ is a minion with the same minor operations as the ones for $\operatorname{qPol}(\mathbb{A})$; i.e., those defined in eq.
+2.
+We next define minion homomorphisms.
+Definition 29 ([18, 6]).
+A minion homomorphism $\xi:{\mathscr{M}}\to{\mathscr{N}}$ between two minions $\mathscr{M}$} and $\mathscr{N}$} is a natural transformation from $\mathscr{M}$} to $\mathscr{N}$}.
+I.e., $\xi$ is a map from the underlying set of $\mathscr{M}$} to the underlying set of $\mathscr{N}$} that
+  • preserves the arity—i.e., $\xi(M)\in{\mathscr{N}}^{(n)}$ if $M\in{\mathscr{M}}^{(n)}$;
+  • preserves the minors—i.e., $\xi(M_{/\pi})=\xi(M)_{/\pi}$ if $M\in{\mathscr{M}}^{(n)}$ and $\pi:[n]\to[m]$.
+We denote the existence of a minion homomorphism from $\mathscr{M}$} to $\mathscr{N}$} by the notation ${\mathscr{M}}\to{\mathscr{N}}$.
+The following result shows that minion homomorphisms between polymorphism minions of CSPs exactly capture reducibility via pp-constructions.
+Theorem 30 ([10]).
+Let $\mathbb{A}$ and $\mathbb{B}$ be relational structures. Then the following are equivalent:
+  • $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$;
+  • $\mathbb{A}$ $p$-constructs $\mathbb{B}$.
+In particular, a minion homomorphism $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$ gives rise to a complexity reduction from $\operatorname{CSP}(\mathbb{B})$ to $\operatorname{CSP}(\mathbb{A})$.
+Since inclusions between minions are in particular homomorphisms, we immediately obtain from (3) that $\operatorname{Pol}(\mathbb{A})\to\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{A})$ for each $\mathbb{A}$.
+In addition, we observe the next fact.
+Proposition 31.
+Let $\mathbb{A},\mathbb{B}$ be relational structures such that $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$. Then,
+$\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{\oplus Pol}(\mathbb{B})$.
+Proof.
+Let $\xi\colon\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$.
+Define $\zeta\colon\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{\oplus Pol}(\mathbb{B})$ as follows.
+For any $Q\in\operatorname{\oplus Pol}(\mathbb{A})$, there are by definition 1-dimensional Hilbert spaces $H_{1},\dots,H_{d}$ and classical polymorphisms $f_{1},\dots,f_{d}\in\operatorname{Pol}(\mathbb{A})$ such that $Q=\oplus_{i=1}^{d}f_{i}$.
+Then define $\zeta(Q)$ as $\oplus_{i=1}^{d}\xi(f_{i})$, in the same basis.
+The verification that $\zeta$ is a minion homomorphism is immediate since $Q_{/\pi}=\oplus_{i=1}^{d}(f_{i})_{/\pi}$.
+∎
+Although this bears no consequences for the present work, we remark here that $\operatorname{qPol}(\mathbb{A})$ can also be endowed with the algebraic structure of an operad.
+The $n$-ary elements of the operad are the $n$-ary quantum polymorphisms of $\mathbb{A}$, the identity is the identity map $\operatorname{id}\colon\mathbb{A}\to\mathbb{A}$, and the composition is given by the * operation of Definition 22.
+Proposition 32.
+For each relational structure $\mathbb{A}$, $\operatorname{qPol}(\mathbb{A})$ is an operad.
+Proof.
+We first need to check that $Q*(\operatorname{id},\dots,\operatorname{id})=Q=\operatorname{id}{}*Q$.
+For the first equality, we get $(Q*(\operatorname{id},\dots,\operatorname{id}))_{\mathbf{a},c}=\sum_{\mathbf{b}\in B^{n}}Q_{\mathbf{b},c}\otimes\operatorname{id}_{a_{1},b_{1}}\otimes\dots\otimes\operatorname{id}_{a_{n},b_{n}}=Q_{\mathbf{a},c}\otimes\operatorname{id}^{\otimes n}=Q_{\mathbf{a},c}$.
+(Note that, as in the proof of Proposition 26, we work here with the identification of $H$ and $H\otimes\mathbb{K}^{\otimes n}$.) The verification for the right-hand side is similar.
+We now prove operad-associativity; i.e., we prove that if $Q$ is $n$-ary, $W_{i}$ is $k_{i}$-ary for $i\in[n]$, and $V_{i,j}$ is $\ell_{i,j}$-ary for $i\in[n]$ and $j\in[k_{i}]$, then
+$(Q*(W_{1},\dots,W_{n}))*(V_{1,1},\dots,V_{1,k_{1}},\dots,V_{n,1},\dots,V_{n,k_{n}})$
+and
+$Q*(W_{1}*(V_{1,1},\dots,V_{1,k_{1}}),\dots,W_{n}*(V_{n,1},\dots,V_{n,k_{n}}))$
+coincide.
+The verification is slightly tedious and we omit it to the benefit of the reader; it suffices to remark that, for both sides, the projectors corresponding to the tuples $\mathbf{a}^{i,j}\in A^{\ell_{i,j}}$ (for $i\in[n]$ and $j\in[k_{i}]$) and $d\in A$, are equal to
+$\sum_{\mathbf{c}\in A^{n},\mathbf{b}^{1}\in A^{k_{1}},\dots,\mathbf{b}^{n}\in A^{k_{n}}}Q_{\mathbf{c},d}\otimes\bigotimes_{i\in[n]}(W_{i})_{\mathbf{b}^{i},c_{i}}\otimes\bigotimes_{i\in[n],j\in[k_{i}]}(V_{i,j})_{\mathbf{a}^{i,j},b^{i}_{j}}.\qed$
+It would seem tempting to believe that $\operatorname{qPol}(\mathbb{A})$, being simultaneously an operad and a minion, is a clone.
+The natural definition of an (abstract) clone from an operad-minion is the following.
+Let $Q$ be $n$-ary, and let $W_{1},\dots,W_{n}$ be $k$-ary.
+Define $\sigma\colon[nk]\to[k]$ by $\sigma(i)=i\bmod k$, where we take here the convention that the $\bmod$ operation takes values in $\{1,\dots,k\}$.
+Then define $Q\circ(W_{1},\dots,W_{n})$ to be the $k$-ary operation $(Q*(W_{1},\dots,W_{n}))_{/\sigma}$.
+While this definition is valid, and gives rise to the quantum function with projectors $\sum_{\mathbf{b}}Q_{\mathbf{b},c}\otimes\bigotimes_{i=1}^{n}(W_{i})_{\mathbf{a},b_{i}}$, it fails to be associative.
+4.2 Complexity reductions via minion homomorphisms
+The goal of this subsection is to prove that the existence of a minion homomorphism between the quantum polymorphism minions of two structures gives rise to a logspace reduction between the corresponding quantum CSPs.
+See 3 To prove the above result, we quantise the completeness and soundness analysis for the reduction coming from the Long-Code testing of classical constraint systems [13], see for example [5, §3.3] and [24].
+We observe that, while in the classical case the completeness of the reduction is readily proved, in the quantum setting both the completeness and the soundness cases are quite technical.
+Proof.
+Let $\sigma$ and $\rho$ be the signatures of $\mathbb{A}$ and $\mathbb{B}$, respectively.
+Without loss of generality, we take the domain of $B$ to be $[n]$.
+Take an instance $\mathbb{X}$ of $\operatorname{CSP_{q}}(\mathbb{B})$ (i.e., $\mathbb{X}$ is an arbitrary $\rho$-structure).
+Consider the $\sigma$-structure $\mathbb{Y}$ defined via the following process.
+  • $(i)$
+    For each $x\in X$, add in $\mathbb{Y}$ a copy of $\mathbb{A}^{n}$. Given a tuple $\mathbf{a}\in A^{n}$, we denote the corresponding variable of $\mathbb{Y}$ by $(x,\mathbf{a})$}.
+  • $(ii)$
+    For each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$, add in $\mathbb{Y}$ a copy of $\mathbb{A}^{|R^{\mathbb{B}}|}$.
+Given a tuple $\mathbf{c}\in A^{|R^{\mathbb{B}}|}$, we denote the corresponding variable of $\mathbb{Y}$ by $(R,\mathbf{x},\mathbf{c})$}.
+  • $(iii)$
+    We identify variables in $\mathbb{Y}$ as follows.
+For each symbol $R\in\rho$ (of arity, say, $r$), each tuple $\mathbf{x}\in R^{\mathbb{X}}$, each index $i\in[r]$, and each tuple $\mathbf{a}\in A^{n}$ we identify the variable $(x_{i},\mathbf{a})$} with the variable $(R,\mathbf{x},\mathbf{c})$}, where $\mathbf{c}\in A^{|R^{\mathbb{B}}|}$ is the tuple obtaining by composing $\mathbf{a}$ with the $i$-th projection $\pi_{i}\colon R^{\mathbb{B}}\to B$ mapping each $(b_{1},\dots,b_{r})\in R^{\mathbb{B}}$ to $b_{i}$.
+In more details, $\mathbf{c}$ is the tuple $(a_{\mathbf{r}_{1}(i)},\dots,a_{\mathbf{r}_{N}(i)})$, where $R^{\mathbb{B}}=\{\mathbf{r}_{1},\dots,\mathbf{r}_{N}\}$ is a fixed enumeration of $R^{\mathbb{B}}$.
+Note that the structure $\mathbb{Y}$ resulting from this process is a $\sigma$-structure and, thus, a well-defined instance of $\operatorname{CSP_{q}}(\mathbb{A})$.
+Given some $y\in Y$, we shall say that $y$ is a variable of type $(i)$ if it comes from step $(i)$, and we shall say that $y$ is a variable of type $(ii)$ if it comes from step $(ii)$; observe that, because of step $(iii)$, $y$ may be of type $(i)$ and type $(ii)$, simultaneously.
+We claim that the reduction $\mathbb{X}\mapsto\mathbb{Y}$ is complete and sound.
+Henceforth in this proof, with a slight abuse of notation, we identify tuples in $S^{|T|}$ with maps $T\to S$.
+(Completeness.) Take a quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{B}$.
+We define a quantum function $W\colon Y\to_{\operatorname{q}}A$ as follows.
+For each $x\in X$, each $\mathbf{a}\in A^{n}$, and each $p\in A$, we set
+$W_{{(x,\mathbf{a})},p}=\sum_{\begin{subarray}{c}b\in B\\ \mathbf{a}(b)=p\end{subarray}}Q_{x,b}.$ (4)
+Moreover, for each $R\in\rho$, each $\mathbf{x}\in R^{\mathbb{X}}$, each $\mathbf{c}\in A^{|R^{\mathbb{B}}|}$, and each $p\in A$, we set
+$W_{{(R,\mathbf{x},\mathbf{c})},p}=\sum_{\begin{subarray}{c}\mathbf{b}\in R^{\mathbb{B}}\\ \mathbf{c}(\mathbf{b})=p\end{subarray}}\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}.$ (5)
+Claim 33.
+$W$ preserves the identification step $(iii)$.
+Proof.
+Suppose that the variable $(R,\mathbf{x},\mathbf{c})$} is identified with the variable $(x_{i},\mathbf{a})$} in $\mathbb{Y}$; i.e., $\mathbf{c}=\mathbf{a}\circ\pi_{i}$. We find
+$W_{{(R,\mathbf{x},\mathbf{c})},p}$ $=\sum_{\begin{subarray}{c}\mathbf{b}\in R^{\mathbb{B}}\\ \mathbf{a}(b_{i})=p\end{subarray}}\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}=\sum_{\begin{subarray}{c}\mathbf{b}\in B^{\operatorname{ar}(R)}\\ \mathbf{a}(b_{i})=p\end{subarray}}\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}$
+$=\left(\sum_{b_{1}\in B}Q_{x_{1},b_{1}}\right)\cdot\left(\sum_{b_{2}\in B}Q_{x_{2},b_{2}}\right)\dots\left(\sum_{b_{i-1}\in B}Q_{x_{i-1},b_{i-1}}\right)\cdot\left(\sum_{\begin{subarray}{c}b_{i}\in B\\ \mathbf{a}(b_{i})=p\end{subarray}}Q_{x_{i},b_{i}}\right)$
+$\cdot\left(\sum_{b_{i+1}\in B}Q_{x_{i+1},b_{i+1}}\right)\dots\left(\sum_{b_{\operatorname{ar}(R)}\in B}Q_{x_{\operatorname{ar}(R)},b_{\operatorname{ar}(R)}}\right)$
+$=\sum_{\begin{subarray}{c}b_{i}\in B\\ \mathbf{a}(b_{i})=p\end{subarray}}Q_{x_{i},b_{i}}=W_{{(x_{i},\mathbf{a})},p}$
+as required. Note that the second equality crucially uses that $Q$ satisfies the constraint $\mathbf{x}$, in the sense of $($$\mathbf{QH_{1}}$)). ∎
+Claim 34.
+$W$ is a quantum map from $Y$ to $A$.
+Proof.
+First note that the summands in eq.
+4 are mutually orthogonal and, thus, their sum is a projector.
+Similarly, each product $\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}$ in eq.
+4 consists of mutually commuting projectors (by $($$\mathbf{QH_{2}}$))) and, thus, the product is a projector.
+Moreover, products associated with distinct $\mathbf{b}\in R^{\mathbb{B}}$ are orthogonal by $($$\mathbf{QH_{1}}$)) and $($$\mathbf{QH_{2}}$)), so the sum is a well-defined projector.
+For variables of type $(i)$, we have
+$\sum_{p\in A}W_{{(x,\mathbf{a})},p}=\sum_{p\in A}\sum_{\begin{subarray}{c}b\in B\\ \mathbf{a}(b)=p\end{subarray}}Q_{x,b}=\sum_{b\in B}Q_{x,b}=\operatorname{id}.$
+For variables of type $(ii)$, we have
+$\sum_{p\in A}W_{{(\mathbf{x},\mathbf{c})},p}=\sum_{p\in A}\sum_{\begin{subarray}{c}\mathbf{b}\in R^{\mathbb{B}}\\ \mathbf{c}(\mathbf{b})=p\end{subarray}}\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}=\sum_{\mathbf{b}\in R^{\mathbb{B}}}\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}=\sum_{\mathbf{b}\in B^{\operatorname{ar}(R)}}\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}=\operatorname{id},$
+where we are again using $($$\mathbf{QH_{1}}$)). ∎
+We now proceed to show that the quantum map $Q\colon Y\to_{\operatorname{q}}A$ is in fact a quantum homomorphism $\mathbb{Y}\to_{\operatorname{q}}\mathbb{A}$.
+To that end, we need to check that $W$ satisfies the conditions $($$\mathbf{QH_{1}}$)) and $($$\mathbf{QH_{2}}$)).
+Claim 35.
+Proof.
+Let $S$ be a symbol in $\sigma$ of some arity $s$. Take a tuple $\mathbf{y}\in S^{\mathbb{Y}}$ and a tuple $\mathbf{a}\in A^{s}\setminus S^{\mathbb{A}}$. We need to prove that
+$\prod_{k\in[s]}W_{y_{k},a_{k}}=0.$ (6)
+We now distinguish between two cases.
+  • Case 1: All variables appearing in $\mathbf{y}$ are of type $(i)$.
+In this case, there exists some $x\in X$ and some tuples $\mathbf{z}^{(1)},\dots,\mathbf{z}^{(s)}\in A^{n}$ such that $y_{k}={(x,\mathbf{z}^{(k)})}$ for each $k\in[s]$, and the tuple $(\mathbf{a}^{(1)},\dots,\mathbf{a}^{(s)})$ belongs to $S^{\mathbb{A}^{n}}$ (where, recall, $S^{\mathbb{A}^{n}}$ is the interpretation of the symbol $S\in\sigma$ in the $n$-wise categorical power $\mathbb{A}^{n}$ of $\mathbb{A}$).
+In other words, for each $\ell\in B$ it holds that $(z^{(1)}_{\ell},\dots,z^{(s)}_{\ell})\in S^{\mathbb{A}}$.
+In this case, we find
+    $\prod_{k\in[s]}W_{y_{k},a_{k}}=\prod_{k\in[s]}W_{{(x,\mathbf{z}^{(k)})},a_{k}}=\prod_{k\in[s]}\sum_{\begin{subarray}{c}b_{k}\in B\\ \mathbf{z}^{(k)}(b_{k})=a_{k}\end{subarray}}Q_{x,b_{k}}=\sum_{\begin{subarray}{c}\mathbf{b}\in B^{s}\\ \mathbf{z}^{(k)}(b_{k})=a_{k}\;\forall k\in[s]\end{subarray}}\prod_{k\in[s]}Q_{x,b_{k}}.$
+    Now, because of the idempotency of projectors and the fact that projectors in a PVM having different outcomes are mutually orthogonal, the above expression trivialises unless all $b_{k}$’s are equal.
+Hence, we find
+    $\prod_{k\in[s]}W_{y_{k},a_{k}}=\sum_{\begin{subarray}{c}b\in B\\ \mathbf{z}^{(k)}(b)=a_{k}\;\forall k\in[s]\end{subarray}}Q_{x,b}.$
+    We now argue that the above sum is zero.
+Indeed, for any $b\in B$, the requirement that $\mathbf{z}^{(k)}(b)=a_{k}$ $\forall k\in[s]$ implies that the tuple $(z^{(1)}_{b},\dots,z^{(s)}_{b})$ does not belong to $S^{\mathbb{A}}$, since we are assuming $\mathbf{a}\not\in S^{\mathbb{A}}$.
+This, however, is a contradiction.
+It follows that eq.
+6 holds in this case.
+  • Case 2: All variables appearing in $\mathbf{y}$ are of type $(ii)$.
+The argument is similar to the previous case.
+There exists some symbol $R\in\rho$, some tuple $\mathbf{x}\in R^{\mathbb{X}}$, and some tuples $\mathbf{z}^{(1)},\dots,\mathbf{z}^{(s)}\in A^{|R^{\mathbb{B}}|}$ such that $y_{k}={(R,\mathbf{x},\mathbf{z}^{(k)})}$ for each $k\in[s]$, and the tuple $(\mathbf{a}^{(1)},\dots,\mathbf{a}^{(s)})$ belongs to $S^{\mathbb{A}^{|R^{\mathbb{B}}|}}$.
+Unfolding the definition of relations in categorical powers of structures, this means that, for each $\mathbf{b}\in R^{\mathbb{B}}$, it holds that $(z^{(1)}_{\mathbf{b}},\dots,z^{(s)}_{\mathbf{b}})\in S^{\mathbb{A}}$.
+Letting $r$ be the arity of $R$, we find
+    $\prod_{k\in[s]}W_{y_{k},a_{k}}$ $=\prod_{k\in[s]}W_{{(R,\mathbf{x},\mathbf{z}^{(k)})},a_{k}}=\prod_{k\in[s]}\sum_{\begin{subarray}{c}\mathbf{b}^{(k)}\in R^{\mathbb{B}}\\ \mathbf{z}^{(k)}(\mathbf{b}^{(k)})=a_{k}\end{subarray}}\prod_{j\in[r]}Q_{x_{j},b^{(k)}_{j}}$
+    $=\sum_{\begin{subarray}{c}(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(k)})\in(R^{\mathbb{B}})^{k}\\ \mathbf{z}^{(k)}(\mathbf{b}^{(k)})=a_{k}\;\forall k\in[s]\end{subarray}}\prod_{k\in[s]}\prod_{j\in[r]}Q_{x_{j},b^{(k)}_{j}}$.
+    We now observe that the order in which the products are performed in the expression above is irrelevant, due to the fact that, by assumption, $\mathbf{x}$ belongs to $R^{\mathbb{X}}$ and $Q$ satisfies $($$\mathbf{QH_{2}}$)).
+In particular, we have
+    $\prod_{k\in[s]}W_{y_{k},a_{k}}$ $=\sum_{\begin{subarray}{c}(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(k)})\in(R^{\mathbb{B}})^{k}\\ \mathbf{z}^{(k)}(\mathbf{b}^{(k)})=a_{k}\;\forall k\in[s]\end{subarray}}\prod_{j\in[r]}\prod_{k\in[s]}Q_{x_{j},b^{(k)}_{j}}$.
+    We now proceed similarly to case 1, by observing that, since each $Q_{x_{j}}$ is a PVM, the product $\prod_{k\in[s]}Q_{x_{j},b^{(k)}_{j}}$ trivialises unless all $b_{j}^{(k)}$’s are the same (for any fixed $j\in[r]$).
+This means that
+    $\prod_{k\in[s]}W_{y_{k},a_{k}}$ $=\sum_{\begin{subarray}{c}\mathbf{b}\in R^{\mathbb{B}}\\ \mathbf{z}^{(k)}(\mathbf{b})=a_{k}\;\forall k\in[s]\end{subarray}}\prod_{j\in[r]}Q_{x_{j},b_{j}}.$
+    For each $\mathbf{b}\in R^{\mathbb{B}}$, we know that the tuple $(z^{(1)}_{\mathbf{b}},\dots,z^{(s)}_{\mathbf{b}})$ must belong to $S^{\mathbb{A}}$.
+Hence, such tuple cannot equal $\mathbf{a}$, which is assumed to be in $A^{s}\setminus S^{\mathbb{A}}$.
+It follows that the expression above is zero, thus proving eq.
+6 in this case, too.
+This concludes the proof of the claim.∎
+Claim 36.
+Proof.
+Let $y,y^{\prime}\in Y$ be such that $\{y,y^{\prime}\}\in E(\operatorname{Gaif}(\mathbb{Y}))$.
+This means that there exists some symbol $S\in\sigma$, some tuple $\mathbf{y}\in S^{\mathbb{Y}}$, and two distinct indices $i,j\in[\operatorname{ar}(S)]$ such that $y=y_{i}$ and $y^{\prime}=y_{j}$.
+If all entries of $\mathbf{y}$ are variables of type $(i)$, we see from eq.
+4 that the PVMs $W_{y}$ and $W_{y^{\prime}}$ are both linear combinations of projectors in the PVM $Q_{x}$ for some common $x\in X$, and thus they commute.
+If all entries of $\mathbf{y}$ are variables of type $(ii)$, we see from eq.
+5 that both $W_{y}$ and $W_{y^{\prime}}$ are linear combinations of projectors in the PVM $Q_{\mathbf{x}}$ with outcome set $B^{\operatorname{ar}(R)}$ for some $R\in\rho$ and some $\mathbf{x}\in R^{\mathbb{X}}$, where $Q_{\mathbf{x}}$ is defined by
+$Q_{\mathbf{x},\mathbf{b}}=\prod_{j\in[\operatorname{ar}(R)]}Q_{x_{j},b_{j}}$
+for each $\mathbf{b}\in B^{\operatorname{ar}(R)}$. Hence, $W_{y}$ and $W_{y^{\prime}}$ commute in this case, too. ∎
+In summary, we have shown that, whenever $\mathbb{X}\to_{\operatorname{q}}\mathbb{B}$, it holds that $\mathbb{Y}\to_{\operatorname{q}}\mathbb{A}$, which proves completeness of the reduction.
+(Soundness.) Let $W\colon\mathbb{Y}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism.
+For $x\in X$, let $W^{(x)}$ be the quantum polymorphism of $\mathbb{A}$ defined by $W^{(x)}_{\mathbf{a},b}:=W_{{(x,\mathbf{a})},b}$.
+Similarly, for $R\in\rho$ and $\mathbf{x}\in R^{\mathbb{X}}$, let $W^{(R,\mathbf{x})}_{\mathbf{c},b}:=W_{{(R,\mathbf{x},\mathbf{c})},b}$.
+Recall that for $R\in\rho$, we have fixed an enumeration of $R^{\mathbb{B}}$ as $\{\mathbf{r}_{1},\dots,\mathbf{r}_{N}\}$.
+Note that if $R$ is $r$-ary and $\pi_{i}\colon[|R^{\mathbb{B}}|]\to[n]$ is the map $j\mapsto r_{j}(i)$ for $i\in[r]$, then $W^{(x_{i})}=W^{(R,\mathbf{x})}_{/\pi_{i}}$.
+Indeed, this follows from the equalities
+$(W^{(R,\mathbf{x})}_{/\pi_{i}})_{\mathbf{a},b}=W_{{(R,\mathbf{x},(a_{\mathbf{r}_{1}(i)},\dots,a_{\mathbf{r}_{N}(i)}))},b}=W_{(\mathbf{x},\mathbf{a}),b}=W^{(x)}_{\mathbf{a},b}$,
+where the second equality follows from the fact that the variables $(R,\mathbf{x},a_{\mathbf{r}_{1}(i),\dots,\mathbf{r}_{N}(i)})$ and $(x_{i},\mathbf{a})$ of $\mathbb{Y}$ have been identified.
+Define $Q\colon X\to_{\operatorname{q}}B$ to be the quantum function $Q_{x,b}:=\xi(W^{(x)})_{(1,\dots,n),b}$, where $\xi$ is a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$ whose existence we are assuming.
+We now check that $Q$ is a quantum homomorphism $\mathbb{C}\to_{\operatorname{q}}\mathbb{B}$.
+First, let $(x_{1},\dots,x_{r})\in R^{\mathbb{X}}$ and $(b_{1},\dots,b_{r})\not\in R^{\mathbb{B}}$. We compute
+$Q_{x_{1},b_{1}}\cdots Q_{x_{r},b_{r}}$ $=\prod_{i=1}^{r}\xi(W^{(x_{i})})_{(1,\dots,n),b_{i}}=\prod_{i=1}^{r}\left(\xi(W^{(R,\mathbf{x})})_{/\pi_{i}}\right)_{(1,\dots,n),b_{i}}$
+$=\prod_{i=1}^{r}\xi(W^{(R,\mathbf{x})})_{(\mathbf{r}_{1}(i),\dots,\mathbf{r}_{N}(i)),b_{i}}=0,$
+where the last equality follows from the fact that $\xi(W^{(R,\mathbf{x})})$ is a quantum polymorphism of $\mathbb{B}$, and that the $r$-tuple whose entries are $(\mathbf{r}_{1}(i),\dots,\mathbf{r}_{N}(i))$ for $i\in[r]$ is in $R^{\mathbb{B}^{N}}$, while $\mathbf{b}\not\in R^{\mathbb{B}}$ by assumption.
+Thus, $Q$ satisfies $($$\mathbf{QH_{1}}$)).
+To check $($$\mathbf{QH_{2}}$)), let $(x_{1},\dots,x_{r})\in R^{\mathbb{B}}$ for some $R\in\rho$ and let $i,j\in[r]$. Then for all $b,b^{\prime}\in B$, we have that
+$[Q_{x_{i},b},Q_{x_{j},b}]$ $=\left[\left(\xi(W^{(R,\mathbf{x})})_{/\pi_{i}}\right)_{(1,\dots,n),b},\left(\xi(W^{(R,\mathbf{x})})_{/\pi_{j}}\right)_{(1,\dots,n),b^{\prime}}\right]$
+$=\left[\xi(W^{(R,\mathbf{x})})_{(\mathbf{r}_{1}(i),\dots,\mathbf{r}_{N}(i)),b},\xi(W^{(R,\mathbf{x})})_{(r_{1}(j),\dots,r_{N}(j)),b^{\prime}}\right]=0,$
+where the last equality is because $\xi(W^{(R,\mathbf{x})})$ satisfies $($$\mathbf{QH_{2}}$)). ∎
+Remark 37.
+For the readers familiar with the complexity theory of promise CSPs (in particular, with the reduction in [5]), we remark that the reduction in Theorem 3 can be cast as a reduction from $\operatorname{CSP_{q}}(\mathbb{B})$ to the “quantum minor-condition problem” $\textsc{qMC}(\operatorname{qPol}(\mathbb{B}),N)$, where $N$ is a constant that depends on $\mathbb{A}$.
+If there is a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$, $\textsc{qMC}(\operatorname{qPol}(\mathbb{B}),N)$ trivially reduces to $\textsc{qMC}(\operatorname{qPol}(\mathbb{A}),N)$(where the reduction does nothing).
+Finally, $\textsc{qMC}(\operatorname{qPol}(\mathbb{A}),N)$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ by replacing every symbol $f$ of arity $k\leq N$ by a copy of $\mathbb{A}^{k}$ (steps (i) and (ii) of the reduction above), and identifying vertices according to the minor identities given in the instance (step (iii) of the reduction above).
+Remark 38.
+We point out that for the reduction in Theorem 3 to be implementable in logspace (and, similarly, for all other logspace reductions we give in this work), we implicitly assume that the quantum CSP instances are given via a non-succinct presentation (following the terminology, for example, of [37], see also [32, §1.1]).
+The fact that the identification of vertices in step (iii) can be achieved in logarithmic space is due to [65].
+We remark that in our applications, these reductions are used solely to transfer undecidability, for which it suffices that they be computable.
+Consequently, the distinction between succinct and non-succinct presentations is immaterial for our purposes.
+5 Quantum relational constructions
+In this section, we introduce quantised notions of pp-definitions and pp-constructions that we call q-definitions and q-constructions, respectively, and we link them to the quantum polymorphisms.
+Moreover, we show how these quantum constructions can be systematically built by enriching their classical counterparts with the commutativity gadgets introduced in [47].
 Then, we prove that the existence of commutativity gadgets is entirely determined by a closure property of the quantum polymorphisms of the given language.
-We begin by defining q-definitions.
+5.1 q-definitions and q-constructions
+Recall the notion of pp-definitions given in Section 3.1.
+We now lift it to a quantum version.
+Recall from Section 3.1 that, for a finite set $A$ and a subset $S$ of $A^{r}$, by the notation $(A;S)$ we mean the relational structure having domain $A$ and a single, $r$-ary relation $S$.
 Definition 39.
 Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$.
-Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$.
+Let $R$ be the $r$-ary symbol of the signature of $(A;S)$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$.
 A q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
   • $(\mathbf{q_{1}})$
-    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
+    For every quantum homomorphism $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;S)$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
   • $(\mathbf{q_{2}})$
-    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$.
-We can then define a notion of q-definability, as follows.
+    For every quantum homomorphism $Q\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;S)$.
+Thus, a q-definition is similar to a classical gadget where we impose (in ($(\mathbf{q_{2}})$)) the additional property that quantum homomorphisms $Q\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ must be non-contextual over the special vertices $\{g_{1},\dots,g_{r}\}$ of the gadget.
+In particular, in the case that $r=1$, then q-definitions and pp-definitions coincide.
+Just as in the classical setting, we have a correspondence between the set of relations with a q-definition over a structure $\mathbb{A}$ and the set of relations $R$ such that $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(A;R)$.
+Theorem 40.
+Let $\mathbb{A}$ be a relational structure and let $R\subseteq A^{r}$.
+Then $R$ has a q-definition over $\mathbb{A}$ if, and only if, $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(A;R)$.
+Proof.
+Suppose first that $R$ has a q-definition over $\mathbb{A}$.
+Let $(\mathbb{G},g_{1},\dots,g_{r})$ be a q-definition, and pick a quantum polymorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ of some arity $n$.
+We need to show that $Q$ is an $n$-ary quantum polymorphism of $(A;R)$.
+We first show $($$\mathbf{QH_{1}}$)).
+Let $\mathbf{a}^{1},\dots,\mathbf{a}^{n}\in R$ and $\mathbf{b}\not\in R$.
+For every $i\in[n]$, the classical map $h_{i}\colon[r]\to A$ defined by $j\mapsto a^{i}_{j}$ is a homomorphism $\mathbb{R}\to(A;R)$.
+By ($(\mathbf{q_{1}})$), there exists for all $i\in[n]$ a homomorphism $H_{i}\colon\mathbb{G}\to\mathbb{A}$ such that $H_{i}(g_{j})=h_{i}(j)$ for all $j\in[r]$.
+Define a quantum function $U\colon[r]\to_{\operatorname{q}}A$ by $U_{j,b}:=(Q\bullet(H_{1},\dots,H_{n}))_{g_{j},b}$.
+By ($(\mathbf{q_{2}})$), this is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;R)$ and therefore $\prod_{j=1}^{r}U_{j,b_{j}}=0$.
+But since $U_{j,b_{j}}=Q_{(a^{1}_{j},\dots,a^{n}_{j}),b_{j}}\otimes\operatorname{id}^{\otimes n}$, we obtain that $\prod_{j=1}^{r}Q_{(a^{1}_{j},\dots,a^{n}_{j}),b_{j}}=0$ and therefore $Q$ satisfies $($$\mathbf{QH_{1}}$)).
+We now check $($$\mathbf{QH_{2}}$)).
+Let $\mathbf{a}^{1},\dots,\mathbf{a}^{n}\in R$, and let $i,j\in[r]$.
+Let $a,b\in A$.
+Using the same notation as above, we have $Q_{(a^{1}_{i},\dots,a^{n}_{i}),a},Q_{(a^{1}_{j},\dots,a^{n}_{j}),b}]\otimes\operatorname{id}^{\otimes n}=[U_{i,a},U_{j,b}]=0$.
+We now prove the converse implication.
+Suppose that $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(A;R)$.
+Let $\mathbf{a}^{1},\dots,\mathbf{a}^{n}$ be an enumeration of all the tuples in $R$.
+We define the q-definition $\mathbb{G}$ to be the structure $\mathbb{A}^{n}$, and its $r$ distinguished vertices to be $(a^{1}_{i},\dots,a^{n}_{i})$ for $i\in[r]$.
+We prove that ($(\mathbf{q_{2}})$) holds.
+Let $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ and define $Q^{\prime}_{i,b}=Q_{(r^{1}_{i},\dots,r^{n}_{i}),b}$, which is a quantum function $r]\to A$.
+Let $(b_{1},\dots,b_{r})\not\in R$.
+Then $\prod_{i=1}^{r}Q^{\prime}_{i,b_{i}}=\prod_{i=1}^{r}Q_{(r^{1}_{i},\dots,r^{n}_{i}),b_{i}}=0$ since $Q\in\operatorname{qPol}(A;R)$ by assumption and by $($$\mathbf{QH_{1}}$)).
+Thus, $($$\mathbf{QH_{1}}$)) holds for $Q^{\prime}$.
+Similarly, if $i,j\in[r]$, $a,b\in A$, then $Q^{\prime}_{i,a},Q^{\prime}_{j,b}]=[Q_{(a^{1}_{i},\dots,a^{n}_{i}),a},Q_{(a^{1}_{j},\dots,a^{n}_{j}),b}]=0$, since $($$\mathbf{QH_{2}}$)) holds for $Q$ as a quantum polymorphism of $(A;R)$.
+Finally, we check that ($(\mathbf{q_{1}})$) holds.
+Let $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;R)$ be a quantum homomorphism.
+Since $\mathbb{R}$ consists of a single tuple and $Q$ satisfies $($$\mathbf{QH_{2}}$)), one can write $Q$ as $\oplus_{i=1}^{d}h_{i}$, where each $h_{i}$ is a homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;R)$.
+For each $i$, there exists $i(j)\in[n]$ such that $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{i(j)}$.
+Define $g_{i}\colon\mathbb{A}^{n}\to\mathbb{A}$ to be the $n$-ary projection $\pi_{i(j)}$, and define $Q^{\prime}\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ to be $\oplus g_{i}$, concluding the proof of ($(\mathbf{q_{1}})$).
+∎
 Definition 41.
 Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures.
-We say that $\mathbb{B}$ is q-definable from $\mathbb{A}$ if there exists a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
-Observe, in ($(\mathbf{q_{1}})$), that a quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ is necessarily non-contextual, since all the vertices of $\mathbb{R}$ lie in a common constraint.
-(This will not be true anymore in the non-oracular setting, see Section 7.) Classical gadgets for CSPs that can be defined one from the other via pp-definitions provide logspace reductions between CSPs of different languages.
-The next result shows that the same is true for $\operatorname{CSP^{*}}$, provided that pp-definitions are replaced by the stronger q-definitions.
+We say that $\mathbb{A}$ q-defines $\mathbb{B}$ if there exists a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
+As an immediate corollary of Theorem 40, we obtain the following.
+See 5
+Proof.
+It is enough to apply Theorem 40 to each of the relations of $\mathbb{B}$. ∎
+Classical gadgets for CSPs that can be defined one from the other via pp-definitions provide logspace reductions between CSPs of different languages.
+It follows from the results above that the same holds for q-definitions.
 Proposition 42.
-Let $\mathbb{A}$ and $\mathbb{B}$ be structures such that $\mathbb{B}$ is q-definable from $\mathbb{A}$.
-Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{*}}(\mathbb{A})$ in logspace.
-We prove Proposition 10 later in this section.
-For now, we focus on the following problem: How to construct q-definitions guaranteeing that a given $\operatorname{CSP^{*}}$ can be reduced to a different one?
-It is not hard to show that every q-definition yields a classical pp-definition.
-Indeed, considering ($(\mathbf{q_{1}})$) for quantum homomorphisms with a Hilbert space of dimension 1 (which corresponds to classical homomorphisms), we find that the tuples in $R$ are in one-to-one correspondence with the restrictions of homomorphisms $\mathbb{G}\to\mathbb{A}$ to $\{g_{1},\dots,g_{r}\}$.
-However, not every pp-definition is a q-definition.
-It turns out that, in order for a pp-definition to ensure the existence of a q-definition, we need to pair it with an extra construction making sure that the process of replacing constraints with gadgets preserves the existence of quantum homomorphisms.
-This is the notion of commutativity gadgets we define below, generalising [Ji].
-Definition 11.
-Fix a $\sigma$-structure $\mathbb{A}$. A commutativity gadget for $\mathbb{A}$ is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
-  • $(\mathbf{c_{1}})$
-    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ extending $Q$.
-  • $(\mathbf{c_{2}})$
-    For every quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
-Note that a commutativity gadget as in Definition 11 is not the same thing as a q-definition for $\mathbb{A}$ and $R=A^{2}$ in the sense of Definition 8.
-Remark 12.
-In [Zeman], an algebraic commutativity gadget is defined by replacing ($(\mathbf{c_{1}})$) by the following condition:
-  • $(\mathbf{c_{1}^{\prime}})$
-    For every $a,b\in A$, there exists a quantum homomorphism $Q\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
-Note that this is a special case of $(\mathbf{c_{1}})$ and, thus, any commutativity gadget is an algebraic commutativity gadget.
-While this appears to be a strictly weaker requirement than the one in Definition 11, we shall prove below that the existence of an algebraic commutativity gadget in the sense of [Zeman] is equivalent to the existence of a commutativity gadget in our sense.
-If $\mathbb{A}$ has a commutativity gadget, then we have a natural way of obtaining a sound reduction from $\operatorname{CSP^{*}}(\mathbb{B})$ to $\operatorname{CSP^{*}}(\mathbb{A})$ out of a classical gadget: For every pair of variables $x,y$ in a relation in an instance of $\operatorname{CSP^{*}}(\mathbb{B})$, glue to $x,y$ a copy of the commutativity gadget, identifying $x$ with $u$ and $y$ with $v$.
-The second condition in Definition 11 forces the associated projectors to commute in every quantum homomorphism to $\mathbb{A}$, and the first condition tells us that no additional constraint is forced upon $x,y$ by gluing the commutativity gadget.
-Formally, we prove the following result.
+Let $\mathbb{A}$ and $\mathbb{B}$ be structures such that $\mathbb{A}$ q-defines $\mathbb{B}$. Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace.
+Proof.
+Applying Theorem 5, we know that $\operatorname{qPol}(\mathbb{A})\subseteq\operatorname{qPol}(\mathbb{B})$.
+In particular, the inclusion map gives a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$.
+Then, the result follows from Theorem 3.
+∎
+Next, we show that q-definitions are compatible with quantum homomorphisms, in the following sense.
 Proposition 43.
-Let $\mathbb{A},\mathbb{B}$ be structures such that $(i)$ $\mathbb{A}$ has a commutativity gadget, and $(ii)$ $\mathbb{B}$ is pp-definable from $\mathbb{A}$.
-Then $\mathbb{B}$ is q-definable from $\mathbb{A}$.
-As a consequence of what we have discussed so far, the crucial step in order to quantise classical reductions from CSP theory is to build commutativity gadgets.
-The main result of this section classifies the languages for which commutativity gadgets exist, in terms of a property of their quantum polymorphisms, which we now define.
-Definition 14.
-A quantum polymorphism of arity $n\in\mathbb{N}$ for a structure $\mathbb{A}$ is a quantum homomorphism $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$.
-We denote by $\operatorname{qPol}(\mathbb{A})$ the set of quantum polymorphisms of $\mathbb{A}$.
-Moreover, we define the quantum closure of $\operatorname{Pol}(\mathbb{A})$, denoted by $\operatorname{\oplus Pol}(\mathbb{A})$, as the union of the quantum closures of the set of $n$-ary polymorphisms $\operatorname{Hom}(\mathbb{A}^{n},\mathbb{A})$, for all $n\in\mathbb{N}$.
-Remark 15.
-Observe that $\operatorname{qPol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are both abstract minions (see the definition in [BBKO21, ciardo_quantum_minion]).
-Furthermore, it is not hard to check that $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{\oplus Pol}(\mathbb{A})$ are homomorphically equivalent, and $\operatorname{\oplus Pol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{A})$.
-In particular, it follows that, if $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$, then we have a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{A})$.
-The main result of this section is the following.
-Theorem 16.
-Let $\mathbb{A}$ be a finite structure. The following are equivalent:
-  1. (1)
-    $\mathbb{A}$ has a commutativity gadget.
-  2. (2)
-    $\mathbb{A}$ has an algebraic commutativity gadget.
-  3. (3)
-    $\mathbb{A}^{n}$ is an (algebraic) commutativity gadget for all large enough $n\in\mathbb{N}$.
-  4. (4)
-    $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
-In the rest of this section, we provide the proofs of the results we stated.
-Proof of Proposition 10.
-For every symbol $R\in\rho$ of some arity $r$, let $\mathbb{G}_{R}$ be a q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ (with its distinguished vertices $g_{1},\dots,g_{r}$).
-Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$.
-We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{*}}(\mathbb{A})$ as follows.
-Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$.
-For $R\in\rho$ of some arity $r$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$, and we glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$.
-Call this instance $\mathbb{Y}$.
-(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism.
-For every symbol $R\in\rho$ of some arity $r$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$.
-Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8.
-Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to_{q}\mathbb{A}^{\prime}$.
-Then, ($(\mathbf{q_{1}})$) guarantees the existence of a quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$ such that
-$Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
-for each $i\in[r]$ and each $a\in A$. Overall, this is a quantum homomorphism $\mathbb{Y}\to_{q}\mathbb{A}$.
-(Soundness.) Suppose that $Q\colon\mathbb{Y}\to_{q}\mathbb{A}$ is a quantum homomorphism.
-Consider the quantum function $Q^{\prime}\colon X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$.
-We claim that $Q^{\prime}$ is a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$.
-To that end, take a symbol $R\in\rho$ of some arity $r$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$.
-Observe that the restriction of $Q$ to $\mathbb{G}_{R}^{(\mathbf{x})}$ is a quantum homomorphism $\mathbb{G}_{R}^{(\mathbf{x})}\to_{q}\mathbb{A}$.
-By ($(\mathbf{q_{2}})$), this means that the assignment $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$).
-In particular, recalling that $x_{i}$ and $g_{i}^{(\mathbf{x})}$ are identified in $\mathbb{Y}$, it follows that the PVMs $Q^{\prime}_{x_{i}}=Q_{g_{i}^{(\mathbf{x})}}$ and $Q^{\prime}_{x_{j}}=Q_{g_{j}^{(\mathbf{x})}}$ commute for each choice of $i,j\in[r]$.
-Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)).
-Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$.
-Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
-(2) $\prod_{i\in[r]}Q^{\prime}_{x_{i},b_{i}}=0,$
-which means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. ∎
-Proof of Proposition 13.
-Let $S=R^{\mathbb{B}}$ be a relation of $\mathbb{B}$ of some arity $r$, and let $\mathbb{G}$ be a classical gadget for $R^{\mathbb{B}}$ in $\mathbb{A}$, with distinguished vertices $g_{1},\dots,g_{r}$.
-Let $\mathbb{H}$ be a commutativity gadget for $\mathbb{A}$, with distinguished vertices $u,v$.
-We let $\mathbb{G}^{\prime}$ be the structure that initially consists of the disjoint union of $\mathbb{G}$ and a copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$ for each pair of distinct vertices $x,y$ in $G$, and in which we identify each such pair $x,y$ with the vertices $u,v$ from the copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$.
-We show that $\mathbb{G}^{\prime}$ (with distinguished vertices $g_{1},\dots,g_{r}$) is a q-definition for $S$ in $\mathbb{A}$.
-Let $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$ be a quantum homomorphism.
-In particular, $Q$ is non-contextual and can therefore be written as $h_{1}\oplus\dots\oplus h_{d}$ for some homomorphisms $h_{k}\colon\mathbb{R}\to\mathbb{A}_{S}$.
-Since $\mathbb{G}$ is a gadget for $S$ in $\mathbb{A}$, it follows that there exist homomorphisms $f_{k}\colon\mathbb{G}\to\mathbb{A}$ such that $f_{k}(g_{i})=h_{k}(i)$ for all $i\in[r]$.
-Then $Q^{\prime}=\bigoplus_{k\in[d]}f_{k}$ is a non-contextual quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$.
-In particular, $Q^{\prime}_{x}$ and $Q^{\prime}_{y}$ commute for all $x,y\in G$, and it follows from ($(\mathbf{c_{1}})$) that for all $x,y$ there exists a quantum homomorphism $Q^{(x,y)}\colon\mathbb{H}\to_{q}\mathbb{A}$ such that $Q^{(x,y)}_{u}=Q^{\prime}_{x}$ and $Q^{(x,y)}_{v}=Q^{\prime}_{y}$.
-We thus obtained a quantum homomorphism $\mathbb{G}^{\prime}\to_{q}\mathbb{A}$, which proves ($(\mathbf{q_{1}})$).
-Let now $Q\colon\mathbb{G}^{\prime}\to_{q}\mathbb{A}$ be a quantum homomorphism.
-For each $x,y\in G$, we have by ($(\mathbf{c_{2}})$) that $Q_{x}$ and $Q_{y}$ commute.
-Thus, $Q$ restricts to a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ that is non-contextual and can thus be written as $h_{1}\oplus\dots\oplus h_{d}$, each $h_{k}$ being a classical homomorphism $\mathbb{G}\to\mathbb{A}$.
-Since $\mathbb{G}$ is a classical gadget for $R$, we obtain that $(h_{k}(g_{1}),\dots,h_{k}(g_{r}))\in R^{\mathbb{B}}$ for all $k\in[d]$.
-Thus, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}=\bigoplus_{k}(h_{k})_{g_{i},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{q}\mathbb{A}_{S}$.
-∎
-Next, we show that q-definitions are compatible with quantum homomorphisms.
-Proposition 17.
-Let $\mathbb{A},\mathbb{B}$ be structures, $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively.
-Then every quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$ is a quantum homomorphism $(A;R_{A})\to_{q}(B;R_{B})$.
+Let $\mathbb{A},\mathbb{B}$ be structures, and let $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively.
+Then every quantum homomorphism $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ is a quantum homomorphism $(A;R_{A})\to_{\operatorname{q}}(B;R_{B})$.
 Proof.
@@
 Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$.
-By ($(\mathbf{q_{1}})$), there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in\{1,\dots,r\}$.
-Let $S=Q\bullet Q^{\prime}$, which is a quantum homomorphism $\mathbb{G}\to_{q}\mathbb{B}$.
-By ($(\mathbf{q_{2}})$), we have that the quantum function defined by $Q^{\prime\prime}_{i,b_{i}}:=S_{g_{i},b_{i}}$ for all $i\in\{1,\dots,r\}$ is a quantum homomorphism $Q^{\prime\prime}\colon\mathbb{R}\to(B,R_{B})$.
+By ($(\mathbf{q_{1}})$), there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in[r]$.
+Let $S=Q\bullet Q^{\prime}$, which is a quantum homomorphism $\mathbb{G}\to_{\operatorname{q}}\mathbb{B}$.
+By ($(\mathbf{q_{2}})$), we have that the quantum function $Q^{\prime\prime}$ defined by $Q^{\prime\prime}_{i,b}:=S_{g_{i},b}$ for all $i\in[r]$ and all $b\in B$ is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(B;R_{B})$.
 Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$.
@@
 Let $a_{1},a_{2}$ appear together in a tuple in $R_{A}$ at positions $i$ and $j$ and let $b_{1},b_{2}\in B$ be arbitrary.
-Let $Q^{\prime}\colon\mathbb{G}\to_{q}\mathbb{A}$ be such that $Q^{\prime}_{g_{i},a_{1}}=Q^{\prime}_{g_{j},a_{2}}=\operatorname{id}$, whose existence is asserted by ($(\mathbf{q_{1}})$).
-Again, let $S\colon\mathbb{G}\to_{q}\mathbb{B}$ be the composition $Q\bullet Q^{\prime}$ and $Q^{\prime\prime}\colon\mathbb{R}\to_{q}(B,R_{B})$ be the quantum homomorphism given by ($(\mathbf{q_{2}})$).
-We get $0=[R_{i,b_{1}},R_{j,b_{2}}]=[S_{g_{i},b_{1}},S_{g_{j},b_{2}}]=[Q_{a_{1},b_{1}}\otimes\operatorname{id},Q_{a_{2},b_{2}}\otimes\operatorname{id}]=[Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]\otimes\operatorname{id}$ and therefore $Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]=0$.
+Let $Q^{\prime}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ be such that $Q^{\prime}_{g_{i},a_{1}}=Q^{\prime}_{g_{j},a_{2}}=\operatorname{id}$, whose existence is asserted by ($(\mathbf{q_{1}})$).
+Again, let $S\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{B}$ be the composition $Q\bullet Q^{\prime}$ and $Q^{\prime\prime}\colon\mathbb{R}\to_{\operatorname{q}}(B;R_{B})$ be the quantum homomorphism given by ($(\mathbf{q_{2}})$).
+We get $0=[Q^{\prime\prime}_{i,b_{1}},Q^{\prime\prime}_{j,b_{2}}]=[S_{g_{i},b_{1}},S_{g_{j},b_{2}}]=[Q_{a_{1},b_{1}}\otimes\operatorname{id},Q_{a_{2},b_{2}}\otimes\operatorname{id}]=[Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]\otimes\operatorname{id}$ and therefore $Q_{a_{1},b_{1}},Q_{a_{2},b_{2}}]=0$.
 ∎
-In the rest of this section, we give a proof of Theorem 16. We start by observing the following fact.
+The next step is to make q-definitions more expressive by allowing them to act on powers of the initial structure.
+This results in the notion of q-constructions, which is the quantum counterpart of classical pp-constructions.
+Definition 44.
+Let $\mathbb{A}$ be a $\sigma$-structure and let $\mathbb{B}$ be a $\rho$-structure. We say that $\mathbb{A}$ q-constructs $\mathbb{B}$ if there exists a $\rho$-structure $\mathbb{C}$ such that
+  • $(i)$
+    the domain of $\mathbb{C}$ is $A^{d}$ for some $d\in\mathbb{N}$;
+  • $(ii)$
+    each relation of $\mathbb{C}$ (of arity, say, $r$) is q-definable in $\mathbb{A}$ (identifying $C^{r}=(A^{d})^{r}$ with $A^{dr}$);
+  • $(iii)$
+    $\mathbb{B}$ and $\mathbb{C}$ are quantum-homomorphically equivalent.
+We now show that q-constructions give rise to homomorphisms between quantum polymorphism minions—and thus, via Theorem 3, to reductions between qCSPs. Namely, we prove the following.
+See 4
+It turns out that the key to proving this is to show that the “vectorisation step” in the second condition of Definition 44 yields a minion homomorphism between the corresponding quantum polymorphism minions (and, thus, a reduction between the corresponding quantum CSPs).
+While for classical CSPs this fact is essentially obvious, in the case of $\operatorname{CSP_{q}}$ one might expect that an additional commutativity gadget is needed to quantise this part of the classical reduction.
+We shall prove that, in fact, no commutativity gadget is needed.
+We start by proving a result about abstract minion homomorphisms.
+Given a minion $\mathscr{M}$} and an $n$-ary element $Q\in{\mathscr{M}}^{(n)}$, we say that a coordinate $j\in[n]$ is inessential for $Q$ if there exists an integer $m\in\mathbb{N}$, a map $\sigma\colon[m]\to[n]$, and an element $Q^{\prime}\in{\mathscr{M}}^{(m)}$ such that
+  • $Q=Q^{\prime}_{/\sigma}$, and
+  • $j$ is not in the image of $\sigma$.
+(cfr.
+e.g.
+[28, Remark 41]).
+A coordinate is essential for $Q$ otherwise.
+An essential part of $Q$ is a minor $Q_{/\tau}$ obtained via any function $\tau\colon[n]\to[n^{\prime}]$ that is a bijection when restricted to the essential coordinates of $Q$, and such that all coordinates of $Q_{/\tau}$ are essential.
+This implies in particular that $\tau$ must be surjective.
+Suppose that $Q_{/\tau}$ is an essential part of $Q$, and take a map $\tau^{\prime}\colon[n^{\prime}]\to[n]$ such that $j=(\tau^{\prime}\circ\tau)(j)$ holds for every essential coordinate $j$ of $Q$.
+One readily checks that, in this case, it must hold that $Q=(Q_{/\tau})_{/\tau^{\prime}}$.
+Such a $\tau^{\prime}$ exists given that $\tau$ is bijective when restricted to the essential coordinates of $Q$.
+Lemma 45.
+Let ${\mathscr{M}},{\mathscr{N}}$ be minions and let $\xi\colon{\mathscr{M}}\to{\mathscr{N}}$ be a partial minion homomorphism defined on every $f\in{\mathscr{M}}$ that only has essential coordinates.
+Then $\xi$ extends to a minion homomorphism ${\mathscr{M}}\to{\mathscr{N}}$.
+Proof.
+Suppose $\xi$ is defined for elements with only essential coordinates, and let $f\in{\mathscr{M}}$ be arbitrary.
+Let $f_{/\tau}$ be an essential part of $f$, and let $\tau^{\prime}$ be such that $\tau\circ\tau^{\prime}=\operatorname{id}$.
+Let $\zeta$ be the extension of $\xi$ defined by $\zeta(f):=(\xi(f_{/\tau}))_{/\tau^{\prime}}$.
+Since $f_{/\tau}$ only has essential coordinates, this definition is well posed.
+We prove simultaneously the following two claims: The definition of $\zeta$ does not depend on the choice of an essential part of $f$, and $\zeta$ is a minion homomorphism.
+To this end, suppose that $g=f_{/\sigma}$ holds.
+Let $g_{/\tau_{g}}$ be an essential part of $g$, and let $f_{/\tau_{f}}$ be an essential part of $f$.
+Let $\tau^{\prime}_{g}$ and $\tau^{\prime}_{f}$ be such that $\tau^{\prime}_{g}\circ\tau_{g}=\operatorname{id}$ and $\tau^{\prime}_{f}\circ\tau_{f}=\operatorname{id}$ hold when restricted to the essential coordinates of $g$ and $f$, respectively.
+Then we have that $g_{/\tau_{g}}=f_{/\tau_{g}\circ\sigma}=\left[(f_{/\tau_{f}})_{/\tau^{\prime}_{f}}\right]_{/\tau_{g}\circ\sigma}=(f_{/\tau_{f}})_{/\tau_{g}\circ\sigma\circ\tau^{\prime}_{f}}$.
+It follows that
+$\zeta(g)$ $=(\xi(g_{/\tau_{g}}))_{/\tau^{\prime}_{g}}=\left[\xi(f_{/\tau_{f}})_{/\tau_{g}\circ\sigma\circ\tau^{\prime}_{f}}\right]_{/\tau^{\prime}_{g}}=\xi(f/_{\tau_{f}})_{/\sigma\circ\tau^{\prime}_{f}}=(\xi(f_{/{\tau_{f}}})_{/\tau^{\prime}_{f}})_{/\sigma}=\zeta(f)_{/\sigma}.$
+This proves that $\zeta$ is a minion homomorphism, and that it is well defined by taking $g=f$ and $\sigma=\operatorname{id}$ in this claim, with potentially two different essential parts $\tau_{f},\tau_{g}$.
+∎
+Proof of Theorem 4.
+For ease of notation, we assume here that $\mathbb{B}$ has a single relation with symbol $R$; the proof in the case of multiple relations is entirely analogous.
+Let $r$ be the arity of $R$.
+Let $\mathbb{C}=(A^{d};R^{\mathbb{C}})$ be as in the definition of a q-construction, and let $\tilde{R}\subseteq A^{dr}$ be its vectorisation.
+Thus, $(a_{1,1},\dots,a_{1,d},\dots,a_{r,1},\dots,a_{r,d})\in\tilde{R}$ if, and only if, $(\mathbf{a}_{1},\dots,\mathbf{a}_{r})\in R^{\mathbb{C}}$.
+We will establish the chain of minion homomorphisms
+$\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(A;\tilde{R})\to\operatorname{qPol}(\mathbb{C})\to\operatorname{qPol}(\mathbb{B}).$
+The first arrow in this chain follows from Theorem 40.
+For the last arrow in this chain, let $U\colon\mathbb{C}\to_{\operatorname{q}}\mathbb{B}$ and $W\colon\mathbb{B}\to_{\operatorname{q}}\mathbb{C}$.
+Defining $\xi\colon\operatorname{qPol}(\mathbb{C})\to\operatorname{qPol}(\mathbb{B})$ as $\xi(Q):=(U\bullet Q)*(W,\dots,W)$ gives the required minion homomorphism.
+[…]
+We now show that $Q$ consists of PVMs.
+If $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ are as in the first case, then there is nothing to show.
+So let us assume that $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ all lie in the $i$th projection of $R^{\mathbb{C}}$.
+Let $\mathbb{G}$ be a q-definition for $\tilde{R}$ in $\mathbb{A}$, with its distinguished vertices $g_{i,k}$ for $i\in[r]$ and $k\in[d]$.
+By ($(\mathbf{q_{1}})$), for $j\in[n]$, there exist homomorphisms $h_{j}\colon\mathbb{G}\to\mathbb{A}$ such that $h_{j}(g_{i,k})=a_{j,k}$ for all $k\in[d]$.
+Let $U=\tilde{Q}\bullet(h_{1}\otimes\dots\otimes h_{n})\colon\mathbb{G}\to\mathbb{A}$, which can be explicitly computed by $U_{g_{i,k},b}=Q_{(a_{1,k},\dots,a_{n,k}),b}$ for all $b\in A$.
+By ($(\mathbf{q_{2}})$), we must have $U_{g_{i,k},b},U_{g_{i,k^{\prime}},c}]=0$ for all $k,k^{\prime}\in[d]$ and $b,c\in A$.
+Thus, we get $Q_{(a_{1,k},\dots,a_{n,k}),b},Q_{(a_{1,k^{\prime}},\dots,a_{n,k^{\prime}}),c}]=0$ for all $k,k^{\prime}\in[d]$ and $b,c\in A$.
+It follows that all the factors in the definition of $Q_{(\mathbf{a}_{1},\dots,\mathbf{a}_{n}),\mathbf{b}}$ commute, so that the operator is indeed a projection.
+A simple computation shows that $\sum_{\mathbf{b}}Q_{(\mathbf{a}_{1},\dots,\mathbf{a}_{n}),\mathbf{b}}=\operatorname{id}$ holds for all $\mathbf{a}_{1},\dots,\mathbf{a}_{n}\in A^{d}$, so that $Q$ is indeed a quantum function.
+Next, we show that $Q$ is in $\operatorname{qPol}(\mathbb{C})$, starting with $($$\mathbf{QH_{1}}$)).
+To this end, let $(\mathbf{a}_{1,j},\dots,\mathbf{a}_{r,j})\in R^{\mathbb{C}}$ for all $j\in[n]$ and let $(\mathbf{b}_{1},\dots,\mathbf{b}_{r})\not\in R^{\mathbb{C}}$.
+Note that, for all $i\in[r]$, we have $Q_{(\mathbf{a}_{i,1},\dots,\mathbf{a}_{i,n}),\mathbf{b}_{i}}=\prod_{k=1}^{d}\tilde{Q}_{(a_{i,1,k},\dots,a_{i,n,k}),b_{k}}$ (i.e., corresponding to the second case of the definition of $Q$), since $\mathbf{a}_{i,1},\dots,\mathbf{a}_{i,n}$ all belong to the $i$th projection of $R$.
+Thus, we obtain
+$\prod_{i=1}^{r}Q_{(\mathbf{a}_{i,1},\dots,\mathbf{a}_{i,n}),\mathbf{b}_{i}}$ $=\prod_{i=1}^{r}\prod_{k=1}^{d}\tilde{Q}_{(a_{i,1,k},\dots,a_{i,n,k}),b_{k}}=0,$
+where the second equality follows from $($$\mathbf{QH_{1}}$)) applied to $\tilde{Q}$, since every $dr$-tuple of the form
+$(a_{1,j,1},\dots,a_{1,j,d},\dots,a_{r,j,1},\dots,a_{r,j,d})$
+is in $\tilde{R}$ for $j\in[n]$, while the $dr$-tuple $(b_{1,1},\dots,b_{r,d})$ is not.
+We now turn to $($$\mathbf{QH_{2}}$)). Using the same notation as in the previous paragraph, we aim to show that
+$[Q_{(\mathbf{a}_{i,1},\dots,\mathbf{a}_{i,n}),\mathbf{b}},Q_{(\mathbf{a}_{i^{\prime},1},\dots,\mathbf{a}_{i^{\prime},n}),\mathbf{c}}]=0$ (7)
+for all $i,i^{\prime}\in[r]$ and all $\mathbf{b},\mathbf{c}\in A^{d}$.
+Using the same argument as above, we note that the PVMs for $Q$ are defined as in the second case, since all the tuples belong to the $i$th (resp.
+$i^{\prime}$th) projection of $R^{\mathbb{C}}$.
+Thus, (7) implies
+$\left[\prod_{k=1}^{d}Q_{(a_{i,1,k},\dots,a_{i,n,k}),\mathbf{b}},\prod_{k^{\prime}=1}^{d}Q_{{(a_{i^{\prime},1,k^{\prime}},\dots,a_{i^{\prime},n,k^{\prime}}),\mathbf{c}}}\right]=0.$
+We now show that any two factors in this expression commute, for $i,i^{\prime}\in[r]$ and $k,k^{\prime}\in[d]$.
+Indeed, by ($(\mathbf{q_{1}})$), there is for every $j\in[n]$ a homomorphism $h_{j}\colon\mathbb{G}\to\mathbb{A}$ such that, for all $i^{\prime\prime}\in[r]$ and all $k^{\prime\prime}\in[d]$, it holds that $h_{j}(g_{i^{\prime\prime},k^{\prime\prime}})=a_{i^{\prime\prime},j,k^{\prime\prime}}$.
+Define $U:=\tilde{Q}\bullet(h_{1}\otimes\dots\otimes h_{n})$, so that $U_{g_{i^{\prime\prime},k^{\prime\prime}},c}=\tilde{Q}_{(a_{i^{\prime\prime},1,k^{\prime\prime}},\dots,a_{i^{\prime\prime},n,k^{\prime\prime}}),c}$.
+In particular, for $(i^{\prime\prime},k^{\prime\prime})=(i,k)$ and $(i^{\prime\prime},k^{\prime\prime})=(i^{\prime},k^{\prime})$, we obtain by ($(\mathbf{q_{2}})$) the desired claim.
+This concludes the proof of $($$\mathbf{QH_{2}}$)).
+[…]
+$\zeta(\tilde{W})_{(\mathbf{a}_{1},\dots,\mathbf{a}_{n}),\mathbf{b}}=\prod_{k=1}^{d}\tilde{W}_{(a_{1,k},\dots,a_{n,k}),b_{k}}=\prod_{k=1}^{d}\tilde{Q}_{(a_{\sigma(1),k},\dots,a_{\sigma(m),k}),b_{k}}=\zeta(\tilde{Q})_{(\mathbf{a}_{\sigma(1)},\dots,\mathbf{a}_{\sigma(m)}),\mathbf{b}}.$
+[…]
+As an immediate consequence of the above result, we obtain the following.
+Corollary 46.
+If $\mathbb{A}$ q-constructs $\mathbb{B}$, then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace.
+Proof.
+The result directly follows from Theorem 4 and Theorem 3. ∎
+5.2 Commutativity gadgets
+It is not hard to show that every q-definition yields a classical pp-definition.
+Indeed, considering ($(\mathbf{q_{1}})$) for quantum homomorphisms with a Hilbert space of dimension 1 (which corresponds to classical homomorphisms), we find that the tuples in $R$ are in one-to-one correspondence with the restrictions of homomorphisms $\mathbb{G}\to\mathbb{A}$ to $\{g_{1},\dots,g_{r}\}$.
+However, not every pp-definition is a q-definition.
+It turns out that one can build a q-definition out of a pp-definition in a “uniform” way by pairing it with an extra construction making sure that the process of replacing constraints with gadgets preserves the existence of quantum homomorphisms.
+This is the notion of commutativity gadgets from [47], which we define below.
+Note that, in our terminology, a commutativity gadget is simply a q-definition for the complete binary relation.
+Definition 47.
+Fix a $\sigma$-structure $\mathbb{A}$.
+A commutativity gadget for $\mathbb{A}$ is a q-definition for the relation $A^{2}$.
+In other words, a commutativity gadget is a $\sigma$-structure $\mathbb{G}$ with two distinguished elements $u,v\in G$ satisfying the following properties:
+  • $(\mathbf{c_{1}})$
+    For any quantum function $Q\colon\{u,v\}\to_{\operatorname{q}}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum homomorphism $Q^{\prime}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ extending $Q$.
+  • $(\mathbf{c_{2}})$
+    For every quantum homomorphism $Q\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
+Remark 48.
+In [34], an algebraic commutativity gadget is defined by replacing ($(\mathbf{c_{1}})$) by the following condition:
+  • $(\mathbf{c_{1}^{\prime}})$
+    For every $a,b\in A$, there exists a quantum homomorphism $Q\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
+Note that this is a special case of $(\mathbf{c_{1}})$ and, thus, any commutativity gadget is an algebraic commutativity gadget.
+While this appears to be a strictly weaker requirement than the one in Definition 47, we shall prove below that the existence of an algebraic commutativity gadget in the sense of [34] is equivalent to the existence of a commutativity gadget in our sense.
+If $\mathbb{A}$ has a commutativity gadget, then we have a natural way of obtaining a complete and sound reduction from $\operatorname{CSP_{q}}(\mathbb{B})$ to $\operatorname{CSP_{q}}(\mathbb{A})$ out of a classical gadget: For every pair of variables $x,y$ in a relation in an instance of $\operatorname{CSP_{q}}(\mathbb{B})$, glue to $x,y$ a copy of the commutativity gadget, identifying $x$ with $u$ and $y$ with $v$.
+Condition ($(\mathbf{c_{2}})$) in Definition 47 forces the associated projectors to commute in every quantum homomorphism to $\mathbb{A}$, and condition ($(\mathbf{c_{1}})$) tells us that no additional constraint is forced upon $x,y$ by gluing the commutativity gadget.
+We can view this fact at the relational level, as follows.
 Proposition 49.
-A quantum polymorphism of a structure $\mathbb{A}$ is non-contextual if and only if it belongs to $\operatorname{\oplus Pol}(\mathbb{A})$.
+Let $\mathbb{A},\mathbb{B}$ be structures such that $(i)$ $\mathbb{A}$ has a commutativity gadget, and $(ii)$ $\mathbb{A}$ $p$-defines $\mathbb{B}$. Then $\mathbb{A}$ q-defines $\mathbb{B}$.
 Proof.
-This is a simple consequence of the spectral theorem. ∎
+Let $S=R^{\mathbb{B}}$ be a relation of $\mathbb{B}$ of some arity $r$, and let $\mathbb{G}$ be a classical pp-definition for $R^{\mathbb{B}}$ in $\mathbb{A}$, with distinguished vertices $g_{1},\dots,g_{r}$.
+Let $\mathbb{H}$ be a commutativity gadget for $\mathbb{A}$, with distinguished vertices $u,v$.
+We let $\mathbb{G}^{\prime}$ be the structure that initially consists of the disjoint union of $\mathbb{G}$ and a copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$ for each pair of distinct vertices $x,y$ in $G$, and in which we identify each such pair $x,y$ with the vertices $u,v$ from the copy $\mathbb{H}_{x,y}$ of $\mathbb{H}$.
+We show that $\mathbb{G}^{\prime}$ (with distinguished vertices $g_{1},\dots,g_{r}$) is a q-definition for $S$ in $\mathbb{A}$.
+Let $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;S)$ be a quantum homomorphism.
+In particular, $Q$ is non-contextual and can therefore be written as $h_{1}\oplus\dots\oplus h_{d}$ for some homomorphisms $h_{k}\colon\mathbb{R}\to(A;S)$.
+Since $\mathbb{G}$ is a pp-definition for $S$ in $\mathbb{A}$, it follows that there exist homomorphisms $f_{k}\colon\mathbb{G}\to\mathbb{A}$ such that $f_{k}(g_{i})=h_{k}(i)$ for all $i\in[r]$.
+Then $Q^{\prime}=\bigoplus_{k\in[d]}f_{k}$ is a non-contextual quantum homomorphism $\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$.
+In particular, $Q^{\prime}_{x}$ and $Q^{\prime}_{y}$ commute for all $x,y\in G$, and it follows from ($(\mathbf{c_{1}})$) that for all $x,y$ there exists a quantum homomorphism $Q^{(x,y)}\colon\mathbb{H}\to_{\operatorname{q}}\mathbb{A}$ such that $Q^{(x,y)}_{u}=Q^{\prime}_{x}$ and $Q^{(x,y)}_{v}=Q^{\prime}_{y}$.
+We thus obtained a quantum homomorphism $\mathbb{G}^{\prime}\to_{\operatorname{q}}\mathbb{A}$, which proves ($(\mathbf{q_{1}})$).
+Let now $Q\colon\mathbb{G}^{\prime}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism.
+For each $x,y\in G$, we have by ($(\mathbf{c_{2}})$) that $Q_{x}$ and $Q_{y}$ commute.
+Thus, $Q$ restricts to a quantum homomorphism $\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ that is non-contextual and can thus be written as $h_{1}\oplus\dots\oplus h_{d}$, each $h_{k}$ being a classical homomorphism $\mathbb{G}\to\mathbb{A}$.
+Since $\mathbb{G}$ is a classical pp-definition for $R^{\mathbb{B}}$, we obtain that $(h_{k}(g_{1}),\dots,h_{k}(g_{r}))\in R^{\mathbb{B}}$ for all $k\in[d]$.
+Thus, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}=\bigoplus_{k}(h_{k})_{g_{i},a}$ for $i\in[r],a\in A$ is a quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}(A;S)$.
+∎
+Next, we prove the analogue of Proposition 49 in the more general case of q-constructions.
+Theorem 50.
+Let $\mathbb{A},\mathbb{B}$ be structures such that $(i)$ $\mathbb{A}$ has a commutativity gadget, and $(ii)$ $\mathbb{A}$ $p$-constructs $\mathbb{B}$. Then $\mathbb{A}$ q-constructs $\mathbb{B}$.
+Proof.
+It is well known that condition $(ii)$ is equivalent to the fact that $\mathbb{B}$ is homomorphically equivalent to a pp-power $\mathbb{C}$ of $\mathbb{A}$ (for instance, see [9, 5]).
+Let $\mathbb{D}$ be the structure with domain $D=A$ obtained by identifying $(A^{d})^{r}$ with $A^{dr}$ for each $r$-ary relation of $\mathbb{C}$, as in Definition 44.
+Since $\mathbb{C}$ is a pp-power of $\mathbb{A}$, we have that $\mathbb{D}$ is pp-definable from $\mathbb{A}$.
+Using that $\mathbb{A}$ admits a commutativity gadget, it follows from Proposition 49 that $\mathbb{D}$ is in fact q-definable from $\mathbb{A}$.
+But this precisely means that $\mathbb{A}$ q-constructs $\mathbb{B}$, and we are done.
+∎
+5.3 Complexity reductions via non-contextual quantum polymorphisms
+We now have all ingredients to classify the languages for which commutativity gadgets exist in terms of their quantum polymorphism minion—thus proving the following theorem, which is an extended version of Theorem 6.
+Theorem 51.
+Let $\mathbb{A}$ be a relational structure. The following are equivalent:
+  1. 1.
+    $\mathbb{A}$ has a commutativity gadget.
+  2. 2.
+    $\mathbb{A}$ has an algebraic commutativity gadget.
+  3. 3.
+    $\mathbb{A}^{n}$ is an (algebraic) commutativity gadget for all large enough $n\in\mathbb{N}$.
+  4. 4.
+    $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
+The next definition shall be useful.
 Definition 52.
@@
 The pairs $(a_{1},b_{1}),\dots,(a_{n},b_{n})\in A^{2}$ are called generators of $A^{2}$ in $\operatorname{Pol}(\mathbb{A})$ if for every pair $(c,d)\in A^{2}$, there exists $g\in\operatorname{Pol}(\mathbb{A})$ of arity $n$ such that $g(a_{1},\dots,a_{n})=c$ and $g(b_{1},\dots,b_{n})=d$.
-Note that, if $A$ is finite (as we always assume), $A^{2}$ admits a family of $n=|A|^{2}$ generators.
+Note that, if $A$ is finite (as is always the case in the current paper), $A^{2}$ admits a family of $n=|A|^{2}$ generators.
 Indeed, we can simply let $g$ be the $i$-th projection $\mathbb{A}^{n}\to\mathbb{A}$, with $i\in[n]$ being the index of the pair $(c,d)$.
 However, $n$ could be smaller for certain $\mathbb{A}$.
+This is for example the case for complete graphs (where $n=2$ is enough) or for undirected cycles (where $m$ is enough for an odd cycle of length $2m+1$).
 We point out that there are many cases of infinite structures $\mathbb{A}$ where $A^{2}$ is finitely generated under $\operatorname{Pol}(\mathbb{A})$; this is for example the case for the class of $\omega$-categorical structures, which has been widely studied in the context of classical constraint satisfaction problems.
 Proposition 53.
-Let $\mathbb{A}$ be a structure such that $A^{2}$ admits a family of $n$ generators under $\operatorname{Pol}(\mathbb{A})$.
+Let $\mathbb{A}$ be a (not necessarily finite) structure such that $A^{2}$ admits a family of $n$ generators under $\operatorname{Pol}(\mathbb{A})$.
 Suppose also that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
@@
 Let $\mathbb{G}$ be $\mathbb{A}^{n}$ and the two distinguished elements be $\mathbf{a}=(a_{1},\dots,a_{n})$ and $\mathbf{b}=(b_{1},\dots,b_{n})$.
-Every quantum homomorphism $\mathbb{G}\to_{q}\mathbb{A}$ is a quantum polymorphism of $\mathbb{A}$ and therefore ($(\mathbf{c_{2}})$) holds by the assumption that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
-Moreover, by the spectral theorem, every quantum function $Q\colon\{\mathbf{a},\mathbf{b}\}\to A$ such that $Q(\mathbf{a})$ commutes with $Q(\mathbf{b})$ is of the form $h_{1}\oplus\cdots\oplus h_{d}$, where $h_{i}\colon\{\mathbf{a},\mathbf{b}\}\to A$ is a classical function.
+Every quantum homomorphism $\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ is a quantum polymorphism of $\mathbb{A}$ and therefore ($(\mathbf{c_{2}})$) holds by the assumption that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
+Moreover, by the spectral theorem, every quantum function $Q\colon\{\mathbf{a},\mathbf{b}\}\to_{\operatorname{q}}A$ such that $Q(\mathbf{a})$ commutes with $Q(\mathbf{b})$ is of the form $h_{1}\oplus\cdots\oplus h_{d}$, where $h_{i}\colon\{\mathbf{a},\mathbf{b}\}\to A$ is a classical function.
 Using that $(a_{1},b_{1}),\dots,(a_{n},b_{n})$ are generators for $A^{2}$, we know that, for each $i$, there exists $g_{i}\in\operatorname{Pol}(\mathbb{A})$ such that $g_{i}(\mathbf{a})=h_{i}(\mathbf{a})$ and $g_{i}(\mathbf{b})=h_{i}(\mathbf{b})$.
@@
 Let $\mathbb{G}$ be an algebraic commutativity gadget with distinguished elements $u,v\in G$.
-Let $Q\colon\mathbb{A}^{n}\to_{q}\mathbb{A}$ be a quantum polymorphism, and choose two tuples $\mathbf{a},\mathbf{b}\in A^{n}$ and two elements $c,d\in A$.
+Let $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ be a quantum polymorphism, and choose two tuples $\mathbf{a},\mathbf{b}\in A^{n}$ and two elements $c,d\in A$.
 We want to show that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, which is enough to conclude by Proposition 28.
-Using ($(\mathbf{c_{1}^{\prime}})$) we know that, for each $i\in[n]$, there exists a quantum homomorphism $W^{(i)}\colon\mathbb{G}\to_{q}\mathbb{A}$ such that
+Using ($(\mathbf{c_{1}^{\prime}})$) we know that, for each $i\in[n]$, there exists a quantum homomorphism $W^{(i)}\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}$ such that
 $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}.$
-Consider now the quantum homomorphism $W\colon\mathbb{G}\to_{q}\mathbb{A}^{n}$ given by the tensor product
+Consider now the quantum homomorphism $W\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}^{n}$ given by the tensor product
 $W=\bigotimes_{i\in[n]}W^{(i)}$
 and observe that $W_{u,\mathbf{a}}=W_{v,\mathbf{b}}=\operatorname{id}$. Take now the composition
-$U=Q\bullet W:\mathbb{G}\to_{q}\mathbb{A}.$
+$U=Q\bullet W\colon\mathbb{G}\to_{\operatorname{q}}\mathbb{A}.$
 Using ($(\mathbf{c_{2}})$), we observe that $U$ satisfies $U_{u,c},U_{v,d}]=0$. However, expanding the composition as per Definition 20, we have
@@
 Proof of Theorem 51.
-The result is a direct consequence of Propositions 20 and 21 and the fact that any commutativity gadget is an algebraic commutativity gadget (as observed in Remark 12). ∎
-4. Warm-up: Quantum polymorphisms of cliques
-In this section, we give a first simple example of using quantum polymorphisms to obtain commutativity gadgets, in the setting of cliques. In particular, we prove the following.
+The result is a direct consequence of Propositions 53 and 54 and the fact that any commutativity gadget is an algebraic commutativity gadget (as observed in Remark 48). ∎
+As a straightforward consequence of the above, we obtain the following result on reducibility between quantum CSPs.
+Theorem 55.
+Let $\mathbb{A},\mathbb{B}$ be relational structures such that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$ and $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$.
+Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace.
+Proof.
+The fact that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$ is equivalent to $\mathbb{A}$ admitting a commutativity gadget by Theorem 51, while the fact that $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\mathbb{B})$ is equivalent to $\mathbb{A}$ $p$-constructing $\mathbb{B}$ by Theorem 30.
+Using Theorem 50, we deduce that $\mathbb{A}$ q-constructs $\mathbb{B}$.
+Hence, there exists a minion homomorphism $\operatorname{qPol}(\mathbb{A})\to\operatorname{qPol}(\mathbb{B})$ by Theorem 4, and the result follows from Theorem 3.
+∎
+We immediately derive from the above theorem the following useful undecidability criterion for quantum CSPs. As we shall see later, this generalises the known undecidability criteria for quantum CSPs.
+Corollary 56.
+Let $\mathbb{A}$ be a structure such that $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$ and all quantum polymorphisms of $\mathbb{A}$ are non-contextual.
+Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+Proof.
+Since all quantum polymorphisms of $\mathbb{A}$ are non-contextual, we deduce that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
+Hence, by LABEL:{thm_comm_gadgets_plus_minion_homomorphisms_gives_reduction}, there is a logspace (in particular, computable) reduction from $\operatorname{CSP_{q}}(\operatorname{3SAT})$ to $\operatorname{CSP_{q}}(\mathbb{A})$.
+The result then follows from the undecidability of $\operatorname{CSP_{q}}(\operatorname{3SAT})$ [46, 60].
+∎
+The result above can be rephrased in the following, convenient form.
+See 7
+Proof.
+Since $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$, it follows from the algebraic CSP dichotomy theorem [21, 68] that $\mathbb{A}$ $p$-constructs $\operatorname{3SAT}$ and thus, by Theorem 30, $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$.
+To conclude, we apply Corollary 56.
+∎
+Note that Theorem 7 is in fact equivalent to Corollary 56 if $\mathsf{P}\neq\mathsf{NP}$.
+Indeed, under such hypothesis, $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$ if, and only if, $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$.
+On the other hand, if $\mathsf{P}=\mathsf{NP}$Theorem 7 becomes vacuous (unlike Corollary 56).
+Example 57.
+Recall Figure 1 in Section 2.
+The figure illustrates the undecidability of $\operatorname{CSP_{q}}(\mathbb{C}_{5})$ by reducing from quantum 1-in-3-SAT; i.e., from $\operatorname{CSP_{q}}(\mathbb{A})$ where $\mathbb{A}$ is the language encoding 1-in-3-SAT formulae.
+Formally, $\mathbb{A}$ is the Boolean structure with a single, ternary relation
+$R^{\mathbb{A}}=\{(0,0,1),(0,1,0),(1,0,0)\}.$
+The undecidability of quantum 1-in-3-SAT can be obtained by reducing from quantum 3SAT, using for example that $\operatorname{Pol}(\mathbb{A})$ does not contain the majority polymorphism, see [33] and Section 11.
+5.4 Comparison with known undecidability criteria
+The following undecidability criterion follows by combining [2, Lemma 10] with the undecidability of $\operatorname{CSP_{q}}(\mathbb{XOR})$ from [67], as pointed out in [63].
+(For the formal definition of $\mathbb{XOR}$, see Section 11.)
+Theorem 58 ([2, 67, 63]).
+Let $\mathbb{A}$ be a Boolean relational structure such that $(i)$ its signature contains a binary symbol $R$ for which $R^{\mathbb{A}}=A^{2}$, and $(ii)$ $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$.
+Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+This result was generalised in [33] by replacing the requirement $(i)$ on the existence of a complete binary relation with a weaker structural condition called two-variable (non-)falsifiability, and relaxing the requirement on $\mathbb{A}$ being Boolean.
+Given any relation $S\subseteq A^{r}$ and two indices $i\neq j\in[r]$, we denote by $S_{i,j}$ the binary projection of $S$ onto the coordinates $i,j$; i.e.,
+$S_{i,j}=\bigcup_{(a_{1},\dots,a_{r})\in S}\{(a_{i},a_{j})\}.$
+Definition 59 ([33]).
+Let $\mathbb{A}$ be a relational structure with signature $\sigma$.
+We say that $\mathbb{A}$ is two-variable falsifiable (TVF) if for each symbol $R\in\sigma$ and each pair $i\neq j\in[\operatorname{ar}(R)]$, it holds that $(R^{\mathbb{A}})_{i,j}\neq A^{2}$.
+Theorem 60 ([33, Theorem 4.14]).
+Let $\mathbb{A}$ be a relational structure such that $(i)$ $\mathbb{A}$ is not TVF, and $(ii)$ $\operatorname{CSP}(\mathbb{A})\not\in\mathsf{P}$.
+Then, $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+[^7] We point out that, just like for our Theorem 7, condition $(ii)$ in both Theorem 58 and Theorem 60 can be replaced by the algebraic condition $\operatorname{Pol}(\mathbb{A})\to\operatorname{Pol}(\operatorname{3SAT})$, which does not trivialise if $\mathsf{P}=\mathsf{NP}$.
+Clearly, any structure satisfying the requirements of Theorem 58 also satisfies the requirements of Theorem 60, so the latter is indeed a generalisation of the former.
+We now prove that our undecidability criterion Theorem 7 extends the one in Theorem 60.
+To that end, we prove the following.
+Proposition 61.
+Let $\mathbb{A}$ be a relational structure. If $\mathbb{A}$ is not TVF, then $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
+Note that, in fact, Theorem 7 is strictly more general than Theorem 60.
+Indeed, there are multiple examples of structures $\mathbb{A}$ such that $(i)$ $\operatorname{CSP}(\mathbb{A})$ is NP-complete, $(ii)$ $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$, but $(iii)$ $\mathbb{A}$ is TVF.
+For instance, as we shall see in later sections, the Siggers digraph, any odd cycle, and any clique give examples of such $\mathbb{A}$.
+Proof of Proposition 61.
+Let $\mathbb{A}$ be a $\sigma$-structure that is not TVF, and let $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ be a quantum polymorphism of some arity $n\in\mathbb{N}$.
+Take two tuples $\mathbf{a},\mathbf{b}\in A^{n}$.
+We claim that the PVMs $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute.
+Since $\mathbb{A}$ is not TVF, there exists some symbol $R\in\sigma$ (of arity, say, $r$) and two indices $i\neq j\in[r]$ such that $R^{\mathbb{A}}$ has full binary projection onto the coordinates $i,j$; i.e., $(R^{\mathbb{A}})_{i,j}=A^{2}$.
+Therefore, we can assign to each $\ell\in[n]$ a tuple $\textbf{t}^{(\ell)}=(t_{1}^{(\ell)},\dots,t_{r}^{(\ell)})\in R^{\mathbb{A}}$ such that
+$(t_{i}^{(\ell)},t_{j}^{(\ell)})=(a_{\ell},b_{\ell}).$ (8)
+Let now $T$ be the $n\times r$ matrix whose rows are $\textbf{t}^{(1)},\textbf{t}^{(2)},\dots,\textbf{t}^{(n)}$.
+By definition of the categorical power $\mathbb{A}^{n}$, the $r$-tuple consisting of the list of columns of such matrix belongs to $R^{\mathbb{A}^{n}}$.
+This means that any pair of columns of $T$ having distinct indices is adjacent in $\operatorname{Gaif}(\mathbb{A}^{n})$.
+Note now that, because of (8), the tuples $\mathbf{a}$ and $\mathbf{b}$ occur in $T$ as the $i$-th and $j$-th columns, respectively.
+We deduce that $\mathbf{a}$ and $\mathbf{b}$ are adjacent in $\operatorname{Gaif}(\mathbb{A}^{n})$.
+Using $($$\mathbf{QH_{2}}$)), we conclude that $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ must commute, as claimed.
+∎
+We point out that Proposition 61 can alternatively be proved by observing that any relation having a full binary projection (and, thus, witnessing that $\mathbb{A}$ is not TVF) can be used to directly give a commutativity gadget for $\mathbb{A}$; then, the result follows from Theorem 51.
+6 Quantum polymorphisms of cliques
+In this section, we give a first simple example of using quantum polymorphisms to obtain commutativity gadgets, in the setting of cliques.
+We let $\mathbb{K}_{m}$ denote the $m$-clique.
+We prove the following.
 Theorem 62.
@@
 Proof.
-The proof goes by induction on $m\geq 3$.
-For the case of $m=3$, it follows from [Ji, Lemma 4] that $\mathbb{K}_{3}$ has a commutativity gadget. Then, Theorem 16 implies the result.
+The proof goes by induction on $m$.
+For the case of $m=3$, it follows from [47, Lemma 4] that $\mathbb{K}_{3}$ has a commutativity gadget. Then, Theorem 51 implies the result.
 Take now $m\geq 4$.
-Let $Q\colon(\mathbb{K}_{m})^{n}\to_{q}\mathbb{K}_{m}$ be a quantum homomorphism over a Hilbert space $H$.
+Let $Q\colon(\mathbb{K}_{m})^{n}\to_{\operatorname{q}}\mathbb{K}_{m}$ be a quantum homomorphism over a Hilbert space $H$.
 Pick $\mathbf{u},\mathbf{v}\in[m]^{n}$ and $i,j\in[m]$.
@@
 We first claim that
-(3) $[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},k}]=0.$
+$[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},k}]=0.$ (9)
 This is clearly the case if $Q_{\mathbf{w},k}=0$, so assume this is not the case.
-Define $\mathbb{X}$ to be the subgraph of $(\mathbb{K}_{m})^{n}$ induced by the vertices $\mathbf{z}$ such that $z_{p}\neq w_{p}$ for all $p\in\{1,\dots,n\}$, namely the subgraph induced by the neighborhood of $\mathbf{w}$.
+Define $\mathbb{X}$ to be the subgraph of $(\mathbb{K}_{m})^{n}$ induced by the vertices $\mathbf{z}$ such that $z_{p}\neq w_{p}$ for all $p\in[n]$, namely the subgraph induced by the neighborhood of $\mathbf{w}$.
 Note that $\mathbb{X}$ is isomorphic to $(\mathbb{K}_{m-1})^{n}$.
-Define $R\colon\mathbb{X}\to_{q}\mathbb{K}_{m-1}$ by setting $R_{\mathbf{z},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}$ over the (non-trivial) space $Q_{\mathbf{w},k}(H)$.
+Define $R\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{K}_{m-1}$ by setting $R_{\mathbf{z},\ell}=Q_{\mathbf{w},k}Q_{\mathbf{z},\ell}$ over the (non-trivial) space $Q_{\mathbf{w},k}(H)$.
 We check that this is a quantum homomorphism.
@@
 For $\mathbf{z}=\mathbf{u}$ and $\mathbf{z}^{\prime}=\mathbf{v}$, this gives the claimed eq.
-3.
-We can now already conclude by noting that
-$Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=[Q_{\mathbf{u},i}(\sum_{k}Q_{\mathbf{w},k}),Q_{\mathbf{v},j}(\sum_{\ell}Q_{\mathbf{w},\ell})]=\sum_{k,\ell}[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}],$
-that all the terms with $k=\ell$ in the sum vanish by eq. 3, and that all the terms with $k\neq\ell$ in the sum vanish, since
+9.
+Observe now that
+$Q_{\mathbf{u},i},Q_{\mathbf{v},j}]=[Q_{\mathbf{u},i}(\sum_{k}Q_{\mathbf{w},k}),Q_{\mathbf{v},j}(\sum_{\ell}Q_{\mathbf{w},\ell})]=\sum_{k,\ell}[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}].$
+All terms with $k=\ell$ in the sum vanish by eq. 9. Moreover, all terms with $k\neq\ell$ in the sum vanish as well, since
 $[Q_{\mathbf{u},i}Q_{\mathbf{w},k},Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}]$ $=Q_{\mathbf{u},i}Q_{\mathbf{w},k}Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}-Q_{\mathbf{v},j}Q_{\mathbf{w},\ell}Q_{\mathbf{u},i}Q_{\mathbf{w},k}$
@@
 where the second equality is due to the fact that $Q_{\mathbf{w},k}$ and $Q_{\mathbf{w},\ell}$ commute with $Q_{\mathbf{u},i}$ and $Q_{\mathbf{v},j}$, and the third holds since $Q_{\mathbf{w},k}Q_{\mathbf{w},\ell}=0=Q_{\mathbf{w},\ell}Q_{\mathbf{w},k}$.
+This is enough to conclude.
 ∎
-Note that $\mathbb{K}_{2}$ has quantum polymorphisms that are not in the quantum closure of $\operatorname{Pol}(\mathbb{K}_{2})$.
-Indeed, let $n=2$ and $H=\mathbb{C}^{2}$, and define $Q\colon(\mathbb{K}_{2})^{2}\to_{q}\mathbb{K}_{2}$ by the following matrices:
+Note that $\mathbb{K}_{2}$ does have quantum polymorphisms that are not in the quantum closure of $\operatorname{Pol}(\mathbb{K}_{2})$.
+Indeed, let $n=2$ and $H=\mathbb{C}^{2}$, and define $Q\colon(\mathbb{K}_{2})^{2}\to_{\operatorname{q}}\mathbb{K}_{2}$ by the following matrices:
   • $Q_{(0,0),0}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},Q_{(0,0),1}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$
@@
   • $Q_{(0,1),0}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix},Q_{(0,1),1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$.
-It is easily checked that this is a quantum polymorphism and that $Q_{(0,0),0},Q_{(1,0),0}]\neq 0$. By Theorem 16, we obtain that $\mathbb{K}_{2}$ does not have a commutativity gadget.
+It is easily checked that this is a quantum polymorphism and that $Q_{(0,0),0},Q_{(1,0),0}]\neq 0$.
+By Theorem 51, we deduce that $\mathbb{K}_{2}$ does not have a commutativity gadget.
+In Section 11, we will see that this fact can be alternatively derived via a more general argument, by simply noting that $\mathbb{K}_{2}$ is preserved by the classical majority polymorphism and is not the full binary relation, see Theorem 11.
 Proposition 63.
 Let $m\geq 3$.
-Then $\mathbb{K}_{m}$ has a commutativity gadget.
-In fact, for every $n\geq 2$ and $\mathbf{u},\mathbf{v}\in(\mathbb{K}_{m})^{n}$ such that $u_{i}=v_{i}$ for at least one $i\in[n]$ and $u_{j}\neq v_{j}$ for at least one $j\in[n]$, $(\mathbb{K}_{m})^{n}$ with the distinguished vertices $\mathbf{u},\mathbf{v}$ is a commutativity gadget for $\mathbb{K}_{m}$.
+For every $n\geq 2$ and $\mathbf{u},\mathbf{v}\in(\mathbb{K}_{m})^{n}$ such that $u_{i}=v_{i}$ for at least one $i\in[n]$ and $u_{j}\neq v_{j}$ for at least one $j\in[n]$, $(\mathbb{K}_{m})^{n}$ with the distinguished vertices $\mathbf{u},\mathbf{v}$ is a commutativity gadget for $\mathbb{K}_{m}$.
 Proof.
@@
 ∎
-A consequence of Proposition 23 is the undecidability of the entangled $m$-coloring problem for $m\geq 3$.
-The case of $m=3$ was proved in [culf_re-completeness_2025], and the case for $m>3$ was recently and independently proved in [Zeman, Proposition 5.2, Lemma 4.6].
-We include a simple proof here using commutativity gadgets for the convenience of the reader.
+An immediate consequence of Proposition 63 is the undecidability of the quantum $m$-coloring problem for $m\geq 3$.
 Corollary 64.
-$\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable for all $m\geq 3$.
+$\operatorname{CSP_{q}}(\mathbb{K}_{m})$ is undecidable for all $m\geq 3$.
 Proof.
-We shall reduce from $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, whose undecidability was proved in [Ji, culf2024re].
-Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{K}_{3})$, which we can assume does not contain isolated vertices.
-Let $\mathbb{Y}$ be the graph obtained by adjoining m3m-3 new vertices $y_{4},\dots,y_{m}$ to $\mathbb{X}$ and such that, whenever $x$ is a vertex in $\mathbb{X}$, then $\{x,y_{4},\dots,y_{m}\}$ forms a clique in $\mathbb{Y}$.
-Finally, adjoin a commutativity gadget between every $y_{i}$ and every other vertex in $\mathbb{Y}$.
-Such a commutativity gadget exists by Proposition 23.
-We let $\mathbb{Z}$ be the resulting structure.
-If $\mathbb{X}$ has a quantum homomorphism $Q$ to $\mathbb{K}_{3}$, then $Q^{\prime}\colon\mathbb{Y}\to_{q}\mathbb{K}_{m}$ defined by $Q^{\prime}_{y_{i},i}=\operatorname{id}$ for all $i\in\{4,\dots,m\}$ and $Q_{x,j}=Q^{\prime}_{x,j}$ for all $x\in X$ is a quantum homomorphism.
-This $Q^{\prime}$ can be extended to a quantum homomorphism $\mathbb{Z}\to\mathbb{K}_{m}$ by ($(\mathbf{c_{1}})$).
-Conversely, suppose that there exists a quantum homomorphism $Q\colon\mathbb{Z}\to_{q}\mathbb{K}_{m}$.
-By ($(\mathbf{c_{2}})$), it follows that for every $x\in X$, the restriction of $Q$ to $\{x,y_{2},\dots,y_{4}\}$ is non-contextual.
-Thus, this restriction can be written as $h_{1}\oplus\dots\oplus h_{d}$ for some classical homomorphisms $h_{1},\dots,h_{d}$.
-Up to composing the $h_{i}$ with an automorphism of $\mathbb{K}_{m}$, we can assume without loss of generality that $h_{i}(y_{j})=j$ for all $i\in\{1,\dots,d\}$ and $j\in\{4,\dots,m\}$.
-Thus, for each $i\in\{1,\dots,d\}$ and $x\in X$, we have that $h_{i}(x)\in\{1,2,3\}$.
-It follows that $\sum_{i=1}^{3}Q_{x,i}=\operatorname{id}$ and that $Q$ restricts to a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{K}_{3}$.
-∎
-5. Intermezzo: Contextuality and bifurcations
-In this section, we present a general result on the emergence of a certain non-orthogonality structure in any contextual quantum homomorphism.
-While in the current work we shall only use this result to analyse quantum polymorphisms of odd cycles (see Section 6), we believe this result will be useful in a broader setting.
+The result follows by combining Theorem 62 and Corollary 56. ∎
+We point out that the case of $m=3$ was proved in [33] via the reduction of [47], and the case of $m\geq 4$ was recently and independently proved in [34, Proposition 5.2, Lemma 4.6].
 $Q_{x_{0}}$$Q_{x_{1}}$$Q_{x_{2}}$$a_{0}$$a_{0}^{\prime}$$a_{1}$$a_{2}^{\prime}$$a_{2}$
-Figure 1. A bifurcation of length $2$. Solid lines mean non-orthogonality.
-The high-level idea is that, whenever some non-commutativity occurs in a quantum homomorphism $Q:\mathbb{X}\to_{q}\mathbb{Y}$, a bifurcation must appear in the product $X\times A$, in a sense that we formalise below.
-For simplicity, we restrict the analysis to binary structures (i.e., digraphs), although the same argument applies to quantum homomorphisms between arbitrary structures.
+Figure 3: A bifurcation of length $2$. Solid lines mean non-orthogonality.
+7 Intermezzo: Contextuality and bifurcations
+In this section, we present a general result on the emergence of a certain non-orthogonality pattern in any contextual quantum homomorphism.
+In later sections, we will use this result to analyse the quantum polymorphism minions of odd cycles and the Siggers digraph.
+We believe that the same ideas can be applied in a broader setting to prove non-contextuality of quantum polymorphisms (and, in particular, the existence of commutativity gadgets).
+The high-level idea is that, whenever two PVMs do not commute in a quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$, a bifurcation must appear along some walk in the product $X\times A$, in a sense that we formalise below.
+Recall that the Gaifman graph $\operatorname{Gaif}(\mathbb{X})$ has vertex set $X$, and its edges are all pairs of vertices in $X$ simultaneously appearing in a tuple in some relation of $\mathbb{X}$.
 Definition 65.
-Let $\mathbb{X},\mathbb{A}$ be digraphs, let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a quantum homomorphism, and let $\omega$ be a positive integer. A bifurcation for $Q$ is a set
-$V=\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}\subseteq X\times A$
-such that
-  • $x_{0},x_{1},\dots,x_{\omega}$ is a path in $\mathbb{X}$;
+Let $\mathbb{X},\mathbb{A}$ be relational structures, let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and let $\omega$ be a positive integer.
+Take a $4$-tuple $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$, where
+  • $\mathbf{x}=(x_{0},\dots,x_{\omega})\in X^{\omega+1}$,
+  • $\mathbf{a}=(a_{0},\dots,a_{\omega})\in A^{\omega+1}$, and
+  • $a_{0}^{\prime},a_{\omega}^{\prime}\in A$.
+We say that $\mathcal{B}$ is a (contextuality) bifurcation for $Q$ if it meets the following five conditions:
+  1. $(i)$
+    $x_{0},x_{1},\dots,x_{\omega}$ is a walk in $\operatorname{Gaif}(\mathbb{X})$;
+  2. $(ii)$
     $a_{0}\neq a_{0}^{\prime}$ and $a_{\omega}\neq a_{\omega}^{\prime}$;
-  • $Q_{\alpha,\beta}Q_{\delta,\gamma}\neq 0$ for each $((\alpha,\beta),(\delta,\gamma))\in V^{2}\setminus\{((x_{0},a_{0}^{\prime}),(x_{\omega},a_{\omega}^{\prime})),((x_{\omega},a_{\omega}^{\prime}),(x_{0},a_{0}^{\prime}))\}$ such that $\alpha\neq\delta$.
-We call $\omega$ the length of the bifurcation.
-Recall that $\mathbb{X}^{u}$ denotes the undirected graph underlying a digraph $\mathbb{X}$.
-Theorem 26.
-Let $\mathbb{X}$ and $\mathbb{A}$ be digraphs such that $\mathbb{X}^{u}$ is connected, and let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism.
-Then there exists a bifurcation for $Q$ whose length is at most the diameter of $\mathbb{X}^{u}$.
-We will need the following result from [manvcinska2016quantum], of which we give an alternative proof making use of the machinary of Section 3.
-Proposition 27 ([manvcinska2016quantum]).
-Let $Q\colon\mathbb{X}\to_{q}\mathbb{Y}$ be a quantum homomorphism between two digraphs $\mathbb{X}$ and $\mathbb{Y}$.
-If $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ are such that there exists a directed $\ell$-walk from $x$ to $x^{\prime}$ in $\mathbb{X}$ but there is no directed $\ell$-walk from $y$ to $y^{\prime}$ in $\mathbb{Y}$, then $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$.
+  3. $(iii)$
+    $Q_{x_{i},a_{i}}Q_{x_{j},a_{j}}\neq 0$ for each $i<j\in\{0,\dots,\omega\}$;
+  4. $(iv)$
+    $Q_{x_{0},a_{0}^{\prime}}Q_{x_{j},a_{j}}\neq 0$ for each $j\in\{1,\dots,\omega\}$;
+  5. $(v)$
+    $Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0$ for each $i\in\{0,\dots,\omega-1\}$.
+We call $\omega$ the length of the bifurcation $\mathcal{B}$.
+Figure 3 and Figure 4 illustrate two examples of contextuality bifurcations.
+The following is the main result of this section, showing that bifurcations occur in any contextual quantum homomorphism, provided that the left structure is connected.
+See 8 We shall derive Theorem 8 as a consequence of the next, slightly stronger result.
+Proposition 66.
+Let $\mathbb{X}$ and $\mathbb{A}$ be relational structures and let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism.
+Let $x_{0},x_{1},\dots,x_{\omega}$ be a walk in $\operatorname{Gaif}(\mathbb{X})$, and choose two values $a,\tilde{a}\in A$ in a way that
+  • $(I)$
+    $Q_{x_{0},a},Q_{x_{\omega},\tilde{a}}]\neq 0$, but
+  • $(II)$
+    for all $i,j\in\{0,\dots,\omega\}$ such that $\{i,j\}\neq\{0,\omega\}$, the PVMs $Q_{x_{i}}$ and $Q_{x_{j}}$ commute.
+Then $Q$ has a bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$ for some tuple $\mathbf{a}=(a_{0},\dots,a_{\omega})\in A^{\omega+1}$ such that $a_{0}=a$, $a_{\omega}=\tilde{a}$ and two values $a_{0}^{\prime},a_{\omega}^{\prime}\in A$, where $\mathbf{x}=(x_{0},\dots,x_{\omega})$.
+In the proof of Proposition 66, we shall repeatedly use the following fact about projectors.
+Fact 67.
+Let $P,Q$ be projectors onto some Hilbert space $H$. Then $[P,Q]=0$ if, and only if, $PQP=PQ=QP$.
 Proof.
-Consider the structure $\mathbb{G}$ consisting of a path of length $\ell$, and call $u,v$ the two extremities of this path.
-Since this structure is a tree, Proposition 47 implies that $(\mathbb{G},u,v)$ is a q-definition for certain binary relations $R_{X}\subseteq X^{2},R_{Y}\subseteq Y^{2}$ in $\mathbb{X}$ and $\mathbb{Y}$, respectively.
-It follows from Proposition 17 that $Q$ is a quantum homomorphism $(X,R_{X})\to_{q}(Y,R_{Y})$.
-By assumption, $R_{X}$ contains $(x,x^{\prime})$ and $R_{Y}$ does not contain $(y,y^{\prime})$, and therefore $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$ by $($$\mathbf{QH_{1}}$)).
-∎
-We shall repeatedly use a simple fact about projectors, stated below without proof.
-Fact 28.
-Let $P,Q$ be projectors onto some Hilbert space $H$. Then $[P,Q]=0$ if and only if $PQP=PQ=QP$.
-Proof of Theorem 26.
-Let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism over some Hilbert space $H$ of dimension $h$.
-Let $S$ be the set of pairs $(x,x^{\prime})\in X^{2}$ for which there exist $(a,a^{\prime})\in A^{2}$ such that $Q_{x,a},Q_{x^{\prime},a^{\prime}}]\neq 0$.
-Since $Q$ is contextual, $S$ is non-empty.
-Choose $(x,x^{\prime})\in S$ in a way that $\operatorname{dist}(x,x^{\prime})$ is minimum over all pairs in $S$ (where the distance is computed in $\mathbb{X}^{u}$), and call such distance $\omega$.
-Note that $\omega\geq 2$, as $Q$ satisfies $($$\mathbf{QH_{2}}$)).
-Let $x=x_{0},x_{1},\dots,x_{\omega}=x^{\prime}$ be a shortest oriented path in $\mathbb{X}$ connecting $x$ and $x^{\prime}$.
-By the minimality of $\operatorname{dist}(x,x^{\prime})$, observe that the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\beta}}$ commute for each $\alpha\leq\beta\in\{0,\dots,\omega\}$ such that $\alpha\neq 0$ or $\beta\neq\omega$.
-In particular, we can choose an o.n.
-basis $\mathcal{E}=(e_{1},\dots,e_{h})$ for $H$ that diagonalises all PVMs $Q_{x_{\alpha}}$ for each $\alpha\in\{0,\dots,\omega-1\}$.
-On the other hand, since $Q_{x_{0}}$ and $Q_{x_{\omega}}$ do not commute, there exist $a_{0},a_{\omega}\in A$ such that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
-Take the set $Z=\{i\in[h]:Q_{x_{\omega},b}e_{i}\in\{0,e_{i}\}\}$, and consider the projectors $P_{1}$ onto the space $\operatorname{span}(\{e_{i}:i\in Z\})$ and $P_{2}$ onto the space $\operatorname{span}(\{e_{i}:i\in[h]\setminus Z\})$.
+If $[P,Q]=0$, then $PQ=QP=QP^{2}=PQP$. The other direction is clear. ∎
+[…]
+Figure 4: A bifurcation of length $\omega$. Solid lines mean non-orthogonality.
+Proof of Proposition 66.
+Let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism over some Hilbert space $H$ of dimension $h$, and consider a walk $x_{0},x_{1},\dots,x_{\omega}$ in $\operatorname{Gaif}(\mathbb{X})$ and values $a,\tilde{a}\in A$ that satisfy conditions $(I)$ and $(II)$ of the proposition.
+Note that condition $(I)$ forces $\omega\geq 2$, as $Q$ satisfies $($$\mathbf{QH_{2}}$)).
+Moreover, using condition $(II)$, we can choose an orthonormal basis $\mathcal{E}=(e_{1},\dots,e_{h})$ for $H$ that diagonalises all PVMs $Q_{x_{\alpha}}$ for each $\alpha\in\{0,\dots,\omega-1\}$.
+Take the set $Z=\{i\in[h]:Q_{x_{\omega},{\tilde{a}}}e_{i}\in\{0,e_{i}\}\}$, and consider the following endomorphisms of $H$: the projector $P_{1}$ onto the space $\operatorname{span}(\{e_{i}:i\in Z\})$, and the projector $P_{2}$ onto the space $\operatorname{span}(\{e_{i}:i\in[h]\setminus Z\})$.
 Observe that $P_{1}P_{2}=0$ and $P_{1}+P_{2}=\operatorname{id}$.
-Moreover, we have
-(4) $[P_{1},M]=[P_{2},M]=0\quad\forall M\in\{Q_{x_{\alpha},d}:\alpha\in\{0,\dots,\omega-1\},d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}\}.$
+Moreover, calling $a_{0}=a$ and $a_{\omega}=\tilde{a}$, we have
+$[P_{1},M]=[P_{2},M]=0\quad\forall M\in\{Q_{x_{\alpha},d}:\alpha\in\{0,\dots,\omega-1\},d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}\}.$ (10)
 Indeed, $P_{1}$ is diagonal in the same basis $\mathcal{E}$ that diagonalises each $Q_{x_{\alpha}}$, so it commutes with all projectors in such PVMs. Furthermore,
 $Q_{x_{\omega},a_{\omega}}P_{1}=\sum_{i\in Z}Q_{x_{\omega},a_{\omega}}e_{i}e_{i}^{*}=\sum_{i\in Z}e_{i}e_{i}^{*}Q_{x_{\omega},a_{\omega}}=P_{1}Q_{x_{\omega},a_{\omega}}$
-since, for $i\in Z$, $e_{i}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}$.
+since, for each $i\in Z$, $e_{i}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}$.
 Hence, $P_{1}$ also commutes with $Q_{x_{\omega},a_{\omega}}$.
@@
 10 holds.
-In particular, the operators $Q_{x_{0},a_{0}}P_{1}$, $Q_{x_{0},a_{0}}P_{2}$, $Q_{x_{\omega},a_{\omega}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{2}$ are all projectors.
+In particular, the operators $Q_{x_{0},a_{0}}P_{1}$, $Q_{x_{0},a_{0}}P_{2}$, $Q_{x_{\omega},a_{\omega}}P_{1}$, and $Q_{x_{\omega},a_{\omega}}P_{2}$ are all projectors.
 Observe now that
@@
 $=[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{1}]+[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{2}]+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]$
-(5) $+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}].$
-Combining eq. 4 and 28, we find that the second summand in the right-hand side of section 5 is
+$+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}].$ (11)
+Combining eq. 10 and 67, we find that the second summand in the right-hand side of section 7 is
 $[Q_{x_{0},a_{0}}P_{1},Q_{x_{\omega},a_{\omega}}P_{2}]=[P_{1}Q_{x_{0},a_{0}}P_{1},P_{2}Q_{x_{\omega},a_{\omega}}P_{2}]=0$
@@
 Consider now the first summand in the right-hand side of section 7.
-Both the projectors $Q_{x_{0},a_{0}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{1}$ are diagonal in the basis $\mathcal{E}$.
+Both projectors $Q_{x_{0},a_{0}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{1}$ are diagonal in the basis $\mathcal{E}$.
 Hence, they commute, too.
 We deduce that
-(6) $[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0.$
+$[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0.$ (12)
 We now claim that the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ pairwise commute.
-The fact that $Q_{x_{\alpha},d}P_{2},Q_{x_{\alpha}^{\prime},d^{\prime}}P_{2}]=0$ for each $\alpha,\alpha^{\prime}\in[\omega-1],d,d^{\prime}\in A$ is trivial, since $\mathcal{E}$ diagonalises both.
+The fact that $Q_{x_{\alpha},d}P_{2},Q_{x_{\alpha}^{\prime},d^{\prime}}P_{2}]=0$ for each $\alpha,\alpha^{\prime}\in[\omega-1],d,d^{\prime}\in A$ is clear, since $\mathcal{E}$ diagonalises both.
 Moreover,
@@
 where the first and third equalities come from eq.
-4 and 28, while the second holds since the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\omega}}$ are assumed to commute by the minimality of $\operatorname{dist}(x,x^{\prime})$.
-So, the claim is true, and we can simultaneously diagonalise the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ in some o.n.
-basis $\mathcal{F}=(f_{1},\dots,f_{h})$ of $H$.
-For the next step, it will be useful to define the sets $L,M\subseteq[h]$ such that $Q_{x_{0},a_{0}}P_{2}=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}$ and $Q_{x_{\omega},a_{\omega}}P_{2}=\sum_{m\in M}f_{m}f_{m}^{*}$.
+10 and 67, while the second holds since the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\omega}}$ are assumed to commute by condition $(II)$.
+So, the claim is true, and we can simultaneously diagonalise the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ in some orthonormal basis $\mathcal{F}=(f_{1},\dots,f_{h})$ of $H$.
+For the next step, it is useful to define the sets $L,M\subseteq[h]$ such that $Q_{x_{0},a_{0}}P_{2}=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}$ and $Q_{x_{\omega},a_{\omega}}P_{2}=\sum_{m\in M}f_{m}f_{m}^{*}$.
 For $m\in M$, consider also the set $A_{m}=\{i\in[h]:{\langle e_{i},f_{m}\rangle}\neq 0\}$.
@@
 We now choose $\xi_{1}\in A_{\bar{m}}\cap L$ and $\xi_{2}\in A_{\bar{m}}\cap([h]\setminus L)$. We claim that
-(7) $\xi_{i}\in[h]\setminus Z\quad\mbox{ for both }i=1\mbox{ and }i=2.$
+$\xi_{i}\in[h]\setminus Z\quad\mbox{ for both }i=1\mbox{ and }i=2.$ (13)
 Indeed, suppose for the sake of contradiction that $\xi_{i}\in Z$ for some $i\in[2]$. This would imply that $P_{2}e_{\xi_{i}}=0$, whence in turn we could deduce that
@@
 On the other hand, we have
-(8) $0=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}e_{\xi_{2}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{2}}=P_{2}Q_{x_{0},a_{0}}e_{\xi_{2}}=P_{2}e_{\xi_{2}}=e_{\xi_{2}},$
+$0=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}e_{\xi_{2}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{2}}=P_{2}Q_{x_{0},a_{0}}e_{\xi_{2}}=P_{2}e_{\xi_{2}}=e_{\xi_{2}},$ (14)
 where the first equality is true since $\xi_{2}\not\in L$, the second is true by definition of $L$, the third by eq.
-4, the fourth by the observation above, and the fifth by eq.
-7.
+10, the fourth by the observation above, and the fifth by eq.
+13.
 This is a contradiction, so the claimed disequality $a_{0}^{\prime}\neq a_{0}$ must hold.
@@
 Using eq.
-4 and 28, we deduce that
+10 and 67, we deduce that
 $0=Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}P_{2}=Q_{x_{\alpha},c_{1}}P_{2}Q_{x_{\alpha},c_{2}}P_{2},$
@@
 Call such element $a_{\alpha}$.
-Observe now that
+We now make the following claim.
+Claim 68.
+There exists some $a_{\omega}^{\prime}\in A$ such that $a_{\omega}^{\prime}\neq a_{\omega}$ and
+$\left(\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\right)Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0.$ (15)
+The truth of the claim would be enough to conclude the proof of the proposition.
+Indeed, consider the 4-tuple $(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$, where $\mathbf{x}=(x_{0},\dots,x_{\omega})$ and $\mathbf{a}=(a_{0},\dots,a_{\omega})$.
+Observe first that
 $Q_{x_{0},a_{0}}e_{\xi_{1}}=Q_{x_{0},a_{0}}P_{2}e_{\xi_{1}}=e_{\xi_{1}},$
@@
 Moreover, clearly, $Q_{x_{i},a_{i}}Q_{x_{j},a_{j}}\neq 0$ for each $i\neq j\in[\omega-1]$ since $e_{\xi_{1}}$ is an eigenvector of both with eigenvalue $1$.
-Observe now that
+Notice now that
 $Q_{x_{\omega},a_{\omega}}P_{2}f_{\bar{m}}=f_{\bar{m}}$
@@
 Hence, $Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}\neq 0$.
-Note also that the $Q_{x_{0},a_{0}}Q_{x_{\omega},a_{\omega}}=0$ is trivial from the fact that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
-To conclude, we claim that there exists some $a_{\omega}^{\prime}\in A$ such that $a_{\omega}^{\prime}\neq a_{\omega}$ and
-(9) $\left(\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\right)Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0.$
-Note that this would be enough to finish the proof.
-Indeed, it follows from the minimality of $\operatorname{dist}(x,x^{\prime})$ that (9) guarantees that $Q_{x_{\omega},a_{\omega}^{\prime}}$ commutes with $Q_{x_{i},a_{i}}$ for each $0\leq i\leq\omega-1$.
-Suppose, for the sake of contradiction, that the claim is false. Then, the fact that $Q_{x_{\omega}}$ is a PVM would imply that
+Note also that $Q_{x_{0},a_{0}}Q_{x_{\omega},a_{\omega}}\neq 0$, as we know that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
+All in all, the above argument proves conditions $(iii)$ and $(iv)$ of Definition 65, while condition $(i)$ is clear from the initial assumptions and $(ii)$ follows from the choice of $a_{0}^{\prime}$ and from 68.
+Moreover, $(v)$ also follows from 68 by noting that condition $(II)$ of the current proposition guarantees that $Q_{x_{\omega},a_{\omega}^{\prime}}$ commutes with $Q_{x_{i},a_{i}}$ for each $1\leq i\leq\omega-1$ and, thus, in order for (15) to hold it must be the case that $Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0$ for each $i\in\{0,\dots,\omega-1\}$.
+As a consequence, $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$ is a bifurcation for $Q$, as required.
+We are left to prove 68. Suppose, for the sake of contradiction, that the claim is false. Then, the fact that $Q_{x_{\omega}}$ is a PVM would imply that
 $\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}$ $=\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\operatorname{id}=\sum_{d\in A}\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}Q_{x_{\omega},d}$
@@
 $\|\prod_{i\in[\omega-1]}Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}}e_{\xi_{1}}\|<1,$
-a contradiction. Hence, the claim is true.
-Putting all together, we conclude that the vertices
-$\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}$
-form a bifurcation for $Q$ of length $\omega$, and the proof is complete. ∎
+a contradiction. Hence, 68 is true, and the proof of Proposition 66 is complete. ∎
+Having established Proposition 66, we can now prove the main result of this section.
+Proof of Theorem 8.
 […]
-Figure 2. A bifurcation of length $\omega$. Solid lines mean non-orthogonality.
-6. Quantum polymorphisms of odd cycles
-The goal of this section is to prove the following.
+Remark 69.
+Observe that the proofs of both Proposition 66 and Theorem 8 do not make use of the fact that the given quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ satisfies condition $($$\mathbf{QH_{1}}$)).
+Hence, both statements hold for any quantum function $Q\colon X\to_{\operatorname{q}}A$ satisfying $($$\mathbf{QH_{2}}$)).
+As a simple application of Theorem 8, we can prove the following non-contextuality result.
+Proposition 70.
+Let $\mathbb{X},\mathbb{Y}$ be relational structures such that $\operatorname{Gaif}(\mathbb{X})$ and $\operatorname{Gaif}(\mathbb{Y})$ are odd cycles of the same length.
+Then every quantum homomorphism $\mathbb{X}\to_{\operatorname{q}}\mathbb{Y}$ is non-contextual.
+In particular, it follows that quantum endomorphisms of odd cycles are non-contextual.
+Corollary 71.
+For each odd $m\geq 3$, every quantum homomorphism $Q\colon\mathbb{C}_{m}\to_{\operatorname{q}}\mathbb{C}_{m}$ is non-contextual.
+We point out that a different proof of Corollary 71 was obtained in [34] via a result from [4] on the quantum symmetry of vertex-transitive graphs.
+Proof of Proposition 70.
+Suppose first that $\mathbb{X}$ is not essentially binary—i.e., recalling Definition 15, $\mathbb{X}$ contains some relation $R^{\mathbb{X}}$ such that at least one tuple $\mathbf{x}\in R^{\mathbb{X}}$ contains at least three distinct entries—call them $x,y,z$.
+Then the set $\{x,y,z\}\subseteq X$ induces a $\mathbb{K}_{3}$ in $\operatorname{Gaif}(\mathbb{X})$, and the assumption on $\operatorname{Gaif}(\mathbb{X})$ implies that $X$ cannot contain any other vertex.
+Hence, in this case, the fact that each quantum homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ is non-contextual immediately follows from $($$\mathbf{QH_{2}}$)).
+Suppose now that $\mathbb{X}$ is essentially binary and assume, for the sake of contradiction, that $\mathbb{X},\mathbb{Y}$ admit some contextual quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{Y}$.
+Let $m$ be the number of vertices in $\operatorname{Gaif}(\mathbb{X})=\operatorname{Gaif}(\mathbb{Y})$.
+Applying Theorem 8, we deduce that $Q$ has a bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$ of length $\omega\leq\frac{m-1}{2}$.
+Let as usual $\mathbf{x}=(x_{0},\dots,x_{\omega})$ and $\mathbf{a}=(a_{0},\dots,a_{\omega})$, and consider the pairs $(x_{0},a_{0}),(x_{0},a_{0}^{\prime}),(x_{1},a_{1}),(x_{2},a_{2})\in X\times Y$.
+By Definition 65, we know that $Q_{x_{0},a_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{0},a^{\prime}_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$, $Q_{x_{0},a_{0}}Q_{x_{2},a_{2}}\neq 0$, and $Q_{x_{0},a_{0}^{\prime}}Q_{x_{2},a_{2}}\neq 0$.
+Using Proposition 16, we deduce that $\{a_{0},a_{1}\},\{a_{0}^{\prime},a_{1}\}$, and $\{a_{1},a_{2}\}$ are all edges of $\operatorname{Gaif}(\mathbb{Y})$.
+Note also that $x_{0}$ and $x_{2}$ lie on an (m2)(m-2)-walk in $\operatorname{Gaif}(\mathbb{X})$.
+Since no walk of such length connects a vertex to itself in $\operatorname{Gaif}(\mathbb{Y})$Proposition 16 ensures that $a_{0}\neq a_{2}$ and $a_{0}^{\prime}\neq a_{2}$.
+Finally $a_{0}\neq a_{0}^{\prime}$ again by Definition 65.
+We deduce that $a_{1}$ has degree at least three in $\operatorname{Gaif}(\mathbb{Y})$, a contradiction.
+∎
+8 Quantum polymorphisms of the Siggers digraph
+In this section, we investigate the complexity of the quantum CSP parameterised by the Siggers digraph—i.e., the digraph obtained by taking a directed triangle and making one of its directed edges undirected, see Figure 5.
+The significance of this directed graph is that it is the smallest digraph whose classical CSP is NP-complete.
+021
+Figure 5: The Siggers digraph.
+We prove the following result.
+See 10
+Throughout this section, we denote the Siggers digraph by $\mathbb{A}$ (and, as usual, we denote its vertex set $\{0,1,2\}$ by $A$).
+In order to deduce the undecidability of $\operatorname{CSP_{q}}(\mathbb{A})$ from the algebraic machinery developed in the previous sections, we show that all quantum polymorphisms of the Siggers digraph are non-contextual by making use of contextuality bifurcations.
 Theorem 72.
-For each odd $m\geq 3$, it holds that $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$.
-We begin with a simple consequence of Theorem 26, which handles the case of quantum polymorphisms of arity 11—i.e., quantum endomorphisms.
-Corollary 30.
-$\operatorname{qHom}(\mathbb{C}_{m},\mathbb{C}_{m})=\operatorname{\oplus Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ for each odd $m\geq 3$.
+Every quantum polymorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ is non-contextual.
+The key to establishing this result is the analysis of the properties of quantum homomorphisms from the path of length 5, which we shall denote $\mathbb{P}_{5}$, to $\mathbb{A}$.
+Throughout this section, we label the vertices of $\mathbb{P}_{5}$ as follows.
+uabcdv
+We shall make use of the concept of “modular shift” of a PVM, as defined next.
+Definition 73.
+Let $Q=(Q_{i})_{i\in S}$ and $Q^{\prime}=(Q^{\prime}_{i})_{i\in S}$ be two PVMs whose outcome set is labelled as $S=\{0,1,\dots,n-1\}$.
+We say that $Q^{\prime}$ is a $k$-modular shift of $Q$ (for some $k\in S$) if
+$Q^{\prime}_{i}=Q_{i+_{n}k}$
+for each $i\in S$, where $_{n}$” denotes the addition mod $n$. Note that two PVMs $Q$ and $T$ commute if, and only if, any modular shifts $Q^{\prime}$ and $T^{\prime}$ also commute.
+For the Siggers digraph $\mathbb{A}$, the modular shifts of a PVM with outcome set $A$ are constructed by permuting the projectors associated with the vertices along the directed triangle underlying $\mathbb{A}$.
+The main thrust of our argument is captured by the following theorem.
+Theorem 74.
+Let $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and suppose that every modular shift of the PVMs associated with $u$ and $v$ can be extended to a quantum homomorphism from $\mathbb{P}_{5}$ to $\mathbb{A}$.
+Then, $Q_{u,i},Q_{v,j}]=0$ for each $i,j\in A$.
+To prove Theorem 74, it shall be useful to collect a few simple observations about contextuality bifurcations in quantum homomorphisms from $\mathbb{P}_{5}$ to $\mathbb{A}$.
+Henceforth, given a bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{y},p,q)$ of length $\omega$, we shall denote $\mathbf{x}=(x_{0},\dots,x_{\omega})$ and $\mathbf{y}=(y_{0},\dots,y_{\omega})$.
+Moreover, given a tuple $\mathbf{a}=(a_{0},a_{1},\dots,a_{n})$, we let $\mathbf{a}^{R}=(a_{n},a_{n-1},\dots,a_{0})$ be its reversed tuple.
+Lemma 75.
+Let $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and let $\mathcal{B}=(\mathbf{x},\mathbf{y},p,q)$ be a bifurcation of length $\omega$ for $Q$.
+Then there exists a bifurcation $\mathcal{B}^{\prime}=(\mathbf{x}^{\prime},\mathbf{y}^{\prime},p,q)$ of length $\omega^{\prime}$ such that
+  • $\mathbf{x}^{\prime}$ is either a directed path or a reversed path in $\mathbb{P}_{5}$;
+  • $\mathbf{x}^{\prime}$ is a subtuple of $\mathbf{x}$;
+  • $x^{\prime}_{0}=x_{0}$;
+  • $x^{\prime}_{\omega^{\prime}}=x_{\omega}$.
 Proof.
-Suppose, for the sake of contradiction, that $\operatorname{qHom}(\mathbb{C}_{m},\mathbb{C}_{m})\neq\operatorname{\oplus Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$, and let $Q$ be a contextual quantum homomorphism $\mathbb{C}_{m}\to_{q}\mathbb{C}_{m}$.
-Applying Theorem 26, we deduce that there exists a bifurcation for $Q$ of length $\omega\leq\frac{m-1}{2}$.
-Consider, in particular, the initial elements $(x_{0},a_{0}),(x_{0},a_{0}^{\prime}),(x_{1},a_{1}),(x_{2},a_{2})$ of such bifurcation.
-By Definition 25 we know that $Q_{x_{0},a_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{0},a^{\prime}_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$, $Q_{x_{0},a_{0}}Q_{x_{2},a_{2}}\neq 0$, and $Q_{x_{0},a_{0}^{\prime}}Q_{x_{2},a_{2}}\neq 0$.
-Using $($$\mathbf{QH_{1}}$)), we deduce that $(a_{0},a_{1}),(a_{0}^{\prime},a_{1})$, and $(a_{1},a_{2})$ are all edges of $\mathbb{C}_{m}$.
-Note also that $x_{0}$ and $x_{2}$ lie on an (m2)(m-2)-walk in $\mathbb{X}$.
-Since no walk of such length connects a vertex to itself in $\mathbb{C}_{m}$Proposition 27 ensures that $a_{0}\neq a_{2}$ and $a_{0}^{\prime}\neq a_{2}$.
-Finally $a_{0}\neq a_{0}^{\prime}$ again by Definition 25.
-We deduce that $a_{1}$ has degree at least three in $\mathbb{C}_{m}$, a contradiction.
+We first consider an arbitrary bifurcation $\mathcal{B}$ in a quantum homomorphism $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ and show that $\mathbf{x}$ or $\mathbf{x}^{R}$ is a path is $\mathbb{P}_{5}$ when $\mathbf{x}$ contains no repeated vertices.
+To see why, note that $\operatorname{Gaif}(\mathbb{P}_{5})$ is the undirected path of length 5, so any walk $\mathbf{x}$ that does not repeat vertices is a directed subpath in $\operatorname{Gaif}(\mathbb{P}_{5})$.
+If we then consider $\mathbf{x}$, we can see that either $(x_{i},x_{i+1})$ is an edge of $\mathbb{P}_{5}$ for each $0\leq i\leq\omega-1$, or $(x_{i+1},x_{i})$ is an edge for each such $i$.
+It is thus the case that either $\mathbf{x}$ or $\mathbf{x}^{R}$ is a directed walk in $\mathbb{P}_{5}$.
+If $\mathcal{B}$ repeats some vertices, there exist $i<j$ (with $i\neq 0$ or $j\neq\omega$) such that $x_{i}=x_{j}$ but $x_{k}\neq x_{\ell}$ for each $k\in\{0,\dots,i-1\}$ and $\ell\in\{j+1,\dots,\omega\}$.
+We then define $\mathcal{B}^{\prime}=(\mathbf{x}^{\prime},\mathbf{y}^{\prime},p,q)$ by setting
+  • $\mathbf{x}^{\prime}=(x_{0},\dots,x_{i-1},x_{i},x_{j+1},\dots,x_{\omega})$;
+  • $\mathbf{y}^{\prime}=(y_{0},\dots,y_{i-1},y_{i},y_{j+1},\dots,y_{\omega})$.
+It is immediate that $\mathcal{B}^{\prime}$ inherits properties (ii)(v)(ii)-(v) of Definition 65 from $\mathcal{B}$.
+Furthermore, $\mathbf{x}^{\prime}$ is a walk in $\operatorname{Gaif}(\mathbb{P}_{5})$ since the same holds for $\mathbf{x}$.
+The proof is then complete since $\mathbf{x}^{\prime}$ does not repeat any vertices, so either $\mathbf{x}^{\prime}$ or $(\mathbf{x}^{R})$$^{\prime}$ is a directed walk in $\mathbb{P}_{5}$.
 ∎
-We point out that a different proof of Corollary 30 was obtained in [Zeman] via a result from [banica2007quantum] on the quantum symmetry of vertex-transitive graphs.
-In the following, we fix an odd integer $m\in\mathbb{N}$.
-We start with the following observation.
+Lemma 76.
+Let $Q\colon\mathbb{\mathbb{X}}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and let $\mathcal{B}=(\mathbf{x},\mathbf{y},p,q)$ be a bifurcation of length $\omega$ for $Q$.
+Then $\mathcal{B}^{\prime}=(\mathbf{x}^{R},\mathbf{y}^{R},q,p)$ is also a bifurcation for $Q$.
+Proof.
+We show that $\mathcal{B}^{\prime}$ satisfies the conditions in Definition 65.
+For $(i)$, simply recall that $\operatorname{Gaif}(\mathbb{X})$ is an undirected graph; so, if $\mathbf{x}$ is a walk in $\operatorname{Gaif}(\mathbb{X})$, then the same holds for $\mathbf{x}^{R}$.
+$(ii)$ is immediate from $\mathcal{B}$ being a bifurcation.
+$(iii)$, $(iv)$, and $(v)$ are also clear from the fact that, for $A,B$ self-adjoint, $AB=0$ if, and only if, $BA=0$.
+∎
+Consider the following property:
+  • $(\mathbf{BP})$
+    Bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{y},q,p)$ is such that $\mathbf{x}$ is a walk in $\mathbb{P}_{5}$
+Combining Lemma 75 and Lemma 76, we deduce that we can refine any bifurcation $\mathcal{B}$ for some quantum homomorphism $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ into a (possibly different) bifurcation $\mathcal{B}^{\prime}$ for $Q$ that runs between the same $x_{0}$ and $x_{\omega}$, but has property $(\mathbf{BP})$.
+Conversely, if there is no $(\mathbf{BP})$ bifurcation on a given $\mathbb{P}_{5}$ path, then there is no bifurcation between its start and endpoint.
+Lemma 77.
+Let $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and let $\mathcal{B}=(\mathbf{x},\mathbf{y},p,q)$ be a bifurcation of length $\omega$ for $Q$ with property $(\mathbf{BP})$.
+Then
+  • $y_{1}=2$ and $y_{\omega-1}=0$;
+  • both $y_{0}$ and $p$ are in $\{0,1\}$;
+  • both $y_{\omega}$ and $q$ are in $\{1,2\}$.
+In particular, $\omega\geq 3$.
+Proof.
+Consider $Q$ and $\mathcal{B}$ as described.
+By the definition of a bifurcation, $Q_{x_{0},y_{0}}Q_{x_{1},y_{1}}\neq 0$, $Q_{x_{0},p}Q_{x_{1},y_{1}}\neq 0$, and $p\neq y_{0}$.
+Hence, using $($$\mathbf{QH_{1}}$)), $(y_{0},y_{1})$ and $(p,y_{1})$ are edges in $\mathbb{A}$ since $(x_{0},x_{1})$ is an edge in $\mathbb{P}_{5}$ by $(\mathbf{BP})$.
+It follows that $y_{1}=2$, since 2 is the only vertex in $\mathbb{A}$ with in-degree $\geq 2$.
+This also implies that $y_{0},p\in\{0,1\}$ by the edge relation of $\mathbb{A}$.
+Repeating the same argument for $x_{\omega-1}$ (this time considering the out-degree), we obtain that $y_{\omega-1}=0$ and $y_{\omega},q\in\{1,2\}$.
+Finally, the fact that $y_{1}\neq y_{\omega-1}$ implies that $\omega\geq 3$.
+∎
+The following will also be useful.
+Lemma 78.
+For any quantum homomorphism $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ and any two vertices $x,y\in P_{5}$ such that $\operatorname{dist}_{\operatorname{Gaif}{(\mathbb{P}_{5})}}(x,y)\leq 2$, the PVMs $Q_{x}$ and $Q_{y}$ commute.
+Proof.
+Let $d=\operatorname{dist}_{\operatorname{Gaif}{(\mathbb{P}_{5})}}(x,y)$.
+If $d\leq 1$, the fact is clear from $($$\mathbf{QH_{2}}$)).
+Let then $d=2$, and suppose for the sake of contradiction that $Q_{x}$ and $Q_{y}$ do not commute.
+It would follow from Proposition 66 that $Q$ admits a bifurcation running on a path $x,\dots,y$ in $\operatorname{Gaif}(\mathbb{P}_{5})$ which induces a $(\mathbf{BP})$ bifurcation of length $2$ (or 0), thus contradicting Lemma 77.
+∎
+The next step towards proving Theorem 74 is the following result.
+Lemma 79.
+For any quantum homomorphism $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$, the projector $Q_{b,2}$ commutes with every projector of $Q$.
+Proof.
+Choose $\alpha\in P_{5}$ and $i\in A$, and assume, for the sake of contradiction, that $Q_{b,2},Q_{\alpha,i}]\neq 0$.
+By Lemma 78, it must be that $\alpha=v$.
+Observe that the walk $b,c,d,v$ in $\mathbb{P}_{5}$ and the values $2,i\in A$ satisfy the conditions of Proposition 66.
+Therefore, $Q$ admits a $(\mathbf{BP})$ bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{y},p,q)$ such that $y_{0}=2$.
+By Lemma 77, this is impossible.
+∎
+We can thus consider any quantum homomorphism $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$, and split it into the direct sum $S\oplus T$, where, for each $x\in P_{5}$ and each $i\in A$,
+$S_{x,i}$ $=Q_{b,2}Q_{x,i}\quad\mbox{and}$
+$T_{x,i}$ $=(Q_{b,0}+Q_{b,1})Q_{x,i}=(\operatorname{id}-Q_{b,2})Q_{x,i}.$
+(Note that, formally, the projector, say, $S_{x,i}=Q_{b,2}Q_{x,i}$ acts on the subspace given by the range of $Q_{b,2}$.
+We will leave this implicit to avoid adding extra notation for injections and projections onto such space.)
+Claim 80.
+$S_{u,1}=0$.
+Proof.
+Using Lemma 79, we can expand $S_{u,1}$ as
+$S_{u,1}=Q_{b,2}Q_{u,1}=Q_{u,1}Q_{b,2}=\sum_{i\in A}Q_{u,1}Q_{a,i}Q_{b,2}=0+0+0=0,$
+where the penultimate equality follows from applying $($$\mathbf{QH_{1}}$)) to the products $Q_{u,1}Q_{a,0}$, $Q_{u,1}Q_{a,1}$, and $Q_{a,2}Q_{b,2}$. ∎
+Claim 81.
+$T_{v,0}T_{u,2}=0$.
+Proof.
+Using Lemma 79, we can expand the product as
+$T_{v,0}T_{u,2}$ $=(\operatorname{id}-Q_{b,2})Q_{v,0}(\operatorname{id}-Q_{b,2})Q_{u,2}=Q_{v,0}(\operatorname{id}-Q_{b,2})^{2}Q_{u,2}=Q_{v,0}(\operatorname{id}-Q_{b,2})Q_{u,2}$
+$=Q_{v,0}(Q_{b,0}+Q_{b,1})Q_{u,2}=Q_{v,0}Q_{b,0}Q_{u,2}+Q_{v,0}Q_{b,1}Q_{u,2}.$
+We now argue that both terms of the above sum are zero. First,
+$Q_{b,0}Q_{u,2}=\sum_{i\in A}Q_{b,0}Q_{u,2}Q_{a,i}=Q_{b,0}Q_{u,2}Q_{a,0}=Q_{b,0}Q_{a,0}Q_{u,2}=Q_{a,0}Q_{b,0}Q_{u,2}=0,$
+where the third and fourth equalities follow from $($$\mathbf{QH_{2}}$)); therefore, $Q_{v,0}Q_{b,0}Q_{u,2}=0$. Second, using again $($$\mathbf{QH_{2}}$)),
+$Q_{v,0}Q_{b,1}$ $=\sum_{i,j\in A}Q_{d,i}Q_{v,0}Q_{b,1}Q_{c,j}=Q_{d,2}Q_{v,0}Q_{b,1}Q_{c,2}=Q_{v,0}Q_{d,2}Q_{c,2}Q_{b,1}$
+$=Q_{v,0}Q_{c,2}Q_{d,2}Q_{b,1}=0,$
+so $Q_{v,0}Q_{b,1}Q_{u,2}=0$. It follows that $T_{v,0}T_{u,2}=0$, as required. ∎
+The main conclusion that we draw from these observations is that both $S$ and $T$ satisfy a specific orthogonality property: Letting $W\in\{S,T\}$, it holds that $W_{v,0}W_{u,i}=0$ for some $i\in A$.
+In particular, there exists some modular shift $W^{\prime}_{u}$ of $W_{u}$ such that $W_{v,0}W^{\prime}_{u,0}=0$.
+We now proceed to prove the following.
+Proposition 82.
+Let $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and assume that $Q_{u,0}Q_{v,0}=0$. Then, $Q_{\alpha,i},Q_{v,0}]=0$ for each $\alpha\in P_{5}$ and each $i\in A$.
+Proof.
+First of all, using $($$\mathbf{QH_{1}}$)) we note that
+$Q_{d,2}=Q_{d,2}\sum_{i\in A}Q_{v,i}=Q_{d,2}Q_{v,0}=\sum_{i\in A}Q_{d,i}Q_{v,0}=Q_{v,0}.$
+Hence, it is sufficient to prove that $Q_{\alpha,i},Q_{d,2}]=0$.
+By Lemma 78, $Q_{\alpha},Q_{d,2}]=0$ for all $\alpha\in\{b,c,d,v\}$.
+Therefore, we only need to consider the cases of $\alpha=u$ and $\alpha=a$.
+Observe that
+$Q_{a,1}Q_{d,2}$ $=Q_{a,1}Q_{v,0}=\sum_{i\in A}Q_{u,i}Q_{a,1}Q_{v,0}$
+$=Q_{u,0}Q_{a,1}Q_{v,0}=Q_{a,1}Q_{u,0}Q_{v,0}=0,$
+where the last equality is true by the assumption of this proposition.
+Again, from $($$\mathbf{QH_{1}}$)), we find $Q_{a,2}=Q_{b,0}$, which commutes with $Q_{d,2}$ by Lemma 78.
+Thus, the PVM $Q_{a}$ commutes with $Q_{d,2}$, since $Q_{a,0}=\operatorname{id}-Q_{a,1}-Q_{a,2}$.
+The only remaining case is when $\alpha$ equals $u$.
+$Q_{u}$ commutes with $Q_{d,2}$ since $Q_{u,0}Q_{d,2}=Q_{u,0}Q_{v,0}=0$ and $Q_{u,2}=Q_{a,0}$, which commutes with $Q_{v,0}$.
+The proof is then concluded since $Q_{u,1}=\operatorname{id}-Q_{u,0}-Q_{u,2}$ and so, $Q_{u},Q_{v,0}]=0$.∎
+Using Proposition 82, we can now prove Theorem 74, conditioned on $Q_{v,0}Q_{u,0}=0$.
+Proposition 83.
+Let $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism such that $Q_{v,0}Q_{u,0}=0$, and suppose that every modular shift of the PVMs associated with $u$ and $v$ can be extended to a quantum homomorphism from $\mathbb{P}_{5}$ to $\mathbb{A}$.
+Then, $Q_{u,i},Q_{v,j}]=0$ for each $i,j\in A$.
+Proof.
+Since $Q_{v,0}$ commutes with every projector in $Q$ (from Proposition 82), we can decompose $Q$ as $S\oplus T$ as above, by defining for each $x\in P_{5}$ and each $i\in A$
+$S_{x,i}$ $=Q_{v,0}Q_{x,i}\quad\mbox{and}$
+$T_{x,i}$ $=(\operatorname{id}-Q_{v,0})Q_{x,i}.$
+The argument then follows by showing that, in each component, the $u$- and $v$-indexed PVMs commute when their modular shifts can be extended as stated in the proposition.
+Firstly, $S_{u,i}$ always commutes with $S_{v,i}$ (since $S_{v,0}=\operatorname{id}$), so the implication vacuously holds.
+Now we focus on $T$.
+Assign $T_{u,i}=U_{i}$ and $T_{v,i}=V_{i}$, and assume that every assignment $T^{\prime}\colon\{u,v\}\to_{\operatorname{q}}A$ defined by the modular shifts $T^{\prime}_{u,i}=U_{i+_{3}k}$, $T^{\prime}_{v,i}=V_{i+_{3}l}$ can be completed to a quantum homomorphism $T^{\prime}\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ for every $k,l\in A$.
+Remember that $V_{0}=(\operatorname{id}-Q_{v,0})Q_{v,0}=0$.
+To show that $U_{k}$ must commute with $V_{1}$ and $V_{2}$, consider the modular shifts defined by $T^{\prime}_{u,i}=U_{i+_{3}(2-k)}$ and $T^{\prime}_{v,i}=V_{i+_{3}1}$.
+By our assumption, $T^{\prime}_{u}$ and $T^{\prime}_{v}$ can be extended to a quantum homomorphism $T^{\prime}\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$.
+We will show that there can be no bifurcation between $U_{k}$ and any $v$-indexed projector in this homomorphism.
+Firstly, note that $U_{k}=T^{\prime}_{u,2}$, while $T^{\prime}_{u,2}=T^{\prime}_{a,0}$ using $($$\mathbf{QH_{1}}$)).
+Hence, we consider the path $a,b,c,d,v$ and show $T^{\prime}_{a,0}$ commutes with $T^{\prime}_{v}$.
+By Lemma 78, the PVMs $T^{\prime}_{b}$, $T^{\prime}_{c}$, and $T^{\prime}_{d}$ commute with each other since $\operatorname{dist}_{\operatorname{Gaif}{(\mathbb{P}_{5})}}(x,y)\leq 2$ for all $x,y\in\{b,c,d\}$.
+Moreover, $T^{\prime}_{a}$, $T^{\prime}_{b}$, and $T^{\prime}_{c}$ commute by Lemma 78, and
+$T^{\prime}_{d,1}=T^{\prime}_{d,1}\sum_{i\in A}T^{\prime}_{v,i}=0+0+T^{\prime}_{d,1}T^{\prime}_{v,2}=0$
+since $T^{\prime}_{v,2}=V_{0}=0$.
+By Lemma 77 and Proposition 66, $T^{\prime}_{a,y_{0}}T^{\prime}_{d,1}\neq 0$ for any bifurcation along $a,b,c,d$.
+We conclude that there is no such bifurcation by the fact that $T^{\prime}_{d,1}=0$.
+Thus, there is no $(\mathbf{BP})$ bifurcation between $T^{\prime}_{a}$ and $T^{\prime}_{d}$ and, so, they commute by Proposition 66.
+Similarly, $T^{\prime}_{c}$, $T^{\prime}_{d}$, and $T^{\prime}_{v}$ commute by Lemma 78, and $T^{\prime}_{v,2}=V_{0}=0$.
+Thus, $T^{\prime}_{b}$ and $T^{\prime}_{v}$ commute by the same argument.
+In summary, we have established that $T^{\prime}_{i}$ and $T^{\prime}_{j}$ commute for all $i,j\in\{a,b,c,d,v\}$ with $\{i,j\}\neq\{a,v\}$.
+Lemma 77 shows that there cannot be a bifurcation using $a,b,c,d,v$ since $T^{\prime}_{v,2}=0$.
+Hence, by Proposition 66, there is no $(\mathbf{BP})$ bifurcation using the walk $u,a,b,c,d,v$ and thus $T^{\prime}_{u,2},T^{\prime}_{v,1}]=0$.
+Observe now that $T^{\prime}_{v,2}=0$ implies $T^{\prime}_{v,0}=\operatorname{id}-T^{\prime}_{v,1}$.
+Therefore, $T^{\prime}_{u,2}$ commutes with the entire PVM $T^{\prime}_{v}$, and the proof is complete.
+∎
+We can now prove the full statement of Theorem 74.
+Proof of Theorem 74.
+Consider any quantum homomorphism $Q\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ such that the modular shifts of the $u$ and $v$-indexed PVMs extend.
+Using Lemma 79, we can decompose $Q=S\oplus T$ and by this construction, the modular shifts of the $u$ and $v$ PVMs in $S$ and $T$ can also be extended to quantum homomorphisms.
+Letting $V\in\{S,T\}$, we know by 80 and 81 that there is some $W\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ where
+  1. 1.
+    $W$ is an extension of some modular shifts of $V_{u}$ and $V_{v}$ (so, $W_{u}$ and $W_{v}$ are modular shifts of $V_{u}$ and $V_{v}$) and
+  2. 2.
+    $W_{v,0}W_{u,0}=0$.
+Any modular shifts $W^{\prime}_{u}$ and $W^{\prime}_{v}$ of $W_{u}$ and $W_{v}$ are, in particular, modular shifts of $V_{u}$ and $V_{v}$.
+Thus, they can be extended to a quantum homomorphism $W^{\prime}\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ and we can apply Proposition 83 to $W$.
+We conclude that $W_{u}$ and $W_{v}$ commute.
+Thus, $V_{u}$ and $V_{v}$ commute for $V\in\{S,T\}$, and it follows that $Q_{u}$ and $Q_{v}$ commute.
+∎
+Now consider any quantum polymorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$, and fix two PVMs $Q_{u}$ and $Q_{v}$, associated with arbitrary elements $u,v\in A^{n}$.
+We show that for any modular shifts $Q^{\prime}_{u}$ and $Q^{\prime}_{v}$, there exists an appropriate restriction $\overline{Q}\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ of $Q$ that extends $Q^{\prime}_{u}$ and $Q^{\prime}_{v}$.
+Proposition 84.
+For any quantum polymorphism $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ and $x\in A^{n}$, there exist $y$, $z$ in $A^{n}$ such that $Q_{y}$ and $Q_{z}$ are modular shifts of $Q_{x}$.
+Proof.
+Take some $x\in A^{n}$ and let $y=x+_{3}\{1\}^{n}$ and $z=x+_{3}\{2\}^{n}$, where $_{3}$ is the componentwise addition modulo 3 and $\{i\}^{n}$ is the $n$-tuple that with $i$ in each entry.
+Then the directed triangle $\mathbb{T}_{x}=(\{x,y,z\},\{(x,y),(y,z),(z,x)\})$ is a subgraph of $\mathbb{A}$, so every $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ restricts to $\overline{Q}\colon\mathbb{T}_{x}\to_{\operatorname{q}}\mathbb{A}$.
+We now show that, in any quantum homomorphism $S\colon\mathbb{T}_{x}\to_{\operatorname{q}}\mathbb{A}$, $S_{y}$ and $S_{z}$ are the non-trivial modular shifts of $S_{x}$. To see this, note that
+$S_{x,1}$ $=\sum_{i\in A}S_{z,i}S_{x,1}\sum_{i\in A}S_{y,i}=S_{z,0}S_{x,1}S_{y,2},$
+$S_{y,2}$ $=S_{y,2}\sum_{i\in A}S_{z,i}=S_{y,2}S_{z,0}=\sum_{i\in A}S_{y,i}S_{z,0}=Q_{z,0},$
+and
+$S_{y,2}S_{z,0}$ $=S_{z,0}S_{y,2}=S_{z,0}\sum_{i\in A}S_{x,i}S_{y,2}=S_{z,0}S_{x,1}S_{y,2},$
+from which it follows that $S_{x,1}=S_{y,2}=S_{z,0}$.
+To conclude, we simply repeat the argument for $S_{y,1}$ and $S_{z,1}$ and notice that $S_{x}$, cannot have more than 2 distinct, non-trivial, modular shifts (since $|A|=3$).
+∎
+Proof of Theorem 72.
+We begin by noting that one can find a path of length 5 between any two vertices in $\mathbb{A}$. Thus, one can find a path of length 5 between any two vertices in $\mathbb{A}^{n}$.
+Fix any $u,v\in A^{n}$ and $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$; we consider the PVMs $Q_{u}$ and $Q_{v}$.
+We can find $a,b,c,d\in A^{n}$ such that $u,a,b,c,d,v$ is a walk in $\mathbb{A}^{n}$ thus, we can restrict $Q$ to $\overline{Q}\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ and obtain a quantum homomorphism such that $\overline{Q}_{u}=Q_{u}$ and $\overline{Q}_{v}=Q_{v}$.
+Now consider any modular shifts $Q^{\prime}_{u}$ and $Q^{\prime}_{v}$ of $Q_{u}$ and $Q_{v}$, respectively.
+By Proposition 84, there exist $u^{\prime},v^{\prime}\in A^{n}$ such that $Q^{\prime}_{u}=Q_{u^{\prime}}$ and $Q^{\prime}_{v}=Q_{v^{\prime}}$.
+Hence, we can, again, restrict $Q$ to find $\overline{Q^{\prime}}\colon\mathbb{P}_{5}\to_{\operatorname{q}}\mathbb{A}$ such that $Q_{u^{\prime}}=\overline{Q^{\prime}}_{u}$ and $Q_{v^{\prime}}=\overline{Q^{\prime}}_{v}$.
+An application of Theorem 74 then yields that $Q_{u}$ and $Q_{v}$ commute for every $Q\colon\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ and $u,v\in A^{n}$, which means that $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
+∎
+We can finally conclude the proof of the main result of this section.
+Proof of Theorem 10.
+It is known that $\mathbb{A}$ pp-constructs 3SAT [7].
+Moreover, all quantum polymorphisms of $\mathbb{A}$ are non-contextual by Theorem 72.
+It follows from Corollary 56 that $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+∎
+9 Quantum polymorphisms of odd cycles
+The goal of this section is to prove that the quantum CSP parameterised by an odd cycle $\mathbb{C}_{m}$ is undecidable.
+This resolves an open question raised in [34].
+See 9 To prove this result, we will show that all quantum polymorphisms of odd cycles are non-contextual—and, thus, odd cycles admit commutativity gadgets.
+Theorem 85.
+$\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$ for each odd $m\geq 3$.
+We start with the following observation. Henceforth in this section, we fix an odd integer $m\in\mathbb{N}$.
 Lemma 86.
@@
 Proof.
-If $x=z$ the result is trivial so we may assume that $x\neq z$.
-We know from Corollary 30 that there exists $d\in\mathbb{N}$ and $h_{1},\dots h_{d}\in\operatorname{Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ such that $Q=\bigoplus_{i\in[d]}h_{i}$.
+If $x=z$ the result is trivial, so we may assume that $x\neq z$.
+We know from Corollary 71 that there exist $d\in\mathbb{N}$ and $h_{1},\dots,h_{d}\in\operatorname{Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ such that $Q=\bigoplus_{i\in[d]}h_{i}$.
 Using eq.
 1, we can write
-$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\left(\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right)\left(\sum_{\begin{subarray}{c}j\in[d]\\ h_{j}(y)=b\end{subarray}}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\right)\left(\sum_{\begin{subarray}{c}k\in[d]\\ h_{k}(z)=c\end{subarray}}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)$
-$=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}.$
+$Q_{x,a}Q_{y,b}Q_{z,c}$ $=\left(\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{*}\right)\left(\sum_{\begin{subarray}{c}j\in[d]\\ h_{j}(y)=b\end{subarray}}\mathbf{e}_{j}\mathbf{e}_{j}^{*}\right)\left(\sum_{\begin{subarray}{c}k\in[d]\\ h_{k}(z)=c\end{subarray}}\mathbf{e}_{k}\mathbf{e}_{k}^{*}\right)=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{*}.$
 Let $i\in[d]$ give a non-trivial contribution to the sum above.
@@
 It follows that
-$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}=Q_{y,b}Q_{z,c},$
+$Q_{x,a}Q_{y,b}Q_{z,c}$ $=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{*}=Q_{y,b}Q_{z,c},$
 as needed. ∎
-Next, we focus on quantum polymorphisms of arity at least two.
+Observe that the non-contextuality of unary quantum polymorphisms of odd cycles—i.e., quantum homomorphisms $\mathbb{C}_{m}\to_{\operatorname{q}}\mathbb{C}_{m}$—is clear from Corollary 71.
+Hence, in order to prove Theorem 85, we need to extend non-contextuality to higher arities.
 Henceforth in this section, we fix some arity $n\geq 2$.
@@
 $0$ $=Q_{\mathbf{y}+\mathbf{v}}Q_{\mathbf{x}}=\left(\operatorname{id}-\sum_{i\in\mathbb{Z}_{m}\setminus\{1\}}Q_{\mathbf{y}+i\mathbf{v}}\right)Q_{\mathbf{x}}=Q_{\mathbf{x}}-(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}.$
-Hence, $Q_{\mathbf{x}}=(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}$. Since projectors are Hermitian, it follows that
+Hence, $Q_{\mathbf{x}}=(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}$. Since projectors are self-adjoint, it follows that
 $[Q_{\mathbf{x}},Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}}]=0.$
@@
 $[Q_{\mathbf{x}},Q_{\mathbf{y}}]=[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}].$
-We now apply the same argument after applying the automorphism of $\mathbb{C}_{m}^{n}$ given by $\mathbf{a}\mapsto\mathbf{a}+\mathbf{y}-\mathbf{x}$. We obtain
+We now use the same argument after applying the automorphism of $\mathbb{C}_{m}^{n}$ given by $\mathbf{a}\mapsto\mathbf{a}+\mathbf{y}-\mathbf{x}$. We obtain
 $[Q_{\mathbf{x}},Q_{\mathbf{y}}]=[Q_{\mathbf{x}},Q_{2\mathbf{x}-\mathbf{y}}]$ $=[Q_{3\mathbf{x}-2\mathbf{y}},Q_{2\mathbf{x}-\mathbf{y}}]=[Q_{3\mathbf{x}-2\mathbf{y}},Q_{4\mathbf{x}-3\mathbf{y}}]=\dots$
@@
 $=(Q_{\mathbf{x},a}Q_{\mathbf{y},a+2})^{*}=Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}.$
-In particular, we deduce that $Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}$ is a Hermitian matrix, so
+In particular, we deduce that $Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}$ is self-adjoint, so
 $[Q_{\mathbf{x},a},Q_{\mathbf{y},a+2}]=0,$
 as needed. ∎
-We can now finalise the proof of the main result of this section.
+We can now finalise the proof of Theorem 85.
 Proof of Theorem 85.
@@
 Hence, the inductive hypothesis allows us to freely permute the corresponding projectors in the expression above, thus yielding
-(10) $Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$
+$Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$ (16)
 Multiplying both terms by $Q_{\mathbf{x},a}$ on the right and further rearranging gives
 $Q_{\mathbf{x},a}Q_{\mathbf{y},b}Q_{\mathbf{x},a}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}Q_{\mathbf{x},a}$
-$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}$
+$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$
 Observe now that $\mathbf{x}$, $\mathbf{z}_{1}$, and $\mathbf{z}_{2}$ lie on a common $m$-cycle within $\mathbb{C}_{m}^{n}$ since $\mathbf{z}_{1}-\mathbf{x}=\mathbf{z}_{2}-\mathbf{z}_{1}$.
@@
 $Q_{\mathbf{x},a}Q_{\mathbf{y},b}Q_{\mathbf{x},a}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}=Q_{\mathbf{x},a}Q_{\mathbf{y},b},$
-where the second equality is eq. 10. From this, we conclude that $Q_{\mathbf{x},a}$ and $Q_{\mathbf{y},b}$ commute, as required.
-Corollary 34.
-Let $m\geq 3$ be an odd natural number.
-Then $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$.
-In particular, $\mathbb{C}_{m}$ has a commutativity gadget and $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable.
-Proof.
-$\mathbb{C}_{m}$ has a commutativity gadget by Theorem 29 and Theorem 16.
-Note that $\mathbb{K}_{m}$ is pp-definable in $\mathbb{C}_{m}$ with the gadget consisting of a simple path of length m2m-2.
-Since $\mathbb{C}_{m}$ has a commutativity gadget, it follows that $\mathbb{K}_{m}$ is q-definable in $\mathbb{C}_{m}$ by Proposition 13.
-By Proposition 10, $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ reduces to $\operatorname{CSP^{*}}(\mathbb{C}_{m})$, and $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable by Corollary 24.
-Thus, $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable.
+where the second equality is eq. 16. From this, we conclude that $Q_{\mathbf{x},a}$ and $Q_{\mathbf{y},b}$ commute, as required. ∎
+The undecidability of $\operatorname{CSP_{q}}(\mathbb{C}_{m})$ is now a simple consequence of the machinery developed in Section 4 and Section 5.
+Proof of Theorem 9.
+$\mathbb{K}_{m}$ is pp-definable in $\mathbb{C}_{m}$ with the gadget consisting of a path of length m2m-2; hence, $\operatorname{Pol}(\mathbb{C}_{m})\to\operatorname{Pol}(\mathbb{K}_{m})\to\operatorname{Pol}(\operatorname{3SAT})$, and the result follows by combining Theorem 85 and Corollary 56.
 ∎
-We conclude with the proof that quantum homomorphisms between different odd cycles can be contextual.
+We conclude this section by proving that quantum homomorphisms between different odd cycles can be contextual.
 Proposition 89.
-There is a quantum homomorphism $\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space that is not in $\oplus{\operatorname{Hom}(\mathbb{C}_{7},\mathbb{C}_{5})}$.
+There is a quantum homomorphism $\mathbb{C}_{7}\to_{\operatorname{q}}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space that is not in $\oplus{\operatorname{Hom}(\mathbb{C}_{7},\mathbb{C}_{5})}$.
 Proof.
@@
 Consider the two classical endomorphisms $h,g\colon\mathbb{C}_{5}\to\mathbb{C}_{5}$ such that $h(2)=6,h(5)=0$ and $g(5)=0,g(2)=1$.
-Then $h\oplus g$ is a quantum homomorphism $Q\colon\mathbb{C}_{5}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space.
+Then $h\oplus g$ is a quantum homomorphism $Q\colon\mathbb{C}_{5}\to_{\operatorname{q}}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space.
 Extend $Q$ by setting $Q_{3,a}=Q_{5,a}$ for $a\in\{0,1,2,5,6\}$.
 Choose a basis in which all the projectors are diagonal, and in this basis define $Q_{4,1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$ and $Q_{4,6}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$.
-Then $Q\colon\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ and $Q_{4,0},Q_{2,1}]\neq 0$.
+Then $Q\colon\mathbb{C}_{7}\to_{\operatorname{q}}\mathbb{C}_{5}$ and $Q_{4,1},Q_{2,1}]\neq 0$ since $Q_{2,1}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$.
 ∎
-7. The non-oracular setting
-In this section, we consider the non-oracular version of the results in Section 3.
-A non-oracular quantum homomorphism between two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, denoted by $Q\colon\mathbb{A}\operatorname{\to_{qno}}\mathbb{B}$, is a quantum function $Q\colon A\to_{q}B$ satisfying the condition $($$\mathbf{QH_{1}}$)) but possibly not $($$\mathbf{QH_{2}}$)).
-The following is the non-oracular version of the entangled CSP.
-Definition 36 (Non-oracular entangled CSP).
+10 Non-oracular polymorphisms
+In this section, we consider the non-oracular version of the results described in Section 4 and Section 5.
+A non-oracular quantum homomorphism between two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, denoted by $Q\colon\mathbb{A}\operatorname{\to_{\operatorname{qno}}}\mathbb{B}$, is a quantum function $Q\colon A\to_{\operatorname{q}}B$ satisfying the condition $($$\mathbf{QH_{1}}$)) but possibly not $($$\mathbf{QH_{2}}$)); see Section 3.2.
+The following is the non-oracular version of the quantum CSP.
+Definition 90 (Non-oracular quantum CSP).
 Fix a $\sigma$-structure $\mathbb{A}$.
-$\operatorname{CSP^{**}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a non-oracular quantum homomorphism $\mathbb{X}\operatorname{\to_{qno}}\mathbb{A}$.
-We now define quantum polymorphisms in the non-oracular setting.
+$\operatorname{CSP_{q}^{no}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a non-oracular quantum homomorphism $\mathbb{X}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
+For CSPs parameterised by digraphs, non-oracular quantum homomorphisms capture perfect strategies for the variant of the 2-player CSP game where both cooperating players receive some variable ($x_{A}$ and $x_{B}$, respectively) from the verifier and reply with assignments to their variables with some values $a_{A}$ and $a_{B}$, respectively.
+The answers are deemed correct by the verifier if $a_{A}=a_{B}$ whenever $x_{A}=x_{B}$ (i.e., the game is synchronous), and $(a_{A},a_{B})$ is an edge in the template digraph whenever $(x_{A},x_{B})$ is an edge in the instance.
+The relation between this game and the type of quantum homomorphisms given above was first described in [59] (in the case of undirected graphs) as a generalisation of the quantum colouring game of [23], see also [58].
+Various other works have considered this setting for specific types of binary constraint systems; see, e.g., the non-oracular strategies for Unique Projection Games [51] considered in [50].
+We next define quantum polymorphisms in the non-oracular setting.
 Definition 91.
 Let $\mathbb{A}$ be a relational structure.
-A non-oracular quantum polymorphism is a non-oracular quantum homomorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$.
+A non-oracular quantum polymorphism is a non-oracular quantum homomorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
 We denote by $\operatorname{qPol^{no}}(\mathbb{A})$ the set of non-oracular quantum polymorphisms of $\mathbb{A}$.
-We will obtain a quantisation of the Galois connection between classical polymorphisms and relations. The following is the main result of this section.
+As the main result of this section, we shall prove that, just like in the oracular case Theorem 5, the Galois connection between classical polymorphism clones and pp-definitions [15, 38] can be lifted to a Galois connection for non-oracular quantum CSPs.
+As a consequence, we will prove the following result.
 Theorem 92.
@@
 $\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B}).$
-Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
+Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
+It should be observed that, in this case, we obtain a reduction from the oracular version of the quantum CSP parameterised by $\mathbb{B}$ to the non-oracular version of the quantum CSP parameterised by $\mathbb{A}$.
+As we shall see, this is a consequence of how commutativity in the contexts is handled by the reduction.
 We begin by defining the non-oracular version of q-definitions.
@@
 Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$.
-Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$.
-A non-oracular q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
+Let $R$ be the $r$-ary symbol of the signature of $(A;S)$ (i.e., $R^{(A;S)}=S$), and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$.
+A non-oracular q-definition (in short, a q-no-definition) for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
   1. $(\mathbf{noq_{1}})$
-    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
+    For every quantum homomorphism $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;S)$, there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
   2. $(\mathbf{noq_{2}})$
-    For every non-oracular quantum homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}_{S}$.
+    For every non-oracular quantum homomorphism $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}(A;S)$.
 Note that, in the first item of the definition above, $Q$ is required to be oracular (i.e., to satisfy $($$\mathbf{QH_{2}}$))), while $Q^{\prime}$ need not.
@@
 Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures.
-We say that $\mathbb{B}$ is q-no-definable from $\mathbb{A}$ if there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
+We say that $\mathbb{A}$ q-no-defines $\mathbb{B}$ if there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
 In the following, it shall be useful to observe that Lemma 17, Lemma 24Lemma 19, and Lemma 21 also hold in the non-oracular setting.
 Furthermore, the definition of quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ (Definition 18) does not change in the non-oracular setting.
-The following result is the key to the quantisation of the relation/operation Galois connection in the non-oracular quantum setting.
+The next result is the non-oracular version of Theorem 40, and is the key to the quantisation of the relation/operation Galois connection in the non-oracular quantum setting.
 Theorem 95.
-Let $\mathbb{A}$ be a finite structure and let $S\subseteq A^{r}$ for some $r\in\mathbb{N}$.
-Then there exists a non-oracular q-definition for $\mathbb{A}$ and $S$ if, and only if, $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{A}_{S})$.
+Let $\mathbb{A}$ be a relational structure and let $S\subseteq A^{r}$ for some $r\in\mathbb{N}$.
+Then there exists a non-oracular q-definition for $\mathbb{A}$ and $S$ if, and only if, $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}((A;S))$.
 Proof.
-Henceforth in this proof, we shall denote $\mathbb{A}_{S}$ by $\mathbb{D}$ for the sake of readability.
+Henceforth in this proof, we shall denote $(A;S)$ by $\mathbb{D}$ for the sake of readability.
   • $(\Rightarrow)$
     Let $\mathbb{G}$ be a non-oracular q-definition for $\mathbb{A}$ and $S$.
-This means that every quantum homomorphism $\mathbb{R}\to_{q}\mathbb{D}$ extends to a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, and every non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ restricts to a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{D}$.
-    Fix now a non-oracular quantum polymorphism $Q$ of $\mathbb{A}$; i.e., $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$.
+This means that every quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}\mathbb{D}$ extends to a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$, and every non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ restricts to a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}\mathbb{D}$.
+    Fix now a non-oracular quantum polymorphism $Q$ of $\mathbb{A}$; i.e., $Q\colon\mathbb{A}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
 We want to prove that $Q\in\operatorname{qPol^{no}}(\mathbb{D})$.
-Note that $Q$ is a quantum function $A^{n}\to_{q}A$, so we only need to prove $($$\mathbf{QH_{1}}$)) applied to the relation of $\mathbb{D}^{n}$; i.e., to the relation $R^{\mathbb{D}^{n}}$ containing all $r$-tuples $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})$ where each $\mathbf{b}^{(i)}$ is an element of $A^{n}$, and for each $j\in[n]$ it holds that $(b^{(1)}_{j},\dots,b^{(r)}_{j})\in R^{\mathbb{D}}=S$.
+Note that $Q$ is a quantum function $A^{n}\to_{\operatorname{q}}A$, so we only need to prove $($$\mathbf{QH_{1}}$)) applied to the relation of $\mathbb{D}^{n}$; i.e., to the relation $R^{\mathbb{D}^{n}}$ containing all $r$-tuples $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})$ where each $\mathbf{b}^{(i)}$ is an element of $A^{n}$, and for each $j\in[n]$ it holds that $(b^{(1)}_{j},\dots,b^{(r)}_{j})\in R^{\mathbb{D}}=S$.
     Fix then such a tuple $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})\in R^{\mathbb{D}^{n}}$, and choose a tuple $\mathbf{c}\in A^{r}\setminus S$.
@@
 In particular, $f_{i}$ is a quantum homomorphism.
-Hence, ($(\mathbf{noq_{1}})$) implies that there exists a non-oracular quantum homomorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $f_{i}$.
+Hence, ($(\mathbf{noq_{1}})$) implies that there exists a non-oracular quantum homomorphism $W^{(i)}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ extending $f_{i}$.
 Consider now the tensor product
-    $W=\bigotimes_{i\in[n]}W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}^{n},$
-    and compose it with $Q$ to obtain $Q\bullet W:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$. Fix $\ell\in[r]$ and $a\in A$. Expanding the composition and the tensor product yields
+    $W=\bigotimes_{i\in[n]}W^{(i)}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}^{n},$
+    and compose it with $Q$ to obtain $Q\bullet W\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
+Fix $\ell\in[r]$ and $a\in A$.
+Expanding the composition and the tensor product yields
     $(Q\bullet W)_{g_{\ell},a}=\sum_{\mathbf{a}\in A^{n}}Q_{\mathbf{a},a}\otimes W_{g_{\ell},\mathbf{a}}=\sum_{\mathbf{a}\in A^{n}}Q_{\mathbf{a},a}\otimes\bigotimes_{i\in[n]}W^{(i)}_{g_{\ell},a_{i}}$.
@@
 We show that the two conditions in Items $(\mathbf{noq_{1}})$ and $(\mathbf{noq_{2}})$ hold.
-    Let $Q\colon\mathbb{R}\to_{q}\mathbb{D}$ be a quantum homomorphism over some Hilbert space $H$, and let $d$ be the dimension of $H$.
+    Let $Q\colon\mathbb{R}\to_{\operatorname{q}}\mathbb{D}$ be a quantum homomorphism over some Hilbert space $H$, and let $d$ be the dimension of $H$.
 Since $Q$ satisfies $($$\mathbf{QH_{2}}$)), $Q$ must be classical, so by the spectral theorem we can decompose it as $Q=\bigoplus_{i\in[d]}h_{i}$, where $h_{i}\colon\mathbb{R}\to\mathbb{D}$ is a deterministic homomorphism for each $i\in[d]$.
@@
 In other words, there exists a function $f:[d]\to[n]$ satisfying
-    (11) $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{(f(i))}$
+    $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{(f(i))}$ (17)
     for each $i\in[d]$.
-Consider now the deterministic dictator polymorphisms $\pi^{n}_{j}:\mathbb{A}^{n}\to\mathbb{A}$ defined by $\pi^{n}_{j}(\mathbf{a})=a_{j}$ for each $j\in[n]$ and each $\mathbf{a}\in A^{n}$.
+Consider now the classical projection polymorphisms $\pi^{n}_{j}:\mathbb{A}^{n}\to\mathbb{A}$ defined by $\pi^{n}_{j}(\mathbf{a})=a_{j}$ for each $j\in[n]$ and each $\mathbf{a}\in A^{n}$.
 Define
     $P=\bigoplus_{i\in[d]}\pi^{n}_{f(i)}.$
-    Note that $P$ is a classical (a fortiori, non-oracular quantum) homomorphism from $\mathbb{A}^{n}$ to $\mathbb{A}$. Take now $\ell\in[r]$ and $a\in A$, and observe that
+    Note that $P$ is a classical (so, a fortiori, non-oracular quantum) homomorphism from $\mathbb{A}^{n}$ to $\mathbb{A}$. Take now $\ell\in[r]$ and $a\in A$, and observe that
     $P_{g_{\ell},a}=\bigoplus_{i\in[d]}(\pi^{n}_{f(i)})_{g_{\ell},a}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ \pi^{n}_{f(i)}(g_{\ell})=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ a^{f(i)}_{\ell}=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\\ h_{i}(\ell)=a\end{subarray}}\operatorname{id}=\bigoplus_{\begin{subarray}{c}i\in[d]\end{subarray}}(h_{i})_{\ell,a}=Q_{\ell,a}$
     (where the fourth equality follows from Equation 17). Hence, $P$ witnesses that ($(\mathbf{noq_{1}})$) holds.
-    Finally, let $Q\colon\mathbb{\mathbb{A}}^{n}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism, and consider the quantum function $Q^{\prime}$ given by $Q^{\prime}_{\ell,a}=Q_{g_{\ell},a}$ for $\ell\in[r]$ and $a\in A$.
-Recall now that each $\mathbf{a}^{(i)}$ is an element of $S=R^{\mathbb{D}}$ and, thus, the tuples $(g_{1},\dots,g_{r})$ belongs to $R^{\mathbb{D}^{n}}$ by the definition of categorical power of structures.
+    Finally, let $Q\colon\mathbb{\mathbb{A}}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ be a non-oracular quantum polymorphism, and consider the quantum function $Q^{\prime}$ given by $Q^{\prime}_{\ell,a}=Q_{g_{\ell},a}$ for $\ell\in[r]$ and $a\in A$.
+Recall now that each $\mathbf{a}^{(i)}$ is an element of $S=R^{\mathbb{D}}$ and, thus, the tuple $(g_{1},\dots,g_{r})$ belongs to $R^{\mathbb{D}^{n}}$ by the definition of categorical power of structures.
 Since $Q$ is a non-oracular quantum polymorphism of $\mathbb{A}$, by assumption it is also an $n$-ary non-oracular quantum polymorphism of $\mathbb{D}$.
@@
 The following result is an immediate consequence of Theorem 95.
-Corollary 42.
-Two structures $\mathbb{A},\mathbb{B}$ satisfy $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ if and only if $\mathbb{B}$ is q-no-definable from $\mathbb{A}$.
-Unlike in the oracular setting, non-oracular q-gadgets are not strong enough to guarantee reductions.
+See 13
+Unlike in the oracular setting, non-oracular q-gadgets alone are not strong enough to guarantee reductions.
 To that end, they need to be strengthened via a non-oracular commutativity gadget, which we now define.
@@
   1. $(\mathbf{noc_{1}})$
-    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum non-oracular homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $Q$.
+    For any quantum function $Q\colon\{u,v\}\to_{\operatorname{q}}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum non-oracular homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ extending $Q$.
   2. $(\mathbf{noc_{2}})$
-    For every quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
+    For every quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
+Note that, unlike in the oracular case Definition 47, a non-oracular commutativity gadget is not the same as a q-no-definition for $A^{2}$ in $\mathbb{A}$.
 We now give the non-oracular analogue of Theorem 51.
-Just like in the oracular setting, the characterisation of commutativity gadgets via quantum polymorphisms (stated in the result below) also holds for the weaker algebraic non-oracular commutativity gadgets of [Zeman], which are defined by replacing ($(\mathbf{noc_{1}})$) in Definition 43 with the weaker condition stated below:
+Just like in the oracular setting, the characterisation of commutativity gadgets via quantum polymorphisms (stated in the result below) also holds for the weaker algebraic non-oracular commutativity gadgets of [34], which are defined by replacing ($(\mathbf{noc_{1}})$) in Definition 96 with the weaker condition stated below:
   1. $(\mathbf{noc_{1}^{\prime}})$
-    For every $a,b\in A$, there exists a quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
+    For every $a,b\in A$, there exists a quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
 Theorem 97.
-Let $\mathbb{A}$ be a finite structure. The following are equivalent:
-  1. (1)
+Let $\mathbb{A}$ be a relational structure. The following are equivalent:
+  1. 1.
     $\mathbb{A}$ has a non-oracular commutativity gadget.
-  2. (2)
+  2. 2.
     $\mathbb{A}$ has an algebraic non-oracular commutativity gadget.
-  3. (3)
+  3. 3.
     $\mathbb{A}^{n}$ is an (algebraic) non-oracular commutativity gadget for all large enough $n\in\mathbb{N}$.
-  4. (4)
+  4. 4.
     $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
@@
   • $(4.\Rightarrow 3.)$
-    The proof follows from exactly the same argument as for Proposition 20.
+    The proof follows exactly the same argument as for Proposition 53.
   • $(3.\Rightarrow 1.)$
@@
   • $(2.\Rightarrow 4.)$
-    Take an algebraic non-oracular commutativity gadget $\mathbb{G}$ with distinguished elements $u,v$, a non-oracular quantum polymorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$, two tuples $\mathbf{a},\mathbf{b}\in A^{n}$, and two elements $c,d\in A$.
+    Take an algebraic non-oracular commutativity gadget $\mathbb{G}$ with distinguished elements $u,v$, a non-oracular quantum polymorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$, two tuples $\mathbf{a},\mathbf{b}\in A^{n}$, and two elements $c,d\in A$.
 We proceed as in Proposition 54.
-We first build a non-oracular quantum polymorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ for each $i\in[n]$ with the property that $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}$, using ($(\mathbf{noc_{1}^{\prime}})$).
-Then, we take the tensor product $W=\bigotimes_{i\in[n]}W^{(i)}$ of such homomorphisms, and we define $U:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ as the composition of $Q$ and $W$.
+We first build a non-oracular quantum polymorphism $W^{(i)}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ for each $i\in[n]$ with the property that $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}$, using ($(\mathbf{noc_{1}^{\prime}})$).
+Then, we take the tensor product $W=\bigotimes_{i\in[n]}W^{(i)}$ of such homomorphisms, and we define $U\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ as the composition of $Q$ and $W$.
 On the one hand, ($(\mathbf{noq_{2}})$) guarantees that $U_{u,c},U_{v,d}]=0$.
@@
 Combining the two facts, we conclude that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we deduce that $Q$ is classical.∎
-The next result shows that non-oracular q-definitions, combined with non-oracular commutativity gadgets, give logspace reductions from $\operatorname{CSP^{*}}$s to $\operatorname{CSP^{**}}$s.
+The next result shows that non-oracular q-definitions, combined with non-oracular commutativity gadgets, give logspace reductions from $\operatorname{CSP_{q}}$ to $\operatorname{CSP_{q}^{no}}$.
 Proposition 98.
@@
   • $\mathbb{A}$ has a non-oracular commutativity gadget, and
-  • for every $R\in\rho$, there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$.
-Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
+  • $\mathbb{A}$ q-no-defines $\mathbb{B}$.
+Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
 Proof.
@@
 Let also $\mathbb{H}$ be a non-oracular commutativity gadget for $\mathbb{A}$ with its distinguished vertices $u,v$.
-Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$.
-Recall that $\operatorname{Gaif}(\mathbb{X})$ denotes the Gaifman digraph of $\mathbb{X}$.
-We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{**}}(\mathbb{A})$ as follows.
+Let $\mathbb{X}$ be an instance of $\operatorname{CSP_{q}}(\mathbb{B})$.
+Henceforth in this proof, we shall view $\operatorname{Gaif}(\mathbb{X})$ as the directed graph obtained by viewing each edge $\{x,y\}$ as two directed edges $(x,y)$ and $(y,x)$.
+We define an instance $\mathbb{Y}$ of $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ as follows.
 Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$, as well as a copy $\mathbb{H}^{(x,y)}$ of $\mathbb{H}$ for each edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$.
@@
 Call this instance $\mathbb{Y}$.
-(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism.
-For every symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$.
+(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{B}$ is a quantum homomorphism.
+For every symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=(A;R^{\mathbb{B}})$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$.
 Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 93.
-Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to\mathbb{A}^{\prime}$.
-(Observe that this would not be guaranteed if $Q$ was a non-oracular quantum homomorphism.) Then, ($(\mathbf{noq_{1}})$) guarantees the existence of a non-oracular quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\operatorname{\to_{qno}}\mathbb{A}$ such that
+Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to_{\operatorname{q}}\mathbb{A}^{\prime}$.
+(Observe that this would not be guaranteed if $Q$ was a non-oracular quantum homomorphism.) Then, ($(\mathbf{noq_{1}})$) guarantees the existence of a non-oracular quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that
 $Q^{(\mathbf{x},R)}_{g_{i}^{(\mathbf{x})},a}=Q_{x_{i},a}$
@@
 Hence, we can apply ($(\mathbf{noc_{1}})$) to find a non-oracular quantum homomorphism
-$Q^{(x,y)}\colon\mathbb{H}^{(x,y)}\operatorname{\to_{qno}}\mathbb{A}$
-extending $Q$ in $x$ and $y$. Overall, this is a non-oracular quantum homomorphism $\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$.
-(Soundness.) Suppose that $Q\colon\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$ is a non-oracular quantum homomorphism.
+$Q^{(x,y)}\colon\mathbb{H}^{(x,y)}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$
+extending $Q$ in $x$ and $y$. Overall, this is a non-oracular quantum homomorphism $\mathbb{Y}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
+(Soundness.) Suppose that $Q\colon\mathbb{Y}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ is a non-oracular quantum homomorphism.
 Consider the quantum function $Q^{\prime}:X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$.
@@
 Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)).
-We now apply ($(\mathbf{noq_{2}})$) to deduce that the quantum function $U$ defined by $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}=Q^{\prime}_{x_{i},a}$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$).
+We now apply ($(\mathbf{noq_{2}})$) to deduce that the quantum function $U$ defined by $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}=Q^{\prime}_{x_{i},a}$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}^{\prime}$, where, as before, $\mathbb{A}^{\prime}=(A;R^{\mathbb{B}})$.
 Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$.
 Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
-(12) $\prod_{i\in[r]}Q_{x_{i},b_{i}}=0.$
+$\prod_{i\in[r]}Q_{x_{i},b_{i}}=0.$ (18)
 (Note that for this to be true we do not need $U$ to be oracular.) This means that $Q^{\prime}$ satisfies $($$\mathbf{QH_{1}}$)) and, thus, it is indeed a quantum (oracular) homomorphism from $\mathbb{X}$ to $\mathbb{B}$.
 ∎
-We can finally complete the proof of the main result of this section.
+We can finally complete the proof of Theorem 92.
 Proof of Theorem 92.
 Since $\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})$Theorem 97 guarantees that $\mathbb{A}$ has a non-oracular commutativity gadget.
-Moreover, we know from $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ and Corollary 42 that all the relations of $\mathbb{B}$ have a q-no-definition in $\mathbb{A}$.
+Moreover, we know from $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ and Theorem 13 that $\mathbb{A}$ q-no-defines $\mathbb{B}$.
 Note also that $\mathbb{A}$ and $\mathbb{B}$ have the same domain since $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$.
@@
 Let $\mathbb{T}$ be a tree structure, and let $x,y$ be elements of $\mathbb{T}$.
-Let $Q\colon\mathbb{T}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism.
+Let $Q\colon\mathbb{T}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ be a non-oracular quantum homomorphism.
 If $Q_{x,a}Q_{y,b}\neq 0$, there exists a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ such that $h(x)=a$ and $h(y)=b$.
@@
 We prove this by induction on the size of $\mathbb{T}$.
-Assume first that $\mathbb{T}$ consists of a single constraint $(z_{1},\dots,z_{r})\in R^{\mathbb{T}}$ with $x=z_{i}$ and $y=z_{j}$ for some $i,j\in\{1,\dots,r\}$.
+Assume first that $\mathbb{T}$ consists of a single constraint $(z_{1},\dots,z_{r})\in R^{\mathbb{T}}$ with $x=z_{i}$ and $y=z_{j}$ for some $i,j\in[r]$.
 For the argument below, we can assume without loss of generality that $i<j$ up to swapping $x$ and $y$.
@@
 and since $Q_{x,a}Q_{y,b}\neq 0$, it must be that there exists $(c_{1},\dots,c_{r})\in R^{\mathbb{A}}$ such that $c_{i}=a$ and $c_{j}=b$.
-We obtain a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ by setting $h(z_{k})=c_{k}$ for all $k\in\{1,\dots,r\}$.
+We obtain a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ by setting $h(z_{k})=c_{k}$ for all $k\in[r]$.
 Otherwise, $\mathbb{T}$ consists of the union of some trees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$ that share a single vertex $z$.
 Since $\sum_{c}Q_{x,a}Q_{z,c}Q_{z,c}Q_{y,b}=Q_{x,a}(\sum_{c}Q_{z,c})Q_{y,b}=Q_{x,a}Q_{y,b}\neq 0$, there exists a $c$ such that $Q_{x,a}Q_{z,c}\neq 0$ and $Q_{z,c}Q_{y,b}\neq 0$.
-By induction hypothesis on the subtrees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$, there exist homomorphisms $h_{i}\colon\mathbb{T}_{i}\to\mathbb{A}$ such that $h_{i}(z)=c$ for all $i$, and such that if $x\in T_{i}$ (resp.
-$y\in\mathbb{T}_{i}$) then $h_{i}(x)=a$ (resp.
+By the induction hypothesis on the subtrees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$, there exist homomorphisms $h_{i}\colon\mathbb{T}_{i}\to\mathbb{A}$ such that $h_{i}(z)=c$ for all $i$, and such that if $x\in T_{i}$ (resp.
+$y\in T_{i}$) then $h_{i}(x)=a$ (resp.
 $h_{i}(y)=b$).
@@
 Proof.
-Suppose that $Q\colon\mathbb{R}\to_{q}(A;R)$ is a quantum homomorphism.
-Then we have $Q=h_{1}\oplus\dots\oplus h_{d}$, where each $h_{i}$ is a map $\{1,\dots,r\}\to A$ such that $(h_{i}(1),h_{i}(2))\in R$.
+Suppose that $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;R)$ is a quantum homomorphism.
+Then we have $Q=h_{1}\oplus\dots\oplus h_{d}$, where each $h_{i}$ is a map $r]\to A$ such that $(h_{i}(1),h_{i}(2))\in R$.
 Thus, each $h_{i}$ can be extended to a homomorphism $g_{i}\colon\mathbb{G}\to\mathbb{A}$ since $\mathbb{G}$ is a classical gadget for $R$.
-It follows that $Q^{\prime}:=g_{1}\oplus\dots\oplus g_{d}$ is a (non-oracular) quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ satisfying the conclusion of ($(\mathbf{noq_{1}})$).
-To prove ($(\mathbf{noq_{2}})$), let $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism.
+It follows that $Q^{\prime}:=g_{1}\oplus\dots\oplus g_{d}$ is a (non-oracular) quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ satisfying ($(\mathbf{noq_{1}})$).
+To prove ($(\mathbf{noq_{2}})$), let $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ be a non-oracular quantum homomorphism.
 Let $x_{1},x_{2}$ be the distinguished vertices of $\mathbb{G}$.
@@
 Since $\mathbb{G}$ is a tree, there exists by Lemma 99 a homomorphism $h\colon\mathbb{G}\to\mathbb{A}$ such that $h(x_{i})=a_{i}$ for $i\in\{1,2\}$.
-Thus, $(a_{1},a_{2})\in R$, and therefore the quantum function $Q^{\prime}\colon\{1,2\}\to A$ defined by $Q^{\prime}_{i,a}=Q_{x_{i},a}$ satisfies $($$\mathbf{QH_{1}}$)) and is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}(A;R)$.
+Thus, $(a_{1},a_{2})\in R$, and therefore the quantum function $Q^{\prime}\colon\{1,2\}\to_{\operatorname{q}}A$ defined by $Q^{\prime}_{i,a}=Q_{x_{i},a}$ satisfies $($$\mathbf{QH_{1}}$)) and is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}(A;R)$.
 ∎
-We conclude this section by showing that also the non-oracular version of the CSP parameterised by odd cycles is undecidable. The result is proved in same way as its oracular counterpart.
-Corollary 48.
-For every odd $m\geq 3$, $\operatorname{CSP^{**}}(\mathbb{C}_{m})$ is undecidable.
+The next result is the non-oracular version of Proposition 43.
+Proposition 101.
+Let $\mathbb{A},\mathbb{B}$ be structures, and let $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common non-oracular q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively.
+Then every non-oracular quantum homomorphism $Q\colon\mathbb{A}\operatorname{\to_{\operatorname{qno}}}\mathbb{B}$ is a non-oracular quantum homomorphism $(A;R_{A})\operatorname{\to_{\operatorname{qno}}}(B;R_{B})$.
 Proof.
-Simply observe that the proof of Theorem 29 does not make use of the fact that quantum polymorphisms of $\mathbb{C}_{m}$ are oracular.
-Hence, the same argument establishes that $\operatorname{qPol^{no}}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$ for each odd $m\geq 3$.
-Then, the result follows from Theorem 38 applied to $\mathbb{A}=\mathbb{B}=\mathbb{C}_{m}$ and Corollary 34.
+Let $g_{1},\dots,g_{r}$ be the distinguished vertices of $\mathbb{G}$.
+We only need to prove $($$\mathbf{QH_{1}}$)).
+Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$.
+By ($(\mathbf{noq_{1}})$), there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in[r]$.
+Let $S=Q\bullet Q^{\prime}$, which is a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{B}$.
+By ($(\mathbf{noq_{2}})$), we have that the quantum function $Q^{\prime\prime}$ defined by $Q^{\prime\prime}_{i,b}:=S_{g_{i},b}$ for all $i\in[r]$ and all $b\in B$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}(B;R_{B})$.
+Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$.
+Note that $S_{g_{i},b_{i}}$ is by definition $\sum_{a\in A}Q_{a,b_{i}}\otimes Q^{\prime}_{g_{i},a}=Q_{a_{i},b_{i}}\otimes\operatorname{id}$.
+Thus, $(\prod_{i=1}^{r}Q_{a_{i},b_{i}})\otimes\operatorname{id}=0$, from which we obtain that $\prod_{i=1}^{r}Q_{a_{i},b_{i}}=0$.
 ∎
-8. Quantum polymorphisms of Boolean structures
-In this section, we provide some results regarding the quantum polymorphisms of Boolean structures.
+Remark 102.
+Let $Q\colon\mathbb{X}\operatorname{\to_{\operatorname{qno}}}\mathbb{Y}$ be a non-oracular quantum homomorphism between two digraphs $\mathbb{X}$ and $\mathbb{Y}$.
+Suppose that $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ are such that there exists a directed $\ell$-walk from $x$ to $x^{\prime}$ in $\mathbb{X}$ but there is no directed $\ell$-walk from $y$ to $y^{\prime}$ in $\mathbb{Y}$.
+Consider the structure $\mathbb{G}$ consisting of a path of length $\ell$, and call $u,v$ the two endpoints of this path.
+Since this structure is a tree, Proposition 100 implies that $(\mathbb{G},u,v)$ is a q-no-definition for certain binary relations $R_{X}\subseteq X^{2},R_{Y}\subseteq Y^{2}$ in $\mathbb{X}$ and $\mathbb{Y}$, respectively.
+It follows from Proposition 101 that $Q$ is a non-oracular quantum homomorphism $(X;R_{X})\operatorname{\to_{\operatorname{qno}}}(Y;R_{Y})$.
+By assumption, $R_{X}$ contains $(x,x^{\prime})$ and $R_{Y}$ does not contain $(y,y^{\prime})$, and therefore $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$ by $($$\mathbf{QH_{1}}$)).
+Hence, Proposition 100 provides an alternative proof of [59, Lemma 4.12].
+We now show that our undecidability result for odd cycles, Theorem 9, also holds in the non-oracular setting, by virtue of a result from [34].
+Proposition 103.
+For every odd $m\geq 3$, $\operatorname{CSP_{q}^{no}}(\mathbb{C}_{m})$ is undecidable.
+Proof.
+It was observed in [34, Proposition 5.27 & Lemma 4.9] that, if an undirected graph $\mathbb{A}$ having no 4-cycle admits a commutativity gadget, then it also admits a non-oracular commutativity gadget.
+(The result is phrased therein in terms of algebraic commutativity gadgets; however, as noted in Theorem 51 and Theorem 97, this is equivalent to our definitions.) Hence, Theorem 85 directly implies that $\mathbb{C}_{m}$ has a non-oracular commutativity gadget.
+Note now that $\mathbb{C}_{m}$ trivially pp-defines itself via a tree gadget.
+Hence, Proposition 100 implies that $\mathbb{C}_{m}$ q-no-defines itself.
+Using Proposition 98, we deduce that $\operatorname{CSP_{q}}(\mathbb{C}_{m})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{C}_{m})$ in logspace.
+Hence, it follows from Theorem 9 that $\operatorname{CSP_{q}^{no}}(\mathbb{C}_{m})$ is undecidable.
+∎
+We conclude this section by observing that the notion of q-no-definitions can be extended to a notion of q-no-constructions in a similar way as in the oracular case considered in Section 5.1.
+Definition 104.
+Let $\mathbb{A}$ be a $\sigma$-structure and let $\mathbb{B}$ be a $\rho$-struture. We say that $\mathbb{A}$ q-no-constructs $\mathbb{B}$ if there exists a $\rho$-structure $\mathbb{C}$ such that
+  • $(i)$
+    the domain of $\mathbb{C}$ is $A^{d}$ for some $d\in\mathbb{N}$;
+  • $(ii)$
+    each relation of $\mathbb{C}$ (of arity, say, $r$) is q-no-definable in $\mathbb{A}$ (identifying $C^{r}=(A^{d})^{r}$ with $A^{dr}$);
+  • $(iii)$
+    $\mathbb{B}\to_{\operatorname{q}}\mathbb{C}$ and $\mathbb{C}\to_{\operatorname{q}}\mathbb{B}$.
+Note that, in part $(iii)$, we require $\mathbb{B}$ and $\mathbb{C}$ to be quantum oracular homomorphically equivalent.
+As we shall see, non-oracular homomorphic equivalence is not sufficient in order to obtain a q-no-construction version of Proposition 98.
+Proposition 105.
+Let $\mathbb{A}$ and $\mathbb{B}$ be structures (with potentially different domains and signatures), and suppose that
+  • $\mathbb{A}$ has a non-oracular commutativity gadget, and
+  • $\mathbb{A}$ q-no-constructs $\mathbb{B}$.
+Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
+Proof.
+Let $\sigma$ and $\rho$ be the signatures of $\mathbb{A}$ and $\mathbb{B}$, respectively, and let $\mathbb{C}$ be the $\rho$-structure with domain $A^{d}$ witnessing that $\mathbb{A}$ q-no-constructs $\mathbb{B}$, as per Definition 104.
+We now construct a structure $\mathbb{D}$ in the same way as in the proof of Theorem 4 in Section 5.1.
+Given a tuple $\mathbf{a}=((a_{1,1},\dots,a_{1,d}),\dots,(a_{r,1},\dots,a_{r,d}))\in(A^{d})^{r}$, let $\bar{\mathbf{a}}\in A^{dr}$ be the tuple defined by
+$\bar{\mathbf{a}}=(a_{1,1},\dots,a_{1,d},\dots,a_{r,1},\dots,a_{r,d}).$
+Consider the signature $\tau$ having the same symbols as $\rho$ such that, if a symbol has arity $r$ in $\rho$, it has arity $dr$ in $\tau$.
+We let $\mathbb{D}$ be the $\tau$-structure with domain $D=A$ and relations defined as follows: For any $R\in\rho$, $R^{\mathbb{D}}=\{\bar{\mathbf{c}}:\mathbf{c}\in R^{\mathbb{C}}\}$.
+Using part $(ii)$ of Definition 104, we know that $\mathbb{A}$ q-no-defines $\mathbb{D}$.
+Since, by assumption, $\mathbb{A}$ has a non-oracular commutativity gadget, we deduce from Proposition 98 that $\operatorname{CSP_{q}}(\mathbb{D})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
+On the other hand, we know from the proof of Theorem 4 that $\operatorname{qPol}(\mathbb{D})\to\operatorname{qPol}(\mathbb{C})$ and thus, by Theorem 3, $\operatorname{CSP_{q}}(\mathbb{C})$ reduces to $\operatorname{CSP_{q}}(\mathbb{D})$ in logspace.
+In addition, by part $(iii)$ of Definition 104, $\operatorname{CSP_{q}}(\mathbb{B})$ and $\operatorname{CSP_{q}}(\mathbb{C})$ are trivially interreducible in logspace.
+Combining the three reductions yields the required logspace reduction from $\operatorname{CSP_{q}}(\mathbb{B})$ to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$.
+∎
+Remark 106.
+We point out that the Galois connection for q-no-definitions of Theorem 13 does not appear to lift to q-no-constructions.
+In particular, Theorem 4 does not immediately extends to the non-oracular setting.
+The reason is that the order we choose while identifying the elements of $A^{dr}$ with elements of $C^{r}$ (see the proof of Proposition 105) is now important, as in non-oracular homomorphisms we are not allowed to permute projectors associated with the same constraint.
+This, in turns, prevents from building a minion homomorphism $\operatorname{qPol^{no}}(\mathbb{D})\to\operatorname{qPol^{no}}(\mathbb{C})$, which is needed in order to conclude that q-no-constructions yield minion homomorphisms between the minions of quantum non-oracular polymorphisms.
+11 Quantum polymorphisms of Boolean structures
+In this section, we consider the case of quantum polymorphisms of Boolean structures, and we prove the following result.
+See 11
+As a direct consequence of this result, we obtain a $\mathsf{P}$ vs.
+undecidable complexity dichotomy for quantum CSPs parameterised by Boolean structures.
+We let $\mathbb{XOR}$ be the Boolean structure having two ternary relation symbols $R_{+}$ and $R_{-}$, whose interpretations are
+$R_{+}^{\mathbb{XOR}}$ $=\{(1,0,0),(0,1,0),(0,0,1),(1,1,1)\},$
+$R_{-}^{\mathbb{XOR}}$ $=\{(0,1,1),(1,0,1),(1,1,0),(0,0,0)\}.$
+Note that $\operatorname{CSP}(\mathbb{XOR})$ encodes the satisfiability problem for systems of linear equations mod $2$.
+See 12
+Proof.
+If $\mathbb{A}$ does not pp-define $\mathbb{XOR}$, it was shown in [2] that $\operatorname{CSP_{q}}(\mathbb{A})=\operatorname{CSP}(\mathbb{A})$, and the latter problem is solvable in polynomial time via the bounded-width algorithm [8].
+Suppose now that $\mathbb{A}$ $p$-defines $\mathbb{XOR}$.
+This implies that $\operatorname{Pol}(\mathbb{A})\subseteq\operatorname{Pol}(\mathbb{XOR})$.
+Since $\mathbb{XOR}$ is not preserved by majority, it follows that $\mathbb{A}$ has no majority polymorphism.
+Then, it follows from Theorem 11 that $\mathbb{A}$ admits a commutativity gadget.
+Applying Proposition 49, we deduce that $\mathbb{A}$ q-defines $\mathbb{XOR}$.
+This implies via Proposition 42 that $\operatorname{CSP_{q}}(\mathbb{XOR})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace.
+Since $\operatorname{CSP_{q}}(\mathbb{XOR})$ is undecidable (as proved in [67]), we conclude that $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable.
+∎
+Remark 107.
+Corollary 12 should be compared with [63, Theorem 5.11 (a)].
+The result established therein concerns Boolean constraint systems with contexts, where each context is a subset of variables of the CSP instance whose associated projectors are required to commute, see [63, Definition 3.1].
+By modifying the contexts in the reductions used to prove undecidability, classical pp-definitions can be employed to obtain sound and complete reductions between such problems.
+Syntactically, this can be enforced by requiring that the language admits a full binary relation, see for example [2, §6.1].
+To extend the result to arbitrary Boolean languages, the use of commutativity gadgets appears to be necessary; see also [33].
+On the other hand, we point out that the dichotomy for Boolean constraint systems with contexts in [63] is incomparable to ours, in that it also addresses other variants of quantum satisfiability.
+In contrast, Corollary 12 applies only to finite-dimensional quantum strategies in the tensor-product model (see also the discussion in Section 2.8).
+11.1 The case of structures without majority polymorphism
+In this subsection, we study the quantum polymorphism minions of Boolean structures that are not preserved under majority.
+As a consequence, we give an alternative proof of the fact, originally proved in [33, Section 7], that any Boolean relational structure that is not invariant under the majority polymorphism has a commutativity gadget.
 If a quantum homomorphism $Q$ has range $\{0,1\}$, then $Q_{x,0}$ is determined by $Q_{x,1}$ since $Q_{x,0}=\operatorname{id}-Q_{x,1}$.
-Moreover, the set $\{0,1\}^{n}$ is naturally in bijection with the powerset of $\{1,\dots,n\}$.
-Given $S\subseteq[n]$, we let $s\in\{0,1\}^{n}$ be the tuple such that $s_{i}=1$ iff $i\in S$.
-In the following, if $\mathbb{A}$ is a Boolean structure, we write the projectors of a quantum homomorphism $\mathbb{A}^{n}\to_{q}\mathbb{A}$ by $Q_{S}$ where $S\subseteq[n]$.
-8.1. The case of structures without a majority polymorphism
-In this subsection we show that any Boolean relational structure which is not invariant under the majority polymorphism has a commutativity gadget.
-This specific statement has previously been proven in [culf_re-completeness_2025, Section 7], our contribution is to simplify the proof by using the characterization in Theorem 44 to assert the non-existence of commutativity gadgets through the study of quantum polymorphisms.
-This allows us to make use of standard CSP methods in the proof.
-We focus our study on the following class of Boolean relational structures:
+Moreover, the set $\{0,1\}^{n}$ is naturally in bijection with the powerset of $[n]$.
+Given $S\subseteq[n]$, we let $s\in\{0,1\}^{n}$ be the tuple such that $s_{i}=1$ if, and only if, $i\in S$.
+In the following, if $\mathbb{A}$ is a Boolean structure, we write the projectors of a quantum homomorphism $\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ by $Q_{S}$ where $S\subseteq[n]$.
+We begin our analysis by focussing on the following class of Boolean structures.
 Definition 108.
-Let $t\in\{0,1\}^{k}$ be an arbitrary Boolean tuple. Then the $t$-translate of 11-in-$k$ is the relational structure $\mathbb{O}_{t}=(\{0,1\};\{x\oplus_{2}t\mid x\in R_{1/k}\})$, where
-  1. (1)
+Let $t\in\{0,1\}^{k}$ be an arbitrary Boolean tuple. Then the $t$-translate of 11-in-$k$ is the relational structure $\mathbb{O}_{t}=(\{0,1\};\{x+_{2}t\mid x\in R_{1/k}\})$, where
+  1. 1.
     $R_{1/k}=\{x\in\{0,1\}^{k}\mid\sum_{i=1}^{k}x_{i}=1\}$ is the relation of 11-in-$k$ SAT and,
-  2. (2)
-    $\oplus_{2}$ is componentwise addition modulo 2.
+  2. 2.
+    $_{2}$ is componentwise addition modulo 2.
 The relation of $\mathbb{O}_{t}$ is named $R_{t}$.
-One can easily see that the ternary majority function on $\{0,1\}$ does not preserve $R_{t}$; the majority of any 3 distinct tuples in the relation yields $t\not\in R_{t}$, witnessing the non-preservation.
+One can easily see that the ternary majority function on $\{0,1\}$ does not preserve $R_{t}$; the majority of any three distinct tuples in the relation yields some $t$ that is not in $R_{t}$, witnessing the non-preservation.
 We show that if we translate 11-in-$k$ by a tuple $t\in R_{1/k}$, then $\operatorname{\oplus Pol}(\mathbb{O}_{t})=\operatorname{qPol^{no}}(\mathbb{O}_{t})$.
@@
 Proposition 109.
-For every non-oracular polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$, the following properties hold:
-  1. (1)
+For every non-oracular polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{(1,0,0)}$, the following properties hold:
+  1. 1.
     If $S\cap T=\emptyset$, then $Q_{S}Q_{T}=0$.
-  2. (2)
+  2. 2.
     If $S\cap T=\emptyset$, then $Q_{S\cup T}=Q_{S\cup T}(Q_{S}+Q_{T})$.
-  3. (3)
+  3. 3.
     $Q_{S}=Q_{S\cup T}Q_{S}$ for all $S$
-  4. (4)
+  4. 4.
     $Q_{S}=\sum_{i\in S}Q_{\{i\}}$.
@@
 we can derive (4) by inductively building any $Q_{S}$ from sums of the individual elements of $S$. ∎
-In particular, it follows that in every non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$ and every $S,T\subseteq[n]$, we have $Q_{S},Q_{T}]=0$, which implies that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})$ is the quantum closure of $\operatorname{Pol}(\mathbb{O}_{(1,0,0)})$.
+In particular, it follows that for every non-oracular quantum polymorphism
+$Q\colon\mathbb{O}_{(1,0,0)}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{(1,0,0)}$
+and every $S,T\subseteq[n]$, we have $Q_{S},Q_{T}]=0$, which implies that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})$ is the quantum closure of $\operatorname{Pol}(\mathbb{O}_{(1,0,0)})$.
 Lemma 110.
@@
 Proof.
-Since $(0,0,\dots,0)=(1,0,\dots,0)\oplus_{2}(1,0,\dots,0)\in R_{(1,0,\dots,0)}$ and $(1,1,\dots,1)\notin R_{(1,0,\dots,0)}$, we can pp-define 0 in $\mathbb{O}_{(1,0,\dots,0)}$ by
+Since $(0,0,\dots,0)=(1,0,\dots,0)+_{2}(1,0,\dots,0)\in R_{(1,0,\dots,0)}$ and $(1,1,\dots,1)\notin R_{(1,0,\dots,0)}$, we can pp-define 0 in $\mathbb{O}_{(1,0,\dots,0)}$ by
 $a=0\Leftrightarrow R_{(1,0,\dots,0)}(a,\dots,a)$
@@
 Then $(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{0},\dots,\mathbf{0})$ is in the $n$th power $R_{(1,0,\dots,0)}$ and $(x,y,z,0,\dots,0)$ is not an element of $R_{(1,0,\dots,0)}$
-Now assume we have some non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to_{q}\mathbb{O}_{(1,0,\dots,0)}$ and consider:
+Now let $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{(1,0,\dots,0)}$ be a non-oracular quantum polymorphism and consider:
 $Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}$ $=Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}\sum_{w\in\{0,1\}}Q_{\mathbf{0},w}$
@@
 $=\sum_{w\in\{0,1\}}(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},w}\dots Q_{\mathbf{0},w})(Q_{\mathbf{0},w}\dots Q_{\mathbf{0},w})$
-$=(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})(Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})$
+$=(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})$
+$+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})(Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})$
 $=0(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})0=0$
@@
 Lemma 112.
-Assume $k\geq 3$, and neither $t$ nor $\overline{t}$ ($(1,\dots,1)\oplus_{2}t$) are in $R_{1/k}$. Then $\operatorname{qPol^{no}}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$.
+Assume $k\geq 3$, and neither $t$ nor $\overline{t}$ ($(1,\dots,1)+_{2}t$) are in $R_{1/k}$. Then $\operatorname{qPol^{no}}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$.
 Proof.
 Without loss of generality, assume that $t=t^{l}$ is the tuple with $l$ 1-entries at the start, one can do this from an arbitrary $t$ by just permuting the entries.
-If $l\notin\{1,k-1\}$ (i.e., neither $t$ nor $\overline{t}$ are in $R_{1/k}$) then $(0,\dots,0)=t\oplus_{2}t$ and $(1,\dots,1)=t\oplus_{2}\overline{t}$ are not in $R_{t}$.
+If $l\notin\{1,k-1\}$ (i.e., neither $t$ nor $\overline{t}$ are in $R_{1/k}$) then $(0,\dots,0)=t+_{2}t$ and $(1,\dots,1)=t+_{2}\overline{t}$ are not in $R_{t}$.
 We first pp-define $\neq$ in $\mathbb{O}_{t}$ via
@@
 where we fill the first $l+1$ entries with $x$’s.
-This relation is preserved by every non-oracular $Q\colon\mathbb{O}_{t}^{n}\to\mathbb{O}_{t}^{n}$ since for every $\mathbf{a}\neq\mathbf{b}\in\{0,1\}^{n}$ and $x\in\{0,1\}$ we have
+This relation is preserved by every non-oracular $Q\colon\mathbb{O}_{t}^{n}\to\mathbb{O}_{t}$ since for every $\mathbf{a}\neq\mathbf{b}\in\{0,1\}^{n}$ and $x\in\{0,1\}$ we have
 $Q_{\mathbf{a},x}Q_{\mathbf{b},x}=Q_{\mathbf{a},x}\dots Q_{\mathbf{a},x}Q_{\mathbf{b},x}\dots Q_{\mathbf{b},x}=0.$
@@
 Also, there is no $y\in\{0,1\}$ such that both $(y,x_{2},\dots,x_{k})\in R_{t}$ and $y\neq x_{1}$, i.e., we know that $(\neg x_{1},x_{2},\dots,x_{k})\notin R_{t}$.
-Now let $Q\colon\mathbb{O}_{t}^{n}\to_{q}\mathbb{O}_{t}$ be a non-oracular polymorphism of $\mathbb{O}_{t}$. Then we have
+Now let $Q\colon\mathbb{O}_{t}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{t}$ be a non-oracular polymorphism of $\mathbb{O}_{t}$. Then we have
 $Q_{\mathbf{s}_{1},x_{1}}\dots Q_{\mathbf{s}_{k},x_{k}}$ $=Q_{\mathbf{s}_{1},x_{1}}\left(\sum_{y\in\{0,1\}}Q_{\mathbf{q},y}\right)Q_{\mathbf{s}_{2},x_{2}}\dots Q_{\mathbf{s}_{k},x_{k}}$
@@
 ∎
-Next, we argue that the only real cases we need to consider in the full theorem is the $\mathbb{O}_{t}$’s.
-Lemma 54.
-Every relational structure $\mathbb{A}$ with a relation $R$ that has any full binary projection has the property that $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
-Proof.
-Consider a relational structure with a relation $R$ that has a full binary projection on the components $\{i,j\}$.
-If this is the case, then in any power, any two elements appear together in some tuple of the direct power of RR.∎
+Thus $\operatorname{qPol}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$ for every $k$-tuple $t$.
+We now argue that this result covers every Boolean structure that is of interest to us.
+Specifically, we establish the following result, using similar ideas as in [33, Proposition 7.19].
 Theorem 113.
 For any relation $R\subseteq\{0,1\}^{k}$ the following statements are equivalent:
-  1. (1)
+  1. 1.
     $R$ is not invariant under majority, majority preserves every proper projection of $R$, and every binary projection of $R$ is not full.
-  2. (2)
+  2. 2.
     $R=R_{t}$ for some tuple $t$.
 Proof.
-2.
-implies 1.
+  • $2.\Rightarrow 1.$
     Note that translating a relation by a tuple does not change the size of the binary projections and does not change preservation by majority, since majority is a self-dual operation.
-It follows that it suffices to prove the implication in the case of $R_{1/k}$, for which all the properties in 1.
-are clear.
-For 1.
-implies 2.
+It follows that it suffices to prove the implication in the case of $R_{1/k}$, for which all the properties in $1.$ are clear.
+  • $1.\Rightarrow 2.$
     Suppose $R\subseteq\{0,1\}^{k}$ is not invariant under majority but majority does preserve every proper projection and no binary projection is full.
-Suppose $t\in\{0,1\}^{k}$ is a witness to the fact that $R$ is not invariant under majority, we first show that $\{x\oplus_{2}t\mid x\in R_{1/k}\}\subseteq R$.
+Suppose $t\in\{0,1\}^{k}$ is a witness to the fact that $R$ is not invariant under majority.
+We first show that $\{x+_{2}t\mid x\in R_{1/k}\}\subseteq R$.
     For $i\in[k]$, let $x\in R_{1/k}$ be the tuple with a 1 in the $i$th position.
-We know that $\pi_{[k]\setminus\{i\}}(R)$ is invariant under majority so there is an $r_{i}\in R$ with $\pi_{[k]\setminus\{i\}}(r_{i})=\pi_{[k]\setminus\{i\}}(t)$.
-Now, since $t\notin R$; we can conclude that $\pi_{\{i\}}(r_{i})\neq\pi_{\{i\}}(t)$ and therefore $x\oplus_{2}t=r_{i}$, so $x\oplus_{2}t\in R$.
-Finally, we use the non-fullness of binary projections to argue that any $r\in R$ must match $\overline{t}=(1,\dots,1)\oplus_{2}t$ in exactly one argument, placing it in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$.
+We know that $\pi_{[k]\setminus\{i\}}(R)$ is invariant under majority, so there is an $r_{i}\in R$ with $\pi_{[k]\setminus\{i\}}(r_{i})=\pi_{[k]\setminus\{i\}}(t)$.
+Now, since $t\notin R$, we can conclude that $\pi_{\{i\}}(r_{i})\neq\pi_{\{i\}}(t)$ and therefore $x+_{2}t=r_{i}$, so $x+_{2}t\in R$.
+    Finally, we use the non-fullness of binary projections to argue that any $r\in R$ must match $\overline{t}:=(1,\dots,1)+_{2}t$ in exactly one argument, placing it in $\{x+_{2}t\mid x\in R_{1/k}\}$.
     Assume that $r$ matches $\overline{t}$ in two arguments $1\leq i<j\leq k$.
-We know that $\pi_{i,j}(R)$ contains $\pi_{i,j}(t)$, $(\pi_{i}(t),\pi_{j}(\overline{t}))$ and $(\pi_{i}(\overline{t}),\pi_{j}(t))$; then $\pi_{i,j}(R)$ is full, a contradiction.
+We know that $\pi_{i,j}(R)$ contains $\pi_{i,j}(t)$, $(\pi_{i}(t),\pi_{j}(\overline{t}))$, and $(\pi_{i}(\overline{t}),\pi_{j}(t))$; then $\pi_{i,j}(R)$ is full, a contradiction.
 Additionally, $r$ cannot match $\overline{t}$ on zero arguments since then it would equal $t$, which is not in $R$.
-Thus every $r$ is in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$, completing the proof.
-∎
-We can now easily prove the general result.
+Thus every $r$ is in $\{x+_{2}t\mid x\in R_{1/k}\}$, completing the proof.∎
+Since a structure is TVF if, and only if it doesn’t have any full binary projections, Proposition 61 allows us to apply Theorem 113 and prove the following:
 Theorem 114.
-If $\mathbb{A}$ is a Boolean relational structure such that $\operatorname{Pol}(\mathbb{A})$ does not contain majority, then $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
+If $\mathbb{A}$ is a Boolean relational structure such that $\operatorname{Pol}(\mathbb{A})$ does not contain majority, then $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
 Proof.
@@
 Let $R$ be a relation of $\mathbb{A}$ that is not preserved by majority and of minimal arity.
-By Lemma 54, if a binary projection of $R$ is full then we are done, so assume that this is not the case.
+By Proposition 61, if a binary projection of $R$ is full then we are done, so assume that this is not the case.
 Then Theorem 113 implies that $R=R_{t}$ for some $k$-tuple $t$.
 Finally, we know that $k\geq 3$ since every binary relation of arity 2 is preserved under majority and so, most of the result follows from Corollary 111 and Lemma 112.
-All that remains is for us to argue that Lemma 52 also covers the case where $\overline{t}\in R_{1/k}$.
+All that remains is for us to argue that Corollary 111 also covers the case where $\overline{t}\in R_{1/k}$.
 This is since, $\mathbb{O}_{t}$ has ‘essentially’ the same polymorphisms as $\mathbb{O}_{\overline{t}}$ for any $t$.
-The formal fact we use is that $\{Q_{S}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{t}$ if and only if $\{\operatorname{id}-Q_{\overline{S}}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{\overline{t}}$.
-In particular $\mathbb{O}_{t}$ has contextual polymorphisms if and only if $\mathbb{O}_{\overline{t}}$ does.
+The formal fact we use is that $\{Q_{S}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{t}$ if, and only if, $\{\operatorname{id}-Q_{\overline{S}}\mid S\subseteq[n]\}$ is a (non-oracular) quantum polymorphism of $\mathbb{O}_{\overline{t}}$.
+In particular $\mathbb{O}_{t}$ has contextual polymorphisms if, and only if, $\mathbb{O}_{\overline{t}}$ does.
 Thus we can assume that $\overline{t}\notin R_{1/k}$, completing the proof.
 ∎
-8.2. The minimal contextual clone
-In this section, we provide a clone with contextual polymorphisms that is contained within the quantum polymorphism clone of any Boolean relational structure with contextual polymorphisms.
-In the previous section, we have seen that every boolean relational structure with contextual polymorphisms has non-full binary projections and contains majority in its (classical) polymorphism clone.
-Let $\mathbb{B}$ be the structure $\mathbb{B}:=(\{0,1\};S_{00},S_{11},S_{10})$ where $S_{ab}=\{0,1\}^{2}\setminus\{(a,b)\}$.
-In the classical setting, $\operatorname{Pol}(\mathbb{B})$ is the smallest clone that contains the majority operation.
+11.2 The minimal contextual minion
+In this section, we provide a minion with contextual polymorphisms that is contained within the quantum polymorphisms of any Boolean relational structure that has contextual polymorphisms.
+In the previous section, we have seen that every Boolean relational structure with contextual polymorphisms has non-full binary projections (is TVF) and contains majority in the minion of (classical) polymorphisms.
+Let $\operatorname{\mathbb{M}}$ be the structure $\operatorname{\mathbb{M}}:=(\{0,1\};S_{00},S_{11},S_{10})$ where $S_{ab}=\{0,1\}^{2}\setminus\{(a,b)\}$.
+In the classical setting, $\operatorname{Pol}(\operatorname{\mathbb{M}})$ is the minimal clone containing the majority operation.
 We first show that this property extends to oracular quantum polymorphisms.
-Lemma 57.
-Let $\mathbb{A}$ be a Boolean structure that admits majority as a polymorphism, and such that no binary projection of a relation of $\mathbb{A}$ is full.
-Then $\operatorname{qPol}(\mathbb{B})\subseteq\operatorname{qPol}(\mathbb{A})$.
+Theorem 115.
+Let $\mathbb{A}$ be a Boolean TVF structure that admits majority as a polymorphism. Then $\operatorname{qPol}(\operatorname{\mathbb{M}})\subseteq\operatorname{qPol}(\mathbb{A})$.
 Proof.
 Consider a relational structure $\mathbb{A}:=(\{0,1\};R)$ that is preserved by majority, then it is well known that $R$ (of arity $r$) can be pp-defined in $\mathbb{A}_{2}:=(\{0,1\};\{\pi_{i,j}(R)\mid 1\leq i,j\leq r\})$ via the conjunction of its binary projections $\bigwedge_{1\leq i,j\leq r}\pi_{\{i,j\}}(R)(x_{i},x_{j})$.
-We show that $R$ is preserved by every quantum polymorphism $Q\colon\mathbb{A}_{2}^{n}\to\mathbb{A}_{2}$.
+We show that $R$ is preserved by every quantum polymorphism $Q\colon\mathbb{A}_{2}^{n}\to_{\operatorname{q}}\mathbb{A}_{2}$.
 Assume that $(\mathbf{a}_{1},\dots,\mathbf{a}_{r})$ is in the $n$th direct power of $R$ while in $(x_{1},\dots,x_{r})$, we have indices $i,j$ such that $(x_{i},x_{j})\notin\pi_{\{i,j\}}(R)$.
@@
 where the second equality is by duplicating the $Q_{\mathbf{a}_{i},x_{i}}$ and $Q_{\mathbf{a}_{j},x_{j}}$ projectors and then using property $($$\mathbf{QH_{2}}$)) to move them to the front.
-Moreover, if $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}^{n}))$ then $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}_{2}^{n}))$ so, for any $x,x^{\prime}\in\{0,1\}$ $Q_{\mathbf{a},x},Q_{\mathbf{a}^{\prime},x^{\prime}}]=0$ by property $($$\mathbf{QH_{2}}$)).
+Moreover, if $(\mathbf{a},\mathbf{a}^{\prime})\in E(\operatorname{Gaif}(\mathbb{A}^{n}))$ then $(\mathbf{a},\mathbf{a}^{\prime})\in E(\operatorname{Gaif}(\mathbb{A}_{2}^{n}))$ so, for any $x,x^{\prime}\in\{0,1\}$ $Q_{\mathbf{a},x},Q_{\mathbf{a}^{\prime},x^{\prime}}]=0$ by property $($$\mathbf{QH_{2}}$)).
 $Q$ is then an oracular quantum polymorphism of $\mathbb{A}$ and so $\operatorname{qPol}(\mathbb{A}_{2})\subseteq\operatorname{qPol}(\mathbb{A})$.
-By assumption, no binary projection of a relation of $\mathbb{A}$ is full, thus, each relation $R$ in $\mathbb{A}_{2}$ can be defined as the (non-empty!) conjunction $\bigwedge_{(c,d)\not\in R}S_{cd}(x,y)$, where we use the shorthand $S_{01}(x,y)$ for the formula $S_{10}(y,x)$.
-We show that each relation $R$ of $\mathbb{A}_{2}$ is preserved by the quantum polymorphisms of $\mathbb{B}$.
-Let $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ be a quantum polymorphism, $(\mathbf{a},\mathbf{b})$ be in the $n$th direct power of $R$ and $(c,d)\notin R$.
-We then have $Q_{\mathbf{a},c}Q_{\mathbf{b},d}=0$ by $($$\mathbf{QH_{1}}$)) since $(c,d)\not\in S_{cd}$ and $(\mathbf{a},\mathbf{b})$ is in the relation $S_{cd}$ in the $n$th power of $\mathbb{B}$.
+By assumption, no binary projection of a relation of $\mathbb{A}$ is full, thus, each relation $R$ in $\mathbb{A}_{2}$ can be defined as the (non-empty) conjunction $\bigwedge_{(c,d)\not\in R}S_{cd}(x,y)$, where we use the shorthand $S_{01}(x,y)$ for the formula $S_{10}(y,x)$.
+We show that each relation $R$ of $\mathbb{A}_{2}$ is preserved by the quantum polymorphisms of $\operatorname{\mathbb{M}}$.
+Let $Q\colon\operatorname{\mathbb{M}}^{n}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ be a quantum polymorphism, $(\mathbf{a},\mathbf{b})$ be in the $n$th direct power of $R$ and $(c,d)\notin R$.
+We then have $Q_{\mathbf{a},c}Q_{\mathbf{b},d}=0$ by $($$\mathbf{QH_{1}}$)) since $(c,d)\not\in S_{cd}$ and $(\mathbf{a},\mathbf{b})$ is in the relation $S_{cd}$ in the $n$th power of $\operatorname{\mathbb{M}}$.
 Thus, $($$\mathbf{QH_{1}}$)) holds.
-Finally, if $(\mathbf{a},\mathbf{b})$ is a pair in a relation $R$ in the $n$th power of $\mathbb{A}_{2}$, then it is necessarily a pair in one of the relations $S_{cd}$ in the $n$th power of $\mathbb{B}$.
+Finally, if $(\mathbf{a},\mathbf{b})$ is a pair in a relation $R$ in the $n$th power of $\mathbb{A}_{2}$, then it is necessarily a pair in one of the relations $S_{cd}$ in the $n$th power of $\operatorname{\mathbb{M}}$.
 Thus, $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute, so that $($$\mathbf{QH_{2}}$)) holds.
 ∎
-Now we show that this clone actually contains some contextual polymorphisms.
+Now we show that the minion defined here, actually contains some contextual polymorphisms.
 Theorem 116.
-For $n\leq 3$, every quantum polymorphism $Q\colon\mathbb{B}^{n}\to\mathbb{B}$ is non-contextual. For all $n\geq 4$, $\mathbb{B}$ has contextual quantum polymorphisms of arity $n$.
+For $n\leq 3$, every quantum polymorphism $Q\colon\operatorname{\mathbb{M}}^{n}\to\operatorname{\mathbb{M}}$ is non-contextual.
+For all $n\geq 4$, $\operatorname{\mathbb{M}}$ has contextual quantum polymorphisms of arity $n$.
 Proof.
-We start by proving that every quantum polymorphism $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ is non-contextual for $n\leq 3$.
-For this, it suffices by $($$\mathbf{QH_{2}}$)) to prove that for every pair $S,T$ of subsets of $\{1,\dots,n\}$, either $(S,T)$ or $(T,S)$ belongs to one of the relations of $\mathbb{B}^{n}$.
-If $(S,T)$ does not belong to the relation $S_{00}$ in $\mathbb{B}^{n}$, then there exists $i\in\{1,\dots,n\}$ such that $i\in\overline{S}\cap\overline{T}$.
-Similarly, for it to not belong to the relation $S_{11}$ in $\mathbb{B}^{n}$, there must exist $j\in\{1,\dots,n\}$ such that $j\in S\cap T$.
+We start by proving that every quantum polymorphism $Q\colon\operatorname{\mathbb{M}}^{n}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ is non-contextual for $n\leq 3$.
+For this, it suffices by $($$\mathbf{QH_{2}}$)) to prove that for every pair $S,T$ of subsets of $\{1,\dots,n\}$, either $(S,T)$ or $(T,S)$ belongs to one of the relations of $\operatorname{\mathbb{M}}^{n}$.
+If $(S,T)$ does not belong to the relation $S_{00}$ in $\operatorname{\mathbb{M}}^{n}$, then there exists $i\in\{1,\dots,n\}$ such that $i\in\overline{S}\cap\overline{T}$.
+Similarly, for it to not belong to the relation $S_{11}$ in $\operatorname{\mathbb{M}}^{n}$, there must exist $j\in\{1,\dots,n\}$ such that $j\in S\cap T$.
 If $n\leq 3$, there remains at one most other element $k$ in $\{1,\dots,n\}\setminus\{i,j\}$.
-If $k$ is not in $S\cap T$ nor in $\overline{S}\cap\overline{T}$, then it is the case that $(S,T)$ or $(T,S)$ belongs to the relation $S_{10}$ of $\mathbb{B}^{n}$, so we are done.
-For the $n\geq 4$ case we define a concrete quantum polymorphism for $n=4$. We provide Figure 3 as a visual aid.
-0000100001000010000111001010100101100101001111101101101101111111
-Figure 3.
-Simplified illustration of $\mathbb{B}^{4}$.
-Black lines indicate the poset $S_{10}$.
-The blue (resp.
-red) dotted lines show parts of $S_{00}$ (resp.
-$S_{11}$); this relation additionally relates $(0,0,0,0)$ (resp.
-$(1,1,1,1)$) to everything and relates two nodes when they are a bit-flips of each other (see coloring).
-We define $Q\colon\mathbb{B}^{4}\to_{q}\mathbb{B}$ by listing the projectors.
+If $k$ is not in $S\cap T$ nor in $\overline{S}\cap\overline{T}$, then it is the case that $(S,T)$ or $(T,S)$ belongs to the relation $S_{10}$ of $\operatorname{\mathbb{M}}^{n}$, so we are done.
+We now build a concrete class of contextual polymorphisms of arity $n=4$.
+We provide Figure 2 as a visual aid in this example.
+The figure is a simplified presentation of the relatedness of tuples in $\{0,1\}^{4}$ with the arrows indicating the poset $S_{10}$ and the colouring emphasizing that $S_{11}$ (and $S_{00}$) relate all opposite tuples.
+In particular, the only unrelated tuples are those $s,t$ such that $|S|=|T|=2$ and $S\neq\overline{T}$.
+We define $Q\colon\operatorname{\mathbb{M}}^{4}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ by listing the projectors.
 Let $T\subseteq[4]$.
@@
 Note that if at least two of the matrices $A,B,C$ do not commute, then $Q$ is contextual.
-It remains to prove that $Q$ preserves the relations of $\mathbb{B}^{4}$.
-Assume we have $S,T\subseteq[4]$ with $(s,t)$ in the fourth power of $S_{11}$ and $x,y\in\{0,1\}$ with $(x,y)\notin S_{11}$.
-If $S\neq\overline{T}$, we are in one of 2 cases; either $S$ or $T$ is the empty set or $|S|=|T|=1$.
-In both cases at least $Q_{S}$ or $Q_{T}$ equals 0 and we are done immediately.
-If $S=\overline{T}$, then $Q_{S}Q_{T}=Q_{S}Q_{\overline{S}}=Q_{S}(\operatorname{id}-Q_{S})=0$ and we are done.
+It remains to prove that $Q$ preserves the relations of $\operatorname{\mathbb{M}}^{4}$.
+Assume we have $S,T\subseteq[4]$ with $(s,t)$ in the fourth power of $S_{11}$ and consider $Q_{S}Q_{T}$ (since $(1,1)$ is the only Boolean pair not in $S_{11}$).
+$S$ and $T$ are related if, and only if, $S\cap T=\emptyset$; hence, we have two cases: Either $|S|=|T|=2$ and $S=\overline{T}$, or one of $S,T$ has cardinality less than $2$.
+If $S\neq\overline{T}$, then $|S|\leq 1$ or $|T|\leq 1$ in this case at least $Q_{S}$ or $Q_{T}$ equals 0 and we are done immediately.
+If $S=\overline{T}$, then $Q_{S}Q_{T}=Q_{S}Q_{\smash{\overline{S}}}=Q_{S}(\operatorname{id}-Q_{S})=0$ and we are done.
 The case of $S_{00}$ is symmetric.
-Now assume we have $S,T\subseteq[4]$ such that the corresponding pair of tuples $(s,t)$ in the fourth power of $S_{10}$.
-If both $|S|$ and $|T|$ do not equal $2$ we know $Q_{S},Q_{T},(\operatorname{id}-Q_{S}),(\operatorname{id}-Q_{T})\in\{0,\operatorname{id}\}$ and we immediately get that $Q_{S},{\operatorname{id}-Q_{T}}]=0$.
-The final case is when one of $|S|$ or $|T|$ equal $2$, here we get that either
+Now assume we have $S,T\subseteq[4]$ such that the corresponding pair of tuples $(s,t)$ in the fourth power of $S_{10}$ ($S\subseteq T$).
+If both $|S|$ and $|T|$ do not equal $2$, we know that all projectors $Q_{S}$, $Q_{T}$, $(\operatorname{id}-Q_{S})$, and $(\operatorname{id}-Q_{T})$ are either 0 or $\operatorname{id}$; thus, we immediately find that $Q_{S},{\operatorname{id}-Q_{T}}]=0$.
+The final case is when at least one of $|S|$ or $|T|$ equals $2$.
+In this case, we deduce that either
 $Q_{S}(\operatorname{id}-Q_{T})=0(\operatorname{id}-Q_{T})=0$
-or,
+when $|T|=2$ (and $S\neq T$) or,
 $Q_{S}(\operatorname{id}-Q_{T})=Q_{S}(\operatorname{id}-\operatorname{id})=0$
-showing preservation of the relations. ∎
-9. Acknowledgments
+when $|S|=2$ (and $S\neq T$). Thus all relations are preserved, and $Q\colon\operatorname{\mathbb{M}}^{4}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ is a quantum polymorphism.
+Note that one can easily extend $Q$ to a contextual polymorphism of arbitrary arity $\geq 4$ by adding dummy variables.
+For example, define $\tau\colon[4]\to[n]$ by $\tau(i):=i$ for $i\in[4]$.
+By Proposition 26,
+$Q_{/\tau}=Q\bullet\bigotimes_{j\in[4]}\pi_{\tau(j)}^{n}\colon M^{n}\to_{\operatorname{q}}M$
+is in $\operatorname{qPol}(\operatorname{\mathbb{M}})$.
+We just need to check that $Q_{/\tau}$ is non-contextual, but this is immediate from the definition of $Q$, since e.g.
+$(Q_{/\tau})_{\{1,2\}}=Q_{\{1,2\}}=A$, $(Q_{/\tau})_{\{1,3\}}=Q_{\{1,3\}}=B$, and $A,B]\neq 0$.
+∎
+We can now conclude by proving the main result on this section, giving a full classification of the existence of commutativity gadgets for Boolean structures.
+As remarked in Section 2.6, the $(ii)\Rightarrow(i)$ implication was first proved in [33], which described explicit commutativity gadgets for $\mathbb{A}$ having no majority and for any $\mathbb{A}$ that is not TVF.
+Proof of Theorem 11.
+  • $(ii)\Rightarrow(i)$
+    If $\mathbb{A}$ has no majority polymorphism, then it admits a commutativity gadget by Theorem 114 and Theorem 6. Otherwise, the result follows from Proposition 61 and Theorem 6.
+  • $(i)\Rightarrow(ii)$
+    By contrapositive, if $\mathbb{A}$ admits a majority polymorphism and is TVF, then it holds that $\operatorname{qPol}(\operatorname{\mathbb{M}})\subseteq\operatorname{qPol}(\mathbb{A})$ by Theorem 115.
+Thus $\mathbb{A}$ has at least one contextual polymorphism by Theorem 116.
+So, it has no commutativity gadget.∎
+12 Acknowledgments
 The first and last author are thankful for the attentive listening and comments by Peter Zeman, to whom they explained the definition of quantum polymorphisms and the main results in this paper concerning the characterisation of the existence of commutativity gadgets on the occasion of the CSP World Congress in September 2025.